Shishi

Kerberos 5 implementation for the GNU system for version 1.0.3, 7 August 2022

Simon Josefsson

This manual is last updated 7 August 2022 for version 1.0.3 of Shishi. Copyright \begin{matrix}c \\ ◯ \end{matrix} 2002–2022 Simon Josefsson.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Getting Started. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Features and Status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.4 Cryptographic Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.5 Supported Platforms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.6 Getting help. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.7 Downloading and Installing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.8 Bug Reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 1.9 Contributing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

2 User Manual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

2.1 Proxiable and Proxy Tickets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.2 Forwardable and Forwarded Tickets. . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

3 Administration Manual . . . . . . . . . . . . . . . . . . . . . . . . 18

3.1 Introduction to Shisa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.2 Configuring Shisa. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.3 Using Shisa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.4 Starting Shishid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.5 Configuring DNS for KDC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

3.5.1 DNS vs. Kerberos - Case Sensitivity of Realm Names . . . . . 25 3.5.2 Overview - KDC location information . . . . . . . . . . . . . . . . . . . . . 25 3.5.3 Example - KDC location information . . . . . . . . . . . . . . . . . . . . . 26 3.5.4 Security considerations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

3.6 Kerberos via TLS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6.1 Setting up TLS resume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6.2 Setting up Anonymous TLS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.6.3 Setting up X.509 authenticated TLS . . . . . . . . . . . . . . . . . . . . . . 28

3.6.3.1 Create a Kerberos Certificate Authority . . . . . . . . . . . . . . 28 3.6.3.2 Create a Kerberos KDC Certificate . . . . . . . . . . . . . . . . . . 29 3.6.3.3 Create a Kerberos Client Certificate. . . . . . . . . . . . . . . . . . 30 3.6.3.4 Starting KDC with X.509 authentication support. . . . . 31

3.7 Multiple servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 3.8 Developer information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

4 Reference Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

4.1 Environmental Assumptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.2 Glossary of terms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.3 Realm and Principal Naming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

4.3.1 Realm Names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.3.2 Principal Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.3.2.1 Name of server principals . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.3.2.2 Name of the TGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

4.3.3 Choosing a principal with which to communicate . . . . . . . . . . 40 4.3.4 Principal Name Form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

4.4 Shishi Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.4.1 ‘default-realm’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.4.2 ‘default-principal’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.4.3 ‘client-kdc-etypes’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

4.4.4 ‘verbose’, ‘verbose-asn1’, ‘verbose-noise’,
  ‘verbose-crypto’, ‘verbose-crypto-noise’ . . . . . . . . . . . . . . . . . . . 42

4.4.5 ‘realm-kdc’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.4.6 ‘server-realm’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.4.7 ‘kdc-timeout’, ‘kdc-retries’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.4.8 ‘stringprocess’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.4.9 ‘ticket-life’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.4.10 ‘renew-life’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

4.5 Shisa Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

4.5.1 ‘db’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 4.6 Parameters for shishi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 4.7 Parameters for shishid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 4.8 Parameters for shisa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 4.9 Environment variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.10 Date input formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

4.10.1 General date syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.10.2 Calendar date items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 4.10.3 Time of day items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 4.10.4 Time zone items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 4.10.5 Combined date and time of day items . . . . . . . . . . . . . . . . . . . . 52 4.10.6 Day of week items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 4.10.7 Relative items in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 4.10.8 Pure numbers in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.10.9 Seconds since the Epoch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.10.10 Specifying time zone rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 4.10.11 Authors of parse_datetime . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

5 Programming Manual . . . . . . . . . . . . . . . . . . . . . . . . . 57

5.1 Preparation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.1.1 Header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.1.2 Initialization. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.1.3 Version Check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.1.4 Building the source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.1.5 Autoconf tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

5.1.5.1 Autoconf test via ‘pkg-config’ . . . . . . . . . . . . . . . . . . . . . . 58 5.1.5.2 Standalone Autoconf test using Libtool . . . . . . . . . . . . . . 59 5.1.5.3 Standalone Autoconf test . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

5.2 Initialization Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 5.3 Ticket Set Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 5.4 AP-REQ and AP-REP Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 5.5 SAFE and PRIV Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 5.6 Ticket Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 5.7 AS Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 5.8 TGS Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 5.9 Ticket (ASN.1) Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 5.10 AS/TGS Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 5.11 Authenticator Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 5.12 KRB-ERROR Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 5.13 Cryptographic Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 5.14 X.509 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 5.15 Utility Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 5.16 ASN.1 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 5.17 Error Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

5.17.1 Error Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 5.17.2 Error Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

5.18 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 5.19 Kerberos Database Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 5.20 Generic Security Service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233

6 Acknowledgements. . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

Appendix A Criticism of Kerberos . . . . . . . . . . . . 235

Appendix B Protocol Extensions . . . . . . . . . . . . . . 236

B.1 STARTTLS protected KDC exchanges . . . . . . . . . . . . . . . . . . . . . . . 236 B.1.1 TCP/IP transport with TLS upgrade (STARTTLS) . . . . . 236 B.1.2 Extensible typed hole based on reserved high bit. . . . . . . . . 237 B.1.3 STARTTLS requested by client (extension mode 1) . . . . . . 237 B.1.4 STARTTLS request accepted by server (extension mode 2) . . 237 B.1.5 Proceeding after successful TLS negotiation . . . . . . . . . . . . . 237 B.1.6 Proceeding after failed TLS negotiation. . . . . . . . . . . . . . . . . . 238 B.1.7 Interaction with KDC addresses in DNS . . . . . . . . . . . . . . . . . 238 B.1.8 Using TLS authentication logic in Kerberos. . . . . . . . . . . . . . 238 B.1.9 Security considerations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238

B.2 Telnet encryption with AES-CCM. . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 B.2.1 Command Names and Codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 B.2.2 Command Meanings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 B.2.3 Implementation Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 B.2.4 Integration with the AUTHENTICATION telnet option . . 240 B.2.5 Security Considerations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241

B.2.5.1 Telnet Encryption Protocol Security Considerations . . 241 B.2.5.2 AES-CCM Security Considerations . . . . . . . . . . . . . . . . . 241

B.2.6 Acknowledgments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 B.3 Kerberized rsh and rlogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242

B.3.1 Establish connection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 B.3.2 Kerberos identification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 B.3.3 Kerberos authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 B.3.4 Extended authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 B.3.5 Window size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 B.3.6 End of authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 B.3.7 Encryption. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 B.3.8 KCMDV0.3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 B.3.9 MIT/Heimdal authorization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246

B.4 Key as initialization vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 B.5 The Keytab Binary File Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 B.6 The Credential Cache Binary File Format . . . . . . . . . . . . . . . . . . . . 250

Appendix C Copying Information . . . . . . . . . . . . . 253

C.1 GNU Free Documentation License. . . . . . . . . . . . . . . . . . . . . . . . . . . . 253

Function and Data Index . . . . . . . . . . . . . . . . . . . . . . . . 261

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268

1 Introduction

Shishi implements the Kerberos network authentication system. Shishi can be used to authenticate users in distributed systems, and is most often used via GSS-API for SSH or via SASL for IMAP/POP3.

Shishi contains a library (’libshishi’) that can be used by application developers to add support for Kerberos 5. There is also a command line utility (’shishi’) that is used by users to acquire and manage tickets. The server side, a Key Distribution Center, is implemented by ’shishid’. Of course, a manual documenting usage aspects as well as the programming API is included.

Shishi currently supports AS/TGS exchanges for acquiring tickets, pre-authentication, the AP exchange for performing client and server authentication, and SAFE/PRIV for integrity/privacy protected application data exchanges.

Shishi is internationalized; error and status messages can be translated into the users’ language; user name and passwords can be converted into any available character set (normally including ISO-8859-1 and UTF-8) and also be processed using an experimental Stringprep profile.

Most, if not all, of the widely used encryption and checksum types are supported, such as 3DES, AES, ARCFOUR and HMAC-SHA1.

Shishi is developed for the GNU/Linux system but works on many platforms including most major Unix and Windows systems.

Shishi is free software licensed under the GNU General Public License version 3.0 or later.

1.1 Getting Started

This manual documents the Shishi application and library programming interface. All commands, functions and data types provided by Shishi are explained.

The reader is assumed to possess basic familiarity with network security and the Kerberos 5 security system.

This manual can be used in several ways. If read from the beginning to the end, it gives a good introduction into the library and how it can be used in an application. Forward references are included where necessary. Later on, the manual can be used as a reference manual to get just the information needed about any particular interface of the library. Experienced programmers might want to start looking at the examples at the end of the manual, and then only read up on those parts of the interface which are unclear.

1.2 Features and Status

Shishi might have a couple of advantages over other packages doing a similar job.

It’s Free Software Anybody can use, modify, and redistribute it under the terms of the GNU General Public License version 3.0 or later.

It’s thread-safe The library uses no global variables.

It’s internationalized It handles non-ASCII username and passwords, and user visible strings used in the library (error messages) can be translated into the users’ language.

It’s portable It should work on all Unix like operating systems, including Windows.

Shishi is far from feature complete, it is not even a full RFC 4120 implementation yet. However, some basic functionality is implemented. A few implemented feature are mentioned below.

• Initial authentication (AS) from raw key or password. This step is typically used to acquire a ticket granting ticket and, less commonly, a server ticket.

• Subsequent authentication (TGS). This step is typically used to acquire a server ticket, by authenticating yourself using the ticket granting ticket.

• Client-Server authentication (AP). This step is used by clients and servers to prove to each other who they are, using negotiated tickets.

• Integrity protected communication (SAFE). This step is used by clients and servers to exchange integrity protected data with each other. The key is typically agreed on using the Client-Server authentication step.

• Ticket cache, supporting multiple principals and realms. As tickets have a life time of typically several hours, they are managed in disk files. There can be multiple ticket caches, and each ticket cache can store tickets for multiple clients (users), servers, encryption types, etc. Functionality is provided for locating the proper ticket for every use.

• Most standard cryptographic primitives. The believed most secure algorithms are supported (see Section 1.4 [Cryptographic Overview], page 5).

• Telnet client and server. This is used to remotely login to other machines, after authenticating yourself with a ticket.

• PAM module. This is used to login locally on a machine.

• KDC addresses located using DNS SRV RRs.

• Modularized low-level crypto interface. Currently Gnulib and Libgcrypt are supported. If you wish to add support for another low-level cryptographic library, you only have to implement a few APIs for DES, AES, MD5, SHA1, HMAC, etc. Look at

gl/gc-gnulib.c or gl/gc-libgcrypt.c as a starting pointer.

The following table summarize what the current objectives are (i.e., the todo list) and an estimate on how long it will take to implement the feature, including some reasonable startup-time to get familiar with Shishi in general. If you like to start working on anything, please let me know so work duplication can be avoided.

• Parse /etc/krb5.keytab to extract keys to use for telnetd etc (week)

• Cross-realm support (week).

• PKINIT (use libksba, weeks)

• Finish GSSAPI support via GSSLib (weeks) Shishi will not support GSSLib natively, but a separate project “GSSLib” is under way to produce a generic GSS implementation, and it will use Shishi to implement the Kerberos 5 mechanism.

• Port to cyclone (cyclone need to mature first)

• Modularize ASN.1 library so it can be replaced (days). Almost done, all ASN.1 functionality is found in lib/asn1.c, although the interface is rather libtasn1 centric.

• KDC (initiated, weeks)

• LDAP backend for Shisa.

• Set/Change password protocol (weeks?)

• Port applications to use Shishi (indefinite)

• Finish server-realm stuff

• Improve documentation

• Improve internationalization

• Add AP-REQ replay cache (week).

• Study benefits by introducing a PA-TGS-REP. This would provide mutual authentication of the KDC in a way that is easier to analyze. Currently the mutual authentication property is only implicit from successful decryption of the KDC-REP and the 4 byte

    nonce.

• GUI applet for managing tickets. This is supported via the ticket-applet, of which a Shishi port is published on the Shishi home page.

• Authorization library (months?) The shishi authorized p() is not a good solution, better would be to have a generic and flexible authorization library. Possibly based on S-EXP’s in tickets? Should support non-Kerberos uses as well, of course.

• Proof read manual.

• X.500 support, including DOMAIN-X500-COMPRESS. I will accept patches that implement this, if it causes minimal changes to the current code.

1.3 Overview

This section describes RFC 1510 from a protocol point of view^{1}.

Kerberos provides a means of verifying the identities of principals, (e.g., a workstation user or a network server) on an open (unprotected) network. This is accomplished without relying on authentication by the host operating system, without basing trust on host addresses, without requiring physical security of all the hosts on the network, and under the assumption that packets traveling along the network can be read, modified, and inserted at will. (Note, however, that many applications use Kerberos’ functions only upon the initiation of a stream-based network connection, and assume the absence of any "hijackers" who might subvert such a connection. Such use implicitly trusts the host addresses involved.) Kerberos performs authentication under these conditions as a trusted third- party authentication service by using conventional cryptography, i.e., shared secret key. (shared secret key - Secret and private are often used interchangeably in the literature. In our usage, it takes two (or more) to share a secret, thus a shared DES key is a secret key. Something is only private when no one but its owner knows it. Thus, in public key cryptosystems, one has a public and a private key.)

1 The text is a lightly adapted version of the introduction section from RFC 1510 by J. Kohl and C.

Neuman, September 1993, copyright likely owned by the RFC 1510 authors or some contributor.

The authentication process proceeds as follows: A client sends a request to the authentication server (AS) requesting "credentials" for a given server. The AS responds with these credentials, encrypted in the client’s key. The credentials consist of 1) a "ticket" for the server and 2) a temporary encryption key (often called a "session key"). The client transmits the ticket (which contains the client’s identity and a copy of the session key, all encrypted in the server’s key) to the server. The session key (now shared by the client and server) is used to authenticate the client, and may optionally be used to authenticate the server. It may also be used to encrypt further communication between the two parties or to exchange a separate sub-session key to be used to encrypt further communication.

The implementation consists of one or more authentication servers running on physically secure hosts. The authentication servers maintain a database of principals (i.e., users and servers) and their secret keys. Code libraries provide encryption and implement the Kerberos protocol. In order to add authentication to its transactions, a typical network application adds one or two calls to the Kerberos library, which results in the transmission of the necessary messages to achieve authentication.

The Kerberos protocol consists of several sub-protocols (or exchanges). There are two methods by which a client can ask a Kerberos server for credentials. In the first approach, the client sends a cleartext request for a ticket for the desired server to the AS. The reply is sent encrypted in the client’s secret key. Usually this request is for a ticket-granting ticket (TGT) which can later be used with the ticket-granting server (TGS). In the second method, the client sends a request to the TGS. The client sends the TGT to the TGS in the same manner as if it were contacting any other application server which requires Kerberos credentials. The reply is encrypted in the session key from the TGT.

Once obtained, credentials may be used to verify the identity of the principals in a transaction, to ensure the integrity of messages exchanged between them, or to preserve privacy of the messages. The application is free to choose whatever protection may be necessary.

To verify the identities of the principals in a transaction, the client transmits the ticket to the server. Since the ticket is sent "in the clear" (parts of it are encrypted, but this encryption doesn’t thwart replay) and might be intercepted and reused by an attacker, additional information is sent to prove that the message was originated by the principal to whom the ticket was issued. This information (called the authenticator) is encrypted in the session key, and includes a timestamp. The timestamp proves that the message was recently generated and is not a replay. Encrypting the authenticator in the session key proves that it was generated by a party possessing the session key. Since no one except the requesting principal and the server know the session key (it is never sent over the network in the clear) this guarantees the identity of the client.

The integrity of the messages exchanged between principals can also be guaranteed using the session key (passed in the ticket and contained in the credentials). This approach provides detection of both replay attacks and message stream modification attacks. It is accomplished by generating and transmitting a collision-proof checksum (elsewhere called a hash or digest function) of the client’s message, keyed with the session key. Privacy and integrity of the messages exchanged between principals can be secured by encrypting the data to be passed using the session key passed in the ticket, and contained in the credentials.

1.4 Cryptographic Overview

Shishi implements several of the standard cryptographic primitives. In this section we give the names of the supported encryption suites, and some notes about them, and their associated checksum suite.

Statements such as “it is weak” should be read as meaning that there is no credible security analysis of the mechanism available, and/or that should an attack be published publicly, few people would likely be surprised. Also keep in mind that the key size mentioned is the actual key size, not the effective key space as far as a brute force attack is concerned.

As you may infer from the descriptions, there is currently no encryption algorithm and only one checksum algorithm that inspire great confidence in its design. Hopefully this will change over time.

NULL

NULL is a dummy encryption suite for debugging. Encryption and decryption are identity functions. No integrity protection. It is weak. It is associated with the NULL checksum.

arcfour-hmac
arcfour-hmac-exp

arcfour-hmac-* are a proprietary stream cipher with 56 bit (arcfour-hmac-exp) or 128 bit (arcfour-hmac) keys, used in a proprietary way described in

an expired IETF draft draft-brezak-win2k-krb-rc4-hmac-04.txt. Deriving

keys from passwords is supported, and is done by computing a message digest (MD4) of a 16-bit Unicode representation of the ASCII password, with no salt. Data is integrity protected with a keyed hash (HMAC-MD5), where the key is derived from the base key in a creative way. It is weak. It is associated with

the arcfour-hmac-md5 checksum.
des-cbc-none

des-cbc-none is DES encryption and decryption with 56 bit keys and 8 byte blocks in CBC mode, using a zero IV. The keys can be derived from passwords by an obscure application specific algorithm. It is weak, because it offers no integrity protection. This is typically only used by RFC 1964 GSS-API implementations (which try to protect integrity using an ad-hoc solution). It is associated with the NULL checksum.

des-cbc-crc

des-cbc-crc is DES encryption and decryption with 56 bit keys and 8 byte blocks in CBC mode, using the key as IV (see Section B.4 [Key as initialization vector], page 247). The keys can be derived from passwords by an obscure application specific algorithm. Data is integrity protected with an unkeyed but encrypted CRC32-like checksum. It is weak. It is associated with the rsa-md5-des checksum.

des-cbc-md4

des-cbc-md4 is DES encryption and decryption with 56 bit keys and 8 byte blocks in CBC mode, using a zero IV. The keys can be derived from passwords by an obscure application specific algorithm. Data is integrity protected with an unkeyed but encrypted MD4 hash. It is weak. It is associated with the

rsa-md4-des checksum.
des-cbc-md5

des-cbc-md5 is DES encryption and decryption with 56 bit keys and 8 byte blocks in CBC mode, using a zero IV. The keys can be derived from passwords by an obscure application specific algorithm. Data is integrity protected with an unkeyed but encrypted MD5 hash. It is weak. It is associated with the rsa-md5-des checksum. This is the strongest RFC 1510 interoperable encryption mechanism.

des3-cbc-none

des3-cbc-none is DES encryption and decryption with three 56 bit keys (effective key size 112 bits) and 8 byte blocks in CBC mode. The keys can be derived from passwords by the same algorithm as des3-cbc-sha1-kd. It is weak, because it offers no integrity protection. This is typically only used by GSS-API implementations (which try to protect integrity using an ad-hoc solution) for interoperability with some existing Kerberos GSS implementations. It is associated with the NULL checksum.

des3-cbc-sha1-kd

des3-cbc-sha1-kd is DES encryption and decryption with three 56 bit keys (effective key size 112 bits) and 8 byte blocks in CBC mode. The keys can be derived from passwords by a algorithm based on the paper "A Better Key Schedule For DES-like Ciphers"^{2} by Uri Blumenthal and Steven M. Bellovin (it is not clear if the algorithm, and the way it is used, is used by any other protocols, although it seems unlikely). Data is integrity protected with a keyed SHA1 hash in HMAC mode. It has no security proof, but is assumed to provide adequate security in the sense that knowledge on how to crack it is not known to the public. Note that the key derivation function is not widely used outside of Kerberos, hence not widely studied. It is associated with the hmac-sha1-des3-kd checksum.

aes128-cts-hmac-sha1-96
aes256-cts-hmac-sha1-96
aes128-cts-hmac-sha1-96 and aes256-cts-hmac-sha1-96 is AES encryption

and decryption with 128 bit and 256 bit key, respectively, and 16 byte blocks in CBC mode with Cipher Text Stealing. Cipher Text Stealing means data length of encrypted data is preserved (pure CBC add up to 7 pad characters). The keys can be derived from passwords with RSA Laboratories PKCS#5 Password Based Key Derivation Function 2^{3}, which is allegedly provably secure in a random oracle model. Data is integrity protected with a keyed SHA1 hash, in HMAC mode, truncated to 96 bits. There is no security proof, but the schemes are assumed to provide adequate security in the sense that knowledge on how to crack them is not known to the public. Note that AES has yet to receive the test of time, and the AES cipher encryption mode (CBC with Ciphertext Stealing, and a non-standard IV output) is not widely standardized (hence not widely studied). It is associated with the hmac-sha1-96-aes128 and hmac-sha1-96-aes256 checksums, respectively.

2 http://www.research.att.com/~smb/papers/ides.pdf 3 http://www.rsasecurity.com/rsalabs/pkcs/pkcs-5/

The protocol do not include any way to negotiate which checksum mechanisms to use, so in most cases the associated checksum will be used. However, checksum mechanisms can be used with other encryption mechanisms, as long as they are compatible in terms of key format etc. Here are the names of the supported checksum mechanisms, with some notes on their status and the compatible encryption mechanisms. They are ordered by increased security as perceived by the author.

NULL

NULL is a dummy checksum suite for debugging. It provides no integrity. It is weak. It is compatible with the NULL encryption mechanism.

arcfour-hmac-md5

arcfour-hmac-md5 is a keyed HMAC-MD5 checksum computed on a MD5 message digest, in turn computed on a four byte message type indicator concatenated with the application data. (The arcfour designation is thus somewhat misleading, but since this checksum mechanism is described in the same document as the arcfour encryption mechanisms, it is not a completely unnatural designation.) It is weak. It is compatible with all encryption mechanisms.

rsa-md4

rsa-md4 is a unkeyed MD4 hash computed over the message. It is weak, because it is unkeyed. However applications can, with care, use it non-weak ways (e.g., by including the hash in other messages that are protected by other means). It is compatible with all encryption mechanisms.

rsa-md4-des

rsa-md4-des is a DES CBC encryption of one block of random data and a unkeyed MD4 hash computed over the random data and the message to integrity protect. The key used is derived from the base protocol key by XOR with a constant. It is weak. It is compatible with the des-cbc-crc, des-cbc-md4, des-cbc-md5 encryption mechanisms.

rsa-md5

rsa-md5 is a unkeyed MD5 hash computed over the message. It is weak, because it is unkeyed. However applications can, with care, use it non-weak ways (e.g., by including the hash in other messages that are protected by other means). It is compatible with all encryption mechanisms.

rsa-md5-des

rsa-md5-des is a DES CBC encryption of one block of random data and a unkeyed MD5 hash computed over the random data and the message to integrity protect. The key used is derived from the base protocol key by XOR with a constant. It is weak. It is compatible with the des-cbc-crc, des-cbc-md4, des-cbc-md5 encryption mechanisms.

hmac-sha1-des3-kd

hmac-sha1-des3-kd is a keyed SHA1 hash in HMAC mode computed over the message. The key is derived from the base protocol by the simplified key derivation function (similar to the password key derivation functions of des3-cbc-sha1-kd, which does not appear to be widely used outside Kerberos and hence not widely studied). It has no security proof, but is assumed to provide good security. The weakest part is likely the proprietary key derivation function. It is compatible with the des3-cbc-sha1-kd encryption mechanism.

hmac-sha1-96-aes128
hmac-sha1-96-aes256

hmac-sha1-96-aes* are keyed SHA1 hashes in HMAC mode computed over the message and then truncated to 96 bits. The key is derived from the base protocol by the simplified key derivation function (similar to the password key derivation functions of aes*-cts-hmac-sha1-96, i.e., PKCS#5). It has no security proof, but is assumed to provide good security. It is compatible with the aes*-cts-hmac-sha1-96 encryption mechanisms.

Several of the cipher suites have long names that can be hard to memorize. For your convenience, the following short-hand aliases exists. They can be used wherever the full encryption names are used.

arcfour
          Alias for arcfour-hmac.
des-crc
          Alias for des-cbc-crc.
des-md4
          Alias for des-cbc-md4.
des-md5
des
          Alias for des-cbc-md5.
des3
3des
          Alias for des3-cbc-sha1-kd.
aes128
          Alias for aes128-cts-hmac-sha1-96.
aes
aes256
          Alias for aes256-cts-hmac-sha1-96.

1.5 Supported Platforms

Shishi has at some point in time been tested on the following platforms. Online build reports for each platforms and Shishi version is available at http://autobuild.josefsson.org/

shishi/.

1. Debian GNU/Linux 3.0 (Woody) GCC 2.95.4 and GNU Make. This is the main development plat-

    form.  alphaev67-unknown-linux-gnu,  alphaev6-unknown-linux-gnu,

arm-unknown-linux-gnu, armv4l-unknown-linux-gnu, hppa-unknown-linux-gnu,
hppa64-unknown-linux-gnu, i686-pc-linux-gnu, ia64-unknown-linux-gnu,
m68k-unknown-linux-gnu, mips-unknown-linux-gnu, mipsel-unknown-linux-gnu,
powerpc-unknown-linux-gnu, s390-ibm-linux-gnu, sparc-unknown-linux-gnu,
sparc64-unknown-linux-gnu.

2. Debian GNU/Linux 2.1

GCC 2.95.4 and GNU Make. armv4l-unknown-linux-gnu.

3. Tru64 UNIX

Tru64 UNIX C compiler and Tru64 Make. alphaev67-dec-osf5.1, alphaev68-dec-
osf5.1.

4. SuSE Linux 7.1

GCC 2.96 and GNU Make. alphaev6-unknown-linux-gnu, alphaev67-unknown-
linux-gnu.

5. SuSE Linux 7.2a

GCC 3.0 and GNU Make. ia64-unknown-linux-gnu.

6. SuSE Linux GCC 3.2.2 and GNU Make. x86_64-unknown-linux-gnu (AMD64 Opteron “Melody”).

7. RedHat Linux 7.2

GCC 2.96 and GNU Make. alphaev6-unknown-linux-gnu, alphaev67-unknown-
linux-gnu, ia64-unknown-linux-gnu.

8. RedHat Linux 8.0 GCC 3.2 and GNU Make. i686-pc-linux-gnu.

9. RedHat Advanced Server 2.1 GCC 2.96 and GNU Make. i686-pc-linux-gnu.

10. Slackware Linux 8.0.01 GCC 2.95.3 and GNU Make. i686-pc-linux-gnu.

11. Mandrake Linux 9.0 GCC 3.2 and GNU Make. i686-pc-linux-gnu.

12. IRIX 6.5 MIPS C compiler, IRIX Make. mips-sgi-irix6.5.

13. AIX 4.3.2 IBM C for AIX compiler, AIX Make. rs6000-ibm-aix4.3.2.0.

14. HP-UX 11

HP-UX C compiler and HP Make. ia64-hp-hpux11.22, hppa2.0w-hp-hpux11.11.

15. SUN Solaris 2.8 Sun WorkShop Compiler C 6.0 and SUN Make. sparc-sun-solaris2.8.

16. NetBSD 1.6

GCC 2.95.3 and GNU Make.  alpha-unknown-netbsd1.6, i386-unknown-
netbsdelf1.6.

17. OpenBSD 3.1 and 3.2

    GCC 2.95.3 and GNU Make.  alpha-unknown-openbsd3.1, i386-unknown-
    openbsd3.1.

18. FreeBSD 4.7 and 4.8

    GCC 2.95.4 and GNU Make. alpha-unknown-freebsd4.7, alpha-unknown-
    freebsd4.8, i386-unknown-freebsd4.7, i386-unknown-freebsd4.8.

19. MacOS X 10.2 Server Edition

    GCC 3.1 and GNU Make. powerpc-apple-darwin6.5.

20. Cross compiled to uClinux/uClibc on Motorola Coldfire. GCC 3.4 and GNU Make m68k-uclinux-elf.

If you use Shishi on, or port Shishi to, a new platform please report it to the author (see Section 1.8 [Bug Reports], page 11).

1.6 Getting help

A mailing list where users of Shishi may help each other exists, and you can reach it by sending e-mail to help-shishi@gnu.org. Archives of the mailing list discussions, and an interface to manage subscriptions, is available through the World Wide Web at http://

lists.gnu.org/mailman/listinfo/help-shishi.

1.7 Downloading and Installing

The package can be downloaded from several places, including:

   ftp://alpha.gnu.org/pub/gnu/shishi/

The latest version is stored in a file, e.g., ‘shishi-1.0.3.tar.gz’ where the ‘1.0.3’ indicate the highest version number.

The package is then extracted, configured and built like many other packages that use Autoconf. For detailed information on configuring and building it, refer to the INSTALL file that is part of the distribution archive.

Here is an example terminal session that download, configure, build and install the package. You will need a few basic tools, such as ‘sh’, ‘make’ and ‘cc’.

     $ wget -q ftp://alpha.gnu.org/pub/gnu/shishi/shishi-1.0.3.tar.gz
     $ tar xfz shishi-1.0.3.tar.gz
     $ cd shishi-1.0.3/
     $ ./configure
     ...
     $ make
     ...
     $ make install
     ...

After this you should be prepared to continue with the user, administration or programming manual, depending on how you want to use Shishi.

A few configure options may be relevant, summarized in the table.

--disable-des
--disable-3des
--disable-aes
--disable-md
--disable-null
--disable-arcfour

Disable a cryptographic algorithm at compile time. Usually it is better to disable algorithms during run-time with the configuration file, but this allows you to reduce the code size slightly.

--disable-starttls

Disable the experimental TLS support for KDC connections. If you do not use a Shishi KDC, this support is of no use so you could safely disable it.

--without-stringprep

Disable internationalized string processing.

For the complete list, refer to the output from configure --help.

1.8 Bug Reports

If you think you have found a bug in Shishi, please investigate it and report it.

• Please make sure that the bug is really in Shishi, and preferably also check that it hasn’t already been fixed in the latest version.

• You have to send us a test case that makes it possible for us to reproduce the bug.

• You also have to explain what is wrong; if you get a crash, or if the results printed are not good and in that case, in what way. Make sure that the bug report includes all information you would need to fix this kind of bug for someone else.

Please make an effort to produce a self-contained report, with something definite that can be tested or debugged. Vague queries or piecemeal messages are difficult to act on and don’t help the development effort.

If your bug report is good, we will do our best to help you to get a corrected version of the software; if the bug report is poor, we won’t do anything about it (apart from asking you to send better bug reports).

If you think something in this manual is unclear, or downright incorrect, or if the language needs to be improved, please also send a note.

Send your bug report to:

                         ‘bug-shishi@josefsson.org’

1.9 Contributing

If you want to submit a patch for inclusion – from solve a typo you discovered, up to adding support for a new feature – you should submit it as a bug report (see Section 1.8 [Bug Reports], page 11). There are some things that you can do to increase the chances for it to be included in the official package.

Unless your patch is very small (say, under 10 lines) we require that you assign the copyright of your work to the Free Software Foundation. This is to protect the freedom of the project. If you have not already signed papers, we will send you the necessary information when you submit your contribution.

For contributions that doesn’t consist of actual programming code, the only guidelines are common sense. Use it.

For code contributions, a number of style guides will help you:

• Coding Style. Follow the GNU Standards document (see Section “top” in standards). If you normally code using another coding standard, there is no problem, but you should use ‘indent’ to reformat the code (see Section “top” in indent) before submitting your work.

• Use the unified diff format ‘diff -u’.

• Return errors. The only valid reason for ever aborting the execution of the program is due to memory allocation errors, but for that you should call ‘shishi_xalloc_die’ to allow the application to recover if it wants to.

• Design with thread safety in mind. Don’t use global variables. Don’t even write to per-handle global variables unless the documented behaviour of the function you write is to write to the per-handle global variable.

• Avoid using the C math library. It causes problems for embedded implementations, and in most situations it is very easy to avoid using it.

• Document your functions. Use comments before each function headers, that, if properly formatted, are extracted into Texinfo manuals and GTK-DOC web pages.

• Supply a ChangeLog and NEWS entries, where appropriate.

2 User Manual

Usually Shishi interacts with you to get some initial authentication information like a password, and then contacts a server to receive a so called ticket granting ticket. From now on, you rarely interact with Shishi directly. Applications that need security services instruct the Shishi library to use the ticket granting ticket to get new tickets for various servers. An example could be if you log on to a host remotely via ‘telnet’. The host usually requires authentication before permitting you in. The ‘telnet’ client uses the ticket granting ticket to get a ticket for the server, and then uses this ticket to authenticate you against the server (typically the server is also authenticated to you). You perform the initial authentication by typing shishi at the prompt. Sometimes it is necessary to supply options telling Shishi what your principal name (user name in the Kerberos realm) or your realm is. In the example, I specify the client name simon@JOSEFSSON.ORG.

$ shishi simon@JOSEFSSON.ORG
Enter password for `simon@JOSEFSSON.ORG':
simon@JOSEFSSON.ORG:
Authtime:       Fri Aug 15 04:44:49 2003
Endtime:        Fri Aug 15 05:01:29 2003
Server:         krbtgt/JOSEFSSON.ORG key des3-cbc-sha1-kd (16)
Ticket key:     des3-cbc-sha1-kd (16) protected by des3-cbc-sha1-kd (16)
Ticket flags:   INITIAL (512)
$

As you can see, Shishi also prints a short description of the ticket received. A logical next step is to display all tickets you have received. By the way, the tickets are

usually stored as text in ~/.shishi/tickets. This is achieved by typing shishi --list.

$ shishi --list
Tickets in `/home/jas/.shishi/tickets':
jas@JOSEFSSON.ORG:
Authtime:       Fri Aug 15 04:49:46 2003
Endtime:        Fri Aug 15 05:06:26 2003
Server:         krbtgt/JOSEFSSON.ORG key des-cbc-md5 (3)
Ticket key:     des-cbc-md5 (3) protected by des-cbc-md5 (3)
Ticket flags:   INITIAL (512)
jas@JOSEFSSON.ORG:
Authtime:       Fri Aug 15 04:49:46 2003
Starttime:      Fri Aug 15 04:49:49 2003
Endtime:        Fri Aug 15 05:06:26 2003
Server:         host/latte.josefsson.org key des-cbc-md5 (3)
Ticket key:     des-cbc-md5 (3) protected by des-cbc-md5 (3)
2 tickets found.
$

As you can see, I had a ticket for the server ‘host/latte.josefsson.org’ which was generated by ‘telnet’:ing to that host.

If, for some reason, you want to manually get a ticket for a specific server, you can use the shishi --server-name command. Normally, however, the application that uses Shishi will take care of getting a ticket for the appropriate server, so you normally wouldn’t need to issue this command.

$ shishi --server-name=user/billg --encryption-type=des-cbc-md4
jas@JOSEFSSON.ORG:
Authtime:       Fri Aug 15 04:49:46 2003
Starttime:      Fri Aug 15 04:54:33 2003
Endtime:        Fri Aug 15 05:06:26 2003
Server:         user/billg key des-cbc-md4 (2)
Ticket key:     des-cbc-md4 (2) protected by des-cbc-md5 (3)
$

As you can see, I acquired a ticket for ‘user/billg’ with a ‘des-cbc-md4’ (see Section 1.4 [Cryptographic Overview], page 5) encryption key specified with the ‘--encryption-type’ parameter.

To wrap up this introduction, let us see how you can remove tickets. You may want to do this if you leave your terminal for lunch or similar, and don’t want someone to be able to copy the file and then use your credentials. Note that this only destroys the tickets locally, it does not contact any server telling that these credentials are no longer valid. So, if someone stole your ticket file, you must still contact your administrator and have them reset your account. Simply using this switch is not sufficient.

$ shishi --server-name=imap/latte.josefsson.org --destroy
1 ticket removed.
$ shishi --server-name=foobar --destroy
No tickets removed.
$ shishi --destroy
3 tickets removed.
$

Since the ‘--server-name’ parameter takes a long string to type, it is possible to type the server name directly, after the client name. The following example demonstrates an AS-REQ followed by a TGS-REQ for a specific server (assuming you did not have any tickets to begin with).

$ src/shishi simon@latte.josefsson.org imap/latte.josefsson.org
Enter password for `simon@latte.josefsson.org':
simon@latte.josefsson.org:
Acquired:       Wed Aug 27 17:21:06 2003
Expires:        Wed Aug 27 17:37:46 2003
Server:         imap/latte.josefsson.org key aes256-cts-hmac-sha1-96 (18)
Ticket key:     aes256-cts-hmac-sha1-96 (18) protected by aes256-cts-hmac-sha1-96 (18)
Ticket flags:   FORWARDED PROXIABLE (12)
$

Refer to the reference manual for all available parameters (see Section 4.6 [Parameters for shishi], page 45). The rest of this section contains descriptions of more specialized usage modes that can be ignored by most users.

2.1 Proxiable and Proxy Tickets

At times it may be necessary for a principal to allow a service to perform an operation on its behalf. The service must be able to take on the identity of the client, but only for a particular purpose. A principal can allow a service to take on the principal’s identity for a particular purpose by granting it a proxy.

The process of granting a proxy using the proxy and proxiable flags is used to provide credentials for use with specific services. Though conceptually also a proxy, users wishing to delegate their identity in a form usable for all purpose MUST use the ticket forwarding mechanism described in the next section to forward a ticket-granting ticket.

The PROXIABLE flag in a ticket is normally only interpreted by the ticket-granting service. It can be ignored by application servers. When set, this flag tells the ticketgranting server that it is OK to issue a new ticket (but not a ticket-granting ticket) with a different network address based on this ticket. This flag is set if requested by the client on initial authentication. By default, the client will request that it be set when requesting a ticket-granting ticket, and reset when requesting any other ticket.

This flag allows a client to pass a proxy to a server to perform a remote request on its behalf (e.g. a print service client can give the print server a proxy to access the client’s files on a particular file server in order to satisfy a print request).

In order to complicate the use of stolen credentials, Kerberos tickets are usually valid from only those network addresses specifically included in the ticket[4]. When granting a proxy, the client MUST specify the new network address from which the proxy is to be used, or indicate that the proxy is to be issued for use from any address.

The PROXY flag is set in a ticket by the TGS when it issues a proxy ticket. Application servers MAY check this flag and at their option they MAY require additional authentication from the agent presenting the proxy in order to provide an audit trail.

Here is how you would acquire a PROXY ticket for the service ‘imap/latte.josefsson.org’:

     $ shishi jas@JOSEFSSON.ORG imap/latte.josefsson.org --proxy
     Enter password for `jas@JOSEFSSON.ORG':
     libshishi: warning: KDC bug: Reply encrypted using wrong key.
     jas@JOSEFSSON.ORG:
     Authtime:       Mon Sep  8 20:02:35 2003
     Starttime:      Mon Sep  8 20:02:36 2003
     Endtime:        Tue Sep  9 04:02:35 2003
     Server:         imap/latte.josefsson.org key des3-cbc-sha1-kd (16)
     Ticket key:     des3-cbc-sha1-kd (16) protected by des3-cbc-sha1-kd (16)
     Ticket flags:   PROXY (16)
     $

As you noticed, this asked for your password. The reason is that proxy tickets must be acquired using a proxiable ticket granting ticket, which was not present. If you often need to get proxy tickets, you may acquire a proxiable ticket granting ticket from the start:

     $ shishi --proxiable
     Enter password for `jas@JOSEFSSON.ORG':
     jas@JOSEFSSON.ORG:
     Authtime:       Mon Sep  8 20:04:27 2003

     Endtime:        Tue Sep  9 04:04:27 2003
     Server:         krbtgt/JOSEFSSON.ORG key des3-cbc-sha1-kd (16)
     Ticket key:     des3-cbc-sha1-kd (16) protected by des3-cbc-sha1-kd (16)
     Ticket flags:   PROXIABLE INITIAL (520)

Then you should be able to acquire proxy tickets based on that ticket granting ticket, as follows:

     $ shishi jas@JOSEFSSON.ORG imap/latte.josefsson.org --proxy
     libshishi: warning: KDC bug: Reply encrypted using wrong key.
     jas@JOSEFSSON.ORG:
     Authtime:       Mon Sep  8 20:04:27 2003
     Starttime:      Mon Sep  8 20:04:32 2003
     Endtime:        Tue Sep  9 04:04:27 2003
     Server:         imap/latte.josefsson.org key des3-cbc-sha1-kd (16)
     Ticket key:     des3-cbc-sha1-kd (16) protected by des3-cbc-sha1-kd (16)
     Ticket flags:   PROXY (16)
     $

2.2 Forwardable and Forwarded Tickets

Authentication forwarding is an instance of a proxy where the service that is granted is complete use of the client’s identity. An example where it might be used is when a user logs in to a remote system and wants authentication to work from that system as if the login were local.

The FORWARDABLE flag in a ticket is normally only interpreted by the ticket-granting service. It can be ignored by application servers. The FORWARDABLE flag has an interpretation similar to that of the PROXIABLE flag, except ticket-granting tickets may also be issued with different network addresses. This flag is reset by default, but users MAY request that it be set by setting the FORWARDABLE option in the AS request when they request their initial ticket-granting ticket.

This flag allows for authentication forwarding without requiring the user to enter a password again. If the flag is not set, then authentication forwarding is not permitted, but the same result can still be achieved if the user engages in the AS exchange specifying the requested network addresses and supplies a password.

The FORWARDED flag is set by the TGS when a client presents a ticket with the FORWARDABLE flag set and requests a forwarded ticket by specifying the FORWARDED KDC option and supplying a set of addresses for the new ticket. It is also set in all tickets issued based on tickets with the FORWARDED flag set. Application servers may choose to process FORWARDED tickets differently than non-FORWARDED tickets.

If addressless tickets are forwarded from one system to another, clients SHOULD still use this option to obtain a new TGT in order to have different session keys on the different systems.

Here is how you would acquire a FORWARDED ticket for the service

‘host/latte.josefsson.org’:
     $ shishi jas@JOSEFSSON.ORG host/latte.josefsson.org --forwarded
     Enter password for `jas@JOSEFSSON.ORG':

libshishi: warning: KDC bug: Reply encrypted using wrong key.
jas@JOSEFSSON.ORG:
Authtime:       Mon Sep  8 20:07:11 2003
Starttime:      Mon Sep  8 20:07:12 2003
Endtime:        Tue Sep  9 04:07:11 2003
Server:         host/latte.josefsson.org key des3-cbc-sha1-kd (16)
Ticket key:     des3-cbc-sha1-kd (16) protected by des3-cbc-sha1-kd (16)
Ticket flags:   FORWARDED (4)
$

As you noticed, this asked for your password. The reason is that forwarded tickets must be acquired using a forwardable ticket granting ticket, which was not present. If you often need to get forwarded tickets, you may acquire a forwardable ticket granting ticket from the start:

$ shishi --forwardable
Enter password for `jas@JOSEFSSON.ORG':
jas@JOSEFSSON.ORG:
Authtime:       Mon Sep  8 20:08:53 2003
Endtime:        Tue Sep  9 04:08:53 2003
Server:         krbtgt/JOSEFSSON.ORG key des3-cbc-sha1-kd (16)
Ticket key:     des3-cbc-sha1-kd (16) protected by des3-cbc-sha1-kd (16)
Ticket flags:   FORWARDABLE INITIAL (514)
$

Then you should be able to acquire forwarded tickets based on that ticket granting ticket, as follows:

$ shishi jas@JOSEFSSON.ORG host/latte.josefsson.org --forwarded
libshishi: warning: KDC bug: Reply encrypted using wrong key.
jas@JOSEFSSON.ORG:
Authtime:       Mon Sep  8 20:08:53 2003
Starttime:      Mon Sep  8 20:08:57 2003
Endtime:        Tue Sep  9 04:08:53 2003
Server:         host/latte.josefsson.org key des3-cbc-sha1-kd (16)
Ticket key:     des3-cbc-sha1-kd (16) protected by des3-cbc-sha1-kd (16)
Ticket flags:   FORWARDED (4)
$

3 Administration Manual

Here you will learn how to set up, run and maintain the Shishi Kerberos server. Kerberos is incompatible with the standard Unix /etc/passwd password database^{1}, therefore the first step will be to create a Kerberos user database. Shishi’s user database system is called Shisa. Once Shisa has been configured, you can then start the server and begin issuing Kerberos tickets to your users. The Shishi server is called shishid. After getting the server up and running, we discuss how you can set up multiple Kerberos servers, to increase availability or offer load-balancing. Finally, we include some information intended for developers, that will enable you to customize Shisa to use an external user database, such as a LDAP server or SQL database.

3.1 Introduction to Shisa

The user database part of Shishi is called Shisa. The Shisa library is independent of the core Shishi library. Shisa is responsible for storing the name of your realms, the name of your principals (users), accounting information for the users (i.e., when each account starts to be valid and when it expires), and the cryptographic keys each user has. Some Kerberos internal data can also be stored, such as the key version number, the last dates for when various ticket requests were made, the cryptographic salt, string-to-key parameters and password for each user. Not all information need to be stored. For example, in some situations it is prudent to leave the password field empty, so that somebody who manages to steal the user database will only be able to compromise your system, and not any other systems were your user may have re-used the same password. On the other hand, you may already be storing the password in your customized database, in which case being able to change it via the Shisa interface can be useful.

Shisa is a small (a few thousand lines of C code) standalone library. Shisa does not depend on the Shishi library. Because a user database with passwords may be useful for other applications as well (e.g., GNU SASL), it might be separated into its own project later on. You should keep this in mind, so that you don’t consider writing a Shisa backend for your own database as a purely Shishi specific project. You can, for example, choose to use the Shisa interface in your own applications to have a simple interface to your user database. Your experience and feedback is appreciated if you have chosen to explore this.

Note that the Shisa database does not expose everything you may want to know about a user, such as its full human name, telephone number or even the user’s login account name or home directory. It only stores what is needed to authenticate a peer claiming to be an entity. Thus it does not make sense to replace your current user database or /etc/passwd with data derived from the Shisa database. Instead, it is intended that you write a Shisa backend that exports some of the information stored in your user database. You may be able to replace some existing functionality, such as the password field in /etc/passwd with a Kerberos PAM module, but there is no requirement for doing so.

3.2 Configuring Shisa

The configuration file for Shisa is typically stored in /usr/local/etc/shishi/shisa.conf. You do not have to modify this file, the defaults should be acceptable to first-time users.

1 And besides, Shishi is intended to work on non-Unix platforms as well.

The file is used to define where your user database resides, and some options such as making the database read-only, or whether errors detected when accessing the database should be ignored. (The latter could be useful if the server is a remote LDAP server that might be unavailable, and then you would want to fall back to a local copy of the database.)

The default will store the user database using directories and files, rooted by default in /usr/local/var/shishi. You can use standard file permission settings to control access to the directory hierarchy. It is strongly recommended to restrict access to the directory. Storing the directory on local storage, i.e., hard disk or removable media, is recommended. We discourage placing the database on a network file system, but realize this can be useful in some situations (see Section 3.7 [Multiple servers], page 32).

See the reference manual (see Section 4.5 [Shisa Configuration], page 44) for the details of the configuration file. Again, you are not expected to need to modify anything unless you are an experienced Shishi administrator.

3.3 Using Shisa

There is a command line interface to the Shisa library, aptly named shisa. You will use this tool to add, remove, and change information stored in the database about realms, principals, and keys. The tool can also be used to “dump” all information in the database, for backup or debugging purposes. (Currently the output format cannot be read by any tool, but functionality to do this will be added in the future, possibly as a read-only file-based Shisa database backend.)

The reference manual (see Section 4.8 [Parameters for shisa], page 47) explains all parameters, but here we will give you a walk-through of the typical uses of the tool.

Installing Shishi usually creates a realm with two principals: one ticket granting ticket for the realm, and one host key for the server. This is what you typically need to get started, but it doesn’t serve our purposes, so we start by removing the principals and the realm. To do that, we need to figure out the name of the realm. The ‘--list’ or ‘--dump’ parameters can be used for this. (Most “long” parameters, like ‘--dump’, have shorter names as well, in this case ‘-d’, Section 4.8 [Parameters for shisa], page 47).

     jas@latte:~$ shisa -d
     latte
             krbtgt/latte
                     Account is enabled.
                     Current key version 0 (0x0).
                     Key 0 (0x0).
                             Etype aes256-cts-hmac-sha1-96 (0x12, 18).
                             Salt lattekrbtgt/latte.
             host/latte
                     Account is enabled.
                     Current key version 0 (0x0).
                     Key 0 (0x0).
                             Etype aes256-cts-hmac-sha1-96 (0x12, 18).
                             Salt lattehost/latte.
     jas@latte:~$

The realm names are printed at column 0, the principal names are indented with one ‘TAB’ character (aka ‘\t’ or ASCII 0x09 Horizontal Tabulation), and the information about each principal is indented with two ‘TAB’ characters. The above output means that there is one realm ‘latte’ with two principals: ‘krbtgt/latte’ (which is used to authenticate Kerberos ticket requests) and ‘host/latte’ (used to authenticate host-based applications like Telnet). They were created during ‘make install’ on a host called ‘latte’.

If the installation did not create a default database for you, you might get an error similar to the following output.

     jas@latte:~$ shisa -d
     shisa: Cannot initialize `file' database backend.
     Location `/usr/local/var/shishi' and options `N/A'.
     shisa: Initialization failed:
     Shisa database could not be opened.
     jas@latte:~$

This indicates that the database does not exist. For a file database, you can create it simply by creating the directory, as follows. Note the access permission change with ‘chmod’. Typically the ‘root’ user would own the files, but as these examples demonstrate, setting up a Kerberos server does not require root access. Indeed, it may be prudent to run all Shishi applications as a special non-‘root’ user, and have all Shishi related files owned by that user, so that any security vulnerabilities do not lead to a system compromise. (However, if the user database is ever stolen, system compromises of other systems may be inoccured, should you use, e.g., a kerberized Telnet.)

     jas@latte:~$ mkdir /usr/local/var/shishi
     jas@latte:~$ chmod go-rwx /usr/local/var/shishi

Back to the first example, where you have a realm ‘latte’ with some principals. We want to remove the realm to demonstrate how you create the realm from scratch. (Of course, you can have more than one realm in the database, but for this example we assume you want to set up a realm named the same as Shishi guessed you would name it, so the existing realm need to be removed first.) The ‘--remove’ (short form ‘-r’) parameter is used for this purpose, as follows.

     jas@latte:~$ shisa -r latte host/latte
     Removing principal `host/latte@latte'...
     Removing principal `host/latte@latte'...done
     jas@latte:~$ shisa -r latte krbtgt/latte
     Removing principal `krbtgt/latte@latte'...
     Removing principal `krbtgt/latte@latte'...done
     jas@latte:~$ shisa -r latte
     Removing realm `latte'...
     Removing realm `latte'...done
     jas@latte:~$

You may be asking yourself “What if the realm has many more principals?”. If you fear manual labor (or a small ‘sed’ script, recall the format of ‘--list’?), don’t worry, there is a ‘--force’ (short form ‘-f’) flag. Use it with care. Here is a faster way to do the above:

     jas@latte:~$ shisa -r latte -f
     Removing principal `krbtgt/latte@latte'...

     Removing principal `krbtgt/latte@latte'...done
     Removing principal `host/latte@latte'...
     Removing principal `host/latte@latte'...done
     Removing realm `latte'...
     Removing realm `latte'...done
     jas@latte:~$

You should now have a working, but empty, Shisa database. Let’s set up the realm manually, step by step. The first step is to decide on a name for your realm. The full story is explained elsewhere (see Section 4.3 [Realm and Principal Naming], page 37), but the short story is to take your DNS domain name and translate it to upper case. For example, if your organization uses example.org it is a good idea to use EXAMPLE.ORG as the name of your Kerberos realm. We’ll use EXAMPLE.ORG as the realm name in these examples. Let’s create the realm.

     jas@latte:~$ shisa -a EXAMPLE.ORG
     Adding realm `EXAMPLE.ORG'...
     Adding realm `EXAMPLE.ORG'...done
     jas@latte:~$

Currently, there are no properties associated with entire realms. In the future, it may be possible to set a default realm-wide password expiry policy or similar. Each realm normally has one principal that is used for authenticating against the “ticket granting service” on the Kerberos server with a ticket instead of using the password. This is used by the user when she acquire a ticket for a server. The principal must look like ‘krbtgt/REALM’ (see [Name of the TGS], page 40). Let’s create it.

     jas@latte:~$ shisa -a EXAMPLE.ORG krbtgt/EXAMPLE.ORG
     Adding principal `krbtgt/EXAMPLE.ORG@EXAMPLE.ORG'...
     Adding principal `krbtgt/EXAMPLE.ORG@EXAMPLE.ORG'...done
     jas@latte:~$

Now that wasn’t difficult, although not very satisfying either. What does adding a principal mean? The name is created, obviously, but it also means setting a few values in the database. Let’s view the entry to find out which values.

     jas@latte:~$ shisa -d
     EXAMPLE.ORG
             krbtgt/EXAMPLE.ORG
                     Account is enabled.
                     Current key version 0 (0x0).
                     Key 0 (0x0).
                             Etype aes256-cts-hmac-sha1-96 (0x12, 18).
                             Salt EXAMPLE.ORGkrbtgt/EXAMPLE.ORG.
     jas@latte:~$

To use host based security services like SSH or Telnet with Kerberos, each host must have a key shared between the host and the KDC. The key is typically stored in /usr/local/etc/shishi/shishi.keys. We assume your server is called ‘mail.example.org’ and we create the principal. To illustrate a new parameter, we also set the specific algorithm to use by using the ‘--encryption-type’ (short form ‘-E’) parameter.

jas@latte:~$ shisa -a EXAMPLE.ORG host/mail.example.org -E des3
Adding principal `host/mail.example.org@EXAMPLE.ORG'...
Adding principal `host/mail.example.org@EXAMPLE.ORG'...done
jas@latte:~$

To export the key, there is another Shisa parameter ‘--keys’ that will print the key in a format that is recognized by Shishi. Let’s use it to print the host key.

jas@latte:~$ shisa -d --keys EXAMPLE.ORG host/mail.example.org
EXAMPLE.ORG
        host/mail.example.org
                Account is enabled.
                Current key version 0 (0x0).
                Key 0 (0x0).
                        Etype des3-cbc-sha1-kd (0x10, 16).
-----BEGIN SHISHI KEY-----
Keytype: 16 (des3-cbc-sha1-kd)
Principal: host/mail.example.org
Realm: EXAMPLE.ORG
iQdA8hxdvOUHZNliZJv7noM02rXHV8gq
-----END SHISHI KEY-----
                        Salt EXAMPLE.ORGhost/mail.example.org.
jas@latte:~$

So to set up the host, simply redirect output to the host key file.

jas@latte:~$ shisa -d --keys EXAMPLE.ORG \
     host/mail.example.org >> /usr/local/etc/shishi/shishi.keys
jas@latte:~$

The next logical step is to create a principal for some user, so you can use your password to get a Ticket Granting Ticket via the Authentication Service (AS) from the KDC, and then use the Ticket Granting Service (TGS) from the KDC to get a ticket for a specific host, and then send that ticket to the host to authenticate yourself. Creating this end-user principle is slightly different from the earlier steps, because you want the key to be derived from a password instead of being a random key. The ‘--password’ parameter indicate this. This make the tool ask you for the password.

jas@latte:~$ shisa -a EXAMPLE.ORG simon --password
Password for `simon@EXAMPLE.ORG':
Adding principal `simon@EXAMPLE.ORG'...
Adding principal `simon@EXAMPLE.ORG'...done
jas@latte:~$

The only special thing about this principal now is that it has a password field set in the database.

jas@latte:~$ shisa -d EXAMPLE.ORG simon --keys
EXAMPLE.ORG
        simon
                Account is enabled.
                Current key version 0 (0x0).

                Key 0 (0x0).
                        Etype aes256-cts-hmac-sha1-96 (0x12, 18).
-----BEGIN SHISHI KEY-----
Keytype: 18 (aes256-cts-hmac-sha1-96)
Principal: simon
Realm: EXAMPLE.ORG
Ja7ciNtrAI3gtodLaVDQ5zhcH58ffk0kS5tGAM7ILvM=
-----END SHISHI KEY-----
                        Salt EXAMPLE.ORGsimon.
                        Password foo.
jas@latte:~$

You should now be ready to start the KDC, which is explained in the next section (see Section 3.4 [Starting Shishid], page 23), and get tickets as explained earlier (see Chapter 2 [User Manual], page 13).

3.4 Starting Shishid

The Shishi server, or Key Distribution Center (KDC), is called Shishid. Shishid is responsible for listening on UDP and TCP ports for Kerberos requests. Currently it can handle initial ticket requests (Authentication Service, or AS), typically authenticated with keys derived from passwords, and subsequent ticket requests (Ticket Granting Service, or TGS), typically authenticated with the key acquired during an AS exchange.

Currently there is very little configuration available, the only variables are which ports the server should listen on and an optional user name to setuid into after successfully listening to the ports.

By default, Shishid listens on the ‘kerberos’ service port (typically translated to 88 via /etc/services) on the UDP and TCP transports via IPv4 and (if your machine support it) IPv6 on all interfaces on your machine. Here is a typical startup.

latte:/home/jas/src/shishi# /usr/local/sbin/shishid
Initializing GNUTLS...
Initializing GNUTLS...done
Listening on IPv4:*:kerberos/udp...done
Listening on IPv4:*:kerberos/tcp...done
Listening on IPv6:*:kerberos/udp...failed
socket: Address family not supported by protocol
Listening on IPv6:*:kerberos/tcp...failed
socket: Address family not supported by protocol
Listening on 2 ports...

Running as root is not recommended. Any security problem in shishid and your host may be compromised. Therefor, we recommend using the ‘--setuid’ parameter, as follows.

latte:/home/jas/src/shishi# /usr/local/sbin/shishid --setuid=jas
Initializing GNUTLS...
Initializing GNUTLS...done
Listening on IPv4:*:kerberos/udp...done
Listening on IPv4:*:kerberos/tcp...done

Listening on IPv6:*:kerberos/udp...failed
socket: Address family not supported by protocol
Listening on IPv6:*:kerberos/tcp...failed
socket: Address family not supported by protocol
Listening on 2 ports...
User identity set to `jas' (22541)...

An alternative is to run shishid on an alternative port as a non-privileged user. To continue the example of setting up the EXAMPLE.ORG realm as a non-privileged user from the preceding section, we start the server listen on port 4711 via UDP on IPv4.

jas@latte:~$ /usr/local/sbin/shishid -l IPv4:*:4711/udp
Initializing GNUTLS...
Initializing GNUTLS...done
Listening on *:4711/tcp...
Listening on 1 ports...
shishid: Starting (GNUTLS `1.0.4')
shishid: Listening on *:4711/tcp socket 4

If you have set up the Shisa database as in the previous example, you can now acquire tickets as follows.

jas@latte:~$ shishi -o 'realm-kdc=EXAMPLE.ORG,localhost:4711' \
   simon@EXAMPLE.ORG
Enter password for `simon@EXAMPLE.ORG':
simon@EXAMPLE.ORG:
Authtime:       Fri Dec 12 01:41:01 2003
Endtime:        Fri Dec 12 01:57:41 2003
Server:         krbtgt/EXAMPLE.ORG key aes256-cts-hmac-sha1-96 (18)
Ticket key:     aes256-cts-hmac-sha1-96 (18) protected by aes256-cts-hmac-sha1-96 (18)
Ticket flags:   FORWARDED PROXIABLE RENEWABLE INITIAL (12)
jas@latte:~$

The output from Shishid on a successful invocation would look like:

shishid: Has 131 bytes from *:4711/udp on socket 4
shishid: Processing 131 from *:4711/udp on socket 4
shishid: Trying AS-REQ
shishid: AS-REQ from simon@EXAMPLE.ORG for krbtgt/EXAMPLE.ORG@EXAMPLE.ORG
shishid: Matching client etype 18 against user key etype 18
shishid: Have 511 bytes for *:4711/udp on socket 4
shishid: Sending 511 bytes to *:4711/udp socket 4 via UDP
shishid: Listening on *:4711/udp socket 4

You may use the ’-v’ parameter for Shishid and Shishi to generate more debugging information.

To illustrate what an application, such as the Shishi patched versions of GNU lsh or Telnet from GNU InetUtils, would do when contacting the host ‘mail.example.org’ we illustrate using the TGS service as well.

jas@latte:~$ shishi -o 'realm-kdc=EXAMPLE.ORG,localhost:4711' \
   simon@EXAMPLE.ORG host/mail.example.org
simon@EXAMPLE.ORG:

     Authtime:       Fri Dec 12 01:46:54 2003
     Endtime:        Fri Dec 12 02:03:34 2003
     Server:         host/mail.example.org key des3-cbc-sha1-kd (16)
     Ticket key:     des3-cbc-sha1-kd (16) protected by aes256-cts-hmac-sha1-96 (18)
     Ticket flags:   FORWARDED PROXIABLE (45398796)
     jas@latte:~$

This conclude our walk-through of setting up a new Kerberos realm using Shishi. It is quite likely that one or more steps failed, and if so we encourage you to debug it and submit a patch, or at least report it as a problem. Heck, even letting us know if you got this far would be of interest. See Section 1.8 [Bug Reports], page 11.

3.5 Configuring DNS for KDC

Making sure the configuration files on all hosts running Shishi clients include the addresses of your server is tedious. If the configuration files do not mention the KDC address for a realm, Shishi will try to look up the information from DNS. In order for Shishi to find that information, you need to add the information to DNS. For this to work well, you need to set up a DNS zone with the same name as your Kerberos realm. The easiest is if you own the publicly visible DNS name, such as ‘example.org’ if your realm is ‘EXAMPLE.ORG’, but you can set up an internal DNS server with the information for your realm only. If this is done, you do not need to keep configuration files updated for the KDC addressing information.

3.5.1 DNS vs. Kerberos - Case Sensitivity of Realm Names

In Kerberos, realm names are case sensitive. While it is strongly encouraged that all realm names be all upper case this recommendation has not been adopted by all sites. Some sites use all lower case names and other use mixed case. DNS on the other hand is case insensitive for queries but is case preserving for responses to TXT queries. Since "MYREALM", "myrealm", and "MyRealm" are all different it is necessary that only one of the possible combinations of upper and lower case characters be used. This restriction may be lifted in the future as the DNS naming scheme is expanded to support non-ASCII names.

3.5.2 Overview - KDC location information

KDC location information is to be stored using the DNS SRV RR [RFC 2052]. The format of this RR is as follows:

Service.Proto.Realm TTL Class SRV Priority Weight Port Target

The Service name for Kerberos is always " kerberos".

The Proto can be either " udp", " tcp", or " tls. tcp". If these SRV records are to be used, a " udp" record MUST be included. If the Kerberos implementation supports TCP transport, a " tcp" record MUST be included. When using " tcp" with " kerberos", this indicates a "raw" TCP connection without any additional encapsulation. A " tls. tcp" record MUST be specified for all Kerberos implementations that support communication with the KDC across TCP sockets encapsulated using TLS [RFC2246] (see Section B.1 [STARTTLS protected KDC exchanges], page 236).

The Realm is the Kerberos realm that this record corresponds to.

TTL, Class, SRV, Priority, Weight, and Target have the standard meaning as defined in RFC 2052.

As per RFC 2052 the Port number should be the value assigned to "kerberos" by the Internet Assigned Number Authority (88).

3.5.3 Example - KDC location information

These are DNS records for a Kerberos realm ASDF.COM. It has two Kerberos servers, kdc1.asdf.com and kdc2.asdf.com. Queries should be directed to kdc1.asdf.com first as per the specified priority. Weights are not used in these records.

     _kerberos._udp.ASDF.COM.        IN      SRV     0 0 88 kdc1.asdf.com.
     _kerberos._udp.ASDF.COM.        IN      SRV     1 0 88 kdc2.asdf.com.
     _kerberos._tcp.ASDF.COM.        IN      SRV     0 0 88 kdc1.asdf.com.
     _kerberos._tcp.ASDF.COM.        IN      SRV     1 0 88 kdc2.asdf.com.
     _kerberos._tls._tcp.ASDF.COM.   IN      SRV     0 0 88 kdc1.asdf.com.
     _kerberos._tls._tcp.ASDF.COM.   IN      SRV     1 0 88 kdc2.asdf.com.

3.5.4 Security considerations

As DNS is deployed today, it is an unsecure service. Thus the infor- mation returned by it cannot be trusted.

Current practice for REALM to KDC mapping is to use hostnames to indicate KDC hosts (stored in some implementation-dependent location, but generally a local config file). These hostnames are vulnerable to the standard set of DNS attacks (denial of service, spoofed entries, etc). The design of the Kerberos protocol limits attacks of this sort to denial of service. However, the use of SRV records does not change this attack in any way. They have the same vulnerabilities that already exist in the common practice of using hostnames for KDC locations.

Implementations SHOULD provide a way of specifying this information locally without the use of DNS. However, to make this feature worthwhile a lack of any configuration information on a client should be interpretted as permission to use DNS.

3.6 Kerberos via TLS

If Shishi is built with support for GNUTLS, the messages exchanged between clients and Shishid can be protected with TLS. TLS is only available over TCP connections. A full discussion of the features TLS have is out of scope here, but in short it means the communication is integrity and privacy protected, and that users can use OpenPGP, X.509 or SRP (i.e., any mechanism supported by TLS) to authenticate themselves to the Kerberos server. For details on the implementation, See Section B.1 [STARTTLS protected KDC exchanges], page 236.

3.6.1 Setting up TLS resume

Resuming earlier TLS session is supported and enabled by default. This improves the speed of the TLS handshake, because results from earlier negotiations can be re-used. Currently the TLS resume database is stored in memory (in constract to storing it on disk), in both the client and in the server. Because the server typically runs for a long time, this is not a problem for that side. The client is typically not a long-running process though; the client usually is invoked as part of applications like ‘telnet’ or ‘login’. However, because each use of the client library typically result in a ticket, which is stored on disk and re-used by later processes, this is likely not a serious problem because the number of different tickets required by a user is usually quite small. For the client, TLS resume is typically only useful when you perform an initial authentication (using a password) followed by a ticket request for a service, in the same process.

You can configure the server, ‘shishid’ to never use TLS resume, or to increase or decrease the number of distinct TLS connections that can be resumed before they are garbage collected, see the ‘--resume-limit’ parameter (see Section 4.7 [Parameters for shishid], page 46).

3.6.2 Setting up Anonymous TLS

Anonymous TLS is the simplest to set up and use. In fact, only the client need to be informed that your KDC support TLS. This can be done in the configuration file with the ‘/tls’ parameter for ‘kdc-realm’ (see [Shishi Configuration], page 42), or by placing the KDC address in DNS using the ‘_tls’ SRV record (see Section 3.5 [Configuring DNS for KDC], page 25).

Let’s start Shishid, listening on a TCP socket. TLS require TCP. TCP sockets are automatically upgraded to TLS if the client request it.

jas@latte:~$ /usr/local/sbin/shishid -l IPv4:*:4711/tcp
Initializing GNUTLS...done
Listening on IPv4:*:4711/tcp...
Listening on 1 ports...
shishid: Starting (GNUTLS `1.0.4')
shishid: Listening on IPv4:*:4711/tcp socket 4

Let’s use the client to talk with it, using TLS.

jas@latte:~$ shishi -o 'realm-kdc=EXAMPLE.ORG,localhost:4711/tls \
    simon@EXAMPLE.ORG
Enter password for `simon@EXAMPLE.ORG':
simon@EXAMPLE.ORG:
Authtime:       Tue Dec 16 05:20:47 2003
Endtime:        Tue Dec 16 05:37:27 2003
Server:         krbtgt/EXAMPLE.ORG key aes256-cts-hmac-sha1-96 (18)
Ticket key:     aes256-cts-hmac-sha1-96 (18) protected by aes256-cts-hmac-sha1-96 (18)
Ticket flags:   FORWARDED PROXIABLE (12)
jas@latte:~$

On success, the server will print the following debug information.

shishid: Accepted socket 6 from socket 4 as IPv4:*:4711/tcp peer 127.0.0.1
shishid: Listening on IPv4:*:4711/tcp socket 4
shishid: Listening on IPv4:*:4711/tcp peer 127.0.0.1 socket 6
shishid: Has 4 bytes from IPv4:*:4711/tcp peer 127.0.0.1 on socket 6
shishid: Trying STARTTLS
shishid: TLS handshake negotiated protocol `TLS 1.0', key exchange `Anon DH', certficate type `X.509', cipher `AES 256 CBC', mac `SHA', compression `NULL', session not resumed
shishid: TLS anonymous authentication with 1024 bit Diffie-Hellman
shishid: Listening on IPv4:*:4711/tcp socket 4
shishid: Listening on IPv4:*:4711/tcp peer 127.0.0.1 socket 6
shishid: Has 131 bytes from IPv4:*:4711/tcp peer 127.0.0.1 on socket 6

shishid: Processing 131 from IPv4:*:4711/tcp peer 127.0.0.1 on socket 6
shishid: Trying AS-REQ
shishid: AS-REQ from simon@EXAMPLE.ORG for krbtgt/EXAMPLE.ORG@EXAMPLE.ORG
shishid: Matching client etype 18 against user key etype 18
shishid: Have 511 bytes for IPv4:*:4711/tcp peer 127.0.0.1 on socket 6
shishid: Sending 511 bytes to IPv4:*:4711/tcp peer 127.0.0.1 socket 6 via TLS
shishid: Listening on IPv4:*:4711/tcp socket 4
shishid: Listening on IPv4:*:4711/tcp peer 127.0.0.1 socket 6
shishid: Peer IPv4:*:4711/tcp peer 127.0.0.1 disconnected on socket 6
shishid: Closing IPv4:*:4711/tcp peer 127.0.0.1 socket 6
shishid: Listening on IPv4:*:4711/tcp socket 4

3.6.3 Setting up X.509 authenticated TLS

Setting up X.509 authentication is slightly more complicated than anonymous authentication. You need a X.509 certificate authority (CA) that can generate certificates for your Kerberos server and Kerberos clients. It is often easiest to setup the CA yourself. Managing a CA can be a daunting task, and we only give the bare essentials to get things up and running. We suggest that you study the relevant literature. As a first step beyond this introduction, you may wish to explore more secure forms of key storage than storing them unencrypted on disk.

The following three sections describe how you create the CA, KDC certificate, and client certificates. You can use any tool you like for this task, as long as they generate X.509 (PKIX) certificates in PEM format and RSA keys in PKCS#1 format. Here we use certtool that come with GNUTLS, which is widely available. We conclude by discussing how you use these certificates in the KDC and in the Shishi client.

3.6.3.1 Create a Kerberos Certificate Authority

First create a CA key.

jas@latte:~$ certtool --generate-privkey \
   --outfile /usr/local/etc/shishi/shishi.key
Generating a private key...
Generating a 1024 bit RSA private key...
jas@latte:~$

Then create the CA certificate. Use whatever details you prefer.

jas@latte:~$ certtool --generate-self-signed \
   --load-privkey /usr/local/etc/shishi/shishi.key \
   --outfile /usr/local/etc/shishi/shishi.cert
Generating a self signed certificate...
Please enter the details of the certificate's distinguished name. \
Just press enter to ignore a field.
Country name (2 chars): SE
Organization name: Shishi Example CA
Organizational unit name:
Locality name:
State or province name:
Common name: CA

This field should not be used in new certificates.
E-mail:
Enter the certificate's serial number (decimal): 0
Activation/Expiration time.
The generated certificate will expire in (days): 180
Extensions.
Does the certificate belong to an authority? (Y/N): y
Is this a web server certificate? (Y/N): n
Enter the e-mail of the subject of the certificate:
X.509 certificate info:
Version: 3
Serial Number (hex): 00
Validity:
        Not Before: Sun Dec 21 10:59:00 2003
        Not After: Fri Jun 18 11:59:00 2004
Subject: C=SE,O=Shishi Example CA,CN=CA
Subject Public Key Info:
        Public Key Algorithm: RSA
X.509 Extensions:
        Basic Constraints: (critical)
                CA:TRUE
Is the above information ok? (Y/N): y
Signing certificate...
jas@latte:~$

3.6.3.2 Create a Kerberos KDC Certificate

First create the key for the KDC.

jas@latte:~$ certtool --generate-privkey \
   --outfile /usr/local/etc/shishi/shishid.key
Generating a private key...
Generating a 1024 bit RSA private key...
jas@latte:~$

Then create actual KDC certificate, signed by the CA certificate created in the previous step.

jas@latte:~$ certtool --generate-certificate \
   --load-ca-certificate /usr/local/etc/shishi/shishi.cert \
   --load-ca-privkey /usr/local/etc/shishi/shishi.key \
   --load-privkey /usr/local/etc/shishi/shishid.key \
   --outfile /usr/local/etc/shishi/shishid.cert
Generating a signed certificate...
Loading CA's private key...
Loading CA's certificate...
Please enter the details of the certificate's distinguished name. \
Just press enter to ignore a field.
Country name (2 chars): SE
Organization name: Shishi Example KDC

Organizational unit name:
Locality name:
State or province name:
Common name: KDC
This field should not be used in new certificates.
E-mail:
Enter the certificate's serial number (decimal): 0
Activation/Expiration time.
The generated certificate will expire in (days): 180
Extensions.
Does the certificate belong to an authority? (Y/N): n
Is this a web server certificate? (Y/N): n
Enter the e-mail of the subject of the certificate:
X.509 certificate info:
Version: 3
Serial Number (hex): 00
Validity:
        Not Before: Sun Dec 21 11:02:00 2003
        Not After: Fri Jun 18 12:02:00 2004
Subject: C=SE,O=Shishi Example KDC,CN=KDC
Subject Public Key Info:
        Public Key Algorithm: RSA
X.509 Extensions:
        Basic Constraints: (critical)
                CA:FALSE
Is the above information ok? (Y/N): y
Signing certificate...
jas@latte:~$

3.6.3.3 Create a Kerberos Client Certificate

First create the key for the client.

jas@latte:~$ certtool --generate-privkey \
   --outfile ~/.shishi/client.key
Generating a private key...
Generating a 1024 bit RSA private key...
jas@latte:~$

Then create the client certificate, signed by the CA. An alternative would be to have the KDC sign the client certificates.

jas@latte:~$ certtool --generate-certificate \
   --load-ca-certificate /usr/local/etc/shishi/shishi.cert \
   --load-ca-privkey /usr/local/etc/shishi/shishi.key \
   --load-privkey ~/.shishi/client.key \
   --outfile ~/.shishi/client.certs
Generating a signed certificate...
Loading CA's private key...
Loading CA's certificate...

Please enter the details of the certificate's distinguished name. \
Just press enter to ignore a field.
Country name (2 chars): SE
Organization name: Shishi Example Client
Organizational unit name:
Locality name:
State or province name:
Common name: Client
This field should not be used in new certificates.
E-mail:
Enter the certificate's serial number (decimal): 0
Activation/Expiration time.
The generated certificate will expire in (days): 180
Extensions.
Does the certificate belong to an authority? (Y/N): n
Is this a web server certificate? (Y/N): n
Enter the e-mail of the subject of the certificate:
X.509 certificate info:
Version: 3
Serial Number (hex): 00
Validity:
        Not Before: Sun Dec 21 11:04:00 2003
        Not After: Fri Jun 18 12:04:00 2004
Subject: C=SE,O=Shishi Example Client,CN=Client
Subject Public Key Info:
        Public Key Algorithm: RSA
X.509 Extensions:
        Basic Constraints: (critical)
                CA:FALSE
Is the above information ok? (Y/N): y
Signing certificate...
jas@latte:~$

3.6.3.4 Starting KDC with X.509 authentication support

The KDC need the CA certificate (to verify client certificates) and the server certificate and key (to authenticate itself to the clients). See elsewhere (see Section 4.7 [Parameters for shishid], page 46) for the entire description of the parameters.

jas@latte:~$ shishid -l *:4711/tcp \
   --x509cafile /usr/local/etc/shishi/shishi.cert \
   --x509certfile /usr/local/etc/shishi/shishid.cert \
   --x509keyfile /usr/local/etc/shishi/shishid.key
Initializing GNUTLS...
Parsed 1 CAs...
Loaded server certificate/key...
Generating Diffie-Hellman parameters...
Initializing GNUTLS...done

Listening on *:4711/tcp...
Listening on 1 ports...
shishid: Starting (GNUTLS `1.0.4')
shishid: Listening on *:4711/tcp socket 4

Then acquire tickets as usual. In case you wonder how shishi finds the client certificate and key, the filenames used above when generating the client certificates happen to be the default filenames for these files. So it pick them up automatically.

jas@latte:~$ shishi -o 'realm-kdc=EXAMPLE.ORG,localhost:4711/tls' \
   simon@EXAMPLE.ORG
Enter password for `simon@EXAMPLE.ORG':
simon@EXAMPLE.ORG:
Authtime:       Sun Dec 21 11:15:47 2003
Endtime:        Sun Dec 21 11:32:27 2003
Server:         krbtgt/EXAMPLE.ORG key aes256-cts-hmac-sha1-96 (18)
Ticket key:     aes256-cts-hmac-sha1-96 (18) protected by aes256-cts-hmac-sha1-96 (18)
Ticket flags:   FORWARDED PROXIABLE RENEWABLE HWAUTHENT TRANSITEDPOLICYCHECKED OKASDELEGATE (12)
jas@latte:~$

Here is what the server would print.

shishid: Accepted socket 6 from socket 4 as *:4711/tcp peer 127.0.0.1
shishid: Listening on *:4711/tcp socket 4
shishid: Listening on *:4711/tcp peer 127.0.0.1 socket 6
shishid: Has 4 bytes from *:4711/tcp peer 127.0.0.1 on socket 6
shishid: Trying STARTTLS
shishid: TLS handshake negotiated protocol `TLS 1.0', key exchange `RSA', certficate type `X.509', cipher `AES 256 CBC', mac `SHA', compression `NULL', session not resumed
shishid: TLS client certificate `C=SE,O=Shishi Example Client,CN=Client', issued by `C=SE,O=Shishi Example CA,CN=CA', serial number `00', MD5 fingerprint `a5:d3:1f:58:76:e3:58:cd:2d:eb:f7:45:a2:4b:52:f9:', activated `Sun Dec 21 11:04:00 2003', expires `Fri Jun 18 12:04:00 2004', version #3, key RSA modulus 1024 bits, currently EXPIRED
shishid: Listening on *:4711/tcp socket 4
shishid: Listening on *:4711/tcp peer 127.0.0.1 socket 6
shishid: Has 131 bytes from *:4711/tcp peer 127.0.0.1 on socket 6
shishid: Processing 131 from *:4711/tcp peer 127.0.0.1 on socket 6
shishid: Trying AS-REQ
shishid: AS-REQ from simon@EXAMPLE.ORG for krbtgt/EXAMPLE.ORG@EXAMPLE.ORG
shishid: Matching client etype 18 against user key etype 18
shishid: Have 511 bytes for *:4711/tcp peer 127.0.0.1 on socket 6
shishid: Sending 511 bytes to *:4711/tcp peer 127.0.0.1 socket 6 via TLS
shishid: Listening on *:4711/tcp socket 4
shishid: Listening on *:4711/tcp peer 127.0.0.1 socket 6
shishid: Peer *:4711/tcp peer 127.0.0.1 disconnected on socket 6
shishid: Closing *:4711/tcp peer 127.0.0.1 socket 6
shishid: Listening on *:4711/tcp socket 4

3.7 Multiple servers

Setting up multiple servers is as easy as replicating the user database. Since the default ‘file’ user database is stored in the normal file system, you can use any common tools to replicate a file system. Network file system like NFS (properly secured by, e.g., a point-to-point symmetrically encrypted IPSEC connection) and file synchronizing tools like ‘rsync’ are typical choices.

The secondary server should be configured just like the master server. If you use the ‘file’ database over NFS you do not have to make any modifications. If you use, e.g., a cron job to ‘rsync’ the directory every hour or so, you may want to add a ‘--read-only’ flag to the Shisa ‘db’ definition (see Section 4.5 [Shisa Configuration], page 44). That way, nobody will be lured into creating or changing information in the database on the secondary server, which only would be overwritten during the next synchronization.

     db --read-only file /usr/local/var/backup-shishi

The ‘file’ database is designed so it doesn’t require file locking in the file system, which may be unreliable in some network file systems or implementations. It is also designed so that multiple concurrent readers and writers may access the database without causing corruption.

Warning: The last paragraph is currently not completely accurate. There may be race conditions with concurrent writers. None should cause infinite loops or data loss. However, unexpected results might occur if two writers try to update information about a principal simultaneous.

If you use a remote LDAP server or SQL database to store the user database, and access it via a Shisa backend, you have make sure your Shisa backend handle concurrent writers properly. If you use a modern SQL database, this probably is not a concern. If it is a problem, you may be able to work around it by implementing some kind of synchronization or semaphore mechanism. If all else sounds too complicated, you can set up the secondary servers as ‘--read-only’ servers, although you will lose some functionality (like changing passwords via the secondary server, or updating timestamps when the last ticket request occurred).

One function that is of particular use for users with remote databases (be it LDAP or SQL) is the “database override” feature. Using this you can have the security critical principals (such as the ticket granting ticket) stored on local file system storage, but use the remote database for user principals. Of course, you must keep the local file system storage synchronized between all servers, as before. Here is an example configuration.

     db --read-only file /var/local/master
     db ldap kdc.example.org ca=/etc/shisa/kdc-ca.pem

This instruct the Shisa library to access the two databases sequentially, for each query using the first database that know about the requested principal. If you put the ‘krbtgt/REALM’ principal in the local ‘file’ database, this will override the LDAP interface. Naturally, you can have as many ‘db’ definition lines as you wish.

Users with remote databases can also investigate a so called High Availability mode. This is useful if you wish to have your Kerberos servers be able to continue to operate even when the remote database is offline. This is achieved via the ‘--ignore-errors’ flag in the database definition. Here is a sample configuration.

     db --ignore-errors ldap kdc.example.org ca=/etc/shisa/kdc-ca.pem
     db --read-only file /var/cache/ldap-copy

This instruct the Shisa library to try the LDAP backend first, but if it fails, instead of returning an error, continue to try the operation on a read only local ‘file’ based database. Of course, write requests will still fail, but it may be better than halting the server completely. To make this work, you first need to set up a cron job on a, say, hourly basis, to make a copy of the remote database and store it in the local file database. That way, when the remote server goes away, fairly current information will still be available locally.

If you also wish to experiment with read-write fail over, here is an idea for the configuration.

     db --ignore-errors ldap kdc.example.org ca=/etc/shisa/kdc-ca.pem
     db --ignore-errors --read-only file /var/cache/ldap-copy
     db file /var/cache/local-updates

This is similar to the previous, but it will ignore errors reading and writing from the first two databases, ultimately causing write attempts to end up in the final ‘file’ based database. Of course, you would need to create tools to feed back any local updates made while the remote server was down. It may also be necessary to create a special backend for this purpose, which can auto create principals that are used.

We finish with an example that demonstrate all the ideas presented.

     db --read-only file /var/local/master
     db --ignore-errors ldap kdc.example.org ca=/etc/shisa/kdc-ca.pem
     db --ignore-errors --read-only file /var/cache/ldap-copy
     db file /var/cache/local-updates

3.8 Developer information

The Programming API for Shisa is described below (see Section 5.19 [Kerberos Database Functions], page 225); this section is about extending Shisa, and consequently Shishi, to use your own user database system. You may want to store your Kerberos user information on an LDAP database server, for example.

Adding a new backend is straight forward. You need to implement the backend API function set, add the list of API functions to db/db.c and possibly also add any library dependencies to the Makefile.

The simplest way to write a new backend is to start from the existing ‘file’ based database, in db/file.c, and modify the entry points as needed.

Note that the current backend API will likely change before it is frozen. We may describe it in detail here when it has matured. However, currently it is similar to the external Shisa API (see Section 5.19 [Kerberos Database Functions], page 225).

There should be no need to modify anything else in the Shisa library, and certainly not in the Shishi library or the shishid server.

Naturally, we would appreciate if you would send us your new backend, if you believe it is generally useful (see Section 1.8 [Bug Reports], page 11).

4 Reference Manual

This chapter discuss the underlying assumptions of Kerberos, contain a glossary to Kerberos concepts, give you background information on choosing realm and principal names, and describe all parameters and configuration file syntaxes for the Shishi tools.

4.1 Environmental Assumptions

Kerberos imposes a few assumptions on the environment in which it can properly function:

• "Denial of service" attacks are not solved with Kerberos. There are places in the protocols where an intruder can prevent an application from participating in the proper authentication steps. Detection and solution of such attacks (some of which can appear to be not-uncommon "normal" failure modes for the system) is usually best left to the human administrators and users.

• Principals MUST keep their secret keys secret. If an intruder somehow steals a principal’s key, it will be able to masquerade as that principal or impersonate any server to the legitimate principal.

• "Password guessing" attacks are not solved by Kerberos. If a user chooses a poor password, it is possible for an attacker to successfully mount an offline dictionary attack by repeatedly attempting to decrypt, with successive entries from a dictionary, messages obtained which are encrypted under a key derived from the user’s password.

• Each host on the network MUST have a clock which is "loosely synchronized" to the time of the other hosts; this synchronization is used to reduce the bookkeeping needs of application servers when they do replay detection. The degree of "looseness" can be configured on a per-server basis, but is typically on the order of 5 minutes. If the clocks are synchronized over the network, the clock synchronization protocol MUST itself be secured from network attackers.

• Principal identifiers are not recycled on a short-term basis. A typical mode of access control will use access control lists (ACLs) to grant permissions to particular principals. If a stale ACL entry remains for a deleted principal and the principal identifier is reused, the new principal will inherit rights specified in the stale ACL entry. By not re-using principal identifiers, the danger of inadvertent access is removed.

4.2 Glossary of terms

Authentication Verifying the claimed identity of a principal.

Authentication header A record containing a Ticket and an Authenticator to be presented to a server as part of the authentication process.

Authentication path A sequence of intermediate realms transited in the authentication process when communicating from one realm to another.

Authenticator A record containing information that can be shown to have been recently generated using the session key known only by the client and server.

Authorization The process of determining whether a client may use a service, which objects the client is allowed to access, and the type of access allowed for each.

Capability A token that grants the bearer permission to access an object or service. In Kerberos, this might be a ticket whose use is restricted by the contents of the authorization data field, but which lists no network addresses, together with the session key necessary to use the ticket.

Ciphertext The output of an encryption function. Encryption transforms plaintext into ciphertext.

Client A process that makes use of a network service on behalf of a user. Note that in some cases a Server may itself be a client of some other server (e.g. a print server may be a client of a file server).

Credentials A ticket plus the secret session key necessary to successfully use that ticket in an authentication exchange.

Encryption Type (etype)

When associated with encrypted data, an encryption type identifies the algorithm used to encrypt the data and is used to select the appropriate algorithm for decrypting the data. Encryption type tags are communicated in other messages to enumerate algorithms that are desired, supported, preferred, or allowed to be used for encryption of data between parties. This preference is combined with local information and policy to select an algorithm to be used.

KDC Key Distribution Center, a network service that supplies tickets and temporary session keys; or an instance of that service or the host on which it runs. The KDC services both initial ticket and ticket-granting ticket requests. The initial ticket portion is sometimes referred to as the Authentication Server (or service). The ticket-granting ticket portion is sometimes referred to as the ticket-granting server (or service).

Kerberos The name given to the Project Athena’s authentication service, the protocol used by that service, or the code used to implement the authentication service. The name is adopted from the three-headed dog which guards Hades.

Key Version Number (kvno)

A tag associated with encrypted data identifies which key was used for encryption when a long lived key associated with a principal changes over time. It is used during the transition to a new key so that the party decrypting a message can tell whether the data was encrypted using the old or the new key.

Plaintext The input to an encryption function or the output of a decryption function. Decryption transforms ciphertext into plaintext.

Principal A named client or server entity that participates in a network communication, with one name that is considered canonical.

Principal identifier The canonical name used to uniquely identify each different principal.

Seal To encipher a record containing several fields in such a way that the fields cannot be individually replaced without either knowledge of the encryption key or leaving evidence of tampering.

Secret key An encryption key shared by a principal and the KDC, distributed outside the bounds of the system, with a long lifetime. In the case of a human user’s principal, the secret key MAY be derived from a password.

Server A particular Principal which provides a resource to network clients. The server is sometimes referred to as the Application Server.

Service A resource provided to network clients; often provided by more than one server (for example, remote file service).

Session key A temporary encryption key used between two principals, with a lifetime limited to the duration of a single login "session". In the Kerberos system, a session key is generated by the KDC. The session key is distinct from the sub-session key, described next..

Sub-session key A temporary encryption key used between two principals, selected and exchanged by the principals using the session key, and with a lifetime limited to the duration of a single association. The sub- session key is also referred to as the subkey.

Ticket A record that helps a client authenticate itself to a server; it contains the client’s identity, a session key, a timestamp, and other information, all sealed using the server’s secret key. It only serves to authenticate a client when presented along with a fresh Authenticator.

4.3 Realm and Principal Naming

This section contains the discussion on naming realms and principals from the Kerberos specification.

4.3.1 Realm Names

Although realm names are encoded as GeneralStrings and although a realm can technically select any name it chooses, interoperability across realm boundaries requires agreement on how realm names are to be assigned, and what information they imply.

To enforce these conventions, each realm MUST conform to the conventions itself, and it MUST require that any realms with which inter-realm keys are shared also conform to the conventions and require the same from its neighbors.

Kerberos realm names are case sensitive. Realm names that differ only in the case of the characters are not equivalent. There are presently three styles of realm names: domain, X500, and other. Examples of each style follow:

          domain:   ATHENA.MIT.EDU
            X500:   C=US/O=OSF
           other:   NAMETYPE:rest/of.name=without-restrictions

Domain syle realm names MUST look like domain names: they consist of components separated by periods (.) and they contain neither colons (:) nor slashes (/). Though domain names themselves are case insensitive, in order for realms to match, the case must match as well. When establishing a new realm name based on an internet domain name it is recommended by convention that the characters be converted to upper case.

X.500 names contain an equal (=) and cannot contain a colon (:) before the equal. The realm names for X.500 names will be string representations of the names with components separated by slashes. Leading and trailing slashes will not be included. Note that the slash separator is consistent with Kerberos implementations based on RFC1510, but it is different from the separator recommended in RFC2253.

Names that fall into the other category MUST begin with a prefix that contains no equal (=) or period (.) and the prefix MUST be followed by a colon (:) and the rest of the name. All prefixes must be assigned before they may be used. Presently none are assigned.

The reserved category includes strings which do not fall into the first three categories. All names in this category are reserved. It is unlikely that names will be assigned to this category unless there is a very strong argument for not using the ’other’ category.

These rules guarantee that there will be no conflicts between the various name styles. The following additional constraints apply to the assignment of realm names in the domain and X.500 categories: the name of a realm for the domain or X.500 formats must either be used by the organization owning (to whom it was assigned) an Internet domain name or X.500 name, or in the case that no such names are registered, authority to use a realm name MAY be derived from the authority of the parent realm. For example, if there is no domain name for E40.MIT.EDU, then the administrator of the MIT.EDU realm can authorize the creation of a realm with that name.

This is acceptable because the organization to which the parent is assigned is presumably the organization authorized to assign names to its children in the X.500 and domain name systems as well. If the parent assigns a realm name without also registering it in the domain name or X.500 hierarchy, it is the parent’s responsibility to make sure that there will not in the future exist a name identical to the realm name of the child unless it is assigned to the same entity as the realm name.

4.3.2 Principal Names

As was the case for realm names, conventions are needed to ensure that all agree on what information is implied by a principal name. The name-type field that is part of the principal name indicates the kind of information implied by the name. The name-type SHOULD be treated only as a hint to interpreting the meaning of a name. It is not significant when checking for equivalence. Principal names that differ only in the name-type identify the same principal. The name type does not partition the name space. Ignoring the name type, no two names can be the same (i.e. at least one of the components, or the realm, MUST be different). The following name types are defined:

     name-type      value   meaning
     NT-UNKNOWN        0  Name type not known
     NT-PRINCIPAL      1  Just the name of the principal as in DCE, or for users
     NT-SRV-INST       2  Service and other unique instance (krbtgt)

     NT-SRV-HST        3  Service with host name as instance (telnet, rcommands)
     NT-SRV-XHST       4  Service with host as remaining components
     NT-UID            5  Unique ID
     NT-X500-PRINCIPAL 6  Encoded X.509 Distingished name [RFC 2253]
     NT-SMTP-NAME      7  Name in form of SMTP email name (e.g. user@foo.com)
     NT-ENTERPRISE    10   Enterprise name - may be mapped to principal name

When a name implies no information other than its uniqueness at a particular time the name type PRINCIPAL SHOULD be used. The principal name type SHOULD be used for users, and it might also be used for a unique server. If the name is a unique machine generated ID that is guaranteed never to be reassigned then the name type of UID SHOULD be used (note that it is generally a bad idea to reassign names of any type since stale entries might remain in access control lists).

If the first component of a name identifies a service and the remaining components identify an instance of the service in a server specified manner, then the name type of SRV-INST SHOULD be used. An example of this name type is the Kerberos ticket-granting service whose name has a first component of krbtgt and a second component identifying the realm for which the ticket is valid.

If the first component of a name identifies a service and there is a single component following the service name identifying the instance as the host on which the server is running, then the name type SRV- HST SHOULD be used. This type is typically used for Internet services such as telnet and the Berkeley R commands. If the separate components of the host name appear as successive components following the name of the service, then the name type SRV-XHST SHOULD be used. This type might be used to identify servers on hosts with X.500 names where the slash (/) might otherwise be ambiguous.

A name type of NT-X500-PRINCIPAL SHOULD be used when a name from an X.509 certificate is translated into a Kerberos name. The encoding of the X.509 name as a Kerberos principal shall conform to the encoding rules specified in RFC 2253.

A name type of SMTP allows a name to be of a form that resembles a SMTP email name. This name, including an "@" and a domain name, is used as the one component of the principal name.

A name type of UNKNOWN SHOULD be used when the form of the name is not known. When comparing names, a name of type UNKNOWN will match principals authenticated with names of any type. A principal authenticated with a name of type UNKNOWN, however, will only match other names of type UNKNOWN.

Names of any type with an initial component of ’krbtgt’ are reserved for the Kerberos ticket granting service. See [Name of the TGS], page 40, for the form of such names.

4.3.2.1 Name of server principals

The principal identifier for a server on a host will generally be composed of two parts: (1) the realm of the KDC with which the server is registered, and (2) a two-component name of type NT-SRV-HST if the host name is an Internet domain name or a multi-component name of type NT-SRV-XHST if the name of the host is of a form such as X.500 that allows slash (/) separators. The first component of the two- or multi-component name will identify the service and the latter components will identify the host. Where the name of the host is not case sensitive (for example, with Internet domain names) the name of the host MUST be lower case. If specified by the application protocol for services such as telnet and the Berkeley R commands which run with system privileges, the first component MAY be the string ’host’ instead of a service specific identifier.

4.3.2.2 Name of the TGS

The principal identifier of the ticket-granting service shall be composed of three parts:

(1) the realm of the KDC issuing the TGS ticket (2) a two-part name of type NT-SRV-INST, with the first part "krbtgt" and the second part the name of the realm which will accept the ticket-granting ticket. For example, a ticket-granting ticket issued by the ATHENA.MIT.EDU realm to be used to get tickets from the ATHENA.MIT.EDU KDC has a principal identifier of "ATHENA.MIT.EDU" (realm), ("krbtgt", "ATHENA.MIT.EDU") (name). A ticket-granting ticket issued by the ATHENA.MIT.EDU realm to be used to get tickets from the MIT.EDU realm has a principal identifier of "ATHENA.MIT.EDU" (realm), ("krbtgt", "MIT.EDU") (name).

4.3.3 Choosing a principal with which to communicate

The Kerberos protocol provides the means for verifying (subject to the assumptions in Section 4.1 [Environmental Assumptions], page 35) that the entity with which one communicates is the same entity that was registered with the KDC using the claimed identity (principal name). It is still necessary to determine whether that identity corresponds to the entity with which one intends to communicate.

When appropriate data has been exchanged in advance, this determination may be performed syntactically by the application based on the application protocol specification, information provided by the user, and configuration files. For example, the server principal name (including realm) for a telnet server might be derived from the user specified host name (from the telnet command line), the "host/" prefix specified in the application protocol specification, and a mapping to a Kerberos realm derived syntactically from the domain part of the specified hostname and information from the local Kerberos realms database.

One can also rely on trusted third parties to make this determination, but only when the data obtained from the third party is suitably integrity protected while resident on the third party server and when transmitted. Thus, for example, one should not rely on an unprotected domain name system record to map a host alias to the primary name of a server, accepting the primary name as the party one intends to contact, since an attacker can modify the mapping and impersonate the party with which one intended to communicate.

Implementations of Kerberos and protocols based on Kerberos MUST NOT use insecure DNS queries to canonicalize the hostname components of the service principal names. In an environment without secure name service, application authors MAY append a statically configured domain name to unqualified hostnames before passing the name to the security mechanisms, but should do no more than that. Secure name service facilities, if available, might be trusted for hostname canonicalization, but such canonicalization by the client SHOULD NOT be required by KDC implementations.

Implementation note: Many current implementations do some degree of canonicalization of the provided service name, often using DNS even though it creates security problems. However there is no consistency among implementations about whether the service name is case folded to lower case or whether reverse resolution is used. To maximize interoperability and security, applications SHOULD provide security mechanisms with names which result from folding the user-entered name to lower case, without performing any other modifications or canonicalization.

4.3.4 Principal Name Form

Principal names consist of a sequence of strings, which is often tedious to parse. Therefor, Shishi often uses a “printed” form of principal which embed the entire principal name string sequence, and optionally also the realm, into one string. The format is taken from the Kerberos 5 GSS-API mechanism (RFC 1964).

The elements included within this name representation are as follows, proceeding from the beginning of the string:

1. One or more principal name components; if more than one principal name component is included, the components are separated by ‘/‘. Arbitrary octets may be included within principal name components, with the following constraints and special considerations: a. Any occurrence of the characters ‘@‘ or ‘/‘ within a name component must be immediately preceded by the ‘\‘ quoting character, to prevent interpretation as a component or realm separator.

b. The ASCII newline, tab, backspace, and null characters may occur directly within the component or may be represented, respectively, by ‘\n‘, ‘\t‘, ‘\b‘, or ‘\0‘. c. If the ‘\‘ quoting character occurs outside the contexts described in (1a) and (1b) above, the following character is interpreted literally. As a special case, this allows the doubled representation ‘\\‘ to represent a single occurrence of the quoting character.

d. An occurrence of the ‘\‘ quoting character as the last character of a component is

illegal.

2. Optionally, a ‘@‘ character, signifying that a realm name immediately follows. If no realm name element is included, the local realm name is assumed. The ‘/‘ , ‘:‘, and null characters may not occur within a realm name; the ‘@‘, newline, tab, and backspace characters may be included using the quoting conventions described in (1a), (1b), and (1c) above.

4.4 Shishi Configuration

The valid configuration file tokens are described here. The user configuration file is typically

located in ~/.shishi/shishi.conf (compare ‘shishi --configuration-file’) and the

system configuration is typically located in /usr/local/etc/shishi/shishi.conf (compare ‘shishi --system-configuration-file’). If the first non white space character of a line is a ’#’, the line is ignored. Empty lines are also ignored.

All tokens are valid in both the system and the user configuration files, and have the same meaning. However, as the system file is supposed to apply to all users on a system, it would not make sense to use some tokens in that file. For example, the ‘default-principal’ is rarely useful in a system configuration file.

4.4.1 ‘default-realm’

Specify the default realm, by default the hostname of the host is used. E.g.,

     default-realm JOSEFSSON.ORG

4.4.2 ‘default-principal’

Specify the default principal, by default the login username is used. E.g.,

     default-principal jas
4.4.3 ‘client-kdc-etypes’

Specify which encryption types client asks server to respond in during AS/TGS exchanges. List valid encryption types, in preference order. Supported algorithms include aes256cts-hmac-sha1-96, aes128-cts-hmac-sha1-96, des3-cbc-sha1-kd, des-cbc-md5, des-cbc-md4, des-cbc-crc and null. This option also indicates which encryption types are accepted by the client when receiving the response. Note that the preference order is not cryptographically protected, so a man in the middle can modify the order without being detected. Thus, only specify encryption types you trust completely here. The default only includes aes256-cts-hmac-sha1-96, as suggested by RFC1510bis. E.g.,

     client-kdc-etypes=aes256-cts-hmac-sha1-96 des3-cbc-sha1-kd des-cbc-md5
4.4.4 ‘verbose’, ‘verbose-asn1’, ‘verbose-noise’, ‘verbose-crypto’,
     ‘verbose-crypto-noise’

Enable verbose library messages. E.g.,

     verbose
     verbose-noise
4.4.5 ‘realm-kdc’
Specify KDC addresses for realms. Value is ‘REALM,KDCADDRESS[/TRANSPORT][,KDCADDRESS[/TRANSPORT]...]’.

KDCADDRESS is the hostname or IP address of KDC.

Optional TRANSPORT is “udp” for UDP, “tcp” for TCP, and “tls” for TLS connections. By default UDP is tried first, and TCP used as a fallback if the KRB ERR RESPONSE TOO BIG error is received.

If not specified, Shishi tries to locate the KDC using SRV RRs, which is recommended. This option should normally only be used during experiments, or to access badly maintained realms.

     realm-kdc=JOSEFSSON.ORG,ristretto.josefsson.org
4.4.6 ‘server-realm’
Specify realm for servers. Value is ‘REALM,SERVERREGEXP[,SERVERREGEXP...]’.

SERVERREGEXP is a regular expression matching servers in the realm. The first match is used. E.g.,

     server-realm=JOSEFSSON.ORG,.josefsson.org

Note: currently not used.

4.4.7 ‘kdc-timeout’, ‘kdc-retries’

How long shishi waits for a response from a KDC before continuing to next KDC for realm. The default is 5 seconds. E.g.,

     kdc-timeout=10

How many times shishi sends a request to a KDC before giving up. The default is 3 times. E.g.,

     kdc-retries=5
4.4.8 ‘stringprocess’

How username and passwords entered from the terminal, or taken from the command line, are processed.

"none": no processing is used.

"stringprep": convert from locale charset to UTF-8 and process using experimental RFC 1510 stringprep profile.

It can also be a string indicating a character set supported by iconv via libstringprep, in which case data is converted from locale charset into the indicated character set. E.g., UTF-8, ISO-8859-1, KOI-8, EBCDIC-IS-FRISS are supported on GNU systems. On some systems you can use "locale -m" to list available character sets. By default, the "none" setting is used which is consistent with RFC 1510 that is silent on the issue. In practice, however, converting to UTF-8 improves interoperability.

E.g.,

     stringprocess=UTF-8
4.4.9 ‘ticket-life’

Specify default ticket life time.

The string can be in almost any common format. It can contain month names, time zones, ‘am’ and ‘pm’, ‘yesterday’, ‘ago’, ‘next’, etc. See Section 4.10 [Date input formats], page 49, for the long story.

As an extra feature, if the time specified by your string correspond to a time during the last 24 hours, an extra day is added to it. This allows you to specify relative times such as "17:00" to always mean the next 17:00, even if your system clock happens to be 17:30.

The default is 8 hours.

E.g.,

     #ticket-life=8 hours
     #ticket-life=1 day
     ticket-life=17:00
4.4.10 ‘renew-life’

Specify how long a renewable ticket should remain renewable.

See ticket-life for the syntax. The extra feature that handles negative values within the last 2 hours is not active here.

The default is 7 days.

E.g.,

     #renew-life=1 week
     #renew-life=friday 17:00
     renew-life=sunday

4.5 Shisa Configuration

The configuration file for Shisa is typically stored in /usr/local/etc/shishi/shisa.conf. If the first non white space character of a line is a ’#’, the line is ignored. Empty lines are also ignored.

4.5.1 ‘db’

Currently the only configuration options available is the db token that define the databases to use. The syntax is:

     db [OPTIONS] <TYPE> [LOCATION] [PARAMETERS ...]

Specify the data sources for Kerberos 5 data. Multiple entries, even of the same data source type, are allowed. The data sources are accessed in the same sequence as they are defined here. If an entry is found in one data source, it will be used for the operations, without searching the remaining data sources. Valid OPTIONS include:

     --read-only       No data is written to this data source.
     --ignore-errors   Ignore failures in this backend.

The default (when the configuration file is empty) uses one "file" data source (see below), but for a larger installation you may want to combine several data sources. Here is an example.

     db --read-only file /var/local/master
     db --ignore-errors ldap kdc.example.org ca=/etc/shisa/kdc-ca.pem
     db --read-only file /var/cache/ldap-copy

This demonstrate how you can store critical principals on local disk (the first entry, /var/local/master) that will always be found without looking in the LDAP directory. The critical principals could be, e.g., krbtgt/EXAMPLE.ORG. The second entry denote a LDAP server that could hold user principals. As you can see, Shisa will not let the caller know about errors with the LDAP source (they will be logged, however). Instead, if for instance the LDAP server has crashed, Shisa would continue and read from the /var/cache/ldapcopy file source. That file source may have been set up to contain a copy of the data in the LDAP server, perhaps made on an hourly basis, so that your server will be able to serve recent data even in case of a crash. Any updates or passwords change requests will however not be possible while the LDAP server is inaccessible, to reduce the problem of synchronizing data back into the LDAP server once it is online again.

Currently only the "file" data source is supported, and denote a data source that use the standard file system for storage.

Valid syntaxes for the "file" database:

     db file PATH

Examples:

     db file /var/shishi
     db file /usr/share/shishi read-only

If no ‘db’ tokens are present, the default will be:

     db file /usr/local/var/shishi

4.6 Parameters for shishi

If no command is given, Shishi try to make sure you have a ticket granting ticket for the default realm, and then display it.

Mandatory arguments to long options are mandatory for short options too.

Usage: shishi [OPTIONS]... [CLIENT [SERVER]]...
-h, --help                              Print help and exit
-V, --version                           Print version and exit
Commands:
-d, --destroy                           Destroy tickets in local cache,
limited by any --client-name or
--server-name.  (default=off)
-l, --list                              List tickets in local cache, limited
by any --client-name and
--server-name.  (default=off)
-r, --renew                             Renew ticket.  Use --server-name to
specify ticket, default is the
most recent renewable ticket
granting ticket for the default
realm.  (default=off)
Flags:
--forwardable                       Get a forwardable ticket, i.e., one
that can be used to get forwarded
tickets.  (default=off)
--forwarded                         Get a forwarded ticket.  (default=
off)
--proxiable                         Get a proxiable ticket, i.e., one
that can be used to get proxy
tickets.  (default=off)
--proxy                             Get a proxy ticket.  (default=off)
--renewable                         Get a renewable ticket.  (default=
off)
Options:
--client-name=NAME                  Client name. Default is login
username.
-E, --encryption-type=ETYPE,[ETYPE...]  Encryption types to use.  ETYPE is
either registered name or integer.
 Valid values include 'aes128',
'aes256', 'aes' (same as
'aes256'), '3des', 'des-md5',
'des-md4', 'des-crc', 'des' (same
as 'des-md5'), and 'arcfour'.
-e, --endtime=STRING                    Specify when ticket validity should

expire.  The time syntax may be
relative (to the start time), such
as '20 hours', or absolute, such
as '2001-02-03 04:05:06 CET'. The
default is 8 hours after the start
time.
--realm=STRING                      Set default realm.
--renew-till=STRING                 Specify renewable life of ticket.
Implies --renewable.  Accepts same
time syntax as --endtime.  If
--renewable is specified, the
default is 1 week after the start
time.
--server-name=NAME                  Server name. Default is
'krbtgt/REALM' where REALM is
client realm.
-s, --starttime=STRING                  Specify when ticket should start to
be valid.  Accepts same time
syntax as --endtime. The default
is to become valid immediately.
--ticket-granter=NAME               Service name in ticket to use for
authenticating request. Only for
TGS. Defaults to
'krbtgt/REALM@REALM' where REALM
is client realm.
Other options:
--configuration-file=FILE           Read user configuration from FILE.
-c, --ticket-file=FILE                  Read tickets from FILE.
-o, --library-options=STRING            Parse STRING as a configuration file
statement.
-q, --quiet                             Don't produce any diagnostic output.
  (default=off)
--system-configuration-file=FILE    Read system configuration from FILE.
--ticket-write-file=FILE            Write tickets from FILE.  Default is
to write them back to where they
were read from.
-v, --verbose                           Produce verbose output.
  (default=off)

4.7 Parameters for shishid

If no parameters are specified, ‘shishid’ listens on the defaults interfaces and answers incoming requests using the keys in the default key file.

Mandatory arguments to long options are mandatory for short options too.

Usage: shishid [OPTIONS]...

-h, --help                    Print help and exit
-V, --version                 Print version and exit
Commands:
-l, --listen=[FAMILY:]ADDR:PORT/TYPE
                              Sockets to listen for queries on.  Family is
                                `IPv4' or `IPv6', if absent the family is
                                decided by gethostbyname(ADDR). An address of
                                `*' indicates all addresses on the local
                                host. The default is `*:kerberos/udp,
                                *:kerberos/tcp'.
-u, --setuid=NAME             After binding socket, set user identity.
TLS settings:
    --no-tls                  Disable TLS support  (default=off)
    --x509cafile=FILE         X.509 certificate authorities used to verify
                                client certificates, in PEM format.
    --x509certfile=FILE       X.509 server certificate, in PEM format.
    --x509crlfile=FILE        X.509 certificate revocation list to check for
                                revoked client certificates, in PEM format.
    --x509keyfile=FILE        X.509 server certificate key, in PEM format.
    --resume-limit=SHORT      Keep track of up to this many TLS sessions for
                                resume purposes (0 to disable TLS resume).
                                (default=`50')
Other options:
-c, --configuration-file=FILE Use specified configuration file.
-v, --verbose                 Produce verbose output.
                                Use multiple times to increase amount of
                                information.
-q, --quiet                   Don't produce any diagnostic output.
                                  (default=off)

4.8 Parameters for shisa

The purpose of ‘shisa’ is to manipulate information stored in the Kerberos 5 database used by Shishi.

Mandatory arguments to long options are mandatory for short options too.

Usage: shisa [OPTIONS]... [REALM [PRINCIPAL]]...
-h, --help                         Print help and exit
-V, --version                      Print version and exit
Operations:
-a, --add                          Add realm or principal to database.
-d, --dump                         Dump entries in database.
-n, --key-add                      Add new key to a principal in database.

--key-remove                   Remove a key from a principal in
database.
-l, --list                         List entries in database.
-m, --modify                       Modify principal entry in database.
-r, --remove                       Remove realm or principal from database.
Parameters:
-f, --force                        Allow removal of non-empty realms.
  (default=off)
--enabled                      Only dump or list enabled principals.
  (default=off)
--disabled                     Only dump or list disabled principals.
  (default=off)
--keys                         Print cryptographic key and password in
hostkey format.  (default=off)
Values:
-E, --encryption-type=STRING       Override default key encryption type.
Valid values include 'aes128',
'aes256', 'aes' (same as 'aes256'),
'3des', 'des-md5', 'des-md4',
'des-crc', 'des' (same as 'des-md5'),
and 'arcfour'.
--key-version=NUMBER           Version of key.
--password[=STRING]            Derive key from this password.
--random                       Use a random key.  (default)
--salt=STRING                  Use specified salt for deriving key.
Defaults to concatenation of realm and
(unwrapped) principal name.
--string-to-key-parameter=HEX  Encryption algorithm specific parameter
for password derivation.  Currently
only the AES algorithm can utilize
this, where it is interpreted as the
iteration count of the PKCS#5 PBKDF2
key deriver.
Other options:
-c, --configuration-file=FILE      Use specified configuration file.
-o, --library-options=STRING       Parse string as configuration file
statement.
-v, --verbose                      Produce verbose output.
  (default=off)
-q, --quiet                        Don't produce any diagnostic output.
  (default=off)

4.9 Environment variables

A few of the compile-time defaults may be overridden at run-time by using environment variables. The following variables are supported.

• SHISHI_CONFIG Specify the location of the default system configuration file. Used by the Shishi library. If not specified, the default is specified at compile-time and is usually

$prefix/etc/shishi.conf.

• SHISHI_HOME Specify the user specific directory for configuration files, ticket cache, etc. Used by the Shishi library. If not specified, it is computed as $HOME/.shishi.

• SHISHI_USER Specify the default principal user name. Used by the Shishi library. If not specified, it is taken from the environment variable USER.

• SHISHI_TICKETS Specify the file name of the ticket cache. Used by the Shishi li-

brary. If not specified, it will be $SHISHI_HOME/tickets, or $HOME/.shishi/tickets

if $SHISHI_HOME is not specified.

4.10 Date input formats

First, a quote:

Our units of temporal measurement, from seconds on up to months, are so complicated, asymmetrical and disjunctive so as to make coherent mental reckoning in time all but impossible. Indeed, had some tyrannical god contrived to enslave our minds to time, to make it all but impossible for us to escape subjection to sodden routines and unpleasant surprises, he could hardly have done better than handing down our present system. It is like a set of trapezoidal building blocks, with no vertical or horizontal surfaces, like a language in which the simplest thought demands ornate constructions, useless particles and lengthy circumlocutions. Unlike the more successful patterns of language and science, which enable us to face experience boldly or at least level-headedly, our system of temporal calculation silently and persistently encourages our terror of time. … It is as though architects had to measure length in feet, width in meters and height in ells; as though basic instruction manuals demanded a knowledge of five different languages. It is no wonder then that we often look into our own immediate past or future, last Tuesday or a week from Sunday, with feelings of helpless confusion. … —Robert Grudin, Time and the Art of Living.

This section describes the textual date representations that GNU programs accept. These are the strings you, as a user, can supply as arguments to the various programs. The C interface (via the parse_datetime function) is not described here.

4.10.1 General date syntax

A date is a string, possibly empty, containing many items separated by whitespace. The whitespace may be omitted when no ambiguity arises. The empty string means the beginning of today (i.e., midnight). Order of the items is immaterial. A date string may contain many flavors of items:

• calendar date items • time of day items

• time zone items

• combined date and time of day items

• day of the week items

• relative items

• pure numbers.

We describe each of these item types in turn, below.

A few ordinal numbers may be written out in words in some contexts. This is most useful for specifying day of the week items or relative items (see below). Among the most commonly used ordinal numbers, the word ‘last’ stands for −1, ‘this’ stands for 0, and ‘first’ and ‘next’ both stand for 1. Because the word ‘second’ stands for the unit of time there is no way to write the ordinal number 2, but for convenience ‘third’ stands for 3, ‘fourth’ for 4, ‘fifth’ for 5, ‘sixth’ for 6, ‘seventh’ for 7, ‘eighth’ for 8, ‘ninth’ for 9, ‘tenth’ for 10, ‘eleventh’ for 11 and ‘twelfth’ for 12.

When a month is written this way, it is still considered to be written numerically, instead of being “spelled in full”; this changes the allowed strings.

In the current implementation, only English is supported for words and abbreviations

like ‘AM’, ‘DST’, ‘EST’, ‘first’, ‘January’, ‘Sunday’, ‘tomorrow’, and ‘year’.

The output of the date command is not always acceptable as a date string, not only because of the language problem, but also because there is no standard meaning for time zone items like ‘IST’. When using date to generate a date string intended to be parsed later, specify a date format that is independent of language and that does not use time zone items other than ‘UTC’ and ‘Z’. Here are some ways to do this:

     $ LC_ALL=C TZ=UTC0 date
     Tue Jul 21 23:00:37 UTC 2020
     $ TZ=UTC0 date +'%Y-%m-%d %H:%M:%SZ'
     2020-07-21 23:00:37Z
     $ date --rfc-3339=ns  # --rfc-3339 is a GNU extension.
     2020-07-21 19:00:37.692722128-04:00
     $ date --rfc-2822  # a GNU extension
     Tue, 21 Jul 2020 19:00:37 -0400
     $ date +'%Y-%m-%d %H:%M:%S %z'  # %z is a GNU extension.
     2020-07-21 19:00:37 -0400
     $ date +'@%s.%N'  # %s and %N are GNU extensions.
     @1595372437.692722128

Alphabetic case is completely ignored in dates. Comments may be introduced between round parentheses, as long as included parentheses are properly nested. Hyphens not followed by a digit are currently ignored. Leading zeros on numbers are ignored.

Invalid dates like ‘2019-02-29’ or times like ‘24:00’ are rejected. In the typical case of a host that does not support leap seconds, a time like ‘23:59:60’ is rejected even if it corresponds to a valid leap second.

4.10.2 Calendar date items

A calendar date item specifies a day of the year. It is specified differently, depending on whether the month is specified numerically or literally. All these strings specify the same calendar date:

2020-07-20     # ISO 8601.
20-7-20        # Assume 19xx for 69 through 99,
               # 20xx for 00 through 68 (not recommended).
7/20/2020      # Common U.S. writing.
20 July 2020
20 Jul 2020    # Three-letter abbreviations always allowed.
Jul 20, 2020
20-jul-2020
20jul2020

The year can also be omitted. In this case, the last specified year is used, or the current year if none. For example:

7/20
jul 20

Here are the rules.

For numeric months, the ISO 8601 format ‘year-month-day’ is allowed, where year is any positive number, month is a number between 01 and 12, and day is a number between 01 and 31. A leading zero must be present if a number is less than ten. If year is 68 or smaller, then 2000 is added to it; otherwise, if year is less than 100, then 1900 is added to it. The construct ‘month/day/year’, popular in the United States, is accepted. Also ‘month/day’, omitting the year.

Literal months may be spelled out in full: ‘January’, ‘February’, ‘March’, ‘April’, ‘May’,

‘June’, ‘July’, ‘August’, ‘September’, ‘October’, ‘November’ or ‘December’. Literal months

may be abbreviated to their first three letters, possibly followed by an abbreviating dot. It is also permitted to write ‘Sept’ instead of ‘September’.

When months are written literally, the calendar date may be given as any of the following:

day month year
day month
month day year
day-month-year

Or, omitting the year:

month day

4.10.3 Time of day items

A time of day item in date strings specifies the time on a given day. Here are some examples, all of which represent the same time:

20:02:00.000000
20:02
8:02pm
20:02-0500      # In EST (U.S. Eastern Standard Time).

More generally, the time of day may be given as ‘hour:minute:second’, where hour is a number between 0 and 23, minute is a number between 0 and 59, and second is a number between 0 and 59 possibly followed by ‘.’ or ‘,’ and a fraction containing one or more digits. Alternatively, ‘:second’ can be omitted, in which case it is taken to be zero. On the rare hosts that support leap seconds, second may be 60.

If the time is followed by ‘am’ or ‘pm’ (or ‘a.m.’ or ‘p.m.’), hour is restricted to run from 1 to 12, and ‘:minute’ may be omitted (taken to be zero). ‘am’ indicates the first half of the day, ‘pm’ indicates the second half of the day. In this notation, 12 is the predecessor of 1: midnight is ‘12am’ while noon is ‘12pm’. (This is the zero-oriented interpretation of ‘12am’ and ‘12pm’, as opposed to the old tradition derived from Latin which uses ‘12m’ for noon and ‘12pm’ for midnight.)

The time may alternatively be followed by a time zone correction, expressed as ‘shhmm’, where s is ‘+’ or ‘-’, hh is a number of zone hours and mm is a number of zone minutes. The zone minutes term, mm, may be omitted, in which case the one- or two-digit correction is interpreted as a number of hours. You can also separate hh from mm with a colon. When a time zone correction is given this way, it forces interpretation of the time relative to Coordinated Universal Time (UTC), overriding any previous specification for the time zone or the local time zone. For example, ‘+0530’ and ‘+05:30’ both stand for the time zone 5.5 hours ahead of UTC (e.g., India). This is the best way to specify a time zone correction by fractional parts of an hour. The maximum zone correction is 24 hours.

Either ‘am’/‘pm’ or a time zone correction may be specified, but not both.

4.10.4 Time zone items

A time zone item specifies an international time zone, indicated by a small set of letters, e.g., ‘UTC’ or ‘Z’ for Coordinated Universal Time. Any included periods are ignored. By following a non-daylight-saving time zone by the string ‘DST’ in a separate word (that is, separated by some white space), the corresponding daylight saving time zone may be specified. Alternatively, a non-daylight-saving time zone can be followed by a time zone correction, to add the two values. This is normally done only for ‘UTC’; for example, ‘UTC+05:30’ is equivalent to ‘+05:30’.

Time zone items other than ‘UTC’ and ‘Z’ are obsolescent and are not recommended, because they are ambiguous; for example, ‘EST’ has a different meaning in Australia than in the United States, and ‘A’ has different meaning as a military time zone than as an obsolete RFC 822 time zone. Instead, it’s better to use unambiguous numeric time zone corrections like ‘-0500’, as described in the previous section.

If neither a time zone item nor a time zone correction is supplied, timestamps are interpreted using the rules of the default time zone (see Section 4.10.10 [Specifying time zone rules], page 55).

4.10.5 Combined date and time of day items

The ISO 8601 date and time of day extended format consists of an ISO 8601 date, a ‘T’ character separator, and an ISO 8601 time of day. This format is also recognized if the ‘T’ is replaced by a space.

In this format, the time of day should use 24-hour notation. Fractional seconds are allowed, with either comma or period preceding the fraction. ISO 8601 fractional minutes and hours are not supported. Typically, hosts support nanosecond timestamp resolution; excess precision is silently discarded.

Here are some examples:

     2012-09-24T20:02:00.052-05:00
     2012-12-31T23:59:59,999999999+11:00
     1970-01-01 00:00Z

4.10.6 Day of week items

The explicit mention of a day of the week will forward the date (only if necessary) to reach that day of the week in the future.

Days of the week may be spelled out in full: ‘Sunday’, ‘Monday’, ‘Tuesday’, ‘Wednesday’, ‘Thursday’, ‘Friday’ or ‘Saturday’. Days may be abbreviated to their first three letters, optionally followed by a period. The special abbreviations ‘Tues’ for ‘Tuesday’, ‘Wednes’ for ‘Wednesday’ and ‘Thur’ or ‘Thurs’ for ‘Thursday’ are also allowed.

A number may precede a day of the week item to move forward supplementary weeks. It is best used in expression like ‘third monday’. In this context, ‘last day’ or ‘next day’ is also acceptable; they move one week before or after the day that day by itself would represent.

A comma following a day of the week item is ignored.

4.10.7 Relative items in date strings

Relative items adjust a date (or the current date if none) forward or backward. The effects of relative items accumulate. Here are some examples:

     1 year
     1 year ago
     3 years
     2 days

The unit of time displacement may be selected by the string ‘year’ or ‘month’ for moving by whole years or months. These are fuzzy units, as years and months are not all of equal duration. More precise units are ‘fortnight’ which is worth 14 days, ‘week’ worth 7 days, ‘day’ worth 24 hours, ‘hour’ worth 60 minutes, ‘minute’ or ‘min’ worth 60 seconds, and ‘second’ or ‘sec’ worth one second. An ‘s’ suffix on these units is accepted and ignored.

The unit of time may be preceded by a multiplier, given as an optionally signed number. Unsigned numbers are taken as positively signed. No number at all implies 1 for a multiplier. Following a relative item by the string ‘ago’ is equivalent to preceding the unit by a multiplier with value −1.

The string ‘tomorrow’ is worth one day in the future (equivalent to ‘day’), the string ‘yesterday’ is worth one day in the past (equivalent to ‘day ago’).

The strings ‘now’ or ‘today’ are relative items corresponding to zero-valued time displacement, these strings come from the fact a zero-valued time displacement represents the current time when not otherwise changed by previous items. They may be used to stress other items, like in ‘12:00 today’. The string ‘this’ also has the meaning of a zero-valued time displacement, but is preferred in date strings like ‘this thursday’.

When a relative item causes the resulting date to cross a boundary where the clocks were adjusted, typically for daylight saving time, the resulting date and time are adjusted accordingly.

The fuzz in units can cause problems with relative items. For example, ‘2020-07-31 -1 month’ might evaluate to 2020-07-01, because 2020-06-31 is an invalid date. To determine the previous month more reliably, you can ask for the month before the 15th of the current month. For example:

     $ date -R
     Thu, 31 Jul 2020 13:02:39 -0400
     $ date --date='-1 month' +'Last month was %B?'
     Last month was July?
     $ date --date="$(date +%Y-%m-15) -1 month" +'Last month was %B!'
     Last month was June!

Also, take care when manipulating dates around clock changes such as daylight saving leaps. In a few cases these have added or subtracted as much as 24 hours from the clock, so it is often wise to adopt universal time by setting the TZ environment variable to ‘UTC0’ before embarking on calendrical calculations.

4.10.8 Pure numbers in date strings

The precise interpretation of a pure decimal number depends on the context in the date string.

If the decimal number is of the form yyyymmdd and no other calendar date item (see Section 4.10.2 [Calendar date items], page 51) appears before it in the date string, then yyyy is read as the year, mm as the month number and dd as the day of the month, for the specified calendar date.

If the decimal number is of the form hhmm and no other time of day item appears before it in the date string, then hh is read as the hour of the day and mm as the minute of the hour, for the specified time of day. mm can also be omitted.

If both a calendar date and a time of day appear to the left of a number in the date string, but no relative item, then the number overrides the year.

4.10.9 Seconds since the Epoch

If you precede a number with ‘@’, it represents an internal timestamp as a count of seconds. The number can contain an internal decimal point (either ‘.’ or ‘,’); any excess precision not supported by the internal representation is truncated toward minus infinity. Such a number cannot be combined with any other date item, as it specifies a complete timestamp.

Internally, computer times are represented as a count of seconds since an Epoch—a welldefined point of time. On GNU and POSIX systems, the Epoch is 1970-01-01 00:00:00 UTC, so ‘@0’ represents this time, ‘@1’ represents 1970-01-01 00:00:01 UTC, and so forth. GNU and most other POSIX-compliant systems support such times as an extension to POSIX, using negative counts, so that ‘@-1’ represents 1969-12-31 23:59:59 UTC.

Most modern systems count seconds with 64-bit two’s-complement integers of seconds with nanosecond subcounts, which is a range that includes the known lifetime of the universe with nanosecond resolution. Some obsolescent systems count seconds with 32-bit two’scomplement integers and can represent times from 1901-12-13 20:45:52 through 2038-01-19 03:14:07 UTC. A few systems sport other time ranges.

On most hosts, these counts ignore the presence of leap seconds. For example, on most hosts ‘@1483228799’ represents 2016-12-31 23:59:59 UTC, ‘@1483228800’ represents 2017-01-01 00:00:00 UTC, and there is no way to represent the intervening leap second 2016-12-31 23:59:60 UTC.

4.10.10 Specifying time zone rules

Normally, dates are interpreted using the rules of the current time zone, which in turn are specified by the TZ environment variable, or by a system default if TZ is not set. To specify a different set of default time zone rules that apply just to one date, start the date with a string of the form ‘TZ="rule"’. The two quote characters (‘"’) must be present in the date, and any quotes or backslashes within rule must be escaped by a backslash.

For example, with the GNU date command you can answer the question “What time is it in New York when a Paris clock shows 6:30am on October 31, 2019?” by using a date beginning with ‘TZ="Europe/Paris"’ as shown in the following shell transcript:

     $ export TZ="America/New_York"
     $ date --date='TZ="Europe/Paris" 2019-10-31 06:30'
     Sun Oct 31 01:30:00 EDT 2019

In this example, the --date operand begins with its own TZ setting, so the rest of that operand is processed according to ‘Europe/Paris’ rules, treating the string ‘2019-10-31 06:30’ as if it were in Paris. However, since the output of the date command is processed according to the overall time zone rules, it uses New York time. (Paris was normally six hours ahead of New York in 2019, but this example refers to a brief Halloween period when the gap was five hours.)

A TZ value is a rule that typically names a location in the ‘tz’ database (https:// www.iana.org/time-zones). A recent catalog of location names appears in the TWiki

Date and Time Gateway (https://twiki.org/cgi-bin/xtra/tzdatepick.

html). A few non-GNU hosts require a colon before a location name in a TZ setting,

e.g., ‘TZ=":America/New_York"’.

The ‘tz’ database includes a wide variety of locations ranging from ‘Africa/Abidjan’ to ‘Pacific/Tongatapu’, but if you are at sea and have your own private time zone, or if you are using a non-GNU host that does not support the ‘tz’ database, you may need to use a POSIX rule instead. The previously-mentioned POSIX rule ‘UTC0’ says that the time zone abbreviation is ‘UTC’, the zone is zero hours away from Greenwich, and there is no daylight saving time. POSIX rules can also specify nonzero Greenwich offsets. For example, the following shell transcript answers the question “What time is it five and a half hours east of Greenwich when a clock seven hours west of Greenwich shows 9:50pm on July 12, 2022?”

     $ TZ="<+0530>-5:30" date --date='TZ="<-07>+7" 2022-07-12 21:50'
     Wed Jul 13 10:20:00 +0530 2022

This example uses the somewhat-confusing POSIX convention for rules. ‘TZ="<-07>+7"’ says that the time zone abbreviation is ‘-07’ and the time zone is 7 hours west of Greenwich, and ‘TZ="<+0530>-5:30"’ says that the time zone abbreviation is ‘+0530’ and the time zone is 5 hours 30 minutes east of Greenwich. Although trickier POSIX TZ settings like ‘TZ="<-05>+5<-04>,M3.2.0/2,M11.1.0/2"’ can specify some daylight saving regimes, location-based settings like ‘TZ="America/New_York"’ are typically simpler and more accurate historically. See Section “Specifying the Time Zone with TZ” in The GNU C Library.

4.10.11 Authors of parse_datetime

parse_datetime started life as getdate, as originally implemented by Steven M. Bellovin (smb@research.att.com) while at the University of North Carolina at Chapel Hill. The code was later tweaked by a couple of people on Usenet, then completely overhauled by Rich $alz (rsalz@bbn.com) and Jim Berets (jberets@bbn.com) in August, 1990. Various revisions for the GNU system were made by David MacKenzie, Jim Meyering, Paul Eggert and others, including renaming it to get_date to avoid a conflict with the alternative Posix function getdate, and a later rename to parse_datetime. The Posix function getdate can parse more locale-specific dates using strptime, but relies on an environment variable and external file, and lacks the thread-safety of parse_datetime.

This chapter was originally produced by François Pinard (pinard@iro.umontreal.ca) from the parse_datetime.y source code, and then edited by K. Berry (kb@cs.umb.edu).

5 Programming Manual

This chapter describes all the publicly available functions in the library.

5.1 Preparation

To use ‘Libshishi’, you have to perform some changes to your sources and the build system. The necessary changes are small and explained in the following sections. At the end of this chapter, it is described how the library is initialized, and how the requirements of the library are verified.

A faster way to find out how to adapt your application for use with ‘Libshishi’ may be to look at the examples at the end of this manual (see Section 5.18 [Examples], page 224).

5.1.1 Header

All interfaces (data types and functions) of the library are defined in the header file ‘shishi.h’. You must include this in all programs using the library, either directly or through some other header file, like this:

     #include <shishi.h>

The name space of ‘Libshishi’ is shishi_* for function names, Shishi* for data types and SHISHI_* for other symbols. In addition the same name prefixes with one prepended underscore are reserved for internal use and should never be used by an application.

5.1.2 Initialization

‘Libshishi’ must be initialized before it can be used. The library is initialized by calling shishi_init (see Section 5.2 [Initialization Functions], page 60). The resources allocated by the initialization process can be released if the application no longer has a need to call ‘Libshishi’ functions, this is done by calling shishi_done.

In order to take advantage of the internationalisation features in ‘Libshishi’, such as translated error messages, the application must set the current locale using setlocale before initializing ‘Libshishi’.

5.1.3 Version Check

It is often desirable to check that the version of ‘Libshishi’ used is indeed one which fits all requirements. Even with binary compatibility new features may have been introduced but due to problem with the dynamic linker an old version is actually used. So you may want to check that the version is okay right after program startup.

shishi check version

[Function] const char * shishi_check_version (const char * req_version)

req version: Oldest acceptable version, or NULL.

Description: Checks that the installed library version is at least as recent as the one provided in req version. The version string is formatted like "1.0.2". Whenever NULL is passed to this function, the check is suppressed, but the library version is still returned. Return value: Returns the active library version, or NULL, should the running library be too old.

The normal way to use the function is to put something similar to the following early in your main:

       if (!shishi_check_version (SHISHI_VERSION))
         {
           printf ("shishi_check_version failed:\n"
                   "Header file incompatible with shared library.\n");
           exit (EXIT_FAILURE);
         }

5.1.4 Building the source

If you want to compile a source file including the ‘shishi.h’ header file, you must make sure that the compiler can find it in the directory hierarchy. This is accomplished by adding the path to the directory in which the header file is located to the compilers include file search path (via the -I option).

However, the path to the include file is determined at the time the source is configured. To solve this problem, ‘Libshishi’ uses the external package pkg-config that knows the path to the include file and other configuration options. The options that need to be added to the compiler invocation at compile time are output by the --cflags option to pkg-config shishi. The following example shows how it can be used at the command line:

     gcc -c foo.c `pkg-config shishi --cflags`

Adding the output of ‘pkg-config shishi --cflags’ to the compilers command line will ensure that the compiler can find the ‘Libshishi’ header file.

A similar problem occurs when linking the program with the library. Again, the compiler has to find the library files. For this to work, the path to the library files has to be added to the library search path (via the -L option). For this, the option --libs to pkg-config shishi can be used. For convenience, this option also outputs all other options that are required to link the program with the ‘Libshishi’ libararies (in particular, the ‘-lshishi’ option). The example shows how to link foo.o with the ‘Libshishi’ library to a program

foo.
     gcc -o foo foo.o `pkg-config shishi --libs`

Of course you can also combine both examples to a single command by specifying both

options to pkg-config:
     gcc -o foo foo.c `pkg-config shishi --cflags --libs`

5.1.5 Autoconf tests

If you work on a project that uses Autoconf (see Section “top” in autoconf) to help find installed libraries, the suggestions in the previous section are not the entire story. There are a few methods to detect and incorporate Shishi into your Autoconf based package. The preferred approach, is to use Libtool in your project, and use the normal Autoconf header file and library tests.

5.1.5.1 Autoconf test via ‘pkg-config’

If your audience is a typical GNU/Linux desktop, you can often assume they have the ‘pkg-config’ tool installed, in which you can use its Autoconf M4 macro to find and set up your package for use with Shishi. The following illustrate this scenario.

AC_ARG_ENABLE(kerberos_v5,
AC_HELP_STRING([--disable-kerberos_v5],
                       [don't use the KERBEROS_V5 mechanism]),
kerberos_v5=$enableval)
if test "$kerberos_v5" != "no" ; then
PKG_CHECK_MODULES(SHISHI, shishi >= 0.0.0,
[kerberos_v5=yes],
                        [kerberos_v5=no])
if test "$kerberos_v5" != "yes" ; then
kerberos_v5=no
AC_MSG_WARN([shishi not found, disabling Kerberos 5])
else
kerberos_v5=yes
AC_DEFINE(USE_KERBEROS_V5, 1,
                          [Define to 1 if you want Kerberos 5.])
fi
fi
AC_MSG_CHECKING([if Kerberos 5 should be used])
AC_MSG_RESULT($kerberos_v5)

5.1.5.2 Standalone Autoconf test using Libtool

If your package uses Libtool(see Section “top” in libtool), you can use the normal Autoconf tests to find the Shishi library and rely on the Libtool dependency tracking to include the proper dependency libraries (e.g., Libidn). The following illustrate this scenario.

AC_CHECK_HEADER(shishi.h,
AC_CHECK_LIB(shishi, shishi_check_version,
[kerberos5=yes AC_SUBST(SHISHI_LIBS, -lshishi)],
kerberos5=no),
kerberos5=no)
AC_ARG_ENABLE(kerberos5,
AC_HELP_STRING([--disable-kerberos5],
                       [disable Kerberos 5 unconditionally]),
kerberos5=$enableval)
if test "$kerberos5" != "no" ; then
AC_DEFINE(USE_KERBEROS_V5, 1,
  [Define to 1 if you want Kerberos 5.])
else
AC_MSG_WARN([Shishi not found, disabling Kerberos 5])
fi
AC_MSG_CHECKING([if Kerberos 5 should be used])
AC_MSG_RESULT($kerberos5)

5.1.5.3 Standalone Autoconf test

If your package does not use Libtool, as well as detecting the Shishi library as in the previous case, you must also detect whatever dependencies Shishi requires to work (e.g., libidn). Since the dependencies are in a state of flux, we do not provide an example and we do not recommend this approach, unless you are experienced developer.

5.2 Initialization Functions

shishi

[Function] Shishi * shishi ()

Description: Initializes the Shishi library, and primes logging so that future warnings and informational messages are printed on stderr. If this function fails, it may send its own diagnostic errors to stderr.

Return value: Returns a Shishi library handle, or NULL on error.

shishi server

[Function] Shishi * shishi_server ()

Description: Initializes the Shishi library, and primes logging so that future warnings and informational messages are sent to the syslog system. If this function fails, it may print diagnostic errors in the syslog.

Return value: Returns a Shishi library handle, or NULL on error.

shishi done

[Function] void shishi_done (Shishi * handle)

handle: Shishi handle as allocated by shishi_init().

Description: Deallocates the Shishi library handle. The handle must not be used in any call to a shishi function after an execution of shishi_done().

If there is a default tkts, it is written to the default tkts file. If you do not wish to write the default tkts file, close the default file before calling this function. It is closed with a simple shishi_tkts_done(handle, NULL). For related information, see

shishi_tkts_default_file_set().

shishi init

[Function] int shishi_init (Shishi ** handle)

handle: Pointer to a Shishi handle created by this call.

Description: Creates a Shishi library handle, using shishi(), and reads the system configuration file, user configuration file and user tickets from their default locations. The paths to the system configuration file is decided at compile time, and is sysconfdir/shishi.conf. The user configuration file is HOME/.shishi/config, and the user ticket file is HOME/.shishi/ticket.

The handle is allocated regardless of return value. The single exception being SHISHI_ HANDLE_ERROR, which indicates a problem in allocating the handle. Other error conditions could arise while reading files.

Return value: Returns SHISHI_OK iff successful.

shishi init with paths

[Function] int shishi_init_with_paths (Shishi ** handle,

const char * tktsfile, const char * systemcfgfile,

    const char * usercfgfile)

handle: Pointer to a Shishi handle created by this call. tktsfile: Filename of ticket file, or NULL. systemcfgfile: Filename of system configuration, or NULL. usercfgfile: Filename of user configuration, or NULL. Description: Creates a Shishi library handle, using shishi(), and reads the system configuration file, user configuration file, and user tickets at the specified locations. If any of usercfgfile or systemcfgfile is NULL, the file is read from its default location, which for the system configuration is decided at compile time, and is sysconfdir/shishi.conf, and for the user configuration it is HOME/.shishi/config. If the ticket file name is NULL, a ticket file is not read at all. The handle is allocated regardless of return value. The single exception being SHISHI_ HANDLE_ERROR, which indicates a problem in allocating the handle. Other error conditions could arise while reading files. Return value: Returns SHISHI_OK iff successful.

shishi init server

[Function] int shishi_init_server (Shishi ** handle)

handle: Pointer to a Shishi handle created by this call. Description: Creates a Shishi library handle, using shishi_server(), and reads the system configuration file. The path to the system configuration file is decided at compile time, and is sysconfdir/shishi.conf. The handle is allocated regardless of return value. The single exception being SHISHI_ HANDLE_ERROR, which indicates a problem in allocating the handle. Other error conditions could arise while reading the file. Return value: Returns SHISHI_OK iff successful.

shishi init server with paths

[Function] int shishi_init_server_with_paths (Shishi ** handle,
    const char * systemcfgfile)

handle: Pointer to a Shishi handle created by this call. systemcfgfile: Filename of system configuration, or NULL. Description: Creates a Shishi library handle, using shishi_server(), and reads the system configuration file from the specified location. The path to the system configuration file is decided at compile time, and is sysconfdir/shishi.conf. The handle is allocated regardless of return value. The single exception being SHISHI_ HANDLE_ERROR, which indicates a problem in allocating the handle. Other error conditions could arise while reading the file. Return value: Returns SHISHI_OK iff successful.

shishi cfg

[Function] int shishi_cfg (Shishi * handle, const char * option)

handle: Shishi library handle created by shishi_init(). option: String containing shishi library options. Description: Configures the shishi library according to the options given in option. Return value: Returns SHISHI_OK if option is valid and configuration was successful.

shishi cfg from file

[Function] int shishi_cfg_from_file (Shishi * handle, const char * cfg)

handle: Shishi library handle created by shishi_init(). cfg: Name of configuration file. Description: Configures the shishi library using a configuration file located at cfg. Return value: Returns SHISHI_OK if successful.

shishi cfg print

[Function] int shishi_cfg_print (Shishi * handle, FILE * fh)

handle: Shishi library handle created by shishi_init(). fh: File stream handle opened for writing. Description: Prints library configuration status to fh. This function is mostly intended for debugging purposes. Return value: Always returns SHISHI_OK.

shishi cfg default systemfile

[Function] const char * shishi_cfg_default_systemfile (Shishi * handle)

handle: Shishi library handle created by shishi_init(). Description: The system configuration file name is decided at compile time, but is replaced by assigning another file name to the environment variable SHISHI_CONFIG. This call offers a single interface for determining the file name, to which the library turns for its settings. Return value: Returns file name of present system configuration.

shishi cfg default userdirectory

[Function] const char * shishi_cfg_default_userdirectory

(Shishi * handle) handle: Shishi library handle created by shishi_init(). Description: The default user directory, referred to for Shishi ticket cache and other purposes, is normally computed by appending the fixed string "/.shishi" to the content of the environment variable HOME. This hard coded directory, i.e., "HOME/.shishi/", can be replaced by whatever complete path is stored in the environment variable SHISHI_HOME. Return value: Returns the user’s directory name where the Shishi library will search for configuration files, ticket caches, etcetera.

shishi cfg userdirectory file

[Function] char * shishi_cfg_userdirectory_file (Shishi * handle,

const char * file) handle: Shishi library handle created by shishi_init(). file: Basename of file to use for the user’s configuration settings of the library. Description: Reports the full path to the file where the Shishi library expects to find the user’s library configuration, given that the file itself is named by the parameter file. The answer is composed from the value of file and the directory returned by shishi_ cfg_default_userdirectory(). Typically, the returned string would be expanded from "HOME/.shishi/file". Return value: Returns the absolute filename to the argument file, relative to the user specific Shishi configuration directory.

shishi cfg default userfile

[Function] const char * shishi_cfg_default_userfile (Shishi * handle)

handle: Shishi library handle created by shishi_init(). Description: Reports the absolute filename of the default user configuration file. This is typically "HOME/.shishi/shishi.conf". The value of SHISHI_HOME will change the directory part, as stated regarding shishi_

cfg_default_userdirectory().

Return value: Returns the user’s configuration filename.

shishi cfg clientkdcetype

[Function] int shishi_cfg_clientkdcetype (Shishi * handle,

int32 t ** etypes) handle: Shishi library handle created by shishi_init(). etypes: Pointer to an array of encryption types. Description: Sets the variable etypes to a static array of preferred encryption types applicable to clients. Return value: Returns the number of encryption types referred to by the updated array pointer, or zero, should no type exist.

shishi cfg clientkdcetype fast

[Function] int32_t shishi_cfg_clientkdcetype_fast (Shishi * handle)

handle: Shishi library handle created by shishi_init(). Description: Extracts the default encryption type from the list of preferred encryption types acceptable to the client. When the preferred list is empty, SHISHI_AES256_CTS_HMAC_SHA1_96 is returned as a sensible default type. Return value: Returns the default encryption type.

shishi cfg clientkdcetype set

[Function] int shishi_cfg_clientkdcetype_set (Shishi * handle,

char * value) handle: Shishi library handle created by shishi_init(). value: String naming acceptable encryption types. Description: Sets the configuration option "client-kdc-etypes" from value. The string contains encryption types, integers or names, separated by comma or by whitespace. An example naming three encryption types could be: aes256-cts-hmac-sha1-96 des3-cbc-sha1-kd des-cbc-md5 Return value: Returns SHISHI_OK if successful, and SHISHI_INVALID_ARGUMENT otherwise.

shishi cfg authorizationtype set

[Function] int shishi_cfg_authorizationtype_set (Shishi * handle,

char * value) handle: Shishi library handle created by shishi_init(). value: String listing acceptable authorization types. Description: Sets the configuration option "authorization-types" from value. The string contains authorization types, integers or names, separated by comma or whitespace. As an example, "k5login basic" would first check Kerberos5 authentication based on preset principals, and then fall back to the basic test of identical principal names. Return value: Returns SHISHI_OK if successful, and SHISHI_INVALID_ARGUMENT otherwise.

5.3 Ticket Set Functions

A “ticket set” is, as the name implies, a collection of tickets. Functions are provided to read tickets from file into a ticket set, to query number of tickets in the set, to extract a given ticket from the set, to search the ticket set for tickets matching certain criterium, to write the ticket set to a file, etc. High level functions for performing a initial authentication (see Section 5.7 [AS Functions], page 112) or subsequent authentication (see Section 5.8 [TGS Functions], page 117) and storing the new ticket in the ticket set are also provided.

See Section 5.6 [Ticket Functions], page 102, to manipulate each individual ticket. See Section 5.9 [Ticket (ASN.1) Functions], page 124, for low-level ASN.1 manipulation.

shishi tkts default file guess

[Function] char * shishi_tkts_default_file_guess (Shishi * handle)

handle: Shishi library handle create by shishi_init(). Description: Guesses the default ticket filename; it is SHISHI_TICKETS, SHISHI_ HOME/tickets, or HOME/.shishi/tickets. Return value: Returns default tkts filename as a string that has to be deallocated with free() by the caller.

shishi tkts default file

[Function] const char * shishi_tkts_default_file (Shishi * handle)

handle: Shishi library handle create by shishi_init(). Description: Get filename of default ticket set. Return value: Returns the default ticket set filename used in the library. The string is not a copy, so don’t modify or deallocate it.

shishi tkts default file set

[Function] void shishi_tkts_default_file_set (Shishi * handle,

const char * tktsfile) handle: Shishi library handle create by shishi_init(). tktsfile: string with new default tkts file name, or NULL to reset to default. Description: Set the default ticket set filename used in the library. The string is copied into the library, so you can dispose of the variable immediately after calling this function.

shishi tkts default

[Function] Shishi_tkts * shishi_tkts_default (Shishi * handle)

handle: Shishi library handle create by shishi_init(). Description: Get the default ticket set for library handle. Return value: Return the handle global ticket set.

shishi tkts

[Function] int shishi_tkts (Shishi * handle, Shishi tkts ** tkts)

handle: shishi handle as allocated by shishi_init(). tkts: output pointer to newly allocated tkts handle. Description: Get a new ticket set handle. Return value: Returns SHISHI_OK iff successful.

shishi tkts done

[Function] void shishi_tkts_done (Shishi tkts ** tkts)

tkts: ticket set handle as allocated by shishi_tkts(). Description: Deallocates all resources associated with ticket set. The ticket set handle must not be used in calls to other shishi tkts *() functions after this.

shishi tkts size

[Function] int shishi_tkts_size (Shishi tkts * tkts)

tkts: ticket set handle as allocated by shishi_tkts(). Description: Get size of ticket set. Return value: Returns number of tickets stored in ticket set.

shishi tkts nth

[Function] Shishi_tkt * shishi_tkts_nth (Shishi tkts * tkts, int ticketno)

tkts: ticket set handle as allocated by shishi_tkts(). ticketno: integer indicating requested ticket in ticket set. Get the n: th ticket in ticket set. Return value: Returns a ticket handle to the ticketno:th ticket in the ticket set, or NULL if ticket set is invalid or ticketno is out of bounds. The first ticket is ticketno 0, the second ticketno 1, and so on.

shishi tkts remove

[Function] int shishi_tkts_remove (Shishi tkts * tkts, int ticketno)

tkts: ticket set handle as allocated by shishi_tkts(). ticketno: ticket number of ticket in the set to remove. The first ticket is ticket number 0. Description: Remove a ticket, indexed by ticketno, in ticket set. Return value: SHISHI_OK if successful or if ticketno larger than size of ticket set.

shishi tkts add

[Function] int shishi_tkts_add (Shishi tkts * tkts, Shishi tkt * tkt)

tkts: ticket set handle as allocated by shishi_tkts(). tkt: ticket to be added to ticket set. Description: Add a ticket to the ticket set. Only the pointer is stored, so if you modify tkt, the ticket in the ticket set will also be modified. Return value: Returns SHISHI_OK iff successful.

shishi tkts new

[Function] int shishi_tkts_new (Shishi tkts * tkts, Shishi asn1 ticket,

Shishi asn1 enckdcreppart, Shishi asn1 kdcrep)

tkts: ticket set handle as allocated by shishi_tkts(). ticket: input ticket variable. enckdcreppart: input ticket detail variable. kdcrep: input KDC-REP variable. Description: Allocate a new ticket and add it to the ticket set. Note that ticket, enckdcreppart and kdcrep are stored by reference, so you must not de-allocate them before the ticket is removed from the ticket set and de-allocated. Return value: Returns SHISHI_OK iff successful.

shishi tkts read

[Function] int shishi_tkts_read (Shishi tkts * tkts, FILE * fh)

tkts: ticket set handle as allocated by shishi_tkts(). fh: file descriptor to read from. Description: Read tickets from file descriptor and add them to the ticket set. Return value: Returns SHISHI_OK iff successful.

shishi tkts from file

[Function] int shishi_tkts_from_file (Shishi tkts * tkts,

const char * filename) tkts: ticket set handle as allocated by shishi_tkts(). filename: filename to read tickets from.

Description: Read tickets from file and add them to the ticket set.

Return value: Returns SHISHI_OK iff successful.

shishi tkts write

[Function] int shishi_tkts_write (Shishi tkts * tkts, FILE * fh)

tkts: ticket set handle as allocated by shishi_tkts(). fh: file descriptor to write tickets to.

Description: Write tickets in set to file descriptor.

Return value: Returns SHISHI_OK iff successful.

shishi tkts expire

[Function] int shishi_tkts_expire (Shishi tkts * tkts)

tkts: ticket set handle as allocated by shishi_tkts().

Description: Remove expired tickets from ticket set.

Return value: Returns SHISHI_OK iff successful.

shishi tkts to file

[Function] int shishi_tkts_to_file (Shishi tkts * tkts,

const char * filename) tkts: ticket set handle as allocated by shishi_tkts(). filename: filename to write tickets to.

Description: Write tickets in set to file.

Return value: Returns SHISHI_OK iff successful.

shishi tkts print for service

[Function] int shishi_tkts_print_for_service (Shishi tkts * tkts,

FILE * fh, const char * service)

tkts: ticket set handle as allocated by shishi_tkts(). fh: file descriptor to print to. service: service to limit tickets printed to, or NULL.

Description: Print description of tickets for specified service to file descriptor. If service is NULL, all tickets are printed.

Return value: Returns SHISHI_OK iff successful.

shishi tkts print

[Function] int shishi_tkts_print (Shishi tkts * tkts, FILE * fh)

tkts: ticket set handle as allocated by shishi_tkts(). fh: file descriptor to print to. Description: Print description of all tickets to file descriptor. Return value: Returns SHISHI_OK iff successful.

shishi tkt match p

[Function] int shishi_tkt_match_p (Shishi tkt * tkt, Shishi tkts hint * hint)

tkt: ticket to test hints on. hint: structure with characteristics of ticket to be found. Description: Test if a ticket matches specified hints. Return value: Returns 0 iff ticket fails to match given criteria.

shishi tkts find

[Function] Shishi_tkt * shishi_tkts_find (Shishi tkts * tkts,

Shishi tkts hint * hint) tkts: ticket set handle as allocated by shishi_tkts(). hint: structure with characteristics of ticket to be found. Description: Search the ticketset sequentially (from ticket number 0 through all tickets in the set) for a ticket that fits the given characteristics. If a ticket is found, the hint- >startpos field is updated to point to the next ticket in the set, so this function can be called repeatedly with the same hint argument in order to find all tickets matching a certain criterium. Note that if tickets are added to, or removed from, the ticketset during a query with the same hint argument, the hint->startpos field must be updated appropriately. Here is how you would typically use this function: Shishi tkts hint hint; Shishi tkt tkt; memset(&hint, 0, sizeof(hint)); hint.server = "imap/mail.example.org"; tkt = shishi tkts find (shishi tkts default(handle), &hint); if (!tkt) printf("No ticket found...\n"); else do something with ticket (tkt); Return value: Returns a ticket if found, or NULL if no further matching tickets could be found.

shishi tkts find for clientserver

[Function] Shishi_tkt * shishi_tkts_find_for_clientserver

(Shishi tkts * tkts, const char * client, const char * server)

tkts: ticket set handle as allocated by shishi_tkts(). client: client name to find ticket for. server: server name to find ticket for. Description: Short-hand function for searching the ticket set for a ticket for the given client and server. See shishi_tkts_find(). Return value: Returns a ticket if found, or NULL.

shishi tkts find for server

[Function] Shishi_tkt * shishi_tkts_find_for_server (Shishi tkts * tkts,

const char * server) tkts: ticket set handle as allocated by shishi_tkts(). server: server name to find ticket for. Description: Short-hand function for searching the ticket set for a ticket for the given server using the default client principal. See shishi_tkts_find_for_

clientserver() and shishi_tkts_find().

Return value: Returns a ticket if found, or NULL.

shishi tkts get tgt

[Function] Shishi_tkt * shishi_tkts_get_tgt (Shishi tkts * tkts,

Shishi tkts hint * hint) tkts: ticket set handle as allocated by shishi_tkts(). hint: structure with characteristics of ticket to begot. Description: Get a ticket granting ticket (TGT) suitable for acquiring ticket matching the hint. I.e., get a TGT for the server realm in the hint structure (hint->serverrealm), or the default realm if the serverrealm field is NULL. Can result in AS exchange. Currently this function do not implement cross realm logic. This function is used by shishi_tkts_get(), which is probably what you really want to use unless you have special needs. Return value: Returns a ticket granting ticket if successful, or NULL if this function is unable to acquire on.

shishi tkts get tgs

[Function] Shishi_tkt * shishi_tkts_get_tgs (Shishi tkts * tkts,

Shishi tkts hint * hint, Shishi tkt * tgt)

tkts: ticket set handle as allocated by shishi_tkts(). hint: structure with characteristics of ticket to begot. tgt: ticket granting ticket to use. Description: Get a ticket via TGS exchange using specified ticket granting ticket. This function is used by shishi_tkts_get(), which is probably what you really want to use unless you have special needs.

Return value: Returns a ticket if successful, or NULL if this function is unable to acquire on.

shishi tkts get

[Function] Shishi_tkt * shishi_tkts_get (Shishi tkts * tkts,

Shishi tkts hint * hint) tkts: ticket set handle as allocated by shishi_tkts(). hint: structure with characteristics of ticket to be found. Description: Get a ticket matching given characteristics. This function first looks in the ticket set for a ticket, then tries to find a suitable TGT, possibly via an AS exchange, using shishi_tkts_get_tgt(), and then uses that TGT in a TGS exchange to get the ticket. Currently this function does not implement cross realm logic. Return value: Returns a ticket if found, or NULL if this function is unable to get the ticket.

shishi tkts get for clientserver

[Function] Shishi_tkt * shishi_tkts_get_for_clientserver

(Shishi tkts * tkts, const char * client, const char * server)

tkts: ticket set handle as allocated by shishi_tkts(). client: client name to get ticket for. server: server name to get ticket for. Description: Short-hand function for getting a ticket for the given client and server.

See shishi_tkts_get().

Return value: Returns a ticket if found, or NULL.

shishi tkts get for server

[Function] Shishi_tkt * shishi_tkts_get_for_server (Shishi tkts * tkts,

const char * server) tkts: ticket set handle as allocated by shishi_tkts(). server: server name to get ticket for. Description: Short-hand function for getting a ticket to the given server and for the default principal client. See shishi_tkts_get(). Return value: Returns a ticket if found, or NULL.

shishi tkts get for localservicepasswd

[Function] Shishi_tkt * shishi_tkts_get_for_localservicepasswd

(Shishi tkts * tkts, const char * service, const char * passwd)

tkts: ticket set handle as allocated by shishi_tkts(). service: service name to get ticket for. passwd: password for the default client principal. Description: Short-hand function for getting a ticket to the given local service, and for the default principal client. The latter’s password is given as argument. See

shishi_tkts_get().

Return value: Returns a ticket if found, or NULL otherwise.

5.4 AP-REQ and AP-REP Functions

The “AP-REQ” and “AP-REP” are ASN.1 structures used by application client and servers to prove to each other who they are. The structures contain auxilliary information, together with an authenticator (see Section 5.11 [Authenticator Functions], page 152) which is the real cryptographic proof. The following illustrates the AP-REQ and AP-REP ASN.1 structures.

AP-REQ          ::= [APPLICATION 14] SEQUENCE {
        pvno            [0] INTEGER (5),
        msg-type        [1] INTEGER (14),
        ap-options      [2] APOptions,
        ticket          [3] Ticket,
        authenticator   [4] EncryptedData {Authenticator,

                                { keyuse-pa-TGSReq-authenticator
                                  | keyuse-APReq-authenticator }}
}

AP-REP          ::= [APPLICATION 15] SEQUENCE {
        pvno            [0] INTEGER (5),
        msg-type        [1] INTEGER (15),
        enc-part        [2] EncryptedData {EncAPRepPart,

                                { keyuse-EncAPRepPart }}
}

EncAPRepPart    ::= [APPLICATION 27] SEQUENCE {
        ctime           [0] KerberosTime,
        cusec           [1] Microseconds,
        subkey          [2] EncryptionKey OPTIONAL,
        seq-number      [3] UInt32 OPTIONAL

}

shishi ap

[Function] int shishi_ap (Shishi * handle, Shishi ap ** ap)

handle: shishi handle as allocated by shishi_init().

ap: pointer to new structure that holds information about AP exchange

Description: Create a new AP exchange with a random subkey of the default encryption type from configuration. Note that there is no guarantee that the receiver will understand that key type, you should probably use shishi_ap_etype() or shishi_ ap_nosubkey() instead. In the future, this function will likely behave as shishi_ap_

     nosubkey() and shishi_ap_nosubkey() will be removed.

shishi ap etype

[Function] int shishi_ap_etype (Shishi * handle, Shishi ap ** ap, int etype)

handle: shishi handle as allocated by shishi_init(). ap: pointer to new structure that holds information about AP exchange etype: encryption type of newly generated random subkey. Description: Create a new AP exchange with a random subkey of indicated encryption type. Return value: Returns SHISHI OK iff successful.

shishi ap nosubkey

[Function] int shishi_ap_nosubkey (Shishi * handle, Shishi ap ** ap)

handle: shishi handle as allocated by shishi_init(). ap: pointer to new structure that holds information about AP exchange Description: Create a new AP exchange without subkey in authenticator. Return value: Returns SHISHI OK iff successful.

shishi ap done

[Function] void shishi_ap_done (Shishi ap * ap)

ap: structure that holds information about AP exchange Description: Deallocate resources associated with AP exchange. This should be called by the application when it no longer need to utilize the AP exchange handle.

shishi ap set tktoptions

[Function] int shishi_ap_set_tktoptions (Shishi ap * ap, Shishi tkt * tkt,
    int options)

ap: structure that holds information about AP exchange tkt: ticket to set in AP. options: AP-REQ options to set in AP. Description: Set the ticket (see shishi_ap_tkt_set()) and set the AP-REQ apop-

tions (see shishi_apreq_options_set()).

Return value: Returns SHISHI OK iff successful.

shishi ap set tktoptionsdata

[Function] int shishi_ap_set_tktoptionsdata (Shishi ap * ap,

Shishi tkt * tkt, int options, const char * data, size t len)

ap: structure that holds information about AP exchange tkt: ticket to set in AP. options: AP-REQ options to set in AP. data: input array with data to checksum in Authenticator. len: length of input array with data to checksum in Authenticator. Description: Set the ticket (see shishi_ap_tkt_set()) and set the AP-REQ apoptions (see shishi_apreq_options_set()) and set the Authenticator checksum data. Return value: Returns SHISHI OK iff successful.

shishi ap set tktoptionsraw

[Function] int shishi_ap_set_tktoptionsraw (Shishi ap * ap,

Shishi tkt * tkt, int options, int32 t cksumtype, const char * data,

size t len) ap: structure that holds information about AP exchange tkt: ticket to set in AP. options: AP-REQ options to set in AP. cksumtype: authenticator checksum type to set in AP. data: input array with data to store in checksum field in Authenticator. len: length of input array with data to store in checksum field in Authenticator.

Description: Set the ticket (see shishi_ap_tkt_set()) and set the AP-REQ apoptions (see shishi_apreq_options_set()) and set the raw Authenticator checksum data.

Return value: Returns SHISHI OK iff successful.

shishi ap set tktoptionsasn1usage

[Function] int shishi_ap_set_tktoptionsasn1usage (Shishi ap * ap,

Shishi tkt * tkt, int options, Shishi asn1 node, const char * field,

   int authenticatorcksumkeyusage, int authenticatorkeyusage)

ap: structure that holds information about AP exchange tkt: ticket to set in AP. options: AP-REQ options to set in AP. node: input ASN.1 structure to store as authenticator checksum data. field: field in ASN.1 structure to use. authenticatorcksumkeyusage: key usage for checksum in authenticator. authenticatorkeyusage: key usage for authenticator.

Description: Set ticket, options and authenticator checksum data using shishi_ap_ set_tktoptionsdata(). The authenticator checksum data is the DER encoding of the ASN.1 field provided.

Return value: Returns SHISHI OK iff successful.

shishi ap tktoptions

[Function] int shishi_ap_tktoptions (Shishi * handle, Shishi ap ** ap,

Shishi tkt * tkt, int options)

handle: shishi handle as allocated by shishi_init(). ap: pointer to new structure that holds information about AP exchange tkt: ticket to set in newly created AP. options: AP-REQ options to set in newly created AP.

Description: Create a new AP exchange using shishi_ap(), and set the ticket and AP-REQ apoptions using shishi_ap_set_tktoptions(). A random session key is added to the authenticator, using the same keytype as the ticket.

Return value: Returns SHISHI OK iff successful.

shishi ap tktoptionsdata

[Function] int shishi_ap_tktoptionsdata (Shishi * handle, Shishi ap ** ap,

Shishi tkt * tkt, int options, const char * data, size t len)

handle: shishi handle as allocated by shishi_init(). ap: pointer to new structure that holds information about AP exchange tkt: ticket to set in newly created AP. options: AP-REQ options to set in newly created AP. data: input array with data to checksum in Authenticator. len: length of input array with data to checksum in Authenticator.

Description: Create a new AP exchange using shishi_ap(), and set the ticket, AP-REQ apoptions and the Authenticator checksum data using shishi_ap_set_ tktoptionsdata(). A random session key is added to the authenticator, using the same keytype as the ticket.

Return value: Returns SHISHI OK iff successful.

shishi ap tktoptionsraw

[Function] int shishi_ap_tktoptionsraw (Shishi * handle, Shishi ap ** ap,

Shishi tkt * tkt, int options, int32 t cksumtype, const char * data,

size t len) handle: shishi handle as allocated by shishi_init(). ap: pointer to new structure that holds information about AP exchange tkt: ticket to set in newly created AP. options: AP-REQ options to set in newly created AP. cksumtype: authenticator checksum type to set in AP. data: input array with data to store in checksum field in Authenticator. len: length of input array with data to store in checksum field in Authenticator.

Description: Create a new AP exchange using shishi_ap(), and set the ticket, AP-REQ apoptions and the raw Authenticator checksum data field using shishi_ap_ set_tktoptionsraw(). A random session key is added to the authenticator, using the same keytype as the ticket.

Return value: Returns SHISHI OK iff successful.

shishi ap etype tktoptionsdata

[Function] int shishi_ap_etype_tktoptionsdata (Shishi * handle,

Shishi ap ** ap, int32 t etype, Shishi tkt * tkt, int options, const char * data, size t len)

handle: shishi handle as allocated by shishi_init(). ap: pointer to new structure that holds information about AP exchange etype: encryption type of newly generated random subkey. tkt: ticket to set in newly created AP. options: AP-REQ options to set in newly created AP. data: input array with data to checksum in Authenticator. len: length of input array with data to checksum in Authenticator.

Description: Create a new AP exchange using shishi_ap(), and set the ticket, AP-REQ apoptions and the Authenticator checksum data using shishi_ap_set_ tktoptionsdata(). A random session key is added to the authenticator, using the same keytype as the ticket. Return value: Returns SHISHI OK iff successful.

shishi ap tktoptionsasn1usage

[Function] int shishi_ap_tktoptionsasn1usage (Shishi * handle,

Shishi ap ** ap, Shishi tkt * tkt, int options, Shishi asn1 node,

   const char * field, int authenticatorcksumkeyusage,
    int authenticatorkeyusage)

handle: shishi handle as allocated by shishi_init(). ap: pointer to new structure that holds information about AP exchange tkt: ticket to set in newly created AP. options: AP-REQ options to set in newly created AP. node: input ASN.1 structure to store as authenticator checksum data. field: field in ASN.1 structure to use. authenticatorcksumkeyusage: key usage for checksum in authenticator. authenticatorkeyusage: key usage for authenticator. Description: Create a new AP exchange using shishi_ap(), and set ticket, options and authenticator checksum data from the DER encoding of the ASN.1 field using

shishi_ap_set_tktoptionsasn1usage(). A random session key is added to the

authenticator, using the same keytype as the ticket. Return value: Returns SHISHI OK iff successful.

shishi ap tkt

[Function] Shishi_tkt * shishi_ap_tkt (Shishi ap * ap)

ap: structure that holds information about AP exchange Description: Get Ticket from AP exchange. Return value: Returns the ticket from the AP exchange, or NULL if not yet set or an error occured.

shishi ap tkt set

[Function] void shishi_ap_tkt_set (Shishi ap * ap, Shishi tkt * tkt)

ap: structure that holds information about AP exchange tkt: ticket to store in AP. Description: Set the Ticket in the AP exchange.

shishi ap authenticator cksumdata

[Function] int shishi_ap_authenticator_cksumdata (Shishi ap * ap,

char * out, size t * len)

ap: structure that holds information about AP exchange out: output array that holds authenticator checksum data. len: on input, maximum length of output array that holds authenticator checksum data, on output actual length of output array that holds authenticator checksum data.

Description: Get checksum data from Authenticator.

Return value: Returns SHISHI_OK if successful, or SHISHI_TOO_SMALL_BUFFER if buffer provided was too small (then len will hold necessary buffer size).

shishi ap authenticator cksumdata set

[Function] void shishi_ap_authenticator_cksumdata_set (Shishi ap * ap,
   const char * authenticatorcksumdata,
    size t authenticatorcksumdatalen)

ap: structure that holds information about AP exchange authenticatorcksumdata: input array with data to compute checksum on and store in Authenticator in AP-REQ. authenticatorcksumdatalen: length of input array with data to compute checksum on and store in Authenticator in AP-REQ.

Description: Set the Authenticator Checksum Data in the AP exchange. This is the data that will be checksumed, and the checksum placed in the checksum field. It is not the actual checksum field. See also shishi ap authenticator cksumraw set.

shishi ap authenticator cksumraw set

[Function] void shishi_ap_authenticator_cksumraw_set (Shishi ap * ap,
   int32 t authenticatorcksumtype,
   const char * authenticatorcksumraw,
    size t authenticatorcksumrawlen)

ap: structure that holds information about AP exchange authenticatorcksumtype: authenticator checksum type to set in AP. authenticatorcksumraw: input array with authenticator checksum field value to set in Authenticator in AP-REQ. authenticatorcksumrawlen: length of input array with authenticator checksum field value to set in Authenticator in AP-REQ.

Description: Set the Authenticator Checksum Data in the AP exchange. This is the actual checksum field, not data to compute checksum on and then store in the checksum field. See also shishi ap authenticator cksumdata set.

shishi ap authenticator cksumtype

[Function] int32_t shishi_ap_authenticator_cksumtype (Shishi ap * ap)

ap: structure that holds information about AP exchange

Description: Get the Authenticator Checksum Type in the AP exchange.

Return value: Return the authenticator checksum type.

shishi ap authenticator cksumtype set

[Function] void shishi_ap_authenticator_cksumtype_set (Shishi ap * ap,
    int32 t cksumtype)

ap: structure that holds information about AP exchange cksumtype: authenticator checksum type to set in AP. Description: Set the Authenticator Checksum Type in the AP exchange.

shishi ap authenticator

[Function] Shishi_asn1 shishi_ap_authenticator (Shishi ap * ap)

ap: structure that holds information about AP exchange Description: Get ASN.1 Authenticator structure from AP exchange. Return value: Returns the Authenticator from the AP exchange, or NULL if not yet set or an error occured.

shishi ap authenticator set

[Function] void shishi_ap_authenticator_set (Shishi ap * ap,
    Shishi asn1 authenticator)

ap: structure that holds information about AP exchange authenticator: authenticator to store in AP. Description: Set the Authenticator in the AP exchange.

shishi ap req

[Function] Shishi_asn1 shishi_ap_req (Shishi ap * ap)

ap: structure that holds information about AP exchange Description: Get ASN.1 AP-REQ structure from AP exchange. Return value: Returns the AP-REQ from the AP exchange, or NULL if not yet set or an error occured.

shishi ap req set

[Function] void shishi_ap_req_set (Shishi ap * ap, Shishi asn1 apreq)

ap: structure that holds information about AP exchange apreq: apreq to store in AP. Description: Set the AP-REQ in the AP exchange.

shishi ap req der

[Function] int shishi_ap_req_der (Shishi ap * ap, char ** out,

size t * outlen) ap: structure that holds information about AP exchange out: pointer to output array with der encoding of AP-REQ. outlen: pointer to length of output array with der encoding of AP-REQ. Description: Build AP-REQ using shishi_ap_req_build() and DER encode it. out is allocated by this function, and it is the responsibility of caller to deallocate it. Return value: Returns SHISHI OK iff successful.

shishi ap req der set

[Function] int shishi_ap_req_der_set (Shishi ap * ap, char * der,
    size t derlen)

ap: structure that holds information about AP exchange der: input array with DER encoded AP-REQ. derlen: length of input array with DER encoded AP-REQ. Description: DER decode AP-REQ and set it AP exchange. If decoding fails, the AP-REQ in the AP exchange is lost. Return value: Returns SHISHI OK.

shishi ap req build

[Function] int shishi_ap_req_build (Shishi ap * ap)

ap: structure that holds information about AP exchange Description: Checksum data in authenticator and add ticket and authenticator to AP-REQ. Return value: Returns SHISHI OK iff successful.

shishi ap req decode

[Function] int shishi_ap_req_decode (Shishi ap * ap)

ap: structure that holds information about AP exchange Description: Decode ticket in AP-REQ and set the Ticket fields in the AP exchange. Return value: Returns SHISHI OK iff successful.

shishi ap req process keyusage

[Function] int shishi_ap_req_process_keyusage (Shishi ap * ap,

Shishi key * key, int32 t keyusage)

ap: structure that holds information about AP exchange key: cryptographic key used to decrypt ticket in AP-REQ. keyusage: key usage to use during decryption, for normal AP-REQ’s this is normally SHISHI KEYUSAGE APREQ AUTHENTICATOR, for AP-REQ’s part of TGS-REQ’s, this is normally SHISHI KEYUSAGE TGSREQ APREQ AUTHENTICATOR. Description: Decrypt ticket in AP-REQ using supplied key and decrypt Authenticator in AP-REQ using key in decrypted ticket, and on success set the Ticket and Authenticator fields in the AP exchange. Return value: Returns SHISHI OK iff successful.

shishi ap req process

[Function] int shishi_ap_req_process (Shishi ap * ap, Shishi key * key)

ap: structure that holds information about AP exchange key: cryptographic key used to decrypt ticket in AP-REQ. Description: Decrypt ticket in AP-REQ using supplied key and decrypt Authenticator in AP-REQ using key in decrypted ticket, and on success set the Ticket and Authenticator fields in the AP exchange.

Return value: Returns SHISHI OK iff successful.

shishi ap req asn1

[Function] int shishi_ap_req_asn1 (Shishi ap * ap, Shishi asn1 * apreq)

ap: structure that holds information about AP exchange apreq: output AP-REQ variable.

Description: Build AP-REQ using shishi_ap_req_build() and return it.

Return value: Returns SHISHI OK iff successful.

shishi ap key

[Function] Shishi_key * shishi_ap_key (Shishi ap * ap)

ap: structure that holds information about AP exchange

Description: Extract the application key from AP. If subkeys are used, it is taken from the Authenticator, otherwise the session key is used.

Return value: Return application key from AP.

shishi ap rep

[Function] Shishi_asn1 shishi_ap_rep (Shishi ap * ap)

ap: structure that holds information about AP exchange

Description: Get ASN.1 AP-REP structure from AP exchange.

Return value: Returns the AP-REP from the AP exchange, or NULL if not yet set or an error occured.

shishi ap rep set

[Function] void shishi_ap_rep_set (Shishi ap * ap, Shishi asn1 aprep)

ap: structure that holds information about AP exchange aprep: aprep to store in AP.

Description: Set the AP-REP in the AP exchange.

shishi ap rep der

[Function] int shishi_ap_rep_der (Shishi ap * ap, char ** out,

size t * outlen) ap: structure that holds information about AP exchange out: output array with newly allocated DER encoding of AP-REP. outlen: length of output array with DER encoding of AP-REP.

Description: Build AP-REP using shishi_ap_rep_build() and DER encode it. out is allocated by this function, and it is the responsibility of caller to deallocate it.

Return value: Returns SHISHI OK iff successful.

shishi ap rep der set

[Function] int shishi_ap_rep_der_set (Shishi ap * ap, char * der,
    size t derlen)

ap: structure that holds information about AP exchange der: input array with DER encoded AP-REP. derlen: length of input array with DER encoded AP-REP.

Description: DER decode AP-REP and set it AP exchange. If decoding fails, the AP-REP in the AP exchange remains.

Return value: Returns SHISHI OK.

shishi ap rep build

[Function] int shishi_ap_rep_build (Shishi ap * ap)

ap: structure that holds information about AP exchange

Description: Checksum data in authenticator and add ticket and authenticator to AP-REP.

Return value: Returns SHISHI OK iff successful.

shishi ap rep asn1

[Function] int shishi_ap_rep_asn1 (Shishi ap * ap, Shishi asn1 * aprep)

ap: structure that holds information about AP exchange aprep: output AP-REP variable.

Description: Build AP-REP using shishi_ap_rep_build() and return it.

Return value: Returns SHISHI OK iff successful.

shishi ap rep verify

[Function] int shishi_ap_rep_verify (Shishi ap * ap)

ap: structure that holds information about AP exchange

Description: Verify AP-REP compared to Authenticator.

Return value: Returns SHISHI OK, SHISHI APREP VERIFY FAILED or an error.

shishi ap rep verify der

[Function] int shishi_ap_rep_verify_der (Shishi ap * ap, char * der,
    size t derlen)

ap: structure that holds information about AP exchange der: input array with DER encoded AP-REP. derlen: length of input array with DER encoded AP-REP.

Description: DER decode AP-REP and set it in AP exchange using shishi_ap_rep_

der_set() and verify it using shishi_ap_rep_verify().

Return value: Returns SHISHI OK, SHISHI APREP VERIFY FAILED or an error.

shishi ap rep verify asn1

[Function] int shishi_ap_rep_verify_asn1 (Shishi ap * ap,

Shishi asn1 aprep) ap: structure that holds information about AP exchange aprep: input AP-REP.

Description: Set the AP-REP in the AP exchange using shishi_ap_rep_set() and

verify it using shishi_ap_rep_verify().

Return value: Returns SHISHI OK, SHISHI APREP VERIFY FAILED or an error.

shishi ap encapreppart

[Function] Shishi_asn1 shishi_ap_encapreppart (Shishi ap * ap)

ap: structure that holds information about AP exchange

Description: Get ASN.1 EncAPRepPart structure from AP exchange.

Return value: Returns the EncAPREPPart from the AP exchange, or NULL if not yet set or an error occured.

shishi ap encapreppart set

[Function] void shishi_ap_encapreppart_set (Shishi ap * ap,
    Shishi asn1 encapreppart)

ap: structure that holds information about AP exchange encapreppart: EncAPRepPart to store in AP.

Description: Set the EncAPRepPart in the AP exchange.

shishi ap option2string

[Function] const char * shishi_ap_option2string

(Shishi apoptions option) option: enumerated AP-Option type, see Shishi apoptions.

Description: Convert AP-Option type to AP-Option name string. Note that option must be just one of the AP-Option types, it cannot be an binary ORed indicating several AP-Options.

Return value: Returns static string with name of AP-Option that must not be deallocated, or "unknown" if AP-Option was not understood.

shishi ap string2option

[Function] Shishi_apoptions shishi_ap_string2option (const char * str)

str: zero terminated character array with name of AP-Option, e.g. "use-session-key".

Description: Convert AP-Option name to AP-Option type.

Return value: Returns enumerated type member corresponding to AP-Option, or 0 if string was not understood.

shishi apreq

[Function] Shishi_asn1 shishi_apreq (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: This function creates a new AP-REQ, populated with some default values. Return value: Returns the AP-REQ or NULL on failure.

shishi apreq print

[Function] int shishi_apreq_print (Shishi * handle, FILE * fh,

Shishi asn1 apreq) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. apreq: AP-REQ to print. Description: Print ASCII armored DER encoding of AP-REQ to file. Return value: Returns SHISHI OK iff successful.

shishi apreq save

[Function] int shishi_apreq_save (Shishi * handle, FILE * fh,

Shishi asn1 apreq) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. apreq: AP-REQ to save. Description: Save DER encoding of AP-REQ to file. Return value: Returns SHISHI OK iff successful.

shishi apreq to file

[Function] int shishi_apreq_to_file (Shishi * handle, Shishi asn1 apreq,
   int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ to save. filetype: input variable specifying type of file to be written, see Shishi filetype. filename: input variable with filename to write to. Description: Write AP-REQ to file in specified TYPE. The file will be truncated if it exists. Return value: Returns SHISHI OK iff successful.

shishi apreq parse

[Function] int shishi_apreq_parse (Shishi * handle, FILE * fh,

Shishi asn1 * apreq) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. apreq: output variable with newly allocated AP-REQ.

Description: Read ASCII armored DER encoded AP-REQ from file and populate given variable.

Return value: Returns SHISHI OK iff successful.

shishi apreq read

[Function] int shishi_apreq_read (Shishi * handle, FILE * fh,

Shishi asn1 * apreq) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. apreq: output variable with newly allocated AP-REQ.

Description: Read DER encoded AP-REQ from file and populate given variable.

Return value: Returns SHISHI OK iff successful.

shishi apreq from file

[Function] int shishi_apreq_from_file (Shishi * handle,

Shishi asn1 * apreq, int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). apreq: output variable with newly allocated AP-REQ. filetype: input variable specifying type of file to be read, see Shishi filetype. filename: input variable with filename to read from.

Description: Read AP-REQ from file in specified TYPE.

Return value: Returns SHISHI OK iff successful.

shishi apreq set authenticator

[Function] int shishi_apreq_set_authenticator (Shishi * handle,

Shishi asn1 apreq, int32 t etype, uint32 t kvno, const char * buf,

    size t buflen)

handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ to add authenticator field to. etype: encryption type used to encrypt authenticator. kvno: version of the key used to encrypt authenticator. buf : input array with encrypted authenticator. buflen: size of input array with encrypted authenticator.

Description: Set the encrypted authenticator field in the AP-REP. The encrypted data is usually created by calling shishi_encrypt() on the DER encoded authenticator. To save time, you may want to use shishi_apreq_add_authenticator() instead, which calculates the encrypted data and calls this function in one step.

Return value: Returns SHISHI OK on success.

shishi apreq add authenticator

[Function] int shishi_apreq_add_authenticator (Shishi * handle,

Shishi asn1 apreq, Shishi key * key, int keyusage,

    Shishi asn1 authenticator)

handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ to add authenticator field to. key: key to to use for encryption. keyusage: cryptographic key usage value to use in encryption. authenticator: authenticator as allocated by shishi_authenticator().

Description: Encrypts DER encoded authenticator using key and store it in the AP-REQ.

Return value: Returns SHISHI OK iff successful.

shishi apreq set ticket

[Function] int shishi_apreq_set_ticket (Shishi * handle,

Shishi asn1 apreq, Shishi asn1 ticket)

handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ to add ticket field to. ticket: input ticket to copy into AP-REQ ticket field.

Description: Copy ticket into AP-REQ.

Return value: Returns SHISHI OK iff successful.

shishi apreq options

[Function] int shishi_apreq_options (Shishi * handle, Shishi asn1 apreq,

uint32 t * flags) handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ to get options from. flags: Output integer containing options from AP-REQ.

Description: Extract the AP-Options from AP-REQ into output integer.

Return value: Returns SHISHI OK iff successful.

shishi apreq use session key p

[Function] int shishi_apreq_use_session_key_p (Shishi * handle,

Shishi asn1 apreq) handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ as allocated by shishi_apreq().

Description: Return non-0 iff the "Use session key" option is set in the AP-REQ.

shishi apreq mutual required p

[Function] int shishi_apreq_mutual_required_p (Shishi * handle,

Shishi asn1 apreq) handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ as allocated by shishi_apreq(). Description: Return non-0 iff the "Mutual required" option is set in the AP-REQ. Return value: Returns SHISHI OK iff successful.

shishi apreq options set

[Function] int shishi_apreq_options_set (Shishi * handle,

Shishi asn1 apreq, uint32 t options)

handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ as allocated by shishi_apreq(). options: Options to set in AP-REQ. Description: Set the AP-Options in AP-REQ to indicate integer. Return value: Returns SHISHI OK iff successful.

shishi apreq options add

[Function] int shishi_apreq_options_add (Shishi * handle,

Shishi asn1 apreq, uint32 t option)

handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ as allocated by shishi_apreq(). option: Options to add in AP-REQ. Description: Add the AP-Options in AP-REQ. Options not set in input parameter option are preserved in the AP-REQ. Return value: Returns SHISHI OK iff successful.

shishi apreq options remove

[Function] int shishi_apreq_options_remove (Shishi * handle,

Shishi asn1 apreq, uint32 t option)

handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ as allocated by shishi_apreq(). option: Options to remove from AP-REQ. Description: Remove the AP-Options from AP-REQ. Options not set in input parameter option are preserved in the AP-REQ. Return value: Returns SHISHI OK iff successful.

shishi apreq get authenticator etype

[Function] int shishi_apreq_get_authenticator_etype (Shishi * handle,

Shishi asn1 apreq, int32 t * etype)

handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ variable to get value from. etype: output variable that holds the value.

Description: Extract AP-REQ.authenticator.etype. Return value: Returns SHISHI OK iff successful.

shishi apreq get ticket

[Function] int shishi_apreq_get_ticket (Shishi * handle,

Shishi asn1 apreq, Shishi asn1 * ticket)

handle: shishi handle as allocated by shishi_init(). apreq: AP-REQ variable to get ticket from. ticket: output variable to hold extracted ticket. Description: Extract ticket from AP-REQ. Return value: Returns SHISHI OK iff successful.

shishi aprep

[Function] Shishi_asn1 shishi_aprep (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: This function creates a new AP-REP, populated with some default values. Return value: Returns the authenticator or NULL on failure.

shishi aprep print

[Function] int shishi_aprep_print (Shishi * handle, FILE * fh,

Shishi asn1 aprep) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. aprep: AP-REP to print. Description: Print ASCII armored DER encoding of AP-REP to file. Return value: Returns SHISHI OK iff successful.

shishi aprep save

[Function] int shishi_aprep_save (Shishi * handle, FILE * fh,

Shishi asn1 aprep) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. aprep: AP-REP to save. Description: Save DER encoding of AP-REP to file. Return value: Returns SHISHI OK iff successful.

shishi aprep to file

[Function] int shishi_aprep_to_file (Shishi * handle, Shishi asn1 aprep,
   int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). aprep: AP-REP to save. filetype: input variable specifying type of file to be written, see Shishi filetype. filename: input variable with filename to write to. Description: Write AP-REP to file in specified TYPE. The file will be truncated if it exists. Return value: Returns SHISHI OK iff successful.

shishi aprep parse

[Function] int shishi_aprep_parse (Shishi * handle, FILE * fh,

Shishi asn1 * aprep) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. aprep: output variable with newly allocated AP-REP. Description: Read ASCII armored DER encoded AP-REP from file and populate given variable. Return value: Returns SHISHI OK iff successful.

shishi aprep read

[Function] int shishi_aprep_read (Shishi * handle, FILE * fh,

Shishi asn1 * aprep) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. aprep: output variable with newly allocated AP-REP. Description: Read DER encoded AP-REP from file and populate given variable. Return value: Returns SHISHI OK iff successful.

shishi aprep from file

[Function] int shishi_aprep_from_file (Shishi * handle,

Shishi asn1 * aprep, int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). aprep: output variable with newly allocated AP-REP. filetype: input variable specifying type of file to be read, see Shishi filetype. filename: input variable with filename to read from. Description: Read AP-REP from file in specified TYPE. Return value: Returns SHISHI OK iff successful.

shishi aprep get enc part etype

[Function] int shishi_aprep_get_enc_part_etype (Shishi * handle,

Shishi asn1 aprep, int32 t * etype)

handle: shishi handle as allocated by shishi_init(). aprep: AP-REP variable to get value from. etype: output variable that holds the value. Description: Extract AP-REP.enc-part.etype. Return value: Returns SHISHI OK iff successful.

shishi encapreppart

[Function] Shishi_asn1 shishi_encapreppart (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: This function creates a new EncAPRepPart, populated with some default values. It uses the current time as returned by the system for the ctime and cusec fields. Return value: Returns the encapreppart or NULL on failure.

shishi encapreppart print

[Function] int shishi_encapreppart_print (Shishi * handle, FILE * fh,
    Shishi asn1 encapreppart)

handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. encapreppart: EncAPRepPart to print. Description: Print ASCII armored DER encoding of EncAPRepPart to file. Return value: Returns SHISHI OK iff successful.

shishi encapreppart save

[Function] int shishi_encapreppart_save (Shishi * handle, FILE * fh,
    Shishi asn1 encapreppart)

handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. encapreppart: EncAPRepPart to save. Description: Save DER encoding of EncAPRepPart to file. Return value: Returns SHISHI OK iff successful.

shishi encapreppart to file

[Function] int shishi_encapreppart_to_file (Shishi * handle,
   Shishi asn1 encapreppart, int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). encapreppart: EncAPRepPart to save. filetype: input variable specifying type of file to be written, see Shishi filetype. filename: input variable with filename to write to. Description: Write EncAPRepPart to file in specified TYPE. The file will be truncated if it exists. Return value: Returns SHISHI OK iff successful.

shishi encapreppart parse

[Function] int shishi_encapreppart_parse (Shishi * handle, FILE * fh,
    Shishi asn1 * encapreppart)

handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. encapreppart: output variable with newly allocated EncAPRepPart.

Description: Read ASCII armored DER encoded EncAPRepPart from file and populate given variable.

Return value: Returns SHISHI OK iff successful.

shishi encapreppart read

[Function] int shishi_encapreppart_read (Shishi * handle, FILE * fh,
    Shishi asn1 * encapreppart)

handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. encapreppart: output variable with newly allocated EncAPRepPart.

Description: Read DER encoded EncAPRepPart from file and populate given variable.

Return value: Returns SHISHI OK iff successful.

shishi encapreppart from file

[Function] int shishi_encapreppart_from_file (Shishi * handle,
   Shishi asn1 * encapreppart, int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). encapreppart: output variable with newly allocated EncAPRepPart. filetype: input variable specifying type of file to be read, see Shishi filetype. filename: input variable with filename to read from.

Description: Read EncAPRepPart from file in specified TYPE.

Return value: Returns SHISHI OK iff successful.

shishi encapreppart get key

[Function] int shishi_encapreppart_get_key (Shishi * handle,

Shishi asn1 encapreppart, Shishi key ** key)

handle: shishi handle as allocated by shishi_init(). encapreppart: input EncAPRepPart variable. key: newly allocated key.

Description: Extract the subkey from the encrypted AP-REP part.

Return value: Returns SHISHI_OK iff successful.

shishi encapreppart ctime

[Function] int shishi_encapreppart_ctime (Shishi * handle,

Shishi asn1 encapreppart, char ** t)

handle: shishi handle as allocated by shishi_init(). encapreppart: EncAPRepPart as allocated by shishi_encapreppart(). t: newly allocated zero-terminated character array with client time.

Description: Extract client time from EncAPRepPart.

Return value: Returns SHISHI OK iff successful.

shishi encapreppart ctime set

[Function] int shishi_encapreppart_ctime_set (Shishi * handle,

Shishi asn1 encapreppart, const char * t)

handle: shishi handle as allocated by shishi_init(). encapreppart: EncAPRepPart as allocated by shishi_encapreppart(). t: string with generalized time value to store in EncAPRepPart. Description: Store client time in EncAPRepPart. Return value: Returns SHISHI OK iff successful.

shishi encapreppart cusec get

[Function] int shishi_encapreppart_cusec_get (Shishi * handle,

Shishi asn1 encapreppart, uint32 t * cusec)

handle: shishi handle as allocated by shishi_init(). encapreppart: EncAPRepPart as allocated by shishi_encapreppart(). cusec: output integer with client microseconds field. Description: Extract client microseconds field from EncAPRepPart. Return value: Returns SHISHI OK iff successful.

shishi encapreppart cusec set

[Function] int shishi_encapreppart_cusec_set (Shishi * handle,

Shishi asn1 encapreppart, uint32 t cusec)

handle: shishi handle as allocated by shishi_init(). encapreppart: EncAPRepPart as allocated by shishi_encapreppart(). cusec: client microseconds to set in authenticator, 0-999999. Description: Set the cusec field in the Authenticator. Return value: Returns SHISHI OK iff successful.

shishi encapreppart seqnumber get

[Function] int shishi_encapreppart_seqnumber_get (Shishi * handle,
   Shishi asn1 encapreppart, uint32 t * seqnumber)

handle: shishi handle as allocated by shishi_init(). encapreppart: EncAPRepPart as allocated by shishi_encapreppart(). seqnumber: output integer with sequence number field. Description: Extract sequence number field from EncAPRepPart. Return value: Returns SHISHI OK iff successful.

shishi encapreppart seqnumber remove

[Function] int shishi_encapreppart_seqnumber_remove (Shishi * handle,
    Shishi asn1 encapreppart)

handle: shishi handle as allocated by shishi_init(). encapreppart: encapreppart as allocated by shishi_encapreppart(). Description: Remove sequence number field in EncAPRepPart. Return value: Returns SHISHI_OK iff successful.

shishi encapreppart seqnumber set

[Function] int shishi_encapreppart_seqnumber_set (Shishi * handle,
   Shishi asn1 encapreppart, uint32 t seqnumber)

handle: shishi handle as allocated by shishi_init(). encapreppart: encapreppart as allocated by shishi_encapreppart(). seqnumber: integer with sequence number field to store in encapreppart. Description: Store sequence number field in EncAPRepPart. Return value: Returns SHISHI_OK iff successful.

shishi encapreppart time copy

[Function] int shishi_encapreppart_time_copy (Shishi * handle,
   Shishi asn1 encapreppart, Shishi asn1 authenticator)

handle: shishi handle as allocated by shishi_init(). encapreppart: EncAPRepPart as allocated by shishi_encapreppart(). authenticator: Authenticator to copy time fields from. Description: Copy time fields from Authenticator into EncAPRepPart. Return value: Returns SHISHI OK iff successful.

5.5 SAFE and PRIV Functions

The “KRB-SAFE” is an ASN.1 structure used by application client and servers to exchange integrity protected data. The integrity protection is keyed, usually with a key agreed on via the AP exchange (see Section 5.4 [AP-REQ and AP-REP Functions], page 71). The following illustrates the KRB-SAFE ASN.1 structure.

KRB-SAFE        ::= [APPLICATION 20] SEQUENCE {
      pvno            [0] INTEGER (5),
      msg-type        [1] INTEGER (20),
      safe-body       [2] KRB-SAFE-BODY,
      cksum           [3] Checksum
}
KRB-SAFE-BODY   ::= SEQUENCE {
      user-data       [0] OCTET STRING,
      timestamp       [1] KerberosTime OPTIONAL,
      usec            [2] Microseconds OPTIONAL,
      seq-number      [3] UInt32 OPTIONAL,
      s-address       [4] HostAddress,
      r-address       [5] HostAddress OPTIONAL
}

shishi safe

[Function] int shishi_safe (Shishi * handle, Shishi safe ** safe)

handle: shishi handle as allocated by shishi_init(). safe: pointer to new structure that holds information about SAFE exchange Description: Create a new SAFE exchange. Return value: Returns SHISHI OK iff successful.

shishi safe done

[Function] void shishi_safe_done (Shishi safe * safe)

safe: structure that holds information about SAFE exchange Description: Deallocate resources associated with SAFE exchange. This should be called by the application when it no longer need to utilize the SAFE exchange handle.

shishi safe key

[Function] Shishi_key * shishi_safe_key (Shishi safe * safe)

safe: structure that holds information about SAFE exchange Description: Get key structured from SAFE exchange. Return value: Returns the key used in the SAFE exchange, or NULL if not yet set or an error occured.

shishi safe key set

[Function] void shishi_safe_key_set (Shishi safe * safe, Shishi key * key)

safe: structure that holds information about SAFE exchange key: key to store in SAFE. Description: Set the Key in the SAFE exchange.

shishi safe safe

[Function] Shishi_asn1 shishi_safe_safe (Shishi safe * safe)

safe: structure that holds information about SAFE exchange Description: Get ASN.1 SAFE structured from SAFE exchange. Return value: Returns the ASN.1 safe in the SAFE exchange, or NULL if not yet set or an error occured.

shishi safe safe set

[Function] void shishi_safe_safe_set (Shishi safe * safe,

Shishi asn1 asn1safe) safe: structure that holds information about SAFE exchange asn1safe: KRB-SAFE to store in SAFE exchange. Description: Set the KRB-SAFE in the SAFE exchange.

shishi safe safe der

[Function] int shishi_safe_safe_der (Shishi safe * safe, char ** out,

size t * outlen) safe: safe as allocated by shishi_safe(). out: output array with newly allocated DER encoding of SAFE. outlen: length of output array with DER encoding of SAFE.

Description: DER encode SAFE structure. Typically shishi_safe_build() is used to build the SAFE structure first. out is allocated by this function, and it is the responsibility of caller to deallocate it. Return value: Returns SHISHI OK iff successful.

shishi safe safe der set

[Function] int shishi_safe_safe_der_set (Shishi safe * safe, char * der,
    size t derlen)

safe: safe as allocated by shishi_safe(). der: input array with DER encoded KRB-SAFE. derlen: length of input array with DER encoded KRB-SAFE. Description: DER decode KRB-SAFE and set it SAFE exchange. If decoding fails, the KRB-SAFE in the SAFE exchange remains. Return value: Returns SHISHI OK.

shishi safe print

[Function] int shishi_safe_print (Shishi * handle, FILE * fh,

Shishi asn1 safe) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. safe: SAFE to print. Description: Print ASCII armored DER encoding of SAFE to file. Return value: Returns SHISHI OK iff successful.

shishi safe save

[Function] int shishi_safe_save (Shishi * handle, FILE * fh,

Shishi asn1 safe) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. safe: SAFE to save. Description: Save DER encoding of SAFE to file. Return value: Returns SHISHI OK iff successful.

shishi safe to file

[Function] int shishi_safe_to_file (Shishi * handle, Shishi asn1 safe,
   int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). safe: SAFE to save. filetype: input variable specifying type of file to be written, see Shishi filetype. filename: input variable with filename to write to. Description: Write SAFE to file in specified TYPE. The file will be truncated if it exists. Return value: Returns SHISHI OK iff successful.

shishi safe parse

[Function] int shishi_safe_parse (Shishi * handle, FILE * fh,

Shishi asn1 * safe) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. safe: output variable with newly allocated SAFE.

Description: Read ASCII armored DER encoded SAFE from file and populate given variable.

Return value: Returns SHISHI OK iff successful.

shishi safe read

[Function] int shishi_safe_read (Shishi * handle, FILE * fh,

Shishi asn1 * safe) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. safe: output variable with newly allocated SAFE.

Description: Read DER encoded SAFE from file and populate given variable.

Return value: Returns SHISHI OK iff successful.

shishi safe from file

[Function] int shishi_safe_from_file (Shishi * handle, Shishi asn1 * safe,
   int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). safe: output variable with newly allocated SAFE. filetype: input variable specifying type of file to be read, see Shishi filetype. filename: input variable with filename to read from.

Description: Read SAFE from file in specified TYPE.

Return value: Returns SHISHI OK iff successful.

shishi safe cksum

[Function] int shishi_safe_cksum (Shishi * handle, Shishi asn1 safe,
   int32 t * cksumtype, char ** cksum, size t * cksumlen)

handle: shishi handle as allocated by shishi_init(). safe: safe as allocated by shishi_safe(). cksumtype: output checksum type. cksum: output array with newly allocated checksum data from SAFE. cksumlen: output size of output checksum data buffer.

Description: Read checksum value from KRB-SAFE. cksum is allocated by this function, and it is the responsibility of caller to deallocate it.

Return value: Returns SHISHI OK iff successful.

shishi safe set cksum

[Function] int shishi_safe_set_cksum (Shishi * handle, Shishi asn1 safe,

int32 t cksumtype, const char * cksum, size t cksumlen)

handle: shishi handle as allocated by shishi_init(). safe: safe as allocated by shishi_safe(). cksumtype: input checksum type to store in SAFE. cksum: input checksum data to store in SAFE. cksumlen: size of input checksum data to store in SAFE. Description: Store checksum value in SAFE. A checksum is usually created by calling shishi_checksum() on some application specific data using the key from the ticket that is being used. To save time, you may want to use shishi_safe_build() instead, which calculates the checksum and calls this function in one step. Return value: Returns SHISHI OK iff successful.

shishi safe user data

[Function] int shishi_safe_user_data (Shishi * handle, Shishi asn1 safe,
   char ** userdata, size t * userdatalen)

handle: shishi handle as allocated by shishi_init(). safe: safe as allocated by shishi_safe(). userdata: output array with newly allocated user data from KRB-SAFE. userdatalen: output size of output user data buffer. Description: Read user data value from KRB-SAFE. userdata is allocated by this function, and it is the responsibility of caller to deallocate it. Return value: Returns SHISHI OK iff successful.

shishi safe set user data

[Function] int shishi_safe_set_user_data (Shishi * handle,

Shishi asn1 safe, const char * userdata, size t userdatalen)

handle: shishi handle as allocated by shishi_init(). safe: safe as allocated by shishi_safe(). userdata: input user application to store in SAFE. userdatalen: size of input user application to store in SAFE. Description: Set the application data in SAFE. Return value: Returns SHISHI OK iff successful.

shishi safe build

[Function] int shishi_safe_build (Shishi safe * safe, Shishi key * key)

safe: safe as allocated by shishi_safe(). key: key for session, used to compute checksum. Description: Build checksum and set it in KRB-SAFE. Note that this follows RFC 1510bis and is incompatible with RFC 1510, although presumably few implementations use the RFC1510 algorithm. Return value: Returns SHISHI OK iff successful.

shishi safe verify

[Function] int shishi_safe_verify (Shishi safe * safe, Shishi key * key)

safe: safe as allocated by shishi_safe(). key: key for session, used to verify checksum. Description: Verify checksum in KRB-SAFE. Note that this follows RFC 1510bis and is incompatible with RFC 1510, although presumably few implementations use the RFC1510 algorithm. Return value: Returns SHISHI OK iff successful, SHISHI SAFE BAD KEYTYPE if an incompatible key type is used, or SHISHI SAFE VERIFY FAILED if the actual verification failed.

The “KRB-PRIV” is an ASN.1 structure used by application client and servers to exchange confidential data. The confidentiality is keyed, usually with a key agreed on via the AP exchange (see Section 5.4 [AP-REQ and AP-REP Functions], page 71). The following illustrates the KRB-PRIV ASN.1 structure.

KRB-PRIV        ::= [APPLICATION 21] SEQUENCE {
      pvno            [0] INTEGER (5),
      msg-type        [1] INTEGER (21),
                      -- NOTE: there is no [2] tag
      enc-part        [3] EncryptedData -- EncKrbPrivPart
}
EncKrbPrivPart  ::= [APPLICATION 28] SEQUENCE {
      user-data       [0] OCTET STRING,
      timestamp       [1] KerberosTime OPTIONAL,
      usec            [2] Microseconds OPTIONAL,
      seq-number      [3] UInt32 OPTIONAL,
      s-address       [4] HostAddress -- sender's addr --,
      r-address       [5] HostAddress OPTIONAL -- recip's addr
}

shishi priv

[Function] int shishi_priv (Shishi * handle, Shishi priv ** priv)

handle: shishi handle as allocated by shishi_init(). priv: pointer to new structure that holds information about PRIV exchange Description: Create a new PRIV exchange. Return value: Returns SHISHI OK iff successful.

shishi priv done

[Function] void shishi_priv_done (Shishi priv * priv)

priv: structure that holds information about PRIV exchange Description: Deallocate resources associated with PRIV exchange. This should be called by the application when it no longer need to utilize the PRIV exchange handle.

shishi priv key

[Function] Shishi_key * shishi_priv_key (Shishi priv * priv)

priv: structure that holds information about PRIV exchange

Description: Get key from PRIV exchange.

Return value: Returns the key used in the PRIV exchange, or NULL if not yet set or an error occured.

shishi priv key set

[Function] void shishi_priv_key_set (Shishi priv * priv, Shishi key * key)

priv: structure that holds information about PRIV exchange key: key to store in PRIV.

Description: Set the Key in the PRIV exchange.

shishi priv priv

[Function] Shishi_asn1 shishi_priv_priv (Shishi priv * priv)

priv: structure that holds information about PRIV exchange

Description: Get ASN.1 PRIV structure in PRIV exchange.

Return value: Returns the ASN.1 priv in the PRIV exchange, or NULL if not yet set or an error occured.

shishi priv priv set

[Function] void shishi_priv_priv_set (Shishi priv * priv,

Shishi asn1 asn1priv) priv: structure that holds information about PRIV exchange asn1priv: KRB-PRIV to store in PRIV exchange.

Description: Set the KRB-PRIV in the PRIV exchange.

shishi priv priv der

[Function] int shishi_priv_priv_der (Shishi priv * priv, char ** out,

size t * outlen) priv: priv as allocated by shishi_priv(). out: output array with newly allocated DER encoding of PRIV. outlen: length of output array with DER encoding of PRIV.

Description: DER encode PRIV structure. Typically shishi_priv_build() is used to build the PRIV structure first. out is allocated by this function, and it is the responsibility of caller to deallocate it.

shishi priv priv der set

[Function] int shishi_priv_priv_der_set (Shishi priv * priv, char * der,
    size t derlen)

priv: priv as allocated by shishi_priv(). der: input array with DER encoded KRB-PRIV. derlen: length of input array with DER encoded KRB-PRIV. Description: DER decode KRB-PRIV and set it PRIV exchange. If decoding fails, the KRB-PRIV in the PRIV exchange remains. Return value: Returns SHISHI OK.

shishi priv encprivpart

[Function] Shishi_asn1 shishi_priv_encprivpart (Shishi priv * priv)

priv: structure that holds information about PRIV exchange Description: Get ASN.1 EncPrivPart structure from PRIV exchange. Return value: Returns the ASN.1 encprivpart in the PRIV exchange, or NULL if not yet set or an error occured.

shishi priv encprivpart set

[Function] void shishi_priv_encprivpart_set (Shishi priv * priv,
    Shishi asn1 asn1encprivpart)

priv: structure that holds information about PRIV exchange asn1encprivpart: ENCPRIVPART to store in PRIV exchange. Description: Set the ENCPRIVPART in the PRIV exchange.

shishi priv encprivpart der

[Function] int shishi_priv_encprivpart_der (Shishi priv * priv,

char ** out, size t * outlen)

priv: priv as allocated by shishi_priv(). out: output array with newly allocated DER encoding of ENCPRIVPART. outlen: length of output array with DER encoding of ENCPRIVPART. Description: DER encode ENCPRIVPART structure. out is allocated by this function, and it is the responsibility of caller to deallocate it. Return value: Returns SHISHI OK iff successful.

shishi priv encprivpart der set

[Function] int shishi_priv_encprivpart_der_set (Shishi priv * priv,

char * der, size t derlen)

priv: priv as allocated by shishi_priv(). der: input array with DER encoded ENCPRIVPART. derlen: length of input array with DER encoded ENCPRIVPART. Description: DER decode ENCPRIVPART and set it PRIV exchange. If decoding fails, the ENCPRIVPART in the PRIV exchange remains. Return value: Returns SHISHI OK.

shishi priv print

[Function] int shishi_priv_print (Shishi * handle, FILE * fh,

Shishi asn1 priv) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. priv: PRIV to print.

Description: Print ASCII armored DER encoding of PRIV to file.

Return value: Returns SHISHI OK iff successful.

shishi priv save

[Function] int shishi_priv_save (Shishi * handle, FILE * fh,

Shishi asn1 priv) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. priv: PRIV to save.

Description: Save DER encoding of PRIV to file.

Return value: Returns SHISHI OK iff successful.

shishi priv to file

[Function] int shishi_priv_to_file (Shishi * handle, Shishi asn1 priv,
   int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). priv: PRIV to save. filetype: input variable specifying type of file to be written, see Shishi filetype. filename: input variable with filename to write to.

Description: Write PRIV to file in specified TYPE. The file will be truncated if it exists.

Return value: Returns SHISHI OK iff successful.

shishi priv parse

[Function] int shishi_priv_parse (Shishi * handle, FILE * fh,

Shishi asn1 * priv) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. priv: output variable with newly allocated PRIV.

Description: Read ASCII armored DER encoded PRIV from file and populate given variable.

shishi priv read

[Function] int shishi_priv_read (Shishi * handle, FILE * fh,

Shishi asn1 * priv) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. priv: output variable with newly allocated PRIV. Description: Read DER encoded PRIV from file and populate given variable. Return value: Returns SHISHI OK iff successful.

shishi priv from file

[Function] int shishi_priv_from_file (Shishi * handle, Shishi asn1 * priv,
   int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). priv: output variable with newly allocated PRIV. filetype: input variable specifying type of file to be read, see Shishi filetype. filename: input variable with filename to read from. Description: Read PRIV from file in specified TYPE. Return value: Returns SHISHI OK iff successful.

shishi priv enc part etype

[Function] int shishi_priv_enc_part_etype (Shishi * handle,

Shishi asn1 priv, int32 t * etype)

handle: shishi handle as allocated by shishi_init(). priv: PRIV variable to get value from. etype: output variable that holds the value. Description: Extract PRIV.enc-part.etype. Return value: Returns SHISHI OK iff successful.

shishi priv set enc part

[Function] int shishi_priv_set_enc_part (Shishi * handle,

Shishi asn1 priv, int32 t etype, const char * encpart,

    size t encpartlen)

handle: shishi handle as allocated by shishi_init(). priv: priv as allocated by shishi_priv(). etype: input encryption type to store in PRIV. encpart: input encrypted data to store in PRIV. encpartlen: size of input encrypted data to store in PRIV. Description: Store encrypted data in PRIV. The encrypted data is usually created by calling shishi_encrypt() on some application specific data using the key from the ticket that is being used. To save time, you may want to use shishi_priv_build() instead, which encryptes the data and calls this function in one step. Return value: Returns SHISHI OK iff successful.

shishi encprivpart user data

[Function] int shishi_encprivpart_user_data (Shishi * handle,
   Shishi asn1 encprivpart, char ** userdata, size t * userdatalen)

handle: shishi handle as allocated by shishi_init(). encprivpart: encprivpart as allocated by shishi_priv(). userdata: output array with newly allocated user data from KRB-PRIV. userdatalen: output size of output user data buffer.

Description: Read user data value from KRB-PRIV. userdata is allocated by this function, and it is the responsibility of caller to deallocate it.

Return value: Returns SHISHI OK iff successful.

shishi encprivpart set user data

[Function] int shishi_encprivpart_set_user_data (Shishi * handle,
   Shishi asn1 encprivpart, const char * userdata, size t userdatalen)

handle: shishi handle as allocated by shishi_init(). encprivpart: encprivpart as allocated by shishi_priv(). userdata: input user application to store in PRIV. userdatalen: size of input user application to store in PRIV.

Description: Set the application data in PRIV.

Return value: Returns SHISHI OK iff successful.

shishi priv build

[Function] int shishi_priv_build (Shishi priv * priv, Shishi key * key)

priv: priv as allocated by shishi_priv(). key: key for session, used to encrypt data.

Description: Build checksum and set it in KRB-PRIV. Note that this follows RFC 1510bis and is incompatible with RFC 1510, although presumably few implementations use the RFC1510 algorithm.

Return value: Returns SHISHI OK iff successful.

shishi priv process

[Function] int shishi_priv_process (Shishi priv * priv, Shishi key * key)

priv: priv as allocated by shishi_priv(). key: key to use to decrypt EncPrivPart.

Description: Decrypt encrypted data in KRB-PRIV and set the EncPrivPart in the PRIV exchange.

Return value: Returns SHISHI OK iff successful, SHISHI PRIV BAD KEYTYPE if an incompatible key type is used, or SHISHI CRYPTO ERROR if the actual decryption failed.

5.6 Ticket Functions

A Ticket is an ASN.1 structured that can be used to authenticate the holder to services. It contain an encrypted part, which the ticket holder cannot see, but can be encrypted by the service, and various information about the user and service, including an encryption key to use for the connection. See Section 5.9 [Ticket (ASN.1) Functions], page 124, for more details on the ASN.1 structure of a ticket.

shishi tkt

[Function] int shishi_tkt (Shishi * handle, Shishi tkt ** tkt)

handle: shishi handle as allocated by shishi_init(). tkt: output variable with newly allocated ticket. Description: Create a new ticket handle. Return value: Returns SHISHI OK iff successful.

shishi tkt2

[Function] Shishi_tkt * shishi_tkt2 (Shishi * handle, Shishi asn1 ticket,

Shishi asn1 enckdcreppart, Shishi asn1 kdcrep)

handle: shishi handle as allocated by shishi_init(). ticket: input variable with ticket. enckdcreppart: input variable with auxiliary ticket information. kdcrep: input variable with KDC-REP ticket information. Description: Create a new ticket handle. Return value: Returns new ticket handle, or NULL on error.

shishi tkt done

[Function] void shishi_tkt_done (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Deallocate resources associated with ticket. The ticket must not be used again after this call.

shishi tkt ticket

[Function] Shishi_asn1 shishi_tkt_ticket (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Get ASN.1 Ticket structure from ticket. Return value: Returns actual ticket.

shishi tkt ticket set

[Function] void shishi_tkt_ticket_set (Shishi tkt * tkt,

Shishi asn1 ticket) tkt: input variable with ticket info. ticket: ASN.1 Ticket to store in ticket. Description: Set the ASN.1 Ticket in the Ticket.

shishi tkt enckdcreppart

[Function] Shishi_asn1 shishi_tkt_enckdcreppart (Shishi tkt * tkt)

tkt: input variable with ticket info.

Description: Get ASN.1 EncKDCRepPart structure from ticket.

Return value: Returns auxiliary ticket information.

shishi tkt enckdcreppart set

[Function] void shishi_tkt_enckdcreppart_set (Shishi tkt * tkt,
    Shishi asn1 enckdcreppart)

tkt: structure that holds information about Ticket exchange enckdcreppart: EncKDCRepPart to store in Ticket.

Description: Set the EncKDCRepPart in the Ticket.

shishi tkt kdcrep

[Function] Shishi_asn1 shishi_tkt_kdcrep (Shishi tkt * tkt)

tkt: input variable with ticket info.

Description: Get ASN.1 KDCRep structure from ticket.

Return value: Returns KDC-REP information.

shishi tkt encticketpart

[Function] Shishi_asn1 shishi_tkt_encticketpart (Shishi tkt * tkt)

tkt: input variable with ticket info.

Description: Get ASN.1 EncTicketPart structure from ticket.

Return value: Returns EncTicketPart information.

shishi tkt encticketpart set

[Function] void shishi_tkt_encticketpart_set (Shishi tkt * tkt,
    Shishi asn1 encticketpart)

tkt: input variable with ticket info. encticketpart: encticketpart to store in ticket.

Description: Set the EncTicketPart in the Ticket.

shishi tkt key

[Function] Shishi_key * shishi_tkt_key (Shishi tkt * tkt)

tkt: input variable with ticket info.

Description: Get key used in ticket, by looking first in EncKDCRepPart and then in EncTicketPart. If key is already populated, it is not extracted again.

Return value: Returns key extracted from EncKDCRepPart or EncTicketPart.

shishi tkt key set

[Function] int shishi_tkt_key_set (Shishi tkt * tkt, Shishi key * key)

tkt: input variable with ticket info. key: key to store in ticket.

Description: Set the key in the EncTicketPart.

Return value: Returns SHISHI OK iff successful.

shishi tkt client

[Function] int shishi_tkt_client (Shishi tkt * tkt, char ** client,
    size t * clientlen)

tkt: input variable with ticket info. client: pointer to newly allocated zero terminated string containing principal name. May be NULL (to only populate clientlen). clientlen: pointer to length of client on output, excluding terminating zero. May be NULL (to only populate client).

Description: Represent client principal name in Ticket KDC-REP as zero-terminated string. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length clientlen does not include the terminating zero.

Return value: Returns SHISHI OK iff successful.

shishi tkt client p

[Function] int shishi_tkt_client_p (Shishi tkt * tkt, const char * client)

tkt: input variable with ticket info. client: client name of ticket.

Description: Determine if ticket is for specified client.

Return value: Returns non-0 iff ticket is for specified client.

shishi tkt clientrealm

[Function] int shishi_tkt_clientrealm (Shishi tkt * tkt, char ** client,
    size t * clientlen)

tkt: input variable with ticket info. client: pointer to newly allocated zero terminated string containing principal name and realm. May be NULL (to only populate clientlen). clientlen: pointer to length of client on output, excluding terminating zero. May be NULL (to only populate client).

Description: Convert cname and realm fields from AS-REQ to printable principal name format. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length clientlen does not include the terminating zero.

Return value: Returns SHISHI OK iff successful.

shishi tkt clientrealm p

[Function] int shishi_tkt_clientrealm_p (Shishi tkt * tkt,

const char * client) tkt: input variable with ticket info. client: principal name (client name and realm) of ticket. Description: Determine if ticket is for specified client principal. Return value: Returns non-0 iff ticket is for specified client principal.

shishi tkt realm

[Function] int shishi_tkt_realm (Shishi tkt * tkt, char ** realm,

    size t * realmlen)

tkt: input variable with ticket info. realm: pointer to newly allocated character array with realm name. realmlen: length of newly allocated character array with realm name. Description: Extract realm of server in ticket. Return value: Returns SHISHI OK iff successful.

shishi tkt server

[Function] int shishi_tkt_server (Shishi tkt * tkt, char ** server,
    size t * serverlen)

tkt: input variable with ticket info. server: pointer to newly allocated zero terminated string containing principal name. May be NULL (to only populate serverlen). serverlen: pointer to length of server on output, excluding terminating zero. May be NULL (to only populate server). Description: Represent server principal name in Ticket as zero-terminated string. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length serverlen does not include the terminating zero. Return value: Returns SHISHI OK iff successful.

shishi tkt server p

[Function] int shishi_tkt_server_p (Shishi tkt * tkt, const char * server)

tkt: input variable with ticket info. server: server name of ticket. Description: Determine if ticket is for specified server. Return value: Returns non-0 iff ticket is for specified server.

shishi tkt flags

[Function] int shishi_tkt_flags (Shishi tkt * tkt, uint32 t * flags)

tkt: input variable with ticket info. flags: pointer to output integer with flags.

Description: Extract flags in ticket (i.e., EncKDCRepPart). Return value: Returns SHISHI OK iff successful.

shishi tkt flags set

[Function] int shishi_tkt_flags_set (Shishi tkt * tkt, uint32 t flags)

tkt: input variable with ticket info. flags: integer with flags to store in ticket. Description: Set flags in ticket, i.e., both EncTicketPart and EncKDCRepPart. Note that this reset any already existing flags. Return value: Returns SHISHI OK iff successful.

shishi tkt flags add

[Function] int shishi_tkt_flags_add (Shishi tkt * tkt, uint32 t flag)

tkt: input variable with ticket info. flag: integer with flags to store in ticket. Description: Add ticket flags to Ticket and EncKDCRepPart. This preserves all existing options. Return value: Returns SHISHI OK iff successful.

shishi tkt forwardable p

[Function] int shishi_tkt_forwardable_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket is forwardable. The FORWARDABLE flag in a ticket is normally only interpreted by the ticketgranting service. It can be ignored by application servers. The FORWARDABLE flag has an interpretation similar to that of the PROXIABLE flag, except ticket-granting tickets may also be issued with different network addresses. This flag is reset by default, but users MAY request that it be set by setting the FORWARDABLE option in the AS request when they request their initial ticket-granting ticket. Return value: Returns non-0 iff forwardable flag is set in ticket.

shishi tkt forwarded p

[Function] int shishi_tkt_forwarded_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket is forwarded. The FORWARDED flag is set by the TGS when a client presents a ticket with the FORWARDABLE flag set and requests a forwarded ticket by specifying the FOR-WARDED KDC option and supplying a set of addresses for the new ticket. It is also set in all tickets issued based on tickets with the FORWARDED flag set. Application servers may choose to process FORWARDED tickets differently than non-FORWARDED tickets. Return value: Returns non-0 iff forwarded flag is set in ticket.

shishi tkt proxiable p

[Function] int shishi_tkt_proxiable_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket is proxiable. The PROXIABLE flag in a ticket is normally only interpreted by the ticket-granting service. It can be ignored by application servers. When set, this flag tells the ticketgranting server that it is OK to issue a new ticket (but not a ticket-granting ticket) with a different network address based on this ticket. This flag is set if requested by the client on initial authentication. By default, the client will request that it be set when requesting a ticket-granting ticket, and reset when requesting any other ticket. Return value: Returns non-0 iff proxiable flag is set in ticket.

shishi tkt proxy p

[Function] int shishi_tkt_proxy_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket is proxy ticket. The PROXY flag is set in a ticket by the TGS when it issues a proxy ticket. Application servers MAY check this flag and at their option they MAY require additional authentication from the agent presenting the proxy in order to provide an audit trail. Return value: Returns non-0 iff proxy flag is set in ticket.

shishi tkt may postdate p

[Function] int shishi_tkt_may_postdate_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket may be used to grant postdated tickets. The MAY-POSTDATE flag in a ticket is normally only interpreted by the ticketgranting service. It can be ignored by application servers. This flag MUST be set in a ticket-granting ticket in order to issue a postdated ticket based on the presented ticket. It is reset by default; it MAY be requested by a client by setting the ALLOW-POSTDATE option in the KRB AS REQ message. This flag does not allow a client to obtain a postdated ticket-granting ticket; postdated ticket-granting tickets can only by obtained by requesting the postdating in the KRB AS REQ message. The life (endtime-starttime) of a postdated ticket will be the remaining life of the ticketgranting ticket at the time of the request, unless the RENEWABLE option is also set, in which case it can be the full life (endtime-starttime) of the ticket-granting ticket. The KDC MAY limit how far in the future a ticket may be postdated. Return value: Returns non-0 iff may-postdate flag is set in ticket.

shishi tkt postdated p

[Function] int shishi_tkt_postdated_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket is postdated.

The POSTDATED flag indicates that a ticket has been postdated. The application server can check the authtime field in the ticket to see when the original authentication occurred. Some services MAY choose to reject postdated tickets, or they may only accept them within a certain period after the original authentication. When the KDC issues a POSTDATED ticket, it will also be marked as INVALID, so that the application client MUST present the ticket to the KDC to be validated before use.

Return value: Returns non-0 iff postdated flag is set in ticket.

shishi tkt invalid p

[Function] int shishi_tkt_invalid_p (Shishi tkt * tkt)

tkt: input variable with ticket info.

Description: Determine if ticket is invalid.

The INVALID flag indicates that a ticket is invalid. Application servers MUST reject tickets which have this flag set. A postdated ticket will be issued in this form. Invalid tickets MUST be validated by the KDC before use, by presenting them to the KDC in a TGS request with the VALIDATE option specified. The KDC will only validate tickets after their starttime has passed. The validation is required so that postdated tickets which have been stolen before their starttime can be rendered permanently invalid (through a hot-list mechanism).

Return value: Returns non-0 iff invalid flag is set in ticket.

shishi tkt renewable p

[Function] int shishi_tkt_renewable_p (Shishi tkt * tkt)

tkt: input variable with ticket info.

Description: Determine if ticket is renewable.

The RENEWABLE flag in a ticket is normally only interpreted by the ticket-granting service (discussed below in section 3.3). It can usually be ignored by application servers. However, some particularly careful application servers MAY disallow renewable tickets.

Return value: Returns non-0 iff renewable flag is set in ticket.

shishi tkt initial p

[Function] int shishi_tkt_initial_p (Shishi tkt * tkt)

tkt: input variable with ticket info.

Description: Determine if ticket was issued using AS exchange.

The INITIAL flag indicates that a ticket was issued using the AS protocol, rather than issued based on a ticket-granting ticket. Application servers that want to require the demonstrated knowledge of a client’s secret key (e.g. a password-changing program) can insist that this flag be set in any tickets they accept, and thus be assured that the client’s key was recently presented to the application client.

Return value: Returns non-0 iff initial flag is set in ticket.

shishi tkt pre authent p

[Function] int shishi_tkt_pre_authent_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket was pre-authenticated. The PRE-AUTHENT and HW-AUTHENT flags provide additional information about the initial authentication, regardless of whether the current ticket was issued directly (in which case INITIAL will also be set) or issued on the basis of a ticket-granting ticket (in which case the INITIAL flag is clear, but the PRE-AUTHENT and HW-AUTHENT flags are carried forward from the ticket-granting ticket). Return value: Returns non-0 iff pre-authent flag is set in ticket.

shishi tkt hw authent p

[Function] int shishi_tkt_hw_authent_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket is authenticated using a hardware token. The PRE-AUTHENT and HW-AUTHENT flags provide additional information about the initial authentication, regardless of whether the current ticket was issued directly (in which case INITIAL will also be set) or issued on the basis of a ticket-granting ticket (in which case the INITIAL flag is clear, but the PRE-AUTHENT and HW-AUTHENT flags are carried forward from the ticket-granting ticket). Return value: Returns non-0 iff hw-authent flag is set in ticket.

shishi tkt transited policy checked p

[Function] int shishi_tkt_transited_policy_checked_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket has been policy checked for transit. The application server is ultimately responsible for accepting or rejecting authentication and SHOULD check that only suitably trusted KDCs are relied upon to authenticate a principal. The transited field in the ticket identifies which realms (and thus which KDCs) were involved in the authentication process and an application server would normally check this field. If any of these are untrusted to authenticate the indicated client principal (probably determined by a realm-based policy), the authentication attempt MUST be rejected. The presence of trusted KDCs in this list does not provide any guarantee; an untrusted KDC may have fabricated the list. While the end server ultimately decides whether authentication is valid, the KDC for the end server’s realm MAY apply a realm specific policy for validating the transited field and accepting credentials for cross-realm authentication. When the KDC applies such checks and accepts such cross-realm authentication it will set the TRANSITED-POLICY-CHECKED flag in the service tickets it issues based on the cross-realm TGT. A client MAY request that the KDCs not check the transited field by setting the DISABLE-TRANSITED-CHECK flag. KDCs are encouraged but not required to honor this flag. Application servers MUST either do the transited-realm checks themselves, or reject cross-realm tickets without TRANSITED-POLICY- CHECKED set. Return value: Returns non-0 iff transited-policy-checked flag is set in ticket.

shishi tkt ok as delegate p

[Function] int shishi_tkt_ok_as_delegate_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket is ok as delegated ticket. The copy of the ticket flags in the encrypted part of the KDC reply may have the OK-AS-DELEGATE flag set to indicates to the client that the server specified in the ticket has been determined by policy of the realm to be a suitable recipient of delegation. A client can use the presence of this flag to help it make a decision whether to delegate credentials (either grant a proxy or a forwarded ticket- granting ticket) to this server. It is acceptable to ignore the value of this flag. When setting this flag, an administrator should consider the security and placement of the server on which the service will run, as well as whether the service requires the use of delegated credentials. Return value: Returns non-0 iff ok-as-delegate flag is set in ticket.

shishi tkt keytype

[Function] int shishi_tkt_keytype (Shishi tkt * tkt, int32 t * etype)

tkt: input variable with ticket info. etype: pointer to encryption type that is set, see Shishi etype. Description: Extract encryption type of key in ticket (really EncKDCRepPart). Return value: Returns SHISHI OK iff successful.

shishi tkt keytype fast

[Function] int32_t shishi_tkt_keytype_fast (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Extract encryption type of key in ticket (really EncKDCRepPart). Return value: Returns encryption type of session key in ticket (really EncKDCRep-Part), or -1 on error.

shishi tkt keytype p

[Function] int shishi_tkt_keytype_p (Shishi tkt * tkt, int32 t etype)

tkt: input variable with ticket info. etype: encryption type, see Shishi etype. Description: Determine if key in ticket (really EncKDCRepPart) is of specified key type (really encryption type). Return value: Returns non-0 iff key in ticket is of specified encryption type.

shishi tkt lastreqc

[Function] time_t shishi_tkt_lastreqc (Shishi tkt * tkt,

Shishi lrtype lrtype) tkt: input variable with ticket info. lrtype: lastreq type to extract, see Shishi lrtype. E.g., SHISHI LRTYPE LAST REQUEST. Description: Extract C time corresponding to given lastreq type field in the ticket. Return value: Returns C time interpretation of the specified lastreq field, or (time t) -1.

shishi tkt authctime

[Function] time_t shishi_tkt_authctime (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Extract C time corresponding to the authtime field. The field holds the time when the original authentication took place that later resulted in this ticket. Return value: Returns C time interpretation of the endtime in ticket.

shishi tkt startctime

[Function] time_t shishi_tkt_startctime (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Extract C time corresponding to the starttime field. The field holds the time where the ticket start to be valid (typically in the past). Return value: Returns C time interpretation of the endtime in ticket.

shishi tkt endctime

[Function] time_t shishi_tkt_endctime (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Extract C time corresponding to the endtime field. The field holds the time where the ticket stop being valid. Return value: Returns C time interpretation of the endtime in ticket.

shishi tkt renew tillc

[Function] time_t shishi_tkt_renew_tillc (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Extract C time corresponding to the renew-till field. The field holds the time where the ticket stop being valid for renewal. Return value: Returns C time interpretation of the renew-till in ticket.

shishi tkt valid at time p

[Function] int shishi_tkt_valid_at_time_p (Shishi tkt * tkt, time t now)

tkt: input variable with ticket info. now: time to check for.

Description: Determine if ticket is valid at a specific point in time. Return value: Returns non-0 iff ticket is valid (not expired and after starttime) at specified time.

shishi tkt valid now p

[Function] int shishi_tkt_valid_now_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket is valid now. Return value: Returns 0 iff ticket is invalid (expired or not yet valid).

shishi tkt expired p

[Function] int shishi_tkt_expired_p (Shishi tkt * tkt)

tkt: input variable with ticket info. Description: Determine if ticket has expired (i.e., endtime is in the past). Return value: Returns 0 iff ticket has expired.

shishi tkt lastreq pretty print

[Function] void shishi_tkt_lastreq_pretty_print (Shishi tkt * tkt,

FILE * fh) tkt: input variable with ticket info. fh: file handle open for writing. Description: Print a human readable representation of the various lastreq fields in the ticket (really EncKDCRepPart).

shishi tkt pretty print

[Function] void shishi_tkt_pretty_print (Shishi tkt * tkt, FILE * fh)

tkt: input variable with ticket info. fh: file handle open for writing. Description: Print a human readable representation of a ticket to file handle.

5.7 AS Functions

The Authentication Service (AS) is used to get an initial ticket using e.g. your password. The following illustrates the AS-REQ and AS-REP ASN.1 structures.

-- Request --

AS-REQ          ::= KDC-REQ {10}
KDC-REQ {INTEGER:tagnum}        ::= [APPLICATION tagnum] SEQUENCE {
   pvno            [1] INTEGER (5) -- first tag is [1], not [0] --,
   msg-type        [2] INTEGER (tagnum),
   padata          [3] SEQUENCE OF PA-DATA OPTIONAL,
   req-body        [4] KDC-REQ-BODY

}

KDC-REQ-BODY    ::= SEQUENCE {
kdc-options             [0] KDCOptions,
cname                   [1] PrincipalName OPTIONAL

                            -- Used only in AS-REQ --,

realm                   [2] Realm

                            -- Server's realm
                            -- Also client's in AS-REQ --,

sname                   [3] PrincipalName OPTIONAL,
from                    [4] KerberosTime OPTIONAL,
till                    [5] KerberosTime,
rtime                   [6] KerberosTime OPTIONAL,
nonce                   [7] UInt32,
etype                   [8] SEQUENCE OF Int32 -- EncryptionType

                            -- in preference order --,

addresses               [9] HostAddresses OPTIONAL,
enc-authorization-data  [10] EncryptedData {

                                AuthorizationData,
                                { keyuse-TGSReqAuthData-sesskey
                                  | keyuse-TGSReqAuthData-subkey }
                             } OPTIONAL,

additional-tickets      [11] SEQUENCE OF Ticket OPTIONAL

}
-- Reply --

AS-REP          ::= KDC-REP {11, EncASRepPart, {keyuse-EncASRepPart}}

KDC-REP {INTEGER:tagnum,
 TypeToEncrypt,

 UInt32:KeyUsages}      ::= [APPLICATION tagnum] SEQUENCE {
pvno            [0] INTEGER (5),
msg-type        [1] INTEGER (tagnum),
padata          [2] SEQUENCE OF PA-DATA OPTIONAL,
crealm          [3] Realm,
cname           [4] PrincipalName,
ticket          [5] Ticket,
enc-part        [6] EncryptedData {TypeToEncrypt, KeyUsages}

}

EncASRepPart    ::= [APPLICATION 25] EncKDCRepPart
EncKDCRepPart   ::= SEQUENCE {
key             [0] EncryptionKey,
last-req        [1] LastReq,
nonce           [2] UInt32,

   key-expiration  [3] KerberosTime OPTIONAL,
   flags           [4] TicketFlags,
   authtime        [5] KerberosTime,
   starttime       [6] KerberosTime OPTIONAL,
   endtime         [7] KerberosTime,
   renew-till      [8] KerberosTime OPTIONAL,
   srealm          [9] Realm,
   sname           [10] PrincipalName,
   caddr           [11] HostAddresses OPTIONAL

}

shishi as

[Function] int shishi_as (Shishi * handle, Shishi as ** as)

handle: shishi handle as allocated by shishi_init(). as: holds pointer to newly allocate Shishi as structure. Description: Allocate a new AS exchange variable. Return value: Returns SHISHI OK iff successful.

shishi as done

[Function] void shishi_as_done (Shishi as * as)

as: structure that holds information about AS exchange Description: Deallocate resources associated with AS exchange. This should be called by the application when it no longer need to utilize the AS exchange handle.

shishi as req

[Function] Shishi_asn1 shishi_as_req (Shishi as * as)

as: structure that holds information about AS exchange Description: Get ASN.1 AS-REQ structure from AS exchange. Return value: Returns the generated AS-REQ packet from the AS exchange, or NULL if not yet set or an error occured.

shishi as req build

[Function] int shishi_as_req_build (Shishi as * as)

as: structure that holds information about AS exchange Description: Possibly remove unset fields (e.g., rtime). Return value: Returns SHISHI OK iff successful.

shishi as req set

[Function] void shishi_as_req_set (Shishi as * as, Shishi asn1 asreq)

as: structure that holds information about AS exchange asreq: asreq to store in AS. Description: Set the AS-REQ in the AS exchange.

shishi as req der

[Function] int shishi_as_req_der (Shishi as * as, char ** out,

size t * outlen) as: structure that holds information about AS exchange out: output array with newly allocated DER encoding of AS-REQ. outlen: length of output array with DER encoding of AS-REQ. Description: DER encode AS-REQ. out is allocated by this function, and it is the responsibility of caller to deallocate it. Return value: Returns SHISHI OK iff successful.

shishi as req der set

[Function] int shishi_as_req_der_set (Shishi as * as, char * der,
    size t derlen)

as: structure that holds information about AS exchange der: input array with DER encoded AP-REQ. derlen: length of input array with DER encoded AP-REQ. Description: DER decode AS-REQ and set it AS exchange. If decoding fails, the AS-REQ in the AS exchange remains. Return value: Returns SHISHI OK.

shishi as rep

[Function] Shishi_asn1 shishi_as_rep (Shishi as * as)

as: structure that holds information about AS exchange Description: Get ASN.1 AS-REP structure from AS exchange. Return value: Returns the received AS-REP packet from the AS exchange, or NULL if not yet set or an error occured.

shishi as rep process

[Function] int shishi_as_rep_process (Shishi as * as, Shishi key * key,

const char * password) as: structure that holds information about AS exchange key: user’s key, used to encrypt the encrypted part of the AS-REP. password: user’s password, used if key is NULL. Description: Process new AS-REP and set ticket. The key is used to decrypt the AP-REP. If both key and password is NULL, the user is queried for it. Return value: Returns SHISHI OK iff successful.

shishi as rep build

[Function] int shishi_as_rep_build (Shishi as * as, Shishi key * key)

as: structure that holds information about AS exchange key: user’s key, used to encrypt the encrypted part of the AS-REP. Description: Build AS-REP. Return value: Returns SHISHI OK iff successful.

shishi as rep der

[Function] int shishi_as_rep_der (Shishi as * as, char ** out,

size t * outlen) as: structure that holds information about AS exchange out: output array with newly allocated DER encoding of AS-REP. outlen: length of output array with DER encoding of AS-REP. Description: DER encode AS-REP. out is allocated by this function, and it is the responsibility of caller to deallocate it. Return value: Returns SHISHI OK iff successful.

shishi as rep set

[Function] void shishi_as_rep_set (Shishi as * as, Shishi asn1 asrep)

as: structure that holds information about AS exchange asrep: asrep to store in AS. Description: Set the AS-REP in the AS exchange.

shishi as rep der set

[Function] int shishi_as_rep_der_set (Shishi as * as, char * der,
    size t derlen)

as: structure that holds information about AS exchange der: input array with DER encoded AP-REP. derlen: length of input array with DER encoded AP-REP. Description: DER decode AS-REP and set it AS exchange. If decoding fails, the AS-REP in the AS exchange remains. Return value: Returns SHISHI OK.

shishi as krberror

[Function] Shishi_asn1 shishi_as_krberror (Shishi as * as)

as: structure that holds information about AS exchange Description: Get ASN.1 KRB-ERROR structure from AS exchange. Return value: Returns the received KRB-ERROR packet from the AS exchange, or NULL if not yet set or an error occured.

shishi as krberror der

[Function] int shishi_as_krberror_der (Shishi as * as, char ** out,

size t * outlen) as: structure that holds information about AS exchange out: output array with newly allocated DER encoding of KRB-ERROR. outlen: length of output array with DER encoding of KRB-ERROR. Description: DER encode KRB-ERROR. out is allocated by this function, and it is the responsibility of caller to deallocate it. Return value: Returns SHISHI OK iff successful.

shishi as krberror set

[Function] void shishi_as_krberror_set (Shishi as * as,

Shishi asn1 krberror) as: structure that holds information about AS exchange krberror: krberror to store in AS. Description: Set the KRB-ERROR in the AS exchange.

shishi as tkt

[Function] Shishi_tkt * shishi_as_tkt (Shishi as * as)

as: structure that holds information about AS exchange Description: Get Ticket in AS exchange. Return value: Returns the newly acquired tkt from the AS exchange, or NULL if not yet set or an error occured.

shishi as tkt set

[Function] void shishi_as_tkt_set (Shishi as * as, Shishi tkt * tkt)

as: structure that holds information about AS exchange tkt: tkt to store in AS. Description: Set the Tkt in the AS exchange.

shishi as sendrecv hint

[Function] int shishi_as_sendrecv_hint (Shishi as * as,

Shishi tkts hint * hint) as: structure that holds information about AS exchange hint: additional parameters that modify connection behaviour, or NULL. Description: Send AS-REQ and receive AS-REP or KRB-ERROR. This is the initial authentication, usually used to acquire a Ticket Granting Ticket. The hint structure can be used to set, e.g., parameters for TLS authentication. Return value: Returns SHISHI OK iff successful.

shishi as sendrecv

[Function] int shishi_as_sendrecv (Shishi as * as)

as: structure that holds information about AS exchange Description: Send AS-REQ and receive AS-REP or KRB-ERROR. This is the initial authentication, usually used to acquire a Ticket Granting Ticket. Return value: Returns SHISHI OK iff successful.

5.8 TGS Functions

The Ticket Granting Service (TGS) is used to get subsequent tickets, authenticated by other tickets (so called ticket granting tickets). The following illustrates the TGS-REQ and TGS-REP ASN.1 structures.

-- Request --

TGS-REQ         ::= KDC-REQ {12}
KDC-REQ {INTEGER:tagnum}        ::= [APPLICATION tagnum] SEQUENCE {
pvno            [1] INTEGER (5) -- first tag is [1], not [0] --,
msg-type        [2] INTEGER (tagnum),
padata          [3] SEQUENCE OF PA-DATA OPTIONAL,
req-body        [4] KDC-REQ-BODY

}

KDC-REQ-BODY    ::= SEQUENCE {
kdc-options             [0] KDCOptions,
cname                   [1] PrincipalName OPTIONAL

                            -- Used only in AS-REQ --,

realm                   [2] Realm

                            -- Server's realm
                            -- Also client's in AS-REQ --,

sname                   [3] PrincipalName OPTIONAL,
from                    [4] KerberosTime OPTIONAL,
till                    [5] KerberosTime,
rtime                   [6] KerberosTime OPTIONAL,
nonce                   [7] UInt32,
etype                   [8] SEQUENCE OF Int32 -- EncryptionType

                            -- in preference order --,

addresses               [9] HostAddresses OPTIONAL,
enc-authorization-data  [10] EncryptedData {

                                AuthorizationData,
                                { keyuse-TGSReqAuthData-sesskey
                                  | keyuse-TGSReqAuthData-subkey }
                             } OPTIONAL,

additional-tickets      [11] SEQUENCE OF Ticket OPTIONAL

}
-- Reply --

TGS-REP         ::= KDC-REP {13, EncTGSRepPart,

                { keyuse-EncTGSRepPart-sesskey
                  | keyuse-EncTGSRepPart-subkey }}
KDC-REP {INTEGER:tagnum,
 TypeToEncrypt,

 UInt32:KeyUsages}      ::= [APPLICATION tagnum] SEQUENCE {
pvno            [0] INTEGER (5),
msg-type        [1] INTEGER (tagnum),
padata          [2] SEQUENCE OF PA-DATA OPTIONAL,
crealm          [3] Realm,
cname           [4] PrincipalName,

ticket          [5] Ticket,
enc-part        [6] EncryptedData {TypeToEncrypt, KeyUsages}

}

EncTGSRepPart   ::= [APPLICATION 26] EncKDCRepPart
EncKDCRepPart   ::= SEQUENCE {
key             [0] EncryptionKey,
last-req        [1] LastReq,
nonce           [2] UInt32,
key-expiration  [3] KerberosTime OPTIONAL,
flags           [4] TicketFlags,
authtime        [5] KerberosTime,
starttime       [6] KerberosTime OPTIONAL,
endtime         [7] KerberosTime,
renew-till      [8] KerberosTime OPTIONAL,
srealm          [9] Realm,
sname           [10] PrincipalName,
caddr           [11] HostAddresses OPTIONAL

}

shishi tgs

[Function] int shishi_tgs (Shishi * handle, Shishi tgs ** tgs)

handle: shishi handle as allocated by shishi_init(). tgs: holds pointer to newly allocate Shishi tgs structure. Description: Allocate a new TGS exchange variable. Return value: Returns SHISHI OK iff successful.

shishi tgs done

[Function] void shishi_tgs_done (Shishi tgs * tgs)

tgs: structure that holds information about AS exchange Description: Deallocate resources associated with TGS exchange. This should be called by the application when it no longer need to utilize the TGS exchange handle.

shishi tgs tgtkt

[Function] Shishi_tkt * shishi_tgs_tgtkt (Shishi tgs * tgs)

tgs: structure that holds information about TGS exchange Description: Get Ticket-granting-ticket from TGS exchange. Return value: Returns the ticket-granting-ticket used in the TGS exchange, or NULL if not yet set or an error occured.

shishi tgs tgtkt set

[Function] void shishi_tgs_tgtkt_set (Shishi tgs * tgs, Shishi tkt * tgtkt)

tgs: structure that holds information about TGS exchange tgtkt: ticket granting ticket to store in TGS.

Description: Set the Ticket in the TGS exchange.

shishi tgs ap

[Function] Shishi_ap * shishi_tgs_ap (Shishi tgs * tgs)

tgs: structure that holds information about TGS exchange Description: Get the AP from TGS exchange. Return value: Returns the AP exchange (part of TGS-REQ) from the TGS exchange, or NULL if not yet set or an error occured.

shishi tgs req

[Function] Shishi_asn1 shishi_tgs_req (Shishi tgs * tgs)

tgs: structure that holds information about TGS exchange Description: Get the TGS-REQ from TGS exchange. Return value: Returns the generated TGS-REQ from the TGS exchange, or NULL if not yet set or an error occured.

shishi tgs req set

[Function] void shishi_tgs_req_set (Shishi tgs * tgs, Shishi asn1 tgsreq)

tgs: structure that holds information about TGS exchange tgsreq: tgsreq to store in TGS. Description: Set the TGS-REQ in the TGS exchange.

shishi tgs req der

[Function] int shishi_tgs_req_der (Shishi tgs * tgs, char ** out,

size t * outlen) tgs: structure that holds information about TGS exchange out: output array with newly allocated DER encoding of TGS-REQ. outlen: length of output array with DER encoding of TGS-REQ. Description: DER encode TGS-REQ. out is allocated by this function, and it is the responsibility of caller to deallocate it. Return value: Returns SHISHI OK iff successful.

shishi tgs req der set

[Function] int shishi_tgs_req_der_set (Shishi tgs * tgs, char * der,
    size t derlen)

tgs: structure that holds information about TGS exchange der: input array with DER encoded AP-REQ. derlen: length of input array with DER encoded AP-REQ. Description: DER decode TGS-REQ and set it TGS exchange. If decoding fails, the TGS-REQ in the TGS exchange remains. Return value: Returns SHISHI OK.

shishi tgs req process

[Function] int shishi_tgs_req_process (Shishi tgs * tgs)

tgs: structure that holds information about TGS exchange

Description: Process new TGS-REQ and set ticket. The key to decrypt the TGS-REQ is taken from the EncKDCReqPart of the TGS tgticket.

Return value: Returns SHISHI OK iff successful.

shishi tgs req build

[Function] int shishi_tgs_req_build (Shishi tgs * tgs)

tgs: structure that holds information about TGS exchange

Description: Checksum data in authenticator and add ticket and authenticator to TGS-REQ.

Return value: Returns SHISHI OK iff successful.

shishi tgs rep

[Function] Shishi_asn1 shishi_tgs_rep (Shishi tgs * tgs)

tgs: structure that holds information about TGS exchange

Description: Get TGS-REP from TGS exchange.

Return value: Returns the received TGS-REP from the TGS exchange, or NULL if not yet set or an error occured.

shishi tgs rep der

[Function] int shishi_tgs_rep_der (Shishi tgs * tgs, char ** out,

size t * outlen) tgs: structure that holds information about TGS exchange out: output array with newly allocated DER encoding of TGS-REP. outlen: length of output array with DER encoding of TGS-REP.

Description: DER encode TGS-REP. out is allocated by this function, and it is the responsibility of caller to deallocate it.

Return value: Returns SHISHI OK iff successful.

shishi tgs rep process

[Function] int shishi_tgs_rep_process (Shishi tgs * tgs)

tgs: structure that holds information about TGS exchange

Description: Process new TGS-REP and set ticket. The key to decrypt the TGS-REP is taken from the EncKDCRepPart of the TGS tgticket.

Return value: Returns SHISHI OK iff successful.

shishi tgs rep build

[Function] int shishi_tgs_rep_build (Shishi tgs * tgs, int keyusage,

Shishi key * key) tgs: structure that holds information about TGS exchange keyusage: keyusage integer. key: user’s key, used to encrypt the encrypted part of the TGS-REP.

Description: Build TGS-REP.

Return value: Returns SHISHI OK iff successful.

shishi tgs krberror

[Function] Shishi_asn1 shishi_tgs_krberror (Shishi tgs * tgs)

tgs: structure that holds information about TGS exchange

Description: Get KRB-ERROR from TGS exchange.

Return value: Returns the received TGS-REP from the TGS exchange, or NULL if not yet set or an error occured.

shishi tgs krberror der

[Function] int shishi_tgs_krberror_der (Shishi tgs * tgs, char ** out,

size t * outlen) tgs: structure that holds information about TGS exchange out: output array with newly allocated DER encoding of KRB-ERROR. outlen: length of output array with DER encoding of KRB-ERROR.

Description: DER encode KRB-ERROR. out is allocated by this function, and it is the responsibility of caller to deallocate it.

Return value: Returns SHISHI OK iff successful.

shishi tgs krberror set

[Function] void shishi_tgs_krberror_set (Shishi tgs * tgs,

Shishi asn1 krberror) tgs: structure that holds information about TGS exchange krberror: krberror to store in TGS.

Description: Set the KRB-ERROR in the TGS exchange.

shishi tgs tkt

[Function] Shishi_tkt * shishi_tgs_tkt (Shishi tgs * tgs)

tgs: structure that holds information about TGS exchange

Description: Get Ticket from TGS exchange.

Return value: Returns the newly acquired ticket from the TGS exchange, or NULL if not yet set or an error occured.

shishi tgs tkt set

[Function] void shishi_tgs_tkt_set (Shishi tgs * tgs, Shishi tkt * tkt)

tgs: structure that holds information about TGS exchange tkt: ticket to store in TGS.

Description: Set the Ticket in the TGS exchange.

shishi tgs sendrecv hint

[Function] int shishi_tgs_sendrecv_hint (Shishi tgs * tgs,

Shishi tkts hint * hint) tgs: structure that holds information about TGS exchange hint: additional parameters that modify connection behaviour, or NULL.

Description: Send TGS-REQ and receive TGS-REP or KRB-ERROR. This is the subsequent authentication, usually used to acquire server tickets. The hint structure can be used to set, e.g., parameters for TLS authentication.

Return value: Returns SHISHI OK iff successful.

shishi tgs sendrecv

[Function] int shishi_tgs_sendrecv (Shishi tgs * tgs)

tgs: structure that holds information about TGS exchange

Description: Send TGS-REQ and receive TGS-REP or KRB-ERROR. This is the subsequent authentication, usually used to acquire server tickets.

Return value: Returns SHISHI OK iff successful.

shishi tgs set server

[Function] int shishi_tgs_set_server (Shishi tgs * tgs,

const char * server) tgs: structure that holds information about TGS exchange server: indicates the server to acquire ticket for.

Description: Set the server in the TGS-REQ.

Return value: Returns SHISHI OK iff successful.

shishi tgs set realm

[Function] int shishi_tgs_set_realm (Shishi tgs * tgs, const char * realm)

tgs: structure that holds information about TGS exchange realm: indicates the realm to acquire ticket for.

Description: Set the server in the TGS-REQ.

Return value: Returns SHISHI OK iff successful.

shishi tgs set realmserver

[Function] int shishi_tgs_set_realmserver (Shishi tgs * tgs,

const char * realm, const char * server)

tgs: structure that holds information about TGS exchange realm: indicates the realm to acquire ticket for. server: indicates the server to acquire ticket for.

Description: Set the realm and server in the TGS-REQ.

Return value: Returns SHISHI OK iff successful.

5.9 Ticket (ASN.1) Functions

See Section 5.6 [Ticket Functions], page 102, for an high-level overview of tickets. The following illustrates the Ticket and EncTicketPart ASN.1 structures.

Ticket          ::= [APPLICATION 1] SEQUENCE {
tkt-vno         [0] INTEGER (5),
realm           [1] Realm,
sname           [2] PrincipalName,
enc-part        [3] EncryptedData -- EncTicketPart
}
-- Encrypted part of ticket
EncTicketPart   ::= [APPLICATION 3] SEQUENCE {
flags                   [0] TicketFlags,
key                     [1] EncryptionKey,
crealm                  [2] Realm,
cname                   [3] PrincipalName,
transited               [4] TransitedEncoding,
authtime                [5] KerberosTime,
starttime               [6] KerberosTime OPTIONAL,
endtime                 [7] KerberosTime,
renew-till              [8] KerberosTime OPTIONAL,
caddr                   [9] HostAddresses OPTIONAL,
authorization-data      [10] AuthorizationData OPTIONAL
}

shishi ticket

[Function] Shishi_asn1 shishi_ticket (Shishi * handle)

handle: shishi handle as allocated by shishi_init().

Description: This function creates a new ASN.1 Ticket, populated with some default values.

Return value: Returns the ticket or NULL on failure.

shishi ticket realm get

[Function] int shishi_ticket_realm_get (Shishi * handle,

Shishi asn1 ticket, char ** realm, size t * realmlen)

handle: shishi handle as allocated by shishi_init(). ticket: input variable with ticket info. realm: output array with newly allocated name of realm in ticket. realmlen: size of output array. Description: Extract realm from ticket. Return value: Returns SHISHI OK iff successful.

shishi ticket realm set

[Function] int shishi_ticket_realm_set (Shishi * handle,

Shishi asn1 ticket, const char * realm)

handle: shishi handle as allocated by shishi_init(). ticket: input variable with ticket info. realm: input array with name of realm. Description: Set the realm field in the Ticket. Return value: Returns SHISHI OK iff successful.

shishi ticket server

[Function] int shishi_ticket_server (Shishi * handle, Shishi asn1 ticket,
   char ** server, size t * serverlen)

handle: Shishi library handle create by shishi_init(). ticket: ASN.1 Ticket variable to get server name from. server: pointer to newly allocated zero terminated string containing principal name. May be NULL (to only populate serverlen). serverlen: pointer to length of server on output, excluding terminating zero. May be NULL (to only populate server). Description: Represent server principal name in Ticket as zero-terminated string. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length serverlen does not include the terminating zero. Return value: Returns SHISHI OK iff successful.

shishi ticket sname set

[Function] int shishi_ticket_sname_set (Shishi * handle,

Shishi asn1 ticket, Shishi name type name_type, char * sname[])

handle: shishi handle as allocated by shishi_init(). ticket: Ticket variable to set server name field in. name type: type of principial, see Shishi name type, usually SHISHI NT UNKNOWN.

sname: input array with principal name. Description: Set the server name field in the Ticket. Return value: Returns SHISHI OK iff successful.

shishi ticket get enc part etype

[Function] int shishi_ticket_get_enc_part_etype (Shishi * handle,

Shishi asn1 ticket, int32 t * etype)

handle: shishi handle as allocated by shishi_init(). ticket: Ticket variable to get value from. etype: output variable that holds the value. Description: Extract Ticket.enc-part.etype. Return value: Returns SHISHI OK iff successful.

shishi ticket set enc part

[Function] int shishi_ticket_set_enc_part (Shishi * handle,

Shishi asn1 ticket, int32 t etype, uint32 t kvno, const char * buf,

    size t buflen)

handle: shishi handle as allocated by shishi_init(). ticket: Ticket to add enc-part field to. etype: encryption type used to encrypt enc-part. kvno: key version number. buf : input array with encrypted enc-part. buflen: size of input array with encrypted enc-part. Description: Set the encrypted enc-part field in the Ticket. The encrypted data is usually created by calling shishi_encrypt() on the DER encoded enc-part. To save time, you may want to use shishi_ticket_add_enc_part() instead, which calculates the encrypted data and calls this function in one step. Return value: Returns SHISHI OK iff successful.

shishi ticket add enc part

[Function] int shishi_ticket_add_enc_part (Shishi * handle,

Shishi asn1 ticket, Shishi key * key, Shishi asn1 encticketpart)

handle: shishi handle as allocated by shishi_init(). ticket: Ticket to add enc-part field to. key: key used to encrypt enc-part. encticketpart: EncTicketPart to add. Description: Encrypts DER encoded EncTicketPart using key and stores it in the Ticket. Return value: Returns SHISHI OK iff successful.

shishi encticketpart get key

[Function] int shishi_encticketpart_get_key (Shishi * handle,

Shishi asn1 encticketpart, Shishi key ** key)

handle: shishi handle as allocated by shishi_init(). encticketpart: input EncTicketPart variable. key: newly allocated key. Description: Extract the session key in the Ticket. Return value: Returns SHISHI_OK iff successful.

shishi encticketpart key set

[Function] int shishi_encticketpart_key_set (Shishi * handle,

Shishi asn1 encticketpart, Shishi key * key)

handle: shishi handle as allocated by shishi_init(). encticketpart: input EncTicketPart variable. key: key handle with information to store in encticketpart.

Description: Set the EncTicketPart.key field to key type and value of supplied key.

Return value: Returns SHISHI_OK iff successful.

shishi encticketpart flags set

[Function] int shishi_encticketpart_flags_set (Shishi * handle,
   Shishi asn1 encticketpart, int flags)

handle: shishi handle as allocated by shishi_init(). encticketpart: input EncTicketPart variable. flags: flags to set in encticketpart.

Description: Set the EncTicketPart.flags to supplied value.

Return value: Returns SHISHI_OK iff successful.

shishi encticketpart crealm set

[Function] int shishi_encticketpart_crealm_set (Shishi * handle,

Shishi asn1 encticketpart, const char * realm)

handle: shishi handle as allocated by shishi_init(). encticketpart: input EncTicketPart variable. realm: input array with name of realm.

Description: Set the realm field in the KDC-REQ.

Return value: Returns SHISHI OK iff successful.

shishi encticketpart cname set

[Function] int shishi_encticketpart_cname_set (Shishi * handle,

Shishi asn1 encticketpart, Shishi name type name_type,

const char * principal) handle: shishi handle as allocated by shishi_init(). encticketpart: input EncTicketPart variable. name type: type of principial, see Shishi name type, usually SHISHI NT UNKNOWN.

principal: input array with principal name.

Description: Set the client name field in the EncTicketPart.

shishi encticketpart transited set

[Function] int shishi_encticketpart_transited_set (Shishi * handle,

Shishi asn1 encticketpart, int32 t trtype, const char * trdata,

    size t trdatalen)

handle: shishi handle as allocated by shishi_init(). encticketpart: input EncTicketPart variable. trtype: transitedencoding type, e.g. SHISHI TR DOMAIN X500 COMPRESS. trdata: actual transited realm data. trdatalen: length of actual transited realm data. Description: Set the EncTicketPart.transited field to supplied value. Return value: Returns SHISHI_OK iff successful.

shishi encticketpart authtime set

[Function] int shishi_encticketpart_authtime_set (Shishi * handle,

Shishi asn1 encticketpart, const char * authtime)

handle: shishi handle as allocated by shishi_init(). encticketpart: input EncTicketPart variable. authtime: character buffer containing a generalized time string. Description: Set the EncTicketPart.authtime to supplied value. Return value: Returns SHISHI_OK iff successful.

shishi encticketpart endtime set

[Function] int shishi_encticketpart_endtime_set (Shishi * handle,

Shishi asn1 encticketpart, const char * endtime)

handle: shishi handle as allocated by shishi_init(). encticketpart: input EncTicketPart variable. endtime: character buffer containing a generalized time string. Description: Set the EncTicketPart.endtime to supplied value. Return value: Returns SHISHI_OK iff successful.

shishi encticketpart client

[Function] int shishi_encticketpart_client (Shishi * handle,
   Shishi asn1 encticketpart, char ** client, size t * clientlen)

handle: Shishi library handle create by shishi_init(). encticketpart: EncTicketPart variable to get client name from. client: pointer to newly allocated zero terminated string containing principal name. May be NULL (to only populate clientlen). clientlen: pointer to length of client on output, excluding terminating zero. May be NULL (to only populate client). Description: Represent client principal name in EncTicketPart as zero-terminated string. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length clientlen does not include the terminating zero. Return value: Returns SHISHI OK iff successful.

shishi encticketpart clientrealm

[Function] int shishi_encticketpart_clientrealm (Shishi * handle,
        Shishi asn1 encticketpart, char ** client, size t * clientlen)

handle: Shishi library handle create by shishi_init().

encticketpart: EncTicketPart variable to get client name and realm from.

client: pointer to newly allocated zero terminated string containing principal name and realm. May be NULL (to only populate clientlen). clientlen: pointer to length of client on output, excluding terminating zero. May be NULL (to only populate client). Description: Convert cname and realm fields from EncTicketPart to printable principal name format. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length clientlen does not include the terminating zero. Return value: Returns SHISHI OK iff successful.

5.10 AS/TGS Functions

The Authentication Service (AS) is used to get an initial ticket using e.g. your password. The Ticket Granting Service (TGS) is used to get subsequent tickets using other tickets. Protocol wise the procedures are very similar, which is the reason they are described together. The following illustrates the AS-REQ, TGS-REQ and AS-REP, TGS-REP ASN.1 structures. Most of the functions use the mnemonic “KDC” instead of either AS or TGS, which means the function operates on both AS and TGS types. Only where the distinction between AS and TGS is important are the AS and TGS names used. Remember, these are low-level functions, and normal applications will likely be satisfied with the AS (see Section 5.7 [AS Functions], page 112) and TGS (see Section 5.8 [TGS Functions], page 117) interfaces, or the even more high-level Ticket Set (see Section 5.3 [Ticket Set Functions], page 64) interface.

-- Request --

AS-REQ          ::= KDC-REQ {10}
TGS-REQ         ::= KDC-REQ {12}
KDC-REQ {INTEGER:tagnum}        ::= [APPLICATION tagnum] SEQUENCE {
        pvno            [1] INTEGER (5) -- first tag is [1], not [0] --,
        msg-type        [2] INTEGER (tagnum),
        padata          [3] SEQUENCE OF PA-DATA OPTIONAL,
        req-body        [4] KDC-REQ-BODY

}

KDC-REQ-BODY    ::= SEQUENCE {
        kdc-options             [0] KDCOptions,
        cname                   [1] PrincipalName OPTIONAL

                                    -- Used only in AS-REQ --,

        realm                   [2] Realm

                                    -- Server's realm

                            -- Also client's in AS-REQ --,

sname                   [3] PrincipalName OPTIONAL,
from                    [4] KerberosTime OPTIONAL,
till                    [5] KerberosTime,
rtime                   [6] KerberosTime OPTIONAL,
nonce                   [7] UInt32,
etype                   [8] SEQUENCE OF Int32 -- EncryptionType

                            -- in preference order --,

addresses               [9] HostAddresses OPTIONAL,
enc-authorization-data  [10] EncryptedData {

                                AuthorizationData,
                                { keyuse-TGSReqAuthData-sesskey
                                  | keyuse-TGSReqAuthData-subkey }
                             } OPTIONAL,

additional-tickets      [11] SEQUENCE OF Ticket OPTIONAL

}
-- Reply --

AS-REP          ::= KDC-REP {11, EncASRepPart, {keyuse-EncASRepPart}}
TGS-REP         ::= KDC-REP {13, EncTGSRepPart,

                { keyuse-EncTGSRepPart-sesskey
                  | keyuse-EncTGSRepPart-subkey }}
KDC-REP {INTEGER:tagnum,
 TypeToEncrypt,

 UInt32:KeyUsages}      ::= [APPLICATION tagnum] SEQUENCE {
pvno            [0] INTEGER (5),
msg-type        [1] INTEGER (tagnum),
padata          [2] SEQUENCE OF PA-DATA OPTIONAL,
crealm          [3] Realm,
cname           [4] PrincipalName,
ticket          [5] Ticket,
enc-part        [6] EncryptedData {TypeToEncrypt, KeyUsages}

}

EncASRepPart    ::= [APPLICATION 25] EncKDCRepPart
EncTGSRepPart   ::= [APPLICATION 26] EncKDCRepPart
EncKDCRepPart   ::= SEQUENCE {
key             [0] EncryptionKey,
last-req        [1] LastReq,
nonce           [2] UInt32,
key-expiration  [3] KerberosTime OPTIONAL,
flags           [4] TicketFlags,
authtime        [5] KerberosTime,
starttime       [6] KerberosTime OPTIONAL,

   endtime         [7] KerberosTime,
   renew-till      [8] KerberosTime OPTIONAL,
   srealm          [9] Realm,
   sname           [10] PrincipalName,
   caddr           [11] HostAddresses OPTIONAL

}

shishi as derive salt

[Function] int shishi_as_derive_salt (Shishi * handle, Shishi asn1 asreq,

Shishi asn1 asrep, char ** salt, size t * saltlen)

handle: Shishi handle as allocated by shishi_init(). asreq: Input AS-REQ variable. asrep: Input AS-REP variable. salt: Returned pointer to newly allocated output array. saltlen: Pointer to integer, returning size of output array.

Description: Computes the salt that should be used when deriving a key via shishi_ string_to_key() for an AS exchange. Currently this searches for PA-DATA of type SHISHI_PA_PW_SALT in the AS-REP provided by asrep, and if present returns it. Otherwise the salt is composed from the client name and the realm, both are extracted from the request asreq.

Return value: Returns SHISHI_OK if successful. Failure conditions include various ASN.1 issues.

shishi kdcreq sendrecv hint

[Function] int shishi_kdcreq_sendrecv_hint (Shishi * handle,

Shishi asn1 kdcreq, Shishi asn1 * kdcrep, Shishi tkts hint * hint)

handle: Shishi library handle created by shishi_init(). kdcreq: Input variable with a prepared AS-REQ. kdcrep: Output pointer variable for decoded AS-REP. hint: Input Shishi_tkts_hint structure with flags.

Description: Sends a request to KDC, and receives the response. The provided request kdcreq and the hints structure hint, together determine transmitted data. On reception the reply is decoded as AS-REP into kdcrep.

Return value: Return code is SHISHI_OK on success, SHISHI_KDC_TIMEOUT on time-

outs, SHISHI_ASN1_ERROR on translation errors, and SHISHI_GOT_KRBERROR for other

corruptions.

shishi kdcreq sendrecv

[Function] int shishi_kdcreq_sendrecv (Shishi * handle,

Shishi asn1 kdcreq, Shishi asn1 * kdcrep)

handle: Shishi library handle created by shishi_init(). kdcreq: Input variable with a prepared AS-REQ. kdcrep: Output pointer variable returning received AS-REP.

Description: Sends a request to KDC, and receives the response. The provided AS-REQ, in kdcreq, sets all data for the request. On reception the reply is decoded as AS-REP into kdcrep.

Return value: Return code is SHISHI_OK on success, SHISHI_KDC_TIMEOUT on time-

outs, SHISHI_ASN1_ERROR on translation errors, and SHISHI_GOT_KRBERROR for other

corruptions.

shishi kdc copy crealm

[Function] int shishi_kdc_copy_crealm (Shishi * handle,

Shishi asn1 kdcrep, Shishi asn1 encticketpart)

handle: Shishi handle as allocated by shishi_init(). kdcrep: KDC-REP where the field "crealm" is updated. encticketpart: EncTicketPart providing "crealm" field.

Description: Reads the field "crealm" from the ticket encticketpart and copies the value into the reply kdcrep.

Return value: Returns SHISHI_OK if successful, and ASN.1 failures otherwise.

shishi as check crealm

[Function] int shishi_as_check_crealm (Shishi * handle, Shishi asn1 asreq,

Shishi asn1 asrep) handle: Shishi handle as allocated by shishi_init(). asreq: Request of type AS-REQ. asrep: Reply structure of type AS-REP.

Description: Verifies that the fields asreq.req-body.realm and asrep.crealm contain identical realm names. This is one of the steps that has to be performed when processing an exchange of AS-REQ and AS-REP; see shishi_kdc_process() for more details.

Return value: Returns SHISHI_OK if successful, SHISHI_REALM_MISMATCH whenever the realm names differ, and an error code otherwise.

shishi kdc copy cname

[Function] int shishi_kdc_copy_cname (Shishi * handle, Shishi asn1 kdcrep,
    Shishi asn1 encticketpart)

handle: Shishi handle as allocated by shishi_init(). kdcrep: KDC-REP where the field "cname" is updated. encticketpart: EncTicketPart providing "cname" field.

Description: Reads the field "cname" from the ticket encticketpart and copies the value into the reply kdcrep.

Return value: Returns SHISHI_OK if successful, and ASN.1 failures otherwise.

shishi as check cname

[Function] int shishi_as_check_cname (Shishi * handle, Shishi asn1 asreq,

Shishi asn1 asrep) handle: Shishi handle as allocated by shishi_init(). asreq: Request of type AS-REQ. asrep: Reply structure of type AS-REP. Description: Verifies that the fields asreq.req-body.cname and asrep.cname contain identical names. This is one of the steps that has to be performed when processing an exchange of AS-REQ and AS-REP; see shishi_kdc_process() for more details. Return value: Returns SHISHI_OK if successful, SHISHI_CNAME_MISMATCH if the names differ, and an error code otherwise.

shishi kdc copy nonce

[Function] int shishi_kdc_copy_nonce (Shishi * handle, Shishi asn1 kdcreq,
    Shishi asn1 enckdcreppart)

handle: Shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ providing "nonce" field. enckdcreppart: EncKDCRepPart where "nonce" field is updated. Description: Sets the field "nonce" in enckdcreppart to a value retreived from the corresponding field in kdcreq. Return value: Returns SHISHI_OK if successful.

shishi kdc check nonce

[Function] int shishi_kdc_check_nonce (Shishi * handle,

Shishi asn1 kdcreq, Shishi asn1 enckdcreppart)

handle: Shishi handle as allocated by shishi_init(). kdcreq: Request of type KDC-REQ. enckdcreppart: Encrypted KDC-REP part. Description: Verifies that kdcreq.req-body.nonce and enckdcreppart.nonce contain matching values. This is one of the steps that has to be performed when processing an exchange of KDC-REQ and KDC-REP. Return value: Returns SHISHI_OK if successful, SHISHI_NONCE_MISMATCH whenever the nonces are of differing lengths (usually a sign that a buggy server truncates the nonce to 4 bytes) and the same code if the nonce values differ, or an error code otherwise.

shishi tgs process

[Function] int shishi_tgs_process (Shishi * handle, Shishi asn1 tgsreq,

Shishi asn1 tgsrep, Shishi asn1 authenticator,

   Shishi asn1 oldenckdcreppart, Shishi asn1 * enckdcreppart)

handle: Shishi handle as allocated by shishi_init(). tgsreq: Input variable holding the transmitted KDC-REQ. tgsrep: Input variable holding the received KDC-REP. authenticator: Input variable with an authenticator extracted from the AP-REQ part of tgsreq. oldenckdcreppart: Input variable with EncKDCRepPart used in the request. enckdcreppart: Output variable holding the new EncKDCRepPart. Description: Processes a TGS client exchange and outputs the decrypted EncKD-CRepPart, holding details about the received ticket. This function simply derives the encryption key from the ticket used to construct the original TGS request, and then

calls shishi_kdc_process().

Return value: Returns SHISHI_OK if the TGS client exchange was successful. Failures include ASN.1 and TGS conditions.

shishi as process

[Function] int shishi_as_process (Shishi * handle, Shishi asn1 asreq,

Shishi asn1 asrep, const char * string, Shishi asn1 * enckdcreppart)

handle: Shishi handle as allocated by shishi_init(). asreq: Input variable holding the transmitted KDC-REQ. asrep: Input variable holding the received KDC-REP. string: Input variable with a null terminated password. enckdcreppart: Output variable returning a new EncKDCRepPart. Description: Processes an AS client exchange and returns the decrypted EncKD-CRepPart, holding details about the received ticket. This function simply derives the encryption key from the password, and then calls shishi_kdc_process(). Return value: Returns SHISHI_OK if the AS client exchange was successful. Multiple failure conditions are possible.

shishi kdc process

[Function] int shishi_kdc_process (Shishi * handle, Shishi asn1 kdcreq,

Shishi asn1 kdcrep, Shishi key * key, int keyusage,

    Shishi asn1 * enckdcreppart)

handle: Shishi handle as allocated by shishi_init(). kdcreq: Input variable holding the transmitted KDC-REQ. kdcrep: Input variable holding the received KDC-REP. key: Input pointet to key for decrypting parts of kdcrep. keyusage: Kerberos key usage code. enckdcreppart: Output pointer for the extracted EncKDCRepPart. Description: Processes a KDC client exchange and extracts a decrypted EncKDCRep-Part, holding details about the received ticket. Use shishi_kdcrep_get_ticket() to extract the ticket itself. This function verifies the various conditions that must hold if the response is to be considered valid. In particular, it compares nonces (using shishi_kdc_check_nonce()), and if the exchange was an AS exchange, it also checks

cname and crealm (using shishi_as_check_cname(), shishi_as_check_crealm()).
Usually shishi_as_process() and shishi_tgs_process() should be used instead

of this call, since they simplify computation of the decryption key. Return value: Returns SHISHI_OK if the KDC client exchange was successful. Multiple failure conditions are possible.

shishi asreq

[Function] Shishi_asn1 shishi_asreq (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: This function creates a new AS-REQ, populated with some default values. Return value: Returns the AS-REQ or NULL on failure.

shishi tgsreq

[Function] Shishi_asn1 shishi_tgsreq (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: This function creates a new TGS-REQ, populated with some default values. Return value: Returns the TGS-REQ or NULL on failure.

shishi kdcreq print

[Function] int shishi_kdcreq_print (Shishi * handle, FILE * fh,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. kdcreq: KDC-REQ to print. Description: Print ASCII armored DER encoding of KDC-REQ to file. Return value: Returns SHISHI OK iff successful.

shishi kdcreq save

[Function] int shishi_kdcreq_save (Shishi * handle, FILE * fh,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. kdcreq: KDC-REQ to save. Description: Print DER encoding of KDC-REQ to file. Return value: Returns SHISHI OK iff successful.

shishi kdcreq to file

[Function] int shishi_kdcreq_to_file (Shishi * handle, Shishi asn1 kdcreq,
   int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ to save. filetype: input variable specifying type of file to be written, see Shishi filetype. filename: input variable with filename to write to. Description: Write KDC-REQ to file in specified TYPE. The file will be truncated if it exists. Return value: Returns SHISHI OK iff successful.

shishi kdcreq parse

[Function] int shishi_kdcreq_parse (Shishi * handle, FILE * fh,

Shishi asn1 * kdcreq) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. kdcreq: output variable with newly allocated KDC-REQ.

Description: Read ASCII armored DER encoded KDC-REQ from file and populate given variable.

Return value: Returns SHISHI OK iff successful.

shishi kdcreq read

[Function] int shishi_kdcreq_read (Shishi * handle, FILE * fh,

Shishi asn1 * kdcreq) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. kdcreq: output variable with newly allocated KDC-REQ.

Description: Read DER encoded KDC-REQ from file and populate given variable.

Return value: Returns SHISHI OK iff successful.

shishi kdcreq from file

[Function] int shishi_kdcreq_from_file (Shishi * handle,

Shishi asn1 * kdcreq, int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). kdcreq: output variable with newly allocated KDC-REQ. filetype: input variable specifying type of file to be read, see Shishi filetype. filename: input variable with filename to read from.

Description: Read KDC-REQ from file in specified TYPE.

Return value: Returns SHISHI OK iff successful.

shishi kdcreq nonce set

[Function] int shishi_kdcreq_nonce_set (Shishi * handle,

Shishi asn1 kdcreq, uint32 t nonce)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to set client name field in. nonce: integer nonce to store in KDC-REQ.

Description: Store nonce number field in KDC-REQ.

Return value: Returns SHISHI_OK iff successful.

shishi kdcreq set cname

[Function] int shishi_kdcreq_set_cname (Shishi * handle,

Shishi asn1 kdcreq, Shishi name type name_type,

const char * principal) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to set client name field in. name type: type of principial, see Shishi name type, usually SHISHI NT UNKNOWN.

principal: input array with principal name.

Description: Set the client name field in the KDC-REQ.

Return value: Returns SHISHI OK iff successful.

shishi kdcreq client

[Function] int shishi_kdcreq_client (Shishi * handle, Shishi asn1 kdcreq,
   char ** client, size t * clientlen)

handle: Shishi library handle create by shishi_init(). kdcreq: KDC-REQ variable to get client name from. client: pointer to newly allocated zero terminated string containing principal name. May be NULL (to only populate clientlen). clientlen: pointer to length of client on output, excluding terminating zero. May be NULL (to only populate client).

Description: Represent client principal name in KDC-REQ as zero-terminated string. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length clientlen does not include the terminating zero.

Return value: Returns SHISHI OK iff successful.

shishi asreq clientrealm

[Function] int shishi_asreq_clientrealm (Shishi * handle,

Shishi asn1 asreq, char ** client, size t * clientlen)

handle: Shishi library handle create by shishi_init(). asreq: AS-REQ variable to get client name and realm from. client: pointer to newly allocated zero terminated string containing principal name and realm. May be NULL (to only populate clientlen). clientlen: pointer to length of client on output, excluding terminating zero. May be NULL (to only populate client).

Description: Convert cname and realm fields from AS-REQ to printable principal name format. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length clientlen does not include the terminating zero.

Return value: Returns SHISHI OK iff successful.

shishi kdcreq realm

[Function] int shishi_kdcreq_realm (Shishi * handle, Shishi asn1 kdcreq,

char ** realm, size t * realmlen)

handle: Shishi library handle create by shishi_init(). kdcreq: KDC-REQ variable to get client name from. realm: pointer to newly allocated zero terminated string containing realm. May be NULL (to only populate realmlen). realmlen: pointer to length of realm on output, excluding terminating zero. May be NULL (to only populate realmlen). Description: Get realm field in KDC-REQ as zero-terminated string. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length realmlen does not include the terminating zero. Return value: Returns SHISHI OK iff successful.

shishi kdcreq set realm

[Function] int shishi_kdcreq_set_realm (Shishi * handle,

Shishi asn1 kdcreq, const char * realm)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to set realm field in. realm: input array with name of realm. Description: Set the realm field in the KDC-REQ. Return value: Returns SHISHI OK iff successful.

shishi kdcreq server

[Function] int shishi_kdcreq_server (Shishi * handle, Shishi asn1 kdcreq,
   char ** server, size t * serverlen)

handle: Shishi library handle create by shishi_init(). kdcreq: KDC-REQ variable to get server name from. server: pointer to newly allocated zero terminated string containing principal name. May be NULL (to only populate serverlen). serverlen: pointer to length of server on output, excluding terminating zero. May be NULL (to only populate server). Description: Represent server principal name in KDC-REQ as zero-terminated string. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length serverlen does not include the terminating zero. Return value: Returns SHISHI OK iff successful.

shishi kdcreq set sname

[Function] int shishi_kdcreq_set_sname (Shishi * handle,

Shishi asn1 kdcreq, Shishi name type name_type, const char * sname[])

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to set server name field in. name type: type of principial, see Shishi name type, usually SHISHI NT UNKNOWN.

sname: input array with principal name. Description: Set the server name field in the KDC-REQ. Return value: Returns SHISHI OK iff successful.

shishi kdcreq till

[Function] int shishi_kdcreq_till (Shishi * handle, Shishi asn1 kdcreq,

char ** till, size t * tilllen)

handle: Shishi library handle created by shishi_init(). kdcreq: KDC-REQ variable to get endtime from. till: pointer to newly allocated null terminated string containing "till" field with generalized time. May be passed as NULL to only populate tilllen. tilllen: pointer to length of till for output, excluding the terminating null. Set to NULL, only till is populated. Description: Get "till" field, i.e., "endtime", in KDC-REQ as a null-terminated string. The string is typically 15 characters long and is allocated by this function. It is the responsibility of the caller to deallocate it. Note that the output length tilllen does not include the terminating zero. Return value: Returns SHISHI OK iff successful.

shishi kdcreq tillc

[Function] time_t shishi_kdcreq_tillc (Shishi * handle,

Shishi asn1 kdcreq) handle: Shishi library handle created by shishi_init(). kdcreq: KDC-REQ variable to get "till" field from. Description: Extract C time corresponding to the "till" field. Return value: Returns the C time interpretation of the "till" field in KDC-REQ.

shishi kdcreq etype

[Function] int shishi_kdcreq_etype (Shishi * handle, Shishi asn1 kdcreq,

int32 t * etype, int netype)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get etype field from. etype: output encryption type. netype: element number to return. Return the netype: th encryption type from KDC-REQ. The first etype is number 1. Return value: Returns SHISHI OK iff etype successful set.

shishi kdcreq set etype

[Function] int shishi_kdcreq_set_etype (Shishi * handle,

Shishi asn1 kdcreq, int32 t * etype, int netype)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to set etype field in. etype: input array with encryption types. netype: number of elements in input array with encryption types. Description: Set the list of supported or wanted encryption types in the request. The list should be sorted in priority order. Return value: Returns SHISHI OK iff successful.

shishi kdcreq options

[Function] int shishi_kdcreq_options (Shishi * handle, Shishi asn1 kdcreq,

uint32 t * flags) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from. flags: pointer to output integer with flags. Description: Extract KDC-Options from KDC-REQ. Return value: Returns SHISHI OK iff successful.

shishi kdcreq forwardable p

[Function] int shishi_kdcreq_forwardable_p (Shishi * handle,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from. Description: Determine if KDC-Option forwardable flag is set. The FORWARDABLE option indicates that the ticket to be issued is to have its forwardable flag set. It may only be set on the initial request, or in a subsequent request if the ticket-granting ticket on which it is based is also forwardable. Return value: Returns non-0 iff forwardable flag is set in KDC-REQ.

shishi kdcreq forwarded p

[Function] int shishi_kdcreq_forwarded_p (Shishi * handle,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from. Description: Determine if KDC-Option forwarded flag is set. The FORWARDED option is only specified in a request to the ticket-granting server and will only be honored if the ticket-granting ticket in the request has its FOR-WARDABLE bit set. This option indicates that this is a request for forwarding. The address(es) of the host from which the resulting ticket is to be valid are included in the addresses field of the request. Return value: Returns non-0 iff forwarded flag is set in KDC-REQ.

shishi kdcreq proxiable p

[Function] int shishi_kdcreq_proxiable_p (Shishi * handle,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from.

Description: Determine if KDC-Option proxiable flag is set.

The PROXIABLE option indicates that the ticket to be issued is to have its proxiable flag set. It may only be set on the initial request, or in a subsequent request if the ticket-granting ticket on which it is based is also proxiable.

Return value: Returns non-0 iff proxiable flag is set in KDC-REQ.

shishi kdcreq proxy p

[Function] int shishi_kdcreq_proxy_p (Shishi * handle, Shishi asn1 kdcreq)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from.

Description: Determine if KDC-Option proxy flag is set.

The PROXY option indicates that this is a request for a proxy. This option will only be honored if the ticket-granting ticket in the request has its PROXIABLE bit set. The address(es) of the host from which the resulting ticket is to be valid are included in the addresses field of the request.

Return value: Returns non-0 iff proxy flag is set in KDC-REQ.

shishi kdcreq allow postdate p

[Function] int shishi_kdcreq_allow_postdate_p (Shishi * handle,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from.

Description: Determine if KDC-Option allow-postdate flag is set.

The ALLOW-POSTDATE option indicates that the ticket to be issued is to have its MAY-POSTDATE flag set. It may only be set on the initial request, or in a subsequent request if the ticket-granting ticket on which it is based also has its MAY-POSTDATE flag set.

Return value: Returns non-0 iff allow-postdate flag is set in KDC-REQ.

shishi kdcreq postdated p

[Function] int shishi_kdcreq_postdated_p (Shishi * handle,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from.

Description: Determine if KDC-Option postdated flag is set.

The POSTDATED option indicates that this is a request for a postdated ticket. This option will only be honored if the ticket-granting ticket on which it is based has its MAY-POSTDATE flag set. The resulting ticket will also have its INVALID flag set, and that flag may be reset by a subsequent request to the KDC after the starttime in the ticket has been reached.

Return value: Returns non-0 iff postdated flag is set in KDC-REQ.

shishi kdcreq renewable p

[Function] int shishi_kdcreq_renewable_p (Shishi * handle,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from.

Description: Determine if KDC-Option renewable flag is set.

The RENEWABLE option indicates that the ticket to be issued is to have its RENEW-ABLE flag set. It may only be set on the initial request, or when the ticket-granting ticket on which the request is based is also renewable. If this option is requested, then the rtime field in the request contains the desired absolute expiration time for the ticket.

Return value: Returns non-0 iff renewable flag is set in KDC-REQ.

shishi kdcreq disable transited check p

[Function] int shishi_kdcreq_disable_transited_check_p (Shishi * handle,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from.

Description: Determine if KDC-Option disable-transited-check flag is set.

By default the KDC will check the transited field of a ticket-granting-ticket against the policy of the local realm before it will issue derivative tickets based on the ticketgranting ticket. If this flag is set in the request, checking of the transited field is disabled. Tickets issued without the performance of this check will be noted by the reset (0) value of the TRANSITED-POLICY-CHECKED flag, indicating to the application server that the tranisted field must be checked locally. KDCs are encouraged but not required to honor the DISABLE-TRANSITED-CHECK option.

This flag is new since RFC 1510

Return value: Returns non-0 iff disable-transited-check flag is set in KDC-REQ.

shishi kdcreq renewable ok p

[Function] int shishi_kdcreq_renewable_ok_p (Shishi * handle,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from.

Description: Determine if KDC-Option renewable-ok flag is set.

The RENEWABLE-OK option indicates that a renewable ticket will be acceptable if a ticket with the requested life cannot otherwise be provided. If a ticket with the requested life cannot be provided, then a renewable ticket may be issued with a renew-till equal to the requested endtime. The value of the renew-till field may still be limited by local limits, or limits selected by the individual principal or server.

Return value: Returns non-0 iff renewable-ok flag is set in KDC-REQ.

shishi kdcreq enc tkt in skey p

[Function] int shishi_kdcreq_enc_tkt_in_skey_p (Shishi * handle,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from. Description: Determine if KDC-Option enc-tkt-in-skey flag is set. This option is used only by the ticket-granting service. The ENC-TKT-IN-SKEY option indicates that the ticket for the end server is to be encrypted in the session key from the additional ticket-granting ticket provided. Return value: Returns non-0 iff enc-tkt-in-skey flag is set in KDC-REQ.

shishi kdcreq renew p

[Function] int shishi_kdcreq_renew_p (Shishi * handle, Shishi asn1 kdcreq)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from. Description: Determine if KDC-Option renew flag is set. This option is used only by the ticket-granting service. The RENEW option indicates that the present request is for a renewal. The ticket provided is encrypted in the secret key for the server on which it is valid. This option will only be honored if the ticket to be renewed has its RENEWABLE flag set and if the time in its renew-till field has not passed. The ticket to be renewed is passed in the padata field as part of the authentication header. Return value: Returns non-0 iff renew flag is set in KDC-REQ.

shishi kdcreq validate p

[Function] int shishi_kdcreq_validate_p (Shishi * handle,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to get kdc-options field from. Description: Determine if KDC-Option validate flag is set. This option is used only by the ticket-granting service. The VALIDATE option indicates that the request is to validate a postdated ticket. It will only be honored if the ticket presented is postdated, presently has its INVALID flag set, and would be otherwise usable at this time. A ticket cannot be validated before its starttime. The ticket presented for validation is encrypted in the key of the server for which it is valid and is passed in the padata field as part of the authentication header. Return value: Returns non-0 iff validate flag is set in KDC-REQ.

shishi kdcreq options set

[Function] int shishi_kdcreq_options_set (Shishi * handle,

Shishi asn1 kdcreq, uint32 t options)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to set etype field in. options: integer with flags to store in KDC-REQ.

Description: Set options in KDC-REQ. Note that this reset any already existing flags.

Return value: Returns SHISHI OK iff successful.

shishi kdcreq options add

[Function] int shishi_kdcreq_options_add (Shishi * handle,

Shishi asn1 kdcreq, uint32 t option)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ variable to set etype field in. option: integer with options to add in KDC-REQ.

Description: Add KDC-Option to KDC-REQ. This preserves all existing options.

Return value: Returns SHISHI OK iff successful.

shishi kdcreq clear padata

[Function] int shishi_kdcreq_clear_padata (Shishi * handle,

Shishi asn1 kdcreq) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ to remove PA-DATA from.

Description: Remove the padata field from KDC-REQ.

Return value: Returns SHISHI OK iff successful.

shishi kdcreq get padata

[Function] int shishi_kdcreq_get_padata (Shishi * handle,

Shishi asn1 kdcreq, Shishi padata type padatatype, char ** out,

size t * outlen) handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ to get PA-DATA from. padatatype: type of PA-DATA, see Shishi padata type. out: output array with newly allocated PA-DATA value. outlen: size of output array with PA-DATA value.

Description: Get pre authentication data (PA-DATA) from KDC-REQ. Pre authentication data is used to pass various information to KDC, such as in case of a SHISHI PA TGS REQ padatatype the AP-REQ that authenticates the user to get the ticket.

Return value: Returns SHISHI OK iff successful.

shishi kdcreq get padata tgs

[Function] int shishi_kdcreq_get_padata_tgs (Shishi * handle,

Shishi asn1 kdcreq, Shishi asn1 * apreq)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ to get PA-TGS-REQ from. apreq: Output variable with newly allocated AP-REQ.

Description: Extract TGS pre-authentication data from KDC-REQ. The data is an AP-REQ that authenticates the request. This function call shishi_kdcreq_get_ padata() with a SHISHI PA TGS REQ padatatype and DER decode the result (if any). Return value: Returns SHISHI OK iff successful.

shishi kdcreq add padata

[Function] int shishi_kdcreq_add_padata (Shishi * handle,

Shishi asn1 kdcreq, int padatatype, const char * data, size t datalen)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ to add PA-DATA to. padatatype: type of PA-DATA, see Shishi padata type. data: input array with PA-DATA value. datalen: size of input array with PA-DATA value. Description: Add new pre authentication data (PA-DATA) to KDC-REQ. This is used to pass various information to KDC, such as in case of a SHISHI PA TGS REQ padatatype the AP-REQ that authenticates the user to get the ticket. (But also see

shishi_kdcreq_add_padata_tgs() which takes an AP-REQ directly.)

Return value: Returns SHISHI OK iff successful.

shishi kdcreq add padata tgs

[Function] int shishi_kdcreq_add_padata_tgs (Shishi * handle,

Shishi asn1 kdcreq, Shishi asn1 apreq)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ to add PA-DATA to. apreq: AP-REQ to add as PA-DATA. Description: Add TGS pre-authentication data to KDC-REQ. The data is an AP-REQ that authenticates the request. This functions simply DER encodes the AP-REQ and calls shishi_kdcreq_add_padata() with a SHISHI PA TGS REQ padatatype. Return value: Returns SHISHI OK iff successful.

shishi kdcreq add padata preauth

[Function] int shishi_kdcreq_add_padata_preauth (Shishi * handle,

Shishi asn1 kdcreq, Shishi key * key)

handle: shishi handle as allocated by shishi_init(). kdcreq: KDC-REQ to add pre-authentication data to. key: Key used to encrypt pre-auth data. Description: Add pre-authentication data to KDC-REQ. Return value: Returns SHISHI OK iff successful.

shishi asrep

[Function] Shishi_asn1 shishi_asrep (Shishi * handle)

handle: shishi handle as allocated by shishi_init().

Description: This function creates a new AS-REP, populated with some default values.

Return value: Returns the AS-REP or NULL on failure.

shishi tgsrep

[Function] Shishi_asn1 shishi_tgsrep (Shishi * handle)

handle: shishi handle as allocated by shishi_init().

Description: This function creates a new TGS-REP, populated with some default values.

Return value: Returns the TGS-REP or NULL on failure.

shishi kdcrep print

[Function] int shishi_kdcrep_print (Shishi * handle, FILE * fh,

Shishi asn1 kdcrep) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. kdcrep: KDC-REP to print.

Description: Print ASCII armored DER encoding of KDC-REP to file.

Return value: Returns SHISHI OK iff successful.

shishi kdcrep save

[Function] int shishi_kdcrep_save (Shishi * handle, FILE * fh,

Shishi asn1 kdcrep) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. kdcrep: KDC-REP to save.

Description: Print DER encoding of KDC-REP to file.

Return value: Returns SHISHI OK iff successful.

shishi kdcrep to file

[Function] int shishi_kdcrep_to_file (Shishi * handle, Shishi asn1 kdcrep,
   int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). kdcrep: KDC-REP to save. filetype: input variable specifying type of file to be written, see Shishi filetype. filename: input variable with filename to write to.

Description: Write KDC-REP to file in specified TYPE. The file will be truncated if it exists.

Return value: Returns SHISHI OK iff successful.

shishi kdcrep parse

[Function] int shishi_kdcrep_parse (Shishi * handle, FILE * fh,

Shishi asn1 * kdcrep) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. kdcrep: output variable with newly allocated KDC-REP. Description: Read ASCII armored DER encoded KDC-REP from file and populate given variable. Return value: Returns SHISHI OK iff successful.

shishi kdcrep read

[Function] int shishi_kdcrep_read (Shishi * handle, FILE * fh,

Shishi asn1 * kdcrep) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. kdcrep: output variable with newly allocated KDC-REP. Description: Read DER encoded KDC-REP from file and populate given variable. Return value: Returns SHISHI OK iff successful.

shishi kdcrep from file

[Function] int shishi_kdcrep_from_file (Shishi * handle,

Shishi asn1 * kdcrep, int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). kdcrep: output variable with newly allocated KDC-REP. filetype: input variable specifying type of file to be read, see Shishi filetype. filename: input variable with filename to read from. Description: Read KDC-REP from file in specified TYPE. Return value: Returns SHISHI OK iff successful.

shishi kdcrep crealm set

[Function] int shishi_kdcrep_crealm_set (Shishi * handle,

Shishi asn1 kdcrep, const char * crealm)

handle: shishi handle as allocated by shishi_init(). kdcrep: Kdcrep variable to set realm field in. crealm: input array with name of realm. Description: Set the client realm field in the KDC-REP. Return value: Returns SHISHI OK iff successful.

shishi kdcrep cname set

[Function] int shishi_kdcrep_cname_set (Shishi * handle,

Shishi asn1 kdcrep, Shishi name type name_type, const char * cname[])

handle: shishi handle as allocated by shishi_init(). kdcrep: Kdcrep variable to set server name field in. name type: type of principial, see Shishi name type, usually SHISHI NT UNKNOWN.

cname: input array with principal name. Description: Set the client name field in the KDC-REP. Return value: Returns SHISHI OK iff successful.

shishi kdcrep client set

[Function] int shishi_kdcrep_client_set (Shishi * handle,

Shishi asn1 kdcrep, const char * client)

handle: shishi handle as allocated by shishi_init(). kdcrep: Kdcrep variable to set server name field in. client: zero-terminated string with principal name on RFC 1964 form. Description: Set the client name field in the KDC-REP. Return value: Returns SHISHI OK iff successful.

shishi kdcrep get enc part etype

[Function] int shishi_kdcrep_get_enc_part_etype (Shishi * handle,

Shishi asn1 kdcrep, int32 t * etype)

handle: shishi handle as allocated by shishi_init(). kdcrep: KDC-REP variable to get value from. etype: output variable that holds the value. Description: Extract KDC-REP.enc-part.etype. Return value: Returns SHISHI OK iff successful.

shishi kdcrep get ticket

[Function] int shishi_kdcrep_get_ticket (Shishi * handle,

Shishi asn1 kdcrep, Shishi asn1 * ticket)

handle: shishi handle as allocated by shishi_init(). kdcrep: KDC-REP variable to get ticket from. ticket: output variable to hold extracted ticket. Description: Extract ticket from KDC-REP. Return value: Returns SHISHI OK iff successful.

shishi kdcrep set ticket

[Function] int shishi_kdcrep_set_ticket (Shishi * handle,

Shishi asn1 kdcrep, Shishi asn1 ticket)

handle: shishi handle as allocated by shishi_init(). kdcrep: KDC-REP to add ticket field to. ticket: input ticket to copy into KDC-REP ticket field. Description: Copy ticket into KDC-REP. Return value: Returns SHISHI OK iff successful.

shishi kdcrep set enc part

[Function] int shishi_kdcrep_set_enc_part (Shishi * handle,

Shishi asn1 kdcrep, int32 t etype, uint32 t kvno, const char * buf,

    size t buflen)

handle: shishi handle as allocated by shishi_init(). kdcrep: KDC-REP to add enc-part field to. etype: encryption type used to encrypt enc-part. kvno: key version number. buf : input array with encrypted enc-part. buflen: size of input array with encrypted enc-part.

Description: Set the encrypted enc-part field in the KDC-REP. The encrypted data is usually created by calling shishi_encrypt() on the DER encoded enc-part. To save time, you may want to use shishi_kdcrep_add_enc_part() instead, which calculates the encrypted data and calls this function in one step.

Return value: Returns SHISHI OK iff successful.

shishi kdcrep add enc part

[Function] int shishi_kdcrep_add_enc_part (Shishi * handle,

Shishi asn1 kdcrep, Shishi key * key, int keyusage,

    Shishi asn1 enckdcreppart)

handle: shishi handle as allocated by shishi_init(). kdcrep: KDC-REP to add enc-part field to. key: key used to encrypt enc-part. keyusage: key usage to use, normally SHISHI KEYUSAGE ENCASREPPART, SHISHI KEYUSAGE ENCTGSREPPART SESSION KEY or SHISHI KEYUSAGE ENCTGSREPPART AUTHENTICATOR KEY.

enckdcreppart: EncKDCRepPart to add.

Description: Encrypts DER encoded EncKDCRepPart using key and stores it in the KDC-REP.

Return value: Returns SHISHI OK iff successful.

shishi kdcrep clear padata

[Function] int shishi_kdcrep_clear_padata (Shishi * handle,

Shishi asn1 kdcrep) handle: shishi handle as allocated by shishi_init(). kdcrep: KDC-REP to remove PA-DATA from.

Description: Remove the padata field from KDC-REP.

shishi enckdcreppart get key

[Function] int shishi_enckdcreppart_get_key (Shishi * handle,

Shishi asn1 enckdcreppart, Shishi key ** key)

handle: shishi handle as allocated by shishi_init(). enckdcreppart: input EncKDCRepPart variable. key: newly allocated encryption key handle. Description: Extract the key to use with the ticket sent in the KDC-REP associated with the EncKDCRepPart input variable. Return value: Returns SHISHI_OK iff successful.

shishi enckdcreppart key set

[Function] int shishi_enckdcreppart_key_set (Shishi * handle,

Shishi asn1 enckdcreppart, Shishi key * key)

handle: shishi handle as allocated by shishi_init(). enckdcreppart: input EncKDCRepPart variable. key: key handle with information to store in enckdcreppart. Description: Set the EncKDCRepPart.key field to key type and value of supplied key. Return value: Returns SHISHI_OK iff successful.

shishi enckdcreppart nonce set

[Function] int shishi_enckdcreppart_nonce_set (Shishi * handle,

Shishi asn1 enckdcreppart, uint32 t nonce)

handle: shishi handle as allocated by shishi_init(). enckdcreppart: input EncKDCRepPart variable. nonce: nonce to set in EncKDCRepPart. Description: Set the EncKDCRepPart.nonce field. Return value: Returns SHISHI_OK iff successful.

shishi enckdcreppart flags set

[Function] int shishi_enckdcreppart_flags_set (Shishi * handle,
   Shishi asn1 enckdcreppart, int flags)

handle: shishi handle as allocated by shishi_init(). enckdcreppart: input EncKDCRepPart variable. flags: flags to set in EncKDCRepPart. Description: Set the EncKDCRepPart.flags field. Return value: Returns SHISHI_OK iff successful.

shishi enckdcreppart authtime set

[Function] int shishi_enckdcreppart_authtime_set (Shishi * handle,

Shishi asn1 enckdcreppart, const char * authtime)

handle: shishi handle as allocated by shishi_init(). enckdcreppart: input EncKDCRepPart variable. authtime: character buffer containing a generalized time string.

Description: Set the EncTicketPart.authtime to supplied value.

Return value: Returns SHISHI_OK iff successful.

shishi enckdcreppart starttime set

[Function] int shishi_enckdcreppart_starttime_set (Shishi * handle,
   Shishi asn1 enckdcreppart, const char * starttime)

handle: shishi handle as allocated by shishi_init(). enckdcreppart: input EncKDCRepPart variable. starttime: character buffer containing a generalized time string.

Description: Set the EncTicketPart.starttime to supplied value. Use a NULL value for starttime to remove the field.

Return value: Returns SHISHI_OK iff successful.

shishi enckdcreppart endtime set

[Function] int shishi_enckdcreppart_endtime_set (Shishi * handle,

Shishi asn1 enckdcreppart, const char * endtime)

handle: shishi handle as allocated by shishi_init(). enckdcreppart: input EncKDCRepPart variable. endtime: character buffer containing a generalized time string.

Description: Set the EncTicketPart.endtime to supplied value.

Return value: Returns SHISHI_OK iff successful.

shishi enckdcreppart renew till set

[Function] int shishi_enckdcreppart_renew_till_set (Shishi * handle,
   Shishi asn1 enckdcreppart, const char * renew_till)

handle: shishi handle as allocated by shishi_init(). enckdcreppart: input EncKDCRepPart variable. renew till: character buffer containing a generalized time string.

Description: Set the EncTicketPart.renew-till to supplied value. Use a NULL value for renew till to remove the field.

Return value: Returns SHISHI_OK iff successful.

shishi enckdcreppart srealm set

[Function] int shishi_enckdcreppart_srealm_set (Shishi * handle,

Shishi asn1 enckdcreppart, const char * srealm)

handle: shishi handle as allocated by shishi_init(). enckdcreppart: EncKDCRepPart variable to set realm field in. srealm: input array with name of realm.

Description: Set the server realm field in the EncKDCRepPart.

Return value: Returns SHISHI OK iff successful.

shishi enckdcreppart sname set

[Function] int shishi_enckdcreppart_sname_set (Shishi * handle,

Shishi asn1 enckdcreppart, Shishi name type name_type,

char * sname[]) handle: shishi handle as allocated by shishi_init(). enckdcreppart: EncKDCRepPart variable to set server name field in. name type: type of principial, see Shishi name type, usually SHISHI NT UNKNOWN.

sname: input array with principal name. Description: Set the server name field in the EncKDCRepPart. Return value: Returns SHISHI OK iff successful.

shishi enckdcreppart populate encticketpart

[Function] int shishi_enckdcreppart_populate_encticketpart

(Shishi * handle, Shishi asn1 enckdcreppart,

    Shishi asn1 encticketpart)

handle: shishi handle as allocated by shishi_init(). enckdcreppart: input EncKDCRepPart variable. encticketpart: input EncTicketPart variable. Description: Set the flags, authtime, starttime, endtime, renew-till and caddr fields of the EncKDCRepPart to the corresponding values in the EncTicketPart. Return value: Returns SHISHI_OK iff successful.

5.11 Authenticator Functions

An “Authenticator” is an ASN.1 structure that work as a proof that an entity owns a ticket. It is usually embedded in the AP-REQ structure (see Section 5.4 [AP-REQ and AP-REP Functions], page 71), and you most likely want to use an AP-REQ instead of a Authenticator in normal applications. The following illustrates the Authenticator ASN.1 structure.

Authenticator   ::= [APPLICATION 2] SEQUENCE  {
   authenticator-vno       [0] INTEGER (5),
   crealm                  [1] Realm,
   cname                   [2] PrincipalName,
   cksum                   [3] Checksum OPTIONAL,
   cusec                   [4] Microseconds,
   ctime                   [5] KerberosTime,
   subkey                  [6] EncryptionKey OPTIONAL,
   seq-number              [7] UInt32 OPTIONAL,
   authorization-data      [8] AuthorizationData OPTIONAL

}

shishi authenticator

[Function] Shishi_asn1 shishi_authenticator (Shishi * handle)

handle: shishi handle as allocated by shishi_init().

Description: This function creates a new Authenticator, populated with some default values. It uses the current time as returned by the system for the ctime and cusec fields.

Return value: Returns the authenticator or NULL on failure.

shishi authenticator subkey

[Function] Shishi_asn1 shishi_authenticator_subkey (Shishi * handle)

handle: shishi handle as allocated by shishi_init().

Description: This function creates a new Authenticator, populated with some default values. It uses the current time as returned by the system for the ctime and cusec fields. It adds a random subkey.

Return value: Returns the authenticator or NULL on failure.

shishi authenticator print

[Function] int shishi_authenticator_print (Shishi * handle, FILE * fh,
    Shishi asn1 authenticator)

handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. authenticator: authenticator as allocated by shishi_authenticator().

Description: Print ASCII armored DER encoding of authenticator to file.

Return value: Returns SHISHI OK iff successful.

shishi authenticator save

[Function] int shishi_authenticator_save (Shishi * handle, FILE * fh,
    Shishi asn1 authenticator)

handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. authenticator: authenticator as allocated by shishi_authenticator().

Description: Save DER encoding of authenticator to file.

Return value: Returns SHISHI OK iff successful.

shishi authenticator to file

[Function] int shishi_authenticator_to_file (Shishi * handle,
   Shishi asn1 authenticator, int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). authenticator: Authenticator to save. filetype: input variable specifying type of file to be written, see Shishi filetype. filename: input variable with filename to write to.

Description: Write Authenticator to file in specified TYPE. The file will be truncated if it exists.

Return value: Returns SHISHI OK iff successful.

shishi authenticator parse

[Function] int shishi_authenticator_parse (Shishi * handle, FILE * fh,
    Shishi asn1 * authenticator)

handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. authenticator: output variable with newly allocated authenticator.

Description: Read ASCII armored DER encoded authenticator from file and populate given authenticator variable.

Return value: Returns SHISHI OK iff successful.

shishi authenticator read

[Function] int shishi_authenticator_read (Shishi * handle, FILE * fh,
    Shishi asn1 * authenticator)

handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. authenticator: output variable with newly allocated authenticator.

Description: Read DER encoded authenticator from file and populate given authenticator variable.

Return value: Returns SHISHI OK iff successful.

shishi authenticator from file

[Function] int shishi_authenticator_from_file (Shishi * handle,
   Shishi asn1 * authenticator, int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). authenticator: output variable with newly allocated Authenticator. filetype: input variable specifying type of file to be read, see Shishi filetype. filename: input variable with filename to read from.

Description: Read Authenticator from file in specified TYPE.

Return value: Returns SHISHI OK iff successful.

shishi authenticator set crealm

[Function] int shishi_authenticator_set_crealm (Shishi * handle,

Shishi asn1 authenticator, const char * crealm)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). crealm: input array with realm.

Description: Set realm field in authenticator to specified value.

shishi authenticator set cname

[Function] int shishi_authenticator_set_cname (Shishi * handle,

Shishi asn1 authenticator, Shishi name type name_type,

const char * cname[]) handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). name type: type of principial, see Shishi name type, usually SHISHI NT UNKNOWN.

cname: input array with principal name.

Description: Set principal field in authenticator to specified value.

Return value: Returns SHISHI OK iff successful.

shishi authenticator client set

[Function] int shishi_authenticator_client_set (Shishi * handle,

Shishi asn1 authenticator, const char * client)

handle: shishi handle as allocated by shishi_init(). authenticator: Authenticator to set client name field in. client: zero-terminated string with principal name on RFC 1964 form.

Description: Set the client name field in the Authenticator.

Return value: Returns SHISHI OK iff successful.

shishi authenticator ctime

[Function] int shishi_authenticator_ctime (Shishi * handle,

Shishi asn1 authenticator, char ** t)

handle: shishi handle as allocated by shishi_init(). authenticator: Authenticator as allocated by shishi_authenticator(). t: newly allocated zero-terminated character array with client time.

Description: Extract client time from Authenticator.

Return value: Returns SHISHI OK iff successful.

shishi authenticator ctime set

[Function] int shishi_authenticator_ctime_set (Shishi * handle,

Shishi asn1 authenticator, const char * t)

handle: shishi handle as allocated by shishi_init(). authenticator: Authenticator as allocated by shishi_authenticator(). t: string with generalized time value to store in Authenticator.

Description: Store client time in Authenticator.

shishi authenticator cusec get

[Function] int shishi_authenticator_cusec_get (Shishi * handle,

Shishi asn1 authenticator, uint32 t * cusec)

handle: shishi handle as allocated by shishi_init(). authenticator: Authenticator as allocated by shishi_authenticator(). cusec: output integer with client microseconds field. Description: Extract client microseconds field from Authenticator. Return value: Returns SHISHI OK iff successful.

shishi authenticator cusec set

[Function] int shishi_authenticator_cusec_set (Shishi * handle,

Shishi asn1 authenticator, uint32 t cusec)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). cusec: client microseconds to set in authenticator, 0-999999. Description: Set the cusec field in the Authenticator. Return value: Returns SHISHI OK iff successful.

shishi authenticator seqnumber get

[Function] int shishi_authenticator_seqnumber_get (Shishi * handle,
   Shishi asn1 authenticator, uint32 t * seqnumber)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). seqnumber: output integer with sequence number field. Description: Extract sequence number field from Authenticator. Return value: Returns SHISHI_OK iff successful.

shishi authenticator seqnumber remove

[Function] int shishi_authenticator_seqnumber_remove (Shishi * handle,
    Shishi asn1 authenticator)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). Description: Remove sequence number field in Authenticator. Return value: Returns SHISHI_OK iff successful.

shishi authenticator seqnumber set

[Function] int shishi_authenticator_seqnumber_set (Shishi * handle,
   Shishi asn1 authenticator, uint32 t seqnumber)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). seqnumber: integer with sequence number field to store in Authenticator. Description: Store sequence number field in Authenticator. Return value: Returns SHISHI_OK iff successful.

shishi authenticator client

[Function] int shishi_authenticator_client (Shishi * handle,
   Shishi asn1 authenticator, char ** client, size t * clientlen)

handle: Shishi library handle create by shishi_init(). authenticator: Authenticator variable to get client name from. client: pointer to newly allocated zero terminated string containing principal name. May be NULL (to only populate clientlen). clientlen: pointer to length of client on output, excluding terminating zero. May be NULL (to only populate client).

Description: Represent client principal name in Authenticator as zero-terminated string. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length clientlen does not include the terminating zero.

Return value: Returns SHISHI OK iff successful.

shishi authenticator clientrealm

[Function] int shishi_authenticator_clientrealm (Shishi * handle,
   Shishi asn1 authenticator, char ** client, size t * clientlen)

handle: Shishi library handle create by shishi_init(). authenticator: Authenticator variable to get client name and realm from. client: pointer to newly allocated zero terminated string containing principal name and realm. May be NULL (to only populate clientlen). clientlen: pointer to length of client on output, excluding terminating zero. May be NULL (to only populate client).

Description: Convert cname and realm fields from Authenticator to printable principal name format. The string is allocate by this function, and it is the responsibility of the caller to deallocate it. Note that the output length clientlen does not include the terminating zero.

Return value: Returns SHISHI OK iff successful.

shishi authenticator cksum

[Function] int shishi_authenticator_cksum (Shishi * handle,
   Shishi asn1 authenticator, int32 t * cksumtype, char ** cksum,
    size t * cksumlen)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). cksumtype: output checksum type. cksum: newly allocated output checksum data from authenticator. cksumlen: on output, actual size of allocated output checksum data buffer.

Description: Read checksum value from authenticator. cksum is allocated by this function, and it is the responsibility of caller to deallocate it.

Return value: Returns SHISHI OK iff successful.

shishi authenticator set cksum

[Function] int shishi_authenticator_set_cksum (Shishi * handle,
   Shishi asn1 authenticator, int32 t cksumtype, char * cksum,
    size t cksumlen)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). cksumtype: input checksum type to store in authenticator. cksum: input checksum data to store in authenticator. cksumlen: size of input checksum data to store in authenticator.

Description: Store checksum value in authenticator. A checksum is usually created by calling shishi_checksum() on some application specific data using the key from the ticket that is being used. To save time, you may want to use shishi_authenticator_ add_cksum() instead, which calculates the checksum and calls this function in one step.

Return value: Returns SHISHI OK iff successful.

shishi authenticator add cksum

[Function] int shishi_authenticator_add_cksum (Shishi * handle,

Shishi asn1 authenticator, Shishi key * key, int keyusage, char * data, size t datalen)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). key: key to to use for encryption. keyusage: cryptographic key usage value to use in encryption. data: input array with data to calculate checksum on. datalen: size of input array with data to calculate checksum on.

Description: Calculate checksum for data and store it in the authenticator.

Return value: Returns SHISHI OK iff successful.

shishi authenticator add cksum type

[Function] int shishi_authenticator_add_cksum_type (Shishi * handle,

Shishi asn1 authenticator, Shishi key * key, int keyusage,

   int cksumtype, char * data, size t datalen)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). key: key to to use for encryption. keyusage: cryptographic key usage value to use in encryption. cksumtype: checksum to type to calculate checksum. data: input array with data to calculate checksum on. datalen: size of input array with data to calculate checksum on.

Description: Calculate checksum for data and store it in the authenticator.

Return value: Returns SHISHI OK iff successful.

shishi authenticator clear authorizationdata

[Function] int shishi_authenticator_clear_authorizationdata

(Shishi * handle, Shishi asn1 authenticator) handle: shishi handle as allocated by shishi_init(). authenticator: Authenticator as allocated by shishi_authenticator().

Description: Remove the authorization-data field from Authenticator.

Return value: Returns SHISHI OK iff successful.

shishi authenticator add authorizationdata

[Function] int shishi_authenticator_add_authorizationdata

(Shishi * handle, Shishi asn1 authenticator, int32 t adtype, const char * addata, size t addatalen)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). adtype: input authorization data type to add. addata: input authorization data to add. addatalen: size of input authorization data to add.

Description: Add authorization data to authenticator.

Return value: Returns SHISHI OK iff successful.

shishi authenticator authorizationdata

[Function] int shishi_authenticator_authorizationdata (Shishi * handle,

Shishi asn1 authenticator, int32 t * adtype, char ** addata, size t * addatalen, size t nth)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). adtype: output authorization data type. addata: newly allocated output authorization data. addatalen: on output, actual size of newly allocated authorization data. nth: element number of authorization-data to extract.

Extract n: th authorization data from authenticator. The first field is 1.

Return value: Returns SHISHI OK iff successful.

shishi authenticator remove subkey

[Function] int shishi_authenticator_remove_subkey (Shishi * handle,
    Shishi asn1 authenticator)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator().

Description: Remove subkey from the authenticator.

Return value: Returns SHISHI OK iff successful.

shishi authenticator get subkey

[Function] int shishi_authenticator_get_subkey (Shishi * handle,

Shishi asn1 authenticator, Shishi key ** subkey)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). subkey: output newly allocated subkey from authenticator. Description: Read subkey value from authenticator. Return value: Returns SHISHI OK if successful or SHISHI ASN1 NO ELEMENT if subkey is not present.

shishi authenticator set subkey

[Function] int shishi_authenticator_set_subkey (Shishi * handle,
   Shishi asn1 authenticator, int32 t subkeytype, const char * subkey,
    size t subkeylen)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). subkeytype: input subkey type to store in authenticator. subkey: input subkey data to store in authenticator. subkeylen: size of input subkey data to store in authenticator. Description: Store subkey value in authenticator. A subkey is usually created by calling shishi_key_random() using the default encryption type of the key from the ticket that is being used. To save time, you may want to use shishi_authenticator_ add_subkey() instead, which calculates the subkey and calls this function in one step. Return value: Returns SHISHI OK iff successful.

shishi authenticator add random subkey

[Function] int shishi_authenticator_add_random_subkey (Shishi * handle,
    Shishi asn1 authenticator)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). Description: Generate random subkey, of the default encryption type from configuration, and store it in the authenticator. Return value: Returns SHISHI OK iff successful.

shishi authenticator add random subkey etype

[Function] int shishi_authenticator_add_random_subkey_etype
   (Shishi * handle, Shishi asn1 authenticator, int etype)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). etype: encryption type of random key to generate. Description: Generate random subkey of indicated encryption type, and store it in the authenticator. Return value: Returns SHISHI OK iff successful.

shishi authenticator add subkey

[Function] int shishi_authenticator_add_subkey (Shishi * handle,

Shishi asn1 authenticator, Shishi key * subkey)

handle: shishi handle as allocated by shishi_init(). authenticator: authenticator as allocated by shishi_authenticator(). subkey: subkey to add to authenticator. Description: Store subkey in the authenticator. Return value: Returns SHISHI OK iff successful.

5.12 KRB-ERROR Functions

The “KRB-ERROR” is an ASN.1 structure that can be returned, instead of, e.g., KDC-REP or AP-REP, to indicate various error conditions. Unfortunately, the semantics of several of the fields are ill specified, so the typically procedure is to extract “e-text” and/or “e-data” and show it to the user. The following illustrates the KRB-ERROR ASN.1 structure.

KRB-ERROR       ::= [APPLICATION 30] SEQUENCE {
pvno            [0] INTEGER (5),
msg-type        [1] INTEGER (30),
ctime           [2] KerberosTime OPTIONAL,
cusec           [3] Microseconds OPTIONAL,
stime           [4] KerberosTime,
susec           [5] Microseconds,
error-code      [6] Int32,
crealm          [7] Realm OPTIONAL,
cname           [8] PrincipalName OPTIONAL,
realm           [9] Realm -- service realm --,
sname           [10] PrincipalName -- service name --,
e-text          [11] KerberosString OPTIONAL,
e-data          [12] OCTET STRING OPTIONAL
}

shishi krberror

[Function] Shishi_asn1 shishi_krberror (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: This function creates a new KRB-ERROR, populated with some default values. Return value: Returns the KRB-ERROR or NULL on failure.

shishi krberror print

[Function] int shishi_krberror_print (Shishi * handle, FILE * fh,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. krberror: KRB-ERROR to print.

Description: Print ASCII armored DER encoding of KRB-ERROR to file.

Return value: Returns SHISHI OK iff successful.

shishi krberror save

[Function] int shishi_krberror_save (Shishi * handle, FILE * fh,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). fh: file handle open for writing. krberror: KRB-ERROR to save.

Description: Save DER encoding of KRB-ERROR to file.

Return value: Returns SHISHI OK iff successful.

shishi krberror to file

[Function] int shishi_krberror_to_file (Shishi * handle,

Shishi asn1 krberror, int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). krberror: KRB-ERROR to save. filetype: input variable specifying type of file to be written, see Shishi filetype. filename: input variable with filename to write to.

Description: Write KRB-ERROR to file in specified TYPE. The file will be truncated if it exists.

Return value: Returns SHISHI OK iff successful.

shishi krberror parse

[Function] int shishi_krberror_parse (Shishi * handle, FILE * fh,

Shishi asn1 * krberror) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. krberror: output variable with newly allocated KRB-ERROR.

Description: Read ASCII armored DER encoded KRB-ERROR from file and populate given variable.

Return value: Returns SHISHI OK iff successful.

shishi krberror read

[Function] int shishi_krberror_read (Shishi * handle, FILE * fh,

Shishi asn1 * krberror) handle: shishi handle as allocated by shishi_init(). fh: file handle open for reading. krberror: output variable with newly allocated KRB-ERROR.

Description: Read DER encoded KRB-ERROR from file and populate given variable.

Return value: Returns SHISHI OK iff successful.

shishi krberror from file

[Function] int shishi_krberror_from_file (Shishi * handle,

Shishi asn1 * krberror, int filetype, const char * filename)

handle: shishi handle as allocated by shishi_init(). krberror: output variable with newly allocated KRB-ERROR. filetype: input variable specifying type of file to be read, see Shishi filetype. filename: input variable with filename to read from.

Description: Read KRB-ERROR from file in specified TYPE.

Return value: Returns SHISHI OK iff successful.

shishi krberror build

[Function] int shishi_krberror_build (Shishi * handle,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror().

Description: Finish KRB-ERROR, called before e.g. shishi krberror der. This function removes empty but OPTIONAL fields (such as cname), and

Return value: Returns SHISHI OK iff successful.

shishi krberror der

[Function] int shishi_krberror_der (Shishi * handle, Shishi asn1 krberror,

char ** out, size t * outlen)

handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). out: output array with newly allocated DER encoding of KRB-ERROR. outlen: length of output array with DER encoding of KRB-ERROR.

Description: DER encode KRB-ERROR. The caller must deallocate the OUT buffer.

Return value: Returns SHISHI OK iff successful.

shishi krberror crealm

[Function] int shishi_krberror_crealm (Shishi * handle,

Shishi asn1 krberror, char ** realm, size t * realmlen)

handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). realm: output array with newly allocated name of realm in KRB-ERROR. realmlen: size of output array.

Description: Extract client realm from KRB-ERROR.

shishi krberror remove crealm

[Function] int shishi_krberror_remove_crealm (Shishi * handle,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror().

Description: Remove client realm field in KRB-ERROR.

Return value: Returns SHISHI OK iff successful.

shishi krberror set crealm

[Function] int shishi_krberror_set_crealm (Shishi * handle,

Shishi asn1 krberror, const char * crealm)

handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). crealm: input array with realm.

Description: Set realm field in krberror to specified value.

Return value: Returns SHISHI OK iff successful.

shishi krberror client

[Function] int shishi_krberror_client (Shishi * handle,

Shishi asn1 krberror, char ** client, size t * clientlen)

handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). client: pointer to newly allocated zero terminated string containing principal name. May be NULL (to only populate clientlen). clientlen: pointer to length of client on output, excluding terminating zero. May be NULL (to only populate client).

Description: Return client principal name in KRB-ERROR.

Return value: Returns SHISHI OK iff successful.

shishi krberror set cname

[Function] int shishi_krberror_set_cname (Shishi * handle,

Shishi asn1 krberror, Shishi name type name_type,

const char * cname[]) handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). name type: type of principial, see Shishi name type, usually SHISHI NT UNKNOWN.

cname: input array with principal name.

Description: Set principal field in krberror to specified value.

Return value: Returns SHISHI OK iff successful.

shishi krberror remove cname

[Function] int shishi_krberror_remove_cname (Shishi * handle,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). Description: Remove client realm field in KRB-ERROR. Return value: Returns SHISHI OK iff successful.

shishi krberror client set

[Function] int shishi_krberror_client_set (Shishi * handle,

Shishi asn1 krberror, const char * client)

handle: shishi handle as allocated by shishi_init(). krberror: Krberror to set client name field in. client: zero-terminated string with principal name on RFC 1964 form. Description: Set the client name field in the Krberror. Return value: Returns SHISHI OK iff successful.

shishi krberror realm

[Function] int shishi_krberror_realm (Shishi * handle,

Shishi asn1 krberror, char ** realm, size t * realmlen)

handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). realm: output array with newly allocated name of realm in KRB-ERROR. realmlen: size of output array. Description: Extract (server) realm from KRB-ERROR. Return value: Returns SHISHI OK iff successful.

shishi krberror set realm

[Function] int shishi_krberror_set_realm (Shishi * handle,

Shishi asn1 krberror, const char * realm)

handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). realm: input array with (server) realm. Description: Set (server) realm field in krberror to specified value. Return value: Returns SHISHI OK iff successful.

shishi krberror server

[Function] int shishi_krberror_server (Shishi * handle,

Shishi asn1 krberror, char ** server, size t * serverlen)

handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). server: pointer to newly allocated zero terminated string containing server name.

May be NULL (to only populate serverlen). serverlen: pointer to length of server on output, excluding terminating zero. May be NULL (to only populate server). Description: Return server principal name in KRB-ERROR. Return value: Returns SHISHI OK iff successful.

shishi krberror remove sname

[Function] int shishi_krberror_remove_sname (Shishi * handle,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). krberror: Krberror to set server name field in. Description: Remove server name field in KRB-ERROR. (Since it is not marked OPTIONAL in the ASN.1 profile, what is done is to set the name-type to UNKNOWN and make sure the name-string sequence is empty.) Return value: Returns SHISHI OK iff successful.

shishi krberror set sname

[Function] int shishi_krberror_set_sname (Shishi * handle,

Shishi asn1 krberror, Shishi name type name_type,

const char * sname[]) handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). name type: type of principial, see Shishi name type, usually SHISHI NT UNKNOWN.

sname: input array with principal name. Description: Set principal field in krberror to specified value. Return value: Returns SHISHI OK iff successful.

shishi krberror server set

[Function] int shishi_krberror_server_set (Shishi * handle,

Shishi asn1 krberror, const char * server)

handle: shishi handle as allocated by shishi_init(). krberror: Krberror to set server name field in. server: zero-terminated string with principal name on RFC 1964 form. Description: Set the server name field in the Krberror. Return value: Returns SHISHI OK iff successful.

shishi krberror ctime

[Function] int shishi_krberror_ctime (Shishi * handle,

Shishi asn1 krberror, char ** t)

handle: shishi handle as allocated by shishi_init(). krberror: Krberror to set client name field in. t: newly allocated zero-terminated output array with client time.

Description: Extract client time from KRB-ERROR. Return value: Returns SHISHI OK iff successful.

shishi krberror ctime set

[Function] int shishi_krberror_ctime_set (Shishi * handle,

Shishi asn1 krberror, const char * t)

handle: shishi handle as allocated by shishi_init(). krberror: Krberror as allocated by shishi_krberror(). t: string with generalized time value to store in Krberror. Description: Store client time in Krberror. Return value: Returns SHISHI OK iff successful.

shishi krberror remove ctime

[Function] int shishi_krberror_remove_ctime (Shishi * handle,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). krberror: Krberror as allocated by shishi_krberror(). Description: Remove client time field in Krberror. Return value: Returns SHISHI OK iff successful.

shishi krberror cusec

[Function] int shishi_krberror_cusec (Shishi * handle,

Shishi asn1 krberror, uint32 t * cusec)

handle: shishi handle as allocated by shishi_init(). krberror: Krberror as allocated by shishi_krberror(). cusec: output integer with client microseconds field. Description: Extract client microseconds field from Krberror. Return value: Returns SHISHI OK iff successful.

shishi krberror cusec set

[Function] int shishi_krberror_cusec_set (Shishi * handle,

Shishi asn1 krberror, uint32 t cusec)

handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). cusec: client microseconds to set in krberror, 0-999999. Description: Set the cusec field in the Krberror. Return value: Returns SHISHI OK iff successful.

shishi krberror remove cusec

[Function] int shishi_krberror_remove_cusec (Shishi * handle,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). krberror: Krberror as allocated by shishi_krberror().

Description: Remove client usec field in Krberror.

Return value: Returns SHISHI OK iff successful.

shishi krberror stime

[Function] int shishi_krberror_stime (Shishi * handle,

Shishi asn1 krberror, char ** t)

handle: shishi handle as allocated by shishi_init(). krberror: Krberror to set client name field in. t: newly allocated zero-terminated output array with server time.

Description: Extract server time from KRB-ERROR.

Return value: Returns SHISHI OK iff successful.

shishi krberror stime set

[Function] int shishi_krberror_stime_set (Shishi * handle,

Shishi asn1 krberror, const char * t)

handle: shishi handle as allocated by shishi_init(). krberror: Krberror as allocated by shishi_krberror(). t: string with generalized time value to store in Krberror.

Description: Store server time in Krberror.

Return value: Returns SHISHI OK iff successful.

shishi krberror susec

[Function] int shishi_krberror_susec (Shishi * handle,

Shishi asn1 krberror, uint32 t * susec)

handle: shishi handle as allocated by shishi_init(). krberror: Krberror as allocated by shishi_krberror(). susec: output integer with server microseconds field.

Description: Extract server microseconds field from Krberror.

Return value: Returns SHISHI OK iff successful.

shishi krberror susec set

[Function] int shishi_krberror_susec_set (Shishi * handle,

Shishi asn1 krberror, uint32 t susec)

handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). susec: server microseconds to set in krberror, 0-999999.

Description: Set the susec field in the Krberror.

Return value: Returns SHISHI OK iff successful.

shishi krberror errorcode

[Function] int shishi_krberror_errorcode (Shishi * handle,

Shishi asn1 krberror, int32 t * errorcode)

handle: shishi handle as allocated by shishi_init(). krberror: KRB-ERROR structure with error code. errorcode: output integer KRB-ERROR error code. Description: Extract error code from KRB-ERROR. Return value: Returns SHISHI OK iff successful.

shishi krberror errorcode fast

[Function] int shishi_krberror_errorcode_fast (Shishi * handle,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). krberror: KRB-ERROR structure with error code. Description: Get error code from KRB-ERROR, without error checking. Return value: Return error code (see shishi_krberror_errorcode()) directly, or -1 on error.

shishi krberror errorcode set

[Function] int shishi_krberror_errorcode_set (Shishi * handle,
   Shishi asn1 krberror, int errorcode)

handle: shishi handle as allocated by shishi_init(). krberror: KRB-ERROR structure with error code to set. errorcode: new error code to set in krberror. Description: Set the error-code field to a new error code. Return value: Returns SHISHI OK iff successful.

shishi krberror etext

[Function] int shishi_krberror_etext (Shishi * handle,

Shishi asn1 krberror, char ** etext, size t * etextlen)

handle: shishi handle as allocated by shishi_init(). krberror: KRB-ERROR structure with error code. etext: output array with newly allocated error text. etextlen: output length of error text. Description: Extract additional error text from server (possibly empty). Return value: Returns SHISHI OK iff successful.

shishi krberror set etext

[Function] int shishi_krberror_set_etext (Shishi * handle,

Shishi asn1 krberror, const char * etext)

handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). etext: input array with error text to set.

Description: Set error text (e-text) field in KRB-ERROR to specified value.

Return value: Returns SHISHI OK iff successful.

shishi krberror remove etext

[Function] int shishi_krberror_remove_etext (Shishi * handle,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror().

Description: Remove error text (e-text) field in KRB-ERROR.

Return value: Returns SHISHI OK iff successful.

shishi krberror edata

[Function] int shishi_krberror_edata (Shishi * handle,

Shishi asn1 krberror, char ** edata, size t * edatalen)

handle: shishi handle as allocated by shishi_init(). krberror: KRB-ERROR structure with error code. edata: output array with newly allocated error data. edatalen: output length of error data.

Description: Extract additional error data from server (possibly empty).

Return value: Returns SHISHI OK iff successful.

shishi krberror methoddata

[Function] int shishi_krberror_methoddata (Shishi * handle,

Shishi asn1 krberror, Shishi asn1 * methoddata)

handle: shishi handle as allocated by shishi_init(). krberror: KRB-ERROR structure with error code. methoddata: output ASN.1 METHOD-DATA.

Description: Extract METHOD-DATA ASN.1 object from the e-data field. The edata field will only contain a METHOD-DATA if the krberror error code is SHISHI_

KDC_ERR_PREAUTH_REQUIRED.

Return value: Returns SHISHI OK iff successful.

shishi krberror set edata

[Function] int shishi_krberror_set_edata (Shishi * handle,

Shishi asn1 krberror, const char * edata)

handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). edata: input array with error text to set.

Description: Set error text (e-data) field in KRB-ERROR to specified value.

Return value: Returns SHISHI OK iff successful.

shishi krberror remove edata

[Function] int shishi_krberror_remove_edata (Shishi * handle,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). krberror: krberror as allocated by shishi_krberror(). Description: Remove error text (e-data) field in KRB-ERROR. Return value: Returns SHISHI OK iff successful.

shishi krberror pretty print

[Function] int shishi_krberror_pretty_print (Shishi * handle, FILE * fh,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). fh: file handle opened for writing. krberror: KRB-ERROR structure with error code. Description: Print KRB-ERROR error condition and some explanatory text to file descriptor. Return value: Returns SHISHI OK iff successful.

shishi krberror errorcode message

[Function] const char * shishi_krberror_errorcode_message
    (Shishi * handle, int errorcode)

handle: shishi handle as allocated by shishi_init(). errorcode: integer KRB-ERROR error code. Description: Get human readable string describing KRB-ERROR code. Return value: Return a string describing error code. This function will always return a string even if the error code isn’t known.

shishi krberror message

[Function] const char * shishi_krberror_message (Shishi * handle,

Shishi asn1 krberror) handle: shishi handle as allocated by shishi_init(). krberror: KRB-ERROR structure with error code. Description: Extract error code (see shishi_krberror_errorcode_fast()) and re-

turn error message (see shishi_krberror_errorcode_message()).

Return value: Return a string describing error code. This function will always return a string even if the error code isn’t known.

5.13 Cryptographic Functions

Underneath the high-level functions described earlier, cryptographic operations are happening. If you need to access these cryptographic primitives directly, this section describes the functions available.

Most cryptographic operations need keying material, and cryptographic keys have been isolated into it’s own data structure Shishi_key. The following illustrates it’s contents, but note that you cannot access it’s elements directly but must use the accessor functions described below.

struct Shishi_key
{
int type;    /* RFC 1510 encryption integer type */
char *value; /* Cryptographic key data */
int version; /* RFC 1510 ``kvno'' */
};

All functions that operate on this data structure are described now.

shishi key principal

[Function] const char * shishi_key_principal (const Shishi key * key)

key: structure that holds key information

Description: Get the principal part of the key owner principal name, i.e., except the realm.

Return value: Returns the principal owning the key. (Not a copy of it, so don’t modify or deallocate it.)

shishi key principal set

[Function] void shishi_key_principal_set (Shishi key * key,

const char * principal) key: structure that holds key information principal: string with new principal name.

Description: Set the principal owning the key. The string is copied into the key, so you can dispose of the variable immediately after calling this function.

shishi key realm

[Function] const char * shishi_key_realm (const Shishi key * key)

key: structure that holds key information

Description: Get the realm part of the key owner principal name.

Return value: Returns the realm for the principal owning the key. (Not a copy of it, so don’t modify or deallocate it.)

shishi key realm set

[Function] void shishi_key_realm_set (Shishi key * key, const char * realm)

key: structure that holds key information realm: string with new realm name.

Description: Set the realm for the principal owning the key. The string is copied into the key, so you can dispose of the variable immediately after calling this function.

shishi key type

[Function] int shishi_key_type (const Shishi key * key)

key: structure that holds key information Description: Get key type. Return value: Returns the type of key as an integer as described in the standard.

shishi key type set

[Function] void shishi_key_type_set (Shishi key * key, int32 t type)

key: structure that holds key information type: type to set in key. Description: Set the type of key in key structure.

shishi key value

[Function] const char * shishi_key_value (const Shishi key * key)

key: structure that holds key information Description: Get the raw key bytes. Return value: Returns the key value as a pointer which is valid throughout the lifetime of the key structure.

shishi key value set

[Function] void shishi_key_value_set (Shishi key * key, const char * value)

key: structure that holds key information value: input array with key data. Description: Set the key value and length in key structure. The value is copied into the key (in other words, you can deallocate value right after calling this function without modifying the value inside the key).

shishi key version

[Function] uint32_t shishi_key_version (const Shishi key * key)

key: structure that holds key information Description: Get the "kvno" (key version) of key. It will be UINT32 MAX if the key is not long-lived. Return value: Returns the version of key ("kvno").

shishi key version set

[Function] void shishi_key_version_set (Shishi key * key, uint32 t kvno)

key: structure that holds key information kvno: new version integer. Description: Set the version of key ("kvno") in key structure. Use UINT32 MAX for non-ptermanent keys.

shishi key timestamp

[Function] time_t shishi_key_timestamp (const Shishi key * key)

key: structure that holds key information Description: Get the time the key was established. Typically only present when the key was imported from a keytab format. Return value: Returns the time the key was established, or (time t)-1 if not available. Since: 0.0.42

shishi key timestamp set

[Function] void shishi_key_timestamp_set (Shishi key * key,
    time t timestamp)

key: structure that holds key information timestamp: new timestamp. Description: Set the time the key was established. Typically only relevant when exporting the key to keytab format. Since: 0.0.42

shishi key name

[Function] const char * shishi_key_name (Shishi key * key)

key: structure that holds key information Description: Calls shishi cipher name for key type. Return value: Return name of key.

shishi key length

[Function] size_t shishi_key_length (const Shishi key * key)

key: structure that holds key information Description: Calls shishi cipher keylen for key type. Return value: Returns the length of the key value.

shishi key

[Function] int shishi_key (Shishi * handle, Shishi key ** key)

handle: Shishi library handle create by shishi_init(). key: pointer to structure that will hold newly created key information Description: Create a new Key information structure. Return value: Returns SHISHI OK iff successful.

shishi key done

[Function] void shishi_key_done (Shishi key * key)

key: pointer to structure that holds key information. Description: Deallocates key information structure.

shishi key copy

[Function] void shishi_key_copy (Shishi key * dstkey, Shishi key * srckey)

dstkey: structure that holds destination key information srckey: structure that holds source key information

Description: Copies source key into existing allocated destination key.

shishi key from value

[Function] int shishi_key_from_value (Shishi * handle, int32 t type,

const char * value, Shishi key ** key)

handle: Shishi library handle create by shishi_init(). type: type of key. value: input array with key value, or NULL. key: pointer to structure that will hold newly created key information

Description: Create a new Key information structure, and set the key type and key value. KEY contains a newly allocated structure only if this function is successful.

Return value: Returns SHISHI OK iff successful.

shishi key from base64

[Function] int shishi_key_from_base64 (Shishi * handle, int32 t type,

const char * value, Shishi key ** key)

handle: Shishi library handle create by shishi_init(). type: type of key. value: input string with base64 encoded key value, or NULL. key: pointer to structure that will hold newly created key information

Description: Create a new Key information structure, and set the key type and key value. KEY contains a newly allocated structure only if this function is successful.

Return value: Returns SHISHI INVALID KEY if the base64 encoded key length doesn’t match the key type, and SHISHI OK on success.

shishi key random

[Function] int shishi_key_random (Shishi * handle, int32 t type,

Shishi key ** key) handle: Shishi library handle create by shishi_init(). type: type of key. key: pointer to structure that will hold newly created key information

Description: Create a new Key information structure for the key type and some random data. KEY contains a newly allocated structure only if this function is successful.

Return value: Returns SHISHI OK iff successful.

shishi key from random

[Function] int shishi_key_from_random (Shishi * handle, int32 t type,

const char * rnd, size t rndlen, Shishi key ** outkey)

handle: Shishi library handle create by shishi_init(). type: type of key. rnd: random data. rndlen: length of random data. outkey: pointer to structure that will hold newly created key information Description: Create a new Key information structure, and set the key type and key value using shishi_random_to_key(). KEY contains a newly allocated structure only if this function is successful. Return value: Returns SHISHI OK iff successful.

shishi key from string

[Function] int shishi_key_from_string (Shishi * handle, int32 t type,

const char * password, size t passwordlen, const char * salt, size t saltlen, const char * parameter, Shishi key ** outkey)

handle: Shishi library handle create by shishi_init(). type: type of key. password: input array containing password. passwordlen: length of input array containing password. salt: input array containing salt. saltlen: length of input array containing salt. parameter: input array with opaque encryption type specific information. outkey: pointer to structure that will hold newly created key information Description: Create a new Key information structure, and set the key type and key value using shishi_string_to_key(). KEY contains a newly allocated structure only if this function is successful. Return value: Returns SHISHI OK iff successful.

shishi key from name

[Function] int shishi_key_from_name (Shishi * handle, int32 t type,

const char * name, const char * password, size t passwordlen, const char * parameter, Shishi key ** outkey)

handle: Shishi library handle create by shishi_init(). type: type of key. name: principal name of user. password: input array containing password. passwordlen: length of input array containing password. parameter: input array with opaque encryption type specific information. outkey: pointer to structure that will hold newly created key information Description: Create a new Key information structure, and derive the key from principal name and password using shishi_key_from_name(). The salt is derived from the principal name by concatenating the decoded realm and principal.

Return value: Returns SHISHI OK iff successful.

Applications that run uninteractively may need keying material. In these cases, the keys are stored in a file, a file that is normally stored on the local host. The file should be protected from unauthorized access. The file is in ASCII format and contains keys as outputed by shishi_key_print. All functions that handle these keys sets are described now.

shishi keys

[Function] int shishi_keys (Shishi * handle, Shishi keys ** keys)

handle: shishi handle as allocated by shishi_init(). keys: output pointer to newly allocated keys handle. Description: Get a new key set handle. Return value: Returns SHISHI_OK iff successful.

shishi keys done

[Function] void shishi_keys_done (Shishi keys ** keys)

keys: key set handle as allocated by shishi_keys(). Description: Deallocates all resources associated with key set. The key set handle must not be used in calls to other shishi keys *() functions after this.

shishi keys size

[Function] int shishi_keys_size (Shishi keys * keys)

keys: key set handle as allocated by shishi_keys(). Description: Get size of key set. Return value: Returns number of keys stored in key set.

shishi keys nth

[Function] const Shishi_key * shishi_keys_nth (Shishi keys * keys,
    int keyno)

keys: key set handle as allocated by shishi_keys(). keyno: integer indicating requested key in key set. Get the n: th ticket in key set. Return value: Returns a key handle to the keyno:th key in the key set, or NULL if keys is invalid or keyno is out of bounds. The first key is keyno 0, the second key keyno 1, and so on.

shishi keys remove

[Function] void shishi_keys_remove (Shishi keys * keys, int keyno)

keys: key set handle as allocated by shishi_keys(). keyno: key number of key in the set to remove. The first key is key number 0. Description: Remove a key, indexed by keyno, in given key set.

shishi keys add

[Function] int shishi_keys_add (Shishi keys * keys, Shishi key * key)

keys: key set handle as allocated by shishi_keys(). key: key to be added to key set. Description: Add a key to the key set. A deep copy of the key is stored, so changing key, or deallocating it, will not modify the value stored in the key set. Return value: Returns SHISHI_OK iff successful.

shishi keys print

[Function] int shishi_keys_print (Shishi keys * keys, FILE * fh)

keys: key set to print. fh: file handle, open for writing, to print keys to. Description: Print all keys in set using shishi key print. Returns: Returns SHISHI_OK on success.

shishi keys to file

[Function] int shishi_keys_to_file (Shishi * handle, const char * filename,

Shishi keys * keys) handle: shishi handle as allocated by shishi_init(). filename: filename to append key to. keys: set of keys to print. Description: Print an ASCII representation of a key structure to a file, for each key in the key set. The file is appended to if it exists. See shishi_key_print() for the format of the output. Return value: Returns SHISHI_OK iff successful.

shishi keys from file

[Function] int shishi_keys_from_file (Shishi keys * keys,

const char * filename) keys: key set handle as allocated by shishi_keys(). filename: filename to read keys from. Description: Read zero or more keys from file filename and append them to the keyset keys. See shishi_key_print() for the format of the input. Return value: Returns SHISHI_OK iff successful. Since: 0.0.42

shishi keys for serverrealm in file

[Function] Shishi_key * shishi_keys_for_serverrealm_in_file

(Shishi * handle, const char * filename, const char * server,

const char * realm) handle: Shishi library handle create by shishi_init(). filename: file to read keys from. server: server name to get key for. realm: realm of server to get key for.

Description: Get keys that match specified server and realm from the key set file filename.

Return value: Returns the key for specific server and realm, read from the indicated file, or NULL if no key could be found or an error encountered.

shishi keys for server in file

[Function] Shishi_key * shishi_keys_for_server_in_file (Shishi * handle,

const char * filename, const char * server)

handle: Shishi library handle create by shishi_init(). filename: file to read keys from. server: server name to get key for.

Description: Get key for specified server from filename.

Return value: Returns the key for specific server, read from the indicated file, or NULL if no key could be found or an error encountered.

shishi keys for localservicerealm in file

[Function] Shishi_key * shishi_keys_for_localservicerealm_in_file

(Shishi * handle, const char * filename, const char * service,

const char * realm) handle: Shishi library handle create by shishi_init(). filename: file to read keys from. service: service to get key for. realm: realm of server to get key for, or NULL for default realm.

Description: Get key for specified service and realm from filename.

Return value: Returns the key for the server "SERVICE/HOSTNAME@REALM" (where HOSTNAME is the current system’s hostname), read from the default host keys file (see shishi_hostkeys_default_file()), or NULL if no key could be found or an error encountered.

The previous functions require that the filename is known. For some applications, servers, it makes sense to provide a system default. These key sets used by server applications are known as “hostkeys”. Here are the functions that operate on hostkeys (they are mostly wrappers around generic key sets).

shishi hostkeys default file

[Function] const char * shishi_hostkeys_default_file (Shishi * handle)

handle: Shishi library handle create by shishi_init().

Description: Get file name of default host key file.

Return value: Returns the default host key filename used in the library. (Not a copy of it, so don’t modify or deallocate it.)

shishi hostkeys default file set

[Function] void shishi_hostkeys_default_file_set (Shishi * handle,
    const char * hostkeysfile)

handle: Shishi library handle create by shishi_init(). hostkeysfile: string with new default hostkeys file name, or NULL to reset to default.

Description: Set the default host key filename used in the library. The string is copied into the library, so you can dispose of the variable immediately after calling this function.

shishi hostkeys for server

[Function] Shishi_key * shishi_hostkeys_for_server (Shishi * handle,

const char * server) handle: Shishi library handle create by shishi_init(). server: server name to get key for

Description: Get host key for server.

Return value: Returns the key for specific server, read from the default host keys file (see shishi_hostkeys_default_file()), or NULL if no key could be found or an error encountered.

shishi hostkeys for serverrealm

[Function] Shishi_key * shishi_hostkeys_for_serverrealm

(Shishi * handle, const char * server, const char * realm)

handle: Shishi library handle create by shishi_init(). server: server name to get key for realm: realm of server to get key for.

Description: Get host key for server in realm.

Return value: Returns the key for specific server and realm, read from the default host keys file (see shishi_hostkeys_default_file()), or NULL if no key could be found or an error encountered.

shishi hostkeys for localservicerealm

[Function] Shishi_key * shishi_hostkeys_for_localservicerealm

(Shishi * handle, const char * service, const char * realm)

handle: Shishi library handle create by shishi_init(). service: service to get key for. realm: realm of server to get key for, or NULL for default realm.

Description: Get host key for service on current host in realm.

Return value: Returns the key for the server "SERVICE/HOSTNAME@REALM" (where HOSTNAME is the current system’s hostname), read from the default host keys file (see shishi_hostkeys_default_file()), or NULL if no key could be found or an error encountered.

shishi hostkeys for localservice

[Function] Shishi_key * shishi_hostkeys_for_localservice

(Shishi * handle, const char * service) handle: Shishi library handle create by shishi_init(). service: service to get key for.

Description: Get host key for service on current host in default realm.

Return value: Returns the key for the server "SERVICE/HOSTNAME" (where HOSTNAME is the current system’s hostname), read from the default host keys file (see shishi_hostkeys_default_file()), or NULL if no key could be found or an error encountered.

After creating the key structure, it can be used to encrypt and decrypt data, calculate checksum on data etc. All available functions are described now.

shishi cipher supported p

[Function] int shishi_cipher_supported_p (int32 t type)

type: encryption type, see Shishi etype.

Description: Find out if cipher is supported.

Return value: Return 0 iff cipher is unsupported.

shishi cipher name

[Function] const char * shishi_cipher_name (int32 t type)

type: encryption type, see Shishi etype.

Description: Read humanly readable string for cipher.

Return value: Return name of encryption type, e.g. "des3-cbc-sha1-kd", as defined in the standards.

shishi cipher blocksize

[Function] int shishi_cipher_blocksize (int32 t type)

type: encryption type, see Shishi etype.

Description: Get block size for cipher.

Return value: Return block size for encryption type, as defined in the standards.

shishi cipher confoundersize

[Function] int shishi_cipher_confoundersize (int32 t type)

type: encryption type, see Shishi etype.

Description: Get length of confounder for cipher.

Return value: Returns the size of the confounder (random data) for encryption type, as defined in the standards, or (size t)-1 on error (e.g., unsupported encryption type).

shishi cipher keylen

[Function] size_t shishi_cipher_keylen (int32 t type)

type: encryption type, see Shishi etype. Description: Get key length for cipher. Return value: Return length of key used for the encryption type, as defined in the standards.

shishi cipher randomlen

[Function] size_t shishi_cipher_randomlen (int32 t type)

type: encryption type, see Shishi etype. Description: Get length of random data for cipher. Return value: Return length of random used for the encryption type, as defined in the standards, or (size t)-1 on error (e.g., unsupported encryption type).

shishi cipher defaultcksumtype

[Function] int shishi_cipher_defaultcksumtype (int32 t type)

type: encryption type, see Shishi etype. Description: Get the default checksum associated with cipher. Return value: Return associated checksum mechanism for the encryption type, as defined in the standards.

shishi cipher parse

[Function] int shishi_cipher_parse (const char * cipher)

cipher: name of encryption type, e.g. "des3-cbc-sha1-kd". Description: Get cipher number by parsing string. Return value: Return encryption type corresponding to a string.

shishi checksum supported p

[Function] int shishi_checksum_supported_p (int32 t type)

type: checksum type, see Shishi cksumtype. Description: Find out whether checksum is supported. Return value: Return 0 iff checksum is unsupported.

shishi checksum name

[Function] const char * shishi_checksum_name (int32 t type)

type: checksum type, see Shishi cksumtype. Description: Get name of checksum. Return value: Return name of checksum type, e.g. "hmac-sha1-96-aes256", as defined in the standards.

shishi checksum cksumlen

[Function] size_t shishi_checksum_cksumlen (int32 t type)

type: checksum type, see Shishi cksumtype. Description: Get length of checksum output. Return value: Return length of checksum used for the checksum type, as defined in the standards.

shishi checksum parse

[Function] int shishi_checksum_parse (const char * checksum)

checksum: name of checksum type, e.g. "hmac-sha1-96-aes256". Description: Get checksum number by parsing a string. Return value: Return checksum type, see Shishi cksumtype, corresponding to a string.

shishi string to key

[Function] int shishi_string_to_key (Shishi * handle, int32 t keytype,

const char * password, size t passwordlen, const char * salt, size t saltlen, const char * parameter, Shishi key * outkey)

handle: shishi handle as allocated by shishi_init(). keytype: cryptographic encryption type, see Shishi etype. password: input array with password. passwordlen: length of input array with password. salt: input array with salt. saltlen: length of input array with salt. parameter: input array with opaque encryption type specific information. outkey: allocated key handle that will contain new key. Description: Derive key from a string (password) and salt (commonly concatenation of realm and principal) for specified key type, and set the type and value in the given key to the computed values. The parameter value is specific for each keytype, and can be set if the parameter information is not available. Return value: Returns SHISHI_OK iff successful.

shishi random to key

[Function] int shishi_random_to_key (Shishi * handle, int32 t keytype,

const char * rnd, size t rndlen, Shishi key * outkey)

handle: shishi handle as allocated by shishi_init(). keytype: cryptographic encryption type, see Shishi etype. rnd: input array with random data. rndlen: length of input array with random data. outkey: allocated key handle that will contain new key. Description: Derive key from random data for specified key type, and set the type and value in the given key to the computed values. Return value: Returns SHISHI_OK iff successful.

shishi checksum

[Function] int shishi_checksum (Shishi * handle, Shishi key * key, int keyusage, int cksumtype, const char * in, size t inlen, char ** out,

size t * outlen) handle: shishi handle as allocated by shishi_init(). key: key to compute checksum with. keyusage: integer specifying what this key is used for. cksumtype: the checksum algorithm to use. in: input array with data to integrity protect. inlen: size of input array with data to integrity protect. out: output array with newly allocated integrity protected data. outlen: output variable with length of output array with checksum. Description: Integrity protect data using key, possibly altered by supplied key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. Return value: Returns SHISHI_OK iff successful.

shishi verify

[Function] int shishi_verify (Shishi * handle, Shishi key * key, int keyusage, int cksumtype, const char * in, size t inlen, const char * cksum, size t cksumlen)

handle: shishi handle as allocated by shishi_init(). key: key to verify checksum with. keyusage: integer specifying what this key is used for. cksumtype: the checksum algorithm to use. in: input array with data that was integrity protected. inlen: size of input array with data that was integrity protected. cksum: input array with alleged checksum of data. cksumlen: size of input array with alleged checksum of data. Description: Verify checksum of data using key, possibly altered by supplied key usage. If key usage is 0, no key derivation is used. Return value: Returns SHISHI_OK iff successful.

shishi encrypt ivupdate etype

[Function] int shishi_encrypt_ivupdate_etype (Shishi * handle,

Shishi key * key, int keyusage, int32 t etype, const char * iv, size t ivlen, char ** ivout, size t * ivoutlen, const char * in, size t inlen, char ** out, size t * outlen)

handle: shishi handle as allocated by shishi_init(). key: key to encrypt with. keyusage: integer specifying what this key is encrypting. etype: integer specifying what cipher to use. iv: input array with initialization vector ivlen: size of input array with initialization vector. ivout: output array with newly allocated updated initialization vector. ivoutlen: size of output array with updated initialization vector. in: input array with data to encrypt. inlen: size of input array with data to encrypt. out: output array with newly allocated encrypted data. outlen: output variable with size of newly allocated output array.

Description: Encrypts data as per encryption method using specified initialization vector and key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. If IVOUT or IVOUTLEN is NULL, the updated IV is not saved anywhere.

Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information.

Return value: Returns SHISHI_OK iff successful.

shishi encrypt iv etype

[Function] int shishi_encrypt_iv_etype (Shishi * handle, Shishi key * key,

int keyusage, int32 t etype, const char * iv, size t ivlen, const char * in, size t inlen, char ** out, size t * outlen)

handle: shishi handle as allocated by shishi_init(). key: key to encrypt with. keyusage: integer specifying what this key is encrypting. etype: integer specifying what cipher to use. iv: input array with initialization vector ivlen: size of input array with initialization vector. in: input array with data to encrypt. inlen: size of input array with data to encrypt. out: output array with newly allocated encrypted data. outlen: output variable with size of newly allocated output array.

Description: Encrypts data as per encryption method using specified initialization vector and key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. The next IV is lost, see shishi encrypt ivupdate etype if you need it.

Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information.

Return value: Returns SHISHI_OK iff successful.

shishi encrypt etype

[Function] int shishi_encrypt_etype (Shishi * handle, Shishi key * key,

int keyusage, int32 t etype, const char * in, size t inlen, char ** out,

size t * outlen) handle: shishi handle as allocated by shishi_init(). key: key to encrypt with. keyusage: integer specifying what this key is encrypting. etype: integer specifying what cipher to use. in: input array with data to encrypt. inlen: size of input array with data to encrypt. out: output array with newly allocated encrypted data. outlen: output variable with size of newly allocated output array. Description: Encrypts data as per encryption method using specified initialization vector and key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. The default IV is used, see shishi encrypt iv etype if you need to alter it. The next IV is lost, see shishi encrypt ivupdate etype if you need it. Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information. Return value: Returns SHISHI_OK iff successful.

shishi encrypt ivupdate

[Function] int shishi_encrypt_ivupdate (Shishi * handle, Shishi key * key,

int keyusage, const char * iv, size t ivlen, char ** ivout, size t * ivoutlen, const char * in, size t inlen, char ** out,

size t * outlen) handle: shishi handle as allocated by shishi_init(). key: key to encrypt with. keyusage: integer specifying what this key is encrypting. iv: input array with initialization vector ivlen: size of input array with initialization vector. ivout: output array with newly allocated updated initialization vector. ivoutlen: size of output array with updated initialization vector. in: input array with data to encrypt. inlen: size of input array with data to encrypt. out: output array with newly allocated encrypted data. outlen: output variable with size of newly allocated output array. Description: Encrypts data using specified initialization vector and key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. If IVOUT or IVOUTLEN is NULL, the updated IV is not saved anywhere.

Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information. Return value: Returns SHISHI_OK iff successful.

shishi encrypt iv

[Function] int shishi_encrypt_iv (Shishi * handle, Shishi key * key, int keyusage, const char * iv, size t ivlen, const char * in, size t inlen, char ** out, size t * outlen)

handle: shishi handle as allocated by shishi_init(). key: key to encrypt with. keyusage: integer specifying what this key is encrypting. iv: input array with initialization vector ivlen: size of input array with initialization vector. in: input array with data to encrypt. inlen: size of input array with data to encrypt. out: output array with newly allocated encrypted data. outlen: output variable with size of newly allocated output array. Description: Encrypts data using specified initialization vector and key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. The next IV is lost, see shishi encrypt ivupdate if you need it. Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information. Return value: Returns SHISHI_OK iff successful.

shishi encrypt

[Function] int shishi_encrypt (Shishi * handle, Shishi key * key,

int keyusage, char * in, size t inlen, char ** out, size t * outlen)

handle: shishi handle as allocated by shishi_init(). key: key to encrypt with. keyusage: integer specifying what this key is encrypting. in: input array with data to encrypt. inlen: size of input array with data to encrypt. out: output array with newly allocated encrypted data. outlen: output variable with size of newly allocated output array. Description: Encrypts data using specified key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. The default IV is used, see shishi encrypt iv if you need to alter it. The next IV is lost, see shishi encrypt ivupdate if you need it. Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information. Return value: Returns SHISHI_OK iff successful.

shishi decrypt ivupdate etype

[Function] int shishi_decrypt_ivupdate_etype (Shishi * handle,

Shishi key * key, int keyusage, int32 t etype, const char * iv, size t ivlen, char ** ivout, size t * ivoutlen, const char * in, size t inlen, char ** out, size t * outlen)

handle: shishi handle as allocated by shishi_init(). key: key to decrypt with. keyusage: integer specifying what this key is decrypting. etype: integer specifying what cipher to use. iv: input array with initialization vector ivlen: size of input array with initialization vector. ivout: output array with newly allocated updated initialization vector. ivoutlen: size of output array with updated initialization vector. in: input array with data to decrypt. inlen: size of input array with data to decrypt. out: output array with newly allocated decrypted data. outlen: output variable with size of newly allocated output array. Description: Decrypts data as per encryption method using specified initialization vector and key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. If IVOUT or IVOUTLEN is NULL, the updated IV is not saved anywhere. Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information. Return value: Returns SHISHI_OK iff successful.

shishi decrypt iv etype

[Function] int shishi_decrypt_iv_etype (Shishi * handle, Shishi key * key,

int keyusage, int32 t etype, const char * iv, size t ivlen, const char * in, size t inlen, char ** out, size t * outlen)

handle: shishi handle as allocated by shishi_init(). key: key to decrypt with. keyusage: integer specifying what this key is decrypting. etype: integer specifying what cipher to use. iv: input array with initialization vector ivlen: size of input array with initialization vector. in: input array with data to decrypt. inlen: size of input array with data to decrypt. out: output array with newly allocated decrypted data. outlen: output variable with size of newly allocated output array.

Description: Decrypts data as per encryption method using specified initialization vector and key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. The next IV is lost, see shishi decrypt ivupdate etype if you need it.

Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information.

Return value: Returns SHISHI_OK iff successful.

shishi decrypt etype

[Function] int shishi_decrypt_etype (Shishi * handle, Shishi key * key,

int keyusage, int32 t etype, const char * in, size t inlen, char ** out,

size t * outlen) handle: shishi handle as allocated by shishi_init(). key: key to decrypt with. keyusage: integer specifying what this key is decrypting. etype: integer specifying what cipher to use. in: input array with data to decrypt. inlen: size of input array with data to decrypt. out: output array with newly allocated decrypted data. outlen: output variable with size of newly allocated output array.

Description: Decrypts data as per encryption method using specified key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. The default IV is used, see shishi decrypt iv etype if you need to alter it. The next IV is lost, see shishi decrypt ivupdate etype if you need it.

Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information.

Return value: Returns SHISHI_OK iff successful.

shishi decrypt ivupdate

[Function] int shishi_decrypt_ivupdate (Shishi * handle, Shishi key * key,

int keyusage, const char * iv, size t ivlen, char ** ivout, size t * ivoutlen, const char * in, size t inlen, char ** out,

size t * outlen) handle: shishi handle as allocated by shishi_init(). key: key to decrypt with. keyusage: integer specifying what this key is decrypting. iv: input array with initialization vector ivlen: size of input array with initialization vector. ivout: output array with newly allocated updated initialization vector. ivoutlen: size of output array with updated initialization vector. in: input array with data to decrypt. inlen: size of input array with data to decrypt. out: output array with newly allocated decrypted data. outlen: output variable with size of newly allocated output array. Description: Decrypts data using specified initialization vector and key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. If IVOUT or IVOUTLEN is NULL, the updated IV is not saved anywhere. Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information. Return value: Returns SHISHI_OK iff successful.

shishi decrypt iv

[Function] int shishi_decrypt_iv (Shishi * handle, Shishi key * key, int keyusage, const char * iv, size t ivlen, const char * in, size t inlen, char ** out, size t * outlen)

handle: shishi handle as allocated by shishi_init(). key: key to decrypt with. keyusage: integer specifying what this key is decrypting. iv: input array with initialization vector ivlen: size of input array with initialization vector. in: input array with data to decrypt. inlen: size of input array with data to decrypt. out: output array with newly allocated decrypted data. outlen: output variable with size of newly allocated output array. Description: Decrypts data using specified initialization vector and key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. The next IV is lost, see shishi decrypt ivupdate etype if you need it.

Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information. Return value: Returns SHISHI_OK iff successful.

shishi decrypt

[Function] int shishi_decrypt (Shishi * handle, Shishi key * key, int keyusage, const char * in, size t inlen, char ** out,

size t * outlen) handle: shishi handle as allocated by shishi_init(). key: key to decrypt with. keyusage: integer specifying what this key is decrypting. in: input array with data to decrypt. inlen: size of input array with data to decrypt. out: output array with newly allocated decrypted data. outlen: output variable with size of newly allocated output array. Description: Decrypts data specified key. The key actually used is derived using the key usage. If key usage is 0, no key derivation is used. The OUT buffer must be deallocated by the caller. The default IV is used, see shishi decrypt iv if you need to alter it. The next IV is lost, see shishi decrypt ivupdate if you need it. Note that DECRYPT(ENCRYPT(data)) does not necessarily yield data exactly. Some encryption types add pad to make the data fit into the block size of the encryption algorithm. Furthermore, the pad is not guaranteed to look in any special way, although existing implementations often pad with the zero byte. This means that you may have to "frame" data, so it is possible to infer the original length after decryption. Compare ASN.1 DER which contains such information. Return value: Returns SHISHI_OK iff successful.

shishi n fold

[Function] int shishi_n_fold (Shishi * handle, const char * in, size t inlen,

char * out, size t outlen)

handle: shishi handle as allocated by shishi_init(). in: input array with data to decrypt. inlen: size of input array with data to decrypt ("M"). out: output array with decrypted data. outlen: size of output array ("N"). Description: Fold data into a fixed length output array, with the intent to give each input bit approximately equal weight in determining the value of each output bit. The algorithm is from "A Better Key Schedule For DES-like Ciphers" by Uri Blumenthal and Steven M. Bellovin, http://www.research.att.com/~smb/papers/ides.pdf, although the sample vectors provided by the paper are incorrect. Return value: Returns SHISHI_OK iff successful.

shishi dr

[Function] int shishi_dr (Shishi * handle, Shishi key * key,

   const char * prfconstant, size t prfconstantlen,
   char * derivedrandom, size t derivedrandomlen)

handle: shishi handle as allocated by shishi_init(). key: input array with cryptographic key to use. prfconstant: input array with the constant string. prfconstantlen: size of input array with the constant string. derivedrandom: output array with derived random data. derivedrandomlen: size of output array with derived random data. Description: Derive "random" data from a key and a constant thusly: DR(KEY, PRFCONSTANT) = TRUNCATE(DERIVEDRANDOMLEN, SHISHI ENCRYPT(KEY, PRFCONSTANT)). Return value: Returns SHISHI_OK iff successful.

shishi dk

[Function] int shishi_dk (Shishi * handle, Shishi key * key,

   const char * prfconstant, size t prfconstantlen,

Shishi key * derivedkey) handle: shishi handle as allocated by shishi_init(). key: input cryptographic key to use. prfconstant: input array with the constant string. prfconstantlen: size of input array with the constant string. derivedkey: pointer to derived key (allocated by caller). Derive a key from a key and a constant thusly: DK(KEY, PRFCONSTANT) = SHISHI RANDOM-TO-KEY(SHISHI DR(KEY, PRFCONSTANT)). Return value: Returns SHISHI_OK iff successful.

An easier way to use encryption and decryption if your application repeatedly calls, e.g., shishi_encrypt_ivupdate, is to use the following functions. They store the key, initialization vector, etc, in a context, and the encryption and decryption operations update the IV within the context automatically.

shishi crypto

[Function] Shishi_crypto * shishi_crypto (Shishi * handle,

Shishi key * key, int keyusage, int32 t etype, const char * iv,

size t ivlen) handle: shishi handle as allocated by shishi_init(). key: key to encrypt with. keyusage: integer specifying what this key will encrypt/decrypt. etype: integer specifying what cipher to use. iv: input array with initialization vector ivlen: size of input array with initialization vector. Description: Initialize a crypto context. This store a key, keyusage, encryption type and initialization vector in a "context", and the caller can then use this context to perform encryption via shishi_crypto_encrypt() and decryption via shishi_ crypto_encrypt() without supplying all those details again. The functions also takes care of propagating the IV between calls. When the application no longer need to use the context, it should deallocate resources associated with it by calling shishi_crypto_close(). Return value: Return a newly allocated crypto context.

shishi crypto encrypt

[Function] int shishi_crypto_encrypt (Shishi crypto * ctx, const char * in,

size t inlen, char ** out, size t * outlen)

ctx: crypto context as returned by shishi_crypto(). in: input array with data to encrypt. inlen: size of input array with data to encrypt. out: output array with newly allocated encrypted data. outlen: output variable with size of newly allocated output array. Description: Encrypt data, using information (e.g., key and initialization vector) from context. The IV is updated inside the context after this call. When the application no longer need to use the context, it should deallocate resources associated with it by calling shishi_crypto_close(). Return value: Returns SHISHI_OK iff successful.

shishi crypto decrypt

[Function] int shishi_crypto_decrypt (Shishi crypto * ctx, const char * in,

size t inlen, char ** out, size t * outlen)

ctx: crypto context as returned by shishi_crypto(). in: input array with data to decrypt. inlen: size of input array with data to decrypt. out: output array with newly allocated decrypted data. outlen: output variable with size of newly allocated output array. Description: Decrypt data, using information (e.g., key and initialization vector) from context. The IV is updated inside the context after this call. When the application no longer need to use the context, it should deallocate resources associated with it by calling shishi_crypto_close(). Return value: Returns SHISHI_OK iff successful.

shishi crypto close

[Function] void shishi_crypto_close (Shishi crypto * ctx)

ctx: crypto context as returned by shishi_crypto(). Description: Deallocate resources associated with the crypto context.

Also included in Shishi is an interface to the really low-level cryptographic primitives. They map directly on the underlying cryptographic library used (i.e., Gnulib or Libgcrypt) and is used internally by Shishi.

shishi randomize

[Function] int shishi_randomize (Shishi * handle, int strong, void * data,
    size t datalen)

handle: shishi handle as allocated by shishi_init(). strong: 0 iff operation should not block, non-0 for very strong randomness. data: output array to be filled with random data. datalen: size of output array. Description: Store cryptographically random data of given size in the provided buffer. Return value: Returns SHISHI_OK iff successful.

shishi crc

[Function] int shishi_crc (Shishi * handle, const char * in, size t inlen,

char * out[4]) handle: shishi handle as allocated by shishi_init(). in: input character array of data to checksum. inlen: length of input character array of data to checksum. out: newly allocated character array with checksum of data. Description: Compute checksum of data using CRC32 modified according to RFC 1510. The out buffer must be deallocated by the caller. The modifications compared to standard CRC32 is that no initial and final XOR is performed, and that the output is returned in LSB-first order. Return value: Returns SHISHI OK iff successful.

shishi md4

[Function] int shishi_md4 (Shishi * handle, const char * in, size t inlen,

char * out[16]) handle: shishi handle as allocated by shishi_init(). in: input character array of data to hash. inlen: length of input character array of data to hash. out: newly allocated character array with hash of data. Description: Compute hash of data using MD4. The out buffer must be deallocated by the caller. Return value: Returns SHISHI OK iff successful.

shishi md5

[Function] int shishi_md5 (Shishi * handle, const char * in, size t inlen,

char * out[16]) handle: shishi handle as allocated by shishi_init(). in: input character array of data to hash. inlen: length of input character array of data to hash. out: newly allocated character array with hash of data. Description: Compute hash of data using MD5. The out buffer must be deallocated by the caller. Return value: Returns SHISHI OK iff successful.

shishi hmac md5

[Function] int shishi_hmac_md5 (Shishi * handle, const char * key,

size t keylen, const char * in, size t inlen, char * outhash[16])

handle: shishi handle as allocated by shishi_init(). key: input character array with key to use. keylen: length of input character array with key to use. in: input character array of data to hash. inlen: length of input character array of data to hash. outhash: newly allocated character array with keyed hash of data.

Description: Compute keyed checksum of data using HMAC-MD5. The outhash buffer must be deallocated by the caller.

Return value: Returns SHISHI OK iff successful.

shishi hmac sha1

[Function] int shishi_hmac_sha1 (Shishi * handle, const char * key,

size t keylen, const char * in, size t inlen, char * outhash[20])

handle: shishi handle as allocated by shishi_init(). key: input character array with key to use. keylen: length of input character array with key to use. in: input character array of data to hash. inlen: length of input character array of data to hash. outhash: newly allocated character array with keyed hash of data.

Description: Compute keyed checksum of data using HMAC-SHA1. The outhash buffer must be deallocated by the caller.

Return value: Returns SHISHI OK iff successful.

shishi des cbc mac

[Function] int shishi_des_cbc_mac (Shishi * handle, const char key[8],

const char iv[8], const char * in, size t inlen, char * out[8])

handle: shishi handle as allocated by shishi_init(). key: input character array with key to use. iv: input character array with initialization vector to use, can be NULL. in: input character array of data to hash. inlen: length of input character array of data to hash. out: newly allocated character array with keyed hash of data.

Description: Computed keyed checksum of data using DES-CBC-MAC. The out buffer must be deallocated by the caller.

shishi arcfour

[Function] int shishi_arcfour (Shishi * handle, int decryptp,

const char * key, size t keylen, const char iv[258], char * ivout[258], const char * in, size t inlen, char ** out)

handle: shishi handle as allocated by shishi_init(). decryptp: 0 to indicate encryption, non-0 to indicate decryption. key: input character array with key to use. keylen: length of input key array. iv: input character array with initialization vector to use, or NULL. ivout: output character array with updated initialization vector, or NULL. in: input character array of data to encrypt/decrypt. inlen: length of input character array of data to encrypt/decrypt. out: newly allocated character array with encrypted/decrypted data. Description: Encrypt or decrypt data (depending on decryptp) using ARCFOUR. The out buffer must be deallocated by the caller. The "initialization vector" used here is the concatenation of the sbox and i and j, and is thus always of size 256 + 1 + 1. This is a slight abuse of terminology, and assumes you know what you are doing. Don’t use it if you can avoid to. Return value: Returns SHISHI OK iff successful.

shishi des

[Function] int shishi_des (Shishi * handle, int decryptp, const char key[8], const char iv[8], char * ivout[8], const char * in, size t inlen,

char ** out) handle: shishi handle as allocated by shishi_init(). decryptp: 0 to indicate encryption, non-0 to indicate decryption. key: input character array with key to use. iv: input character array with initialization vector to use, or NULL. ivout: output character array with updated initialization vector, or NULL. in: input character array of data to encrypt/decrypt. inlen: length of input character array of data to encrypt/decrypt. out: newly allocated character array with encrypted/decrypted data. Description: Encrypt or decrypt data (depending on decryptp) using DES in CBC mode. The out buffer must be deallocated by the caller. Return value: Returns SHISHI OK iff successful.

shishi 3des

[Function] int shishi_3des (Shishi * handle, int decryptp, const char key[8], const char iv[8], char * ivout[8], const char * in, size t inlen,

char ** out) handle: shishi handle as allocated by shishi_init(). decryptp: 0 to indicate encryption, non-0 to indicate decryption. key: input character array with key to use. iv: input character array with initialization vector to use, or NULL. ivout: output character array with updated initialization vector, or NULL. in: input character array of data to encrypt/decrypt. inlen: length of input character array of data to encrypt/decrypt. out: newly allocated character array with encrypted/decrypted data. Description: Encrypt or decrypt data (depending on decryptp) using 3DES in CBC mode. The out buffer must be deallocated by the caller. Return value: Returns SHISHI OK iff successful.

shishi aes cts

[Function] int shishi_aes_cts (Shishi * handle, int decryptp,

const char * key, size t keylen, const char iv[16], char * ivout[16], const char * in, size t inlen, char ** out)

handle: shishi handle as allocated by shishi_init(). decryptp: 0 to indicate encryption, non-0 to indicate decryption. key: input character array with key to use. keylen: length of input character array with key to use. iv: input character array with initialization vector to use, or NULL. ivout: output character array with updated initialization vector, or NULL. in: input character array of data to encrypt/decrypt. inlen: length of input character array of data to encrypt/decrypt. out: newly allocated character array with encrypted/decrypted data. Description: Encrypt or decrypt data (depending on decryptp) using AES in CBC-CTS mode. The length of the key, keylen, decide if AES 128 or AES 256 should be used. The out buffer must be deallocated by the caller. Return value: Returns SHISHI OK iff successful.

shishi pbkdf2 sha1

[Function] int shishi_pbkdf2_sha1 (Shishi * handle, const char * P, size t Plen, const char * S, size t Slen, unsigned int c, unsigned int dkLen, char * DK)

handle: shishi handle as allocated by shishi_init(). P: input password, an octet string Plen: length of password, an octet string S: input salt, an octet string Slen: length of salt, an octet string c: iteration count, a positive integer dkLen: intended length in octets of the derived key, a positive integer, at most (2^32 - 1) * hLen. The DK array must have room for this many characters. DK: output derived key, a dkLen-octet string Description: Derive key using the PBKDF2 defined in PKCS5. PBKDF2 applies a pseudorandom function to derive keys. The length of the derived key is essentially unbounded. (However, the maximum effective search space for the derived key may be limited by the structure of the underlying pseudorandom function, which is this function is always SHA1.) Return value: Returns SHISHI OK iff successful.

5.14 X.509 Functions

The functions described in this section are used by the STARTTLS functionality, see Section 3.6 [Kerberos via TLS], page 26.

shishi x509ca default file guess

[Function] char * shishi_x509ca_default_file_guess (Shishi * handle)

handle: Shishi library handle create by shishi_init().

Description: Guesses the default X.509 CA certificate filename; it is HOME/.shishi/client.ca.

Return value: Returns default X.509 client certificate filename as a string that has to be deallocated with free() by the caller.

shishi x509ca default file set

[Function] void shishi_x509ca_default_file_set (Shishi * handle,

const char * x509cafile) handle: Shishi library handle create by shishi_init(). x509cafile: string with new default x509 client certificate file name, or NULL to reset to default.

Description: Set the default X.509 CA certificate filename used in the library. The certificate is used during TLS connections with the KDC to authenticate the KDC. The string is copied into the library, so you can dispose of the variable immediately after calling this function.

shishi x509ca default file

[Function] const char * shishi_x509ca_default_file (Shishi * handle)

handle: Shishi library handle create by shishi_init().

Description: Get filename for default X.509 CA certificate.

Return value: Returns the default X.509 CA certificate filename used in the library. The certificate is used during TLS connections with the KDC to authenticate the KDC. The string is not a copy, so don’t modify or deallocate it.

shishi x509cert default file guess

[Function] char * shishi_x509cert_default_file_guess (Shishi * handle)

handle: Shishi library handle create by shishi_init().

Description: Guesses the default X.509 client certificate filename; it is HOME/.shishi/client.certs.

Return value: Returns default X.509 client certificate filename as a string that has to be deallocated with free() by the caller.

shishi x509cert default file set

[Function] void shishi_x509cert_default_file_set (Shishi * handle,
    const char * x509certfile)

handle: Shishi library handle create by shishi_init(). x509certfile: string with new default x509 client certificate file name, or NULL to reset to default. Description: Set the default X.509 client certificate filename used in the library. The certificate is used during TLS connections with the KDC to authenticate the client. The string is copied into the library, so you can dispose of the variable immediately after calling this function.

shishi x509cert default file

[Function] const char * shishi_x509cert_default_file (Shishi * handle)

handle: Shishi library handle create by shishi_init(). Description: Get filename for default X.509 certificate. Return value: Returns the default X.509 client certificate filename used in the library. The certificate is used during TLS connections with the KDC to authenticate the client. The string is not a copy, so don’t modify or deallocate it.

shishi x509key default file guess

[Function] char * shishi_x509key_default_file_guess (Shishi * handle)

handle: Shishi library handle create by shishi_init(). Description: Guesses the default X.509 client key filename; it is HOME/.shishi/client.key. Return value: Returns default X.509 client key filename as a string that has to be deallocated with free() by the caller.

shishi x509key default file set

[Function] void shishi_x509key_default_file_set (Shishi * handle,
    const char * x509keyfile)

handle: Shishi library handle create by shishi_init(). x509keyfile: string with new default x509 client key file name, or NULL to reset to default. Description: Set the default X.509 client key filename used in the library. The key is used during TLS connections with the KDC to authenticate the client. The string is copied into the library, so you can dispose of the variable immediately after calling this function.

shishi x509key default file

[Function] const char * shishi_x509key_default_file (Shishi * handle)

handle: Shishi library handle create by shishi_init(). Description: Get filename for default X.509 key.

Return value: Returns the default X.509 client key filename used in the library. The key is used during TLS connections with the KDC to authenticate the client. The string is not a copy, so don’t modify or deallocate it.

5.15 Utility Functions

shishi realm default guess

[Function] char * shishi_realm_default_guess ()

Description: Guesses a realm based on getdomainname(), which really responds with a NIS/YP domain, but if set properly, it might be a good first guess. If this NIS query fails, call gethostname(), and on its failure, fall back to returning the artificial string "could-not-guess-default-realm".

Note that the hostname is not trimmed off of the string returned by gethostname(), thus pretending the local host name is a valid realm name. The resulting corner case could merit a check that the suggested realm is distinct from the fully qualifies host, and if not, simply strip the host name from the returned string before it is used in an application. One reason for sticking with the present behaviour, is that some systems respond with a non-qualified host name as reply from gethostname().

Return value: Returns a guessed realm for the running host, containing a string that has to be deallocated with free() by the caller.

shishi realm default

[Function] const char * shishi_realm_default (Shishi * handle)

handle: Shishi library handle created by shishi_init().

Description: Determines name of default realm, i.e., the name of whatever realm the library will use whenever an explicit realm is not stated during a library call.

Return value: Returns the default realm in use by the library. Not a copy, so do not modify or deallocate the returned string.

shishi realm default set

[Function] void shishi_realm_default_set (Shishi * handle,

const char * realm) handle: Shishi library handle created by shishi_init(). realm: String stating a new default realm name, or NULL.

Description: Sets the default realm used by the library; or, with realm set to NULL, resets the library realm setting to that name selected by configuration for default value.

The string is copied into the library, so you can dispose of the content in realm immediately after calling this function.

shishi realm for server file

[Function] char * shishi_realm_for_server_file (Shishi * handle,
    char * server)

handle: Shishi library handle created by shishi_init(). server: Hostname to determine realm for. Description: Finds the realm applicable to a host server, using the standard configuration file. Return value: Returns realm for host, or NULL if not known.

shishi realm for server dns

[Function] char * shishi_realm_for_server_dns (Shishi * handle,
    char * server)

handle: Shishi library handle created by shishi_init(). server: Hostname to find realm for. Description: Finds the realm for a host server using DNS lookup, as is prescribed in "draft-ietf-krb-wg-krb-dns-locate-03.txt". Since DNS lookup can be spoofed, relying on the realm information may result in a redirection attack. In a single-realm scenario, this only achieves a denial of service, but with trust across multiple realms the attack may redirect you to a compromised realm. For this reason, Shishi prints a warning, suggesting that the user should instead add a proper ’server-realm’ configuration token. To illustrate the DNS information used, here is an extract from a zone file for the domain ASDF.COM: kerberos.asdf.com. IN TXT "ASDF.COM" kerberos.mrkserver.asdf.com. IN TXT "MARKETING.ASDF.COM" kerberos.salesserver.asdf.com. IN TXT "SALES.ASDF.COM" Let us suppose that in this case, a client wishes to use a service on the host "foo.asdf.com". It would first query for kerberos.foo.asdf.com. IN TXT Finding no match, it would then query for kerberos.asdf.com. IN TXT With the resource records stated above, the latter query returns a positive answer. Return value: Returns realm for the indicated host, or NULL if no relevant TXT record could be found.

shishi realm for server

[Function] char * shishi_realm_for_server (Shishi * handle, char * server)

handle: Shishi library handle created by shishi_init(). server: Hostname to find realm for. Description: Finds a realm for the host server, using various methods. Currently this includes static configuration files, using the library call shishi_realm_

for_server_file(), and DNS lookup using shishi_realm_for_server_dns().

They are attempted in the stated order. See the documentation of either function for more information.

Return value: Returns realm for the indicated host, or NULL if nothing is known about server.

shishi principal default guess

[Function] char * shishi_principal_default_guess ()

Description: Guesses the principal name for the user, looking at environment variables SHISHI USER, USER and LOGNAME, or if that fails, returns the string "user".

Return value: Returns guessed default principal for user as a string that has to be deallocated by the caller with free().

shishi principal default

[Function] const char * shishi_principal_default (Shishi * handle)

handle: Shishi library handle created by shishi_init().

Description: The default principal name is the name in the environment variable USER, or LOGNAME for some systems, but it can be overridden by specifying the environment variable SHISHI USER.

Return value: Returns the default principal name used by the library. (Not a copy of it, so don’t modify or deallocate it.)

shishi principal default set

[Function] void shishi_principal_default_set (Shishi * handle,

const char * principal) handle: Shishi library handle created by shishi_init(). principal: string with new default principal name, or NULL to reset to default.

Description: Set the default principal used by the library. The string is copied into the library, so you can dispose of the variable immediately after calling this function.

shishi parse name

[Function] int shishi_parse_name (Shishi * handle, const char * name,
   char ** principal, char ** realm)

handle: Shishi library handle created by shishi_init(). name: input principal name string, e.g. imap/mail.gnu.org\GNU.ORG. principal: newly allocated output string with principal name. realm: newly allocated output string with realm name.

Description: Split principal name (e.g., "simon\JOSEFSSON.ORG") into two newly allocated strings, the principal ("simon"), and the realm ("JOSEFSSON.ORG"). If there is no realm part in name, realm is set to NULL.

Return value: Returns SHISHI INVALID PRINCIPAL NAME if name is NULL or ends with the escape character "\", and SHISHI OK if successful.

shishi principal name

[Function] int shishi_principal_name (Shishi * handle,

Shishi asn1 namenode, const char * namefield, char ** out,

size t * outlen) handle: Shishi library handle created by shishi_init(). namenode: ASN.1 structure with principal in namefield. namefield: name of field in namenode containing principal name. out: pointer to newly allocated, null terminated, string containing principal name. May be NULL (to only populate outlen). outlen: pointer to length of out on output, excluding terminating null. May be NULL (to only populate out). Description: Represent principal name in ASN.1 structure as null-terminated string. The string is allocated by this function, and it is the responsibility of the caller to deallocate it. Note that the output length outlen does not include the terminating null. Return value: Returns SHISHI OK if successful.

shishi principal name realm

[Function] int shishi_principal_name_realm (Shishi * handle,

Shishi asn1 namenode, const char * namefield, Shishi asn1 realmnode, const char * realmfield, char ** out, size t * outlen)

handle: Shishi library handle created by shishi_init(). namenode: ASN.1 structure with principal name in namefield. namefield: name of field in namenode containing principal name. realmnode: ASN.1 structure with principal realm in realmfield. realmfield: name of field in realmnode containing principal realm. out: pointer to newly allocated null terminated string containing principal name. May be NULL (to only populate outlen). outlen: pointer to length of out on output, excluding terminating null. May be NULL (to only populate out). Description: Represent principal name and realm in ASN.1 structure as nullterminated string. The string is allocated by this function. It is the responsibility of the caller to deallocate it. Note that the output length outlen does not include the terminating null character. Return value: Returns SHISHI OK if successful.

shishi principal name set

[Function] int shishi_principal_name_set (Shishi * handle,

Shishi asn1 namenode, const char * namefield, Shishi name type name_type, const char * name[])

handle: shishi handle as allocated by shishi_init(). namenode: ASN.1 structure with principal in namefield. namefield: name of field in namenode containing principal name. name type: type of principal, see Shishi name type, usually SHISHI NT UNKNOWN. name: null-terminated input array with principal name. Description: Set the given principal name field to the given name. Return value: Returns SHISHI OK if successful.

shishi principal set

[Function] int shishi_principal_set (Shishi * handle,

Shishi asn1 namenode, const char * namefield, const char * name)

handle: shishi handle as allocated by shishi_init(). namenode: ASN.1 structure with principal in namefield. namefield: name of field in namenode containing principal name. name: null-terminated string with principal name in RFC 1964 form. Description: Set principal name field in an ASN.1 structure to the given name. Return value: Returns SHISHI OK if successful.

shishi derive default salt

[Function] int shishi_derive_default_salt (Shishi * handle,

const char * name, char ** salt)

handle: shishi handle as allocated by shishi_init(). name: principal name of user. salt: output variable with newly allocated salt string. Description: Derive the default salt from a principal. The default salt is the concatenation of the decoded realm and the principal. Return value: Return SHISHI OK if successful.

shishi server for local service

[Function] char * shishi_server_for_local_service (Shishi * handle,

const char * service) handle: shishi handle as allocated by shishi_init(). service: null terminated string with name of service, e.g., "host". Description: Construct a service principal (e.g., "imap/yxa.extuno.com") based on supplied service name (i.e., "imap") and the system’s hostname as returned by hostname() (i.e., "yxa.extundo.com"). The string must be deallocated by the caller. Return value: Return newly allocated service name string.

shishi authorize strcmp

[Function] int shishi_authorize_strcmp (Shishi * handle,

const char * principal, const char * authzname)

handle: shishi handle allocated by shishi_init(). principal: string with desired principal name. authzname: authorization name. Description: Authorization of authzname against desired principal according to "basic" authentication, i.e., testing for identical strings.

Return value: Returns 1 if authzname is authorized for services by the encrypted principal, and 0 otherwise.

shishi authorize k5login

[Function] int shishi_authorize_k5login (Shishi * handle,

const char * principal, const char * authzname)

handle: shishi handle allocated by shishi_init(). principal: string with desired principal name and realm. authzname: authorization name. Description: Authorization of authzname against desired principal in accordance with the MIT/Heimdal authorization method. Return value: Returns 1 if authzname is authorized for services by principal, and returns 0 otherwise.

shishi authorization parse

[Function] int shishi_authorization_parse (const char * authorization)

authorization: name of authorization type, "basic" or "k5login". Description: Parse authorization type name. Return value: Returns authorization type corresponding to a string.

shishi authorized p

[Function] int shishi_authorized_p (Shishi * handle, Shishi tkt * tkt,

const char * authzname) handle: shishi handle allocated by shishi_init(). tkt: input variable with ticket info. authzname: authorization name. Description: Simplistic authorization of authzname against encrypted client principal name inside ticket. For "basic" authentication type, the principal name must coincide with authzname. The "k5login" authentication type attempts the MIT/Heimdal method of parsing the file "~/.k5login" for additional equivalence names. Return value: Returns 1 if authzname is authorized for services by the encrypted principal, and 0 otherwise.

shishi generalize time

[Function] const char * shishi_generalize_time (Shishi * handle,

time t t) handle: Shishi handle as allocated by shishi_init(). t: C time to convert. Description: Converts C time t to a KerberosTime string representation. The returned string must not be deallocated by the caller. Return value: Returns a KerberosTime formatted string corresponding to the input parameter.

shishi generalize now

[Function] const char * shishi_generalize_now (Shishi * handle)

handle: Shishi handle as allocated by shishi_init().

Description: Converts the current time to a KerberosTime string. The returned string must not be deallocated by the caller.

Return value: Returns a KerberosTime formatted string corresponding to the current time.

shishi generalize ctime

[Function] time_t shishi_generalize_ctime (Shishi * handle, const char * t)

handle: Shishi handle as allocated by shishi_init(). t: KerberosTime string to convert.

Description: Converts a KerberosTime formatted string in t to integral C time representation.

Return value: Returns the C time corresponding to the input argument.

shishi time

[Function] int shishi_time (Shishi * handle, Shishi asn1 node,

const char * field, char ** t)

handle: Shishi handle as allocated by shishi_init(). node: ASN.1 structure to get time from. field: Name of the field in the ASN.1 node carrying time. t: Returned pointer to an allocated char array containing a null-terminated time string.

Description: Extracts time information from an ASN.1 structure, and to be precise, does so from the named field field within the structure node.

Return value: Returns SHISHI_OK if successful, or an error.

shishi ctime

[Function] int shishi_ctime (Shishi * handle, Shishi asn1 node,

const char * field, time t * t)

handle: Shishi handle as allocated by shishi_init(). node: ASN.1 structure to read field from. field: Name of field in node to read. t: Pointer to a C-time valued integer, being updated with the time value to be extracted.

Description: Extracts time information from an ASN.1 structure node, and from an arbitrary element field of that structure.

Return value: Returns SHISHI_OK if successful, SHISHI_ASN1_NO_ELEMENT if the element does not exist, SHISHI_ASN1_NO_VALUE if the field has no value. In all other

cases, SHISHI_ASN1_ERROR is returned.

shishi prompt password callback set

[Function] void shishi_prompt_password_callback_set (Shishi * handle,

shishi prompt password func cb) handle: shishi handle as allocated by shishi_init(). cb: function pointer to application password callback, a shishi_prompt_password_ func type. Description: Set a callback function that will be used by shishi_prompt_password() to query the user for a password. The function pointer can be retrieved using shishi_

prompt_password_callback_get().

The cb function should follow the shishi_prompt_password_func prototype: int prompt password (Shishi * handle, char **s, const char *format, va list ap); If the function returns 0, the s variable should contain a newly allocated string with the password read from the user.

shishi prompt password callback get

[Function] shishi_prompt_password_func
   shishi_prompt_password_callback_get (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Get the application password prompt function callback as set by shishi_

prompt_password_callback_set().

Returns: Returns the callback, a shishi_prompt_password_func type, or NULL.

shishi prompt password

[Function] int shishi_prompt_password (Shishi * handle, char ** s,

const char * format, ...)

handle: shishi handle as allocated by shishi_init(). s: pointer to newly allocated output string with read password. format: printf(3) style format string. ...: printf(3) style arguments. Description: Format and print a prompt, and read a password from user. The password is possibly converted (e.g., converted from Latin-1 to UTF-8, or processed using Stringprep profile) following any "stringprocess" keywords in configuration files. Return value: Returns SHISHI OK iff successful.

shishi resolv

[Function] Shishi_dns shishi_resolv (const char * zone,
    uint16 t querytype)

zone: Domain name of authentication zone, e.g. "EXAMPLE.ORG" querytype: Type of domain data to query for. Description: Queries the DNS resolver for data of type querytype about the domain name zone. Currently, the types SHISHI_DNS_TXT and SHISHI_DNS_SRV are the only supported kinds.

After its use, the returned list should be deallocated by a call to shishi_resolv_

free().

Return value: Returns a linked list of DNS resource records, or NULL if the query failed.

shishi resolv free

[Function] void shishi_resolv_free (Shishi dns rrs)

rrs: List of DNS RRs as returned by shishi_resolv(). Description: Deallocates a list of DNS resource records returned by a call to shishi_

resolv().

5.16 ASN.1 Functions

shishi asn1 read inline

[Function] int shishi_asn1_read_inline (Shishi * handle, Shishi asn1 node,

const char * field, char * data, size t * datalen)

handle: shishi handle as allocated by shishi_init(). node: ASN.1 variable to read field from. field: name of field in node to read. data: pre-allocated output buffer that will hold ASN.1 field data. datalen: on input, maximum size of output buffer, on output, actual size of output buffer. Description: Extract data stored in a ASN.1 field into a fixed size buffer allocated by caller. Note that since it is difficult to predict the length of the field, it is often better to use

shishi_asn1_read() instead.

Return value: Returns SHISHI OK if successful, SHISHI ASN1 NO ELEMENT if the element do not exist, SHISHI ASN1 NO VALUE if the field has no value, ot SHISHI ASN1 ERROR otherwise.

shishi asn1 read

[Function] int shishi_asn1_read (Shishi * handle, Shishi asn1 node,

const char * field, char ** data, size t * datalen)

handle: shishi handle as allocated by shishi_init(). node: ASN.1 variable to read field from. field: name of field in node to read. data: newly allocated output buffer that will hold ASN.1 field data. datalen: actual size of output buffer. Description: Extract data stored in a ASN.1 field into a newly allocated buffer. The buffer will always be zero terminated, even though datalen will not include the added zero. Return value: Returns SHISHI OK if successful, SHISHI ASN1 NO ELEMENT if the element do not exist, SHISHI ASN1 NO VALUE if the field has no value, ot SHISHI ASN1 ERROR otherwise.

shishi asn1 read optional

[Function] int shishi_asn1_read_optional (Shishi * handle,

Shishi asn1 node, const char * field, char ** data, size t * datalen)

handle: shishi handle as allocated by shishi_init(). node: ASN.1 variable to read field from. field: name of field in node to read. data: newly allocated output buffer that will hold ASN.1 field data. datalen: actual size of output buffer.

Description: Extract data stored in a ASN.1 field into a newly allocated buffer. If the field does not exist (i.e., SHISHI ASN1 NO ELEMENT), this function set datalen to 0 and succeeds. Can be useful to read ASN.1 fields which are marked OPTIONAL in the grammar, if you want to avoid special error handling in your code.

Return value: Returns SHISHI OK if successful, SHISHI ASN1 NO VALUE if the field has no value, ot SHISHI ASN1 ERROR otherwise.

shishi asn1 done

[Function] void shishi_asn1_done (Shishi * handle, Shishi asn1 node)

handle: shishi handle as allocated by shishi_init(). node: ASN.1 node to deallocate.

Description: Deallocate resources associated with ASN.1 structure. Note that the node must not be used after this call.

shishi asn1 pa enc ts enc

[Function] Shishi_asn1 shishi_asn1_pa_enc_ts_enc (Shishi * handle)

handle: shishi handle as allocated by shishi_init().

Description: Create new ASN.1 structure for PA-ENC-TS-ENC.

Return value: Returns ASN.1 structure.

shishi asn1 encrypteddata

[Function] Shishi_asn1 shishi_asn1_encrypteddata (Shishi * handle)

handle: shishi handle as allocated by shishi_init().

Description: Create new ASN.1 structure for EncryptedData

Return value: Returns ASN.1 structure.

shishi asn1 padata

[Function] Shishi_asn1 shishi_asn1_padata (Shishi * handle)

handle: shishi handle as allocated by shishi_init().

Description: Create new ASN.1 structure for PA-DATA.

Return value: Returns ASN.1 structure.

shishi asn1 methoddata

[Function] Shishi_asn1 shishi_asn1_methoddata (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for METHOD-DATA. Return value: Returns ASN.1 structure.

shishi asn1 etype info

[Function] Shishi_asn1 shishi_asn1_etype_info (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for ETYPE-INFO. Return value: Returns ASN.1 structure.

shishi asn1 etype info2

[Function] Shishi_asn1 shishi_asn1_etype_info2 (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for ETYPE-INFO2. Return value: Returns ASN.1 structure.

shishi asn1 asreq

[Function] Shishi_asn1 shishi_asn1_asreq (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for AS-REQ. Return value: Returns ASN.1 structure.

shishi asn1 asrep

[Function] Shishi_asn1 shishi_asn1_asrep (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for AS-REP. Return value: Returns ASN.1 structure.

shishi asn1 tgsreq

[Function] Shishi_asn1 shishi_asn1_tgsreq (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for TGS-REQ. Return value: Returns ASN.1 structure.

shishi asn1 tgsrep

[Function] Shishi_asn1 shishi_asn1_tgsrep (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for TGS-REP. Return value: Returns ASN.1 structure.

shishi asn1 apreq

[Function] Shishi_asn1 shishi_asn1_apreq (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for AP-REQ. Return value: Returns ASN.1 structure.

shishi asn1 aprep

[Function] Shishi_asn1 shishi_asn1_aprep (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for AP-REP. Return value: Returns ASN.1 structure.

shishi asn1 encapreppart

[Function] Shishi_asn1 shishi_asn1_encapreppart (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for AP-REP. Return value: Returns ASN.1 structure.

shishi asn1 ticket

[Function] Shishi_asn1 shishi_asn1_ticket (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for Ticket. Return value: Returns ASN.1 structure.

shishi asn1 encticketpart

[Function] Shishi_asn1 shishi_asn1_encticketpart (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for EncTicketPart. Return value: Returns ASN.1 structure.

shishi asn1 authenticator

[Function] Shishi_asn1 shishi_asn1_authenticator (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for Authenticator. Return value: Returns ASN.1 structure.

shishi asn1 enckdcreppart

[Function] Shishi_asn1 shishi_asn1_enckdcreppart (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for EncKDCRepPart. Return value: Returns ASN.1 structure.

shishi asn1 encasreppart

[Function] Shishi_asn1 shishi_asn1_encasreppart (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for EncASRepPart. Return value: Returns ASN.1 structure.

shishi asn1 krberror

[Function] Shishi_asn1 shishi_asn1_krberror (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for KRB-ERROR. Return value: Returns ASN.1 structure.

shishi asn1 krbsafe

[Function] Shishi_asn1 shishi_asn1_krbsafe (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for KRB-SAFE. Return value: Returns ASN.1 structure.

shishi asn1 priv

[Function] Shishi_asn1 shishi_asn1_priv (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for KRB-PRIV. Return value: Returns ASN.1 structure.

shishi asn1 encprivpart

[Function] Shishi_asn1 shishi_asn1_encprivpart (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Create new ASN.1 structure for EncKrbPrivPart. Return value: Returns ASN.1 structure.

shishi asn1 to der field

[Function] int shishi_asn1_to_der_field (Shishi * handle,

Shishi asn1 node, const char * field, char ** der, size t * len)

handle: shishi handle as allocated by shishi_init(). node: ASN.1 data that have field to extract. field: name of field in node to extract. der: output array that holds DER encoding of field in node. len: output variable with length of der output array. Description: Extract newly allocated DER representation of specified ASN.1 field. Return value: Returns SHISHI OK if successful, or SHISHI ASN1 ERROR if DER encoding fails (common reasons for this is that the ASN.1 is missing required values).

shishi asn1 to der

[Function] int shishi_asn1_to_der (Shishi * handle, Shishi asn1 node,

char ** der, size t * len)

handle: shishi handle as allocated by shishi_init(). node: ASN.1 data to convert to DER. der: output array that holds DER encoding of node. len: output variable with length of der output array.

Description: Extract newly allocated DER representation of specified ASN.1 data.

Return value: Returns SHISHI OK if successful, or SHISHI ASN1 ERROR if DER encoding fails (common reasons for this is that the ASN.1 is missing required values).

shishi asn1 msgtype

[Function] Shishi_msgtype shishi_asn1_msgtype (Shishi * handle,

Shishi asn1 node) handle: shishi handle as allocated by shishi_init(). node: ASN.1 type to get msg type for.

Description: Determine msg-type of ASN.1 type of a packet. Currently this uses the msg-type field instead of the APPLICATION tag, but this may be changed in the future.

Return value: Returns msg-type of ASN.1 type, 0 on failure.

shishi der msgtype

[Function] Shishi_msgtype shishi_der_msgtype (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Determine msg-type of DER coded data of a packet.

Return value: Returns msg-type of DER data, 0 on failure.

shishi der2asn1

[Function] Shishi_asn1 shishi_der2asn1 (Shishi * handle, const char * der,
    size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Convert arbitrary DER data of a packet to a ASN.1 type.

Return value: Returns newly allocate ASN.1 corresponding to DER data, or NULL on failure.

shishi der2asn1 padata

[Function] Shishi_asn1 shishi_der2asn1_padata (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of PA-DATA and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 methoddata

[Function] Shishi_asn1 shishi_der2asn1_methoddata (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of METHOD-DATA and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 etype info

[Function] Shishi_asn1 shishi_der2asn1_etype_info (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of ETYPE-INFO and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 etype info2

[Function] Shishi_asn1 shishi_der2asn1_etype_info2 (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of ETYPE-INFO2 and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 ticket

[Function] Shishi_asn1 shishi_der2asn1_ticket (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Decode DER encoding of Ticket and create a ASN.1 structure.

Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 encticketpart

[Function] Shishi_asn1 shishi_der2asn1_encticketpart (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Decode DER encoding of EncTicketPart and create a ASN.1 structure.

Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 asreq

[Function] Shishi_asn1 shishi_der2asn1_asreq (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Decode DER encoding of AS-REQ and create a ASN.1 structure.

Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 tgsreq

[Function] Shishi_asn1 shishi_der2asn1_tgsreq (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Decode DER encoding of TGS-REQ and create a ASN.1 structure.

Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 asrep

[Function] Shishi_asn1 shishi_der2asn1_asrep (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Decode DER encoding of AS-REP and create a ASN.1 structure.

Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 tgsrep

[Function] Shishi_asn1 shishi_der2asn1_tgsrep (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of TGS-REP and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 kdcrep

[Function] Shishi_asn1 shishi_der2asn1_kdcrep (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of KDC-REP and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 encasreppart

[Function] Shishi_asn1 shishi_der2asn1_encasreppart (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of EncASRepPart and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 enctgsreppart

[Function] Shishi_asn1 shishi_der2asn1_enctgsreppart (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of EncTGSRepPart and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 enckdcreppart

[Function] Shishi_asn1 shishi_der2asn1_enckdcreppart (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Decode DER encoding of EncKDCRepPart and create a ASN.1 structure.

Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 authenticator

[Function] Shishi_asn1 shishi_der2asn1_authenticator (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Decode DER encoding of Authenticator and create a ASN.1 structure.

Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 krberror

[Function] Shishi_asn1 shishi_der2asn1_krberror (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Decode DER encoding of KRB-ERROR and create a ASN.1 structure.

Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 krbsafe

[Function] Shishi_asn1 shishi_der2asn1_krbsafe (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Decode DER encoding of KRB-SAFE and create a ASN.1 structure.

Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 priv

[Function] Shishi_asn1 shishi_der2asn1_priv (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Decode DER encoding of KRB-PRIV and create a ASN.1 structure.

Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 encprivpart

[Function] Shishi_asn1 shishi_der2asn1_encprivpart (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of EncKrbPrivPart and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 apreq

[Function] Shishi_asn1 shishi_der2asn1_apreq (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of AP-REQ and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 aprep

[Function] Shishi_asn1 shishi_der2asn1_aprep (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of AP-REP and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 encapreppart

[Function] Shishi_asn1 shishi_der2asn1_encapreppart (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding. Description: Decode DER encoding of EncAPRepPart and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi der2asn1 kdcreq

[Function] Shishi_asn1 shishi_der2asn1_kdcreq (Shishi * handle,

const char * der, size t derlen)

handle: shishi handle as allocated by shishi_init(). der: input character array with DER encoding. derlen: length of input character array with DER encoding.

Description: Decode DER encoding of AS-REQ, TGS-REQ or KDC-REQ and create a ASN.1 structure. Return value: Returns ASN.1 structure corresponding to DER data.

shishi asn1 print

[Function] void shishi_asn1_print (Shishi * handle, Shishi asn1 node,

FILE * fh) handle: shishi handle as allocated by shishi_init(). node: ASN.1 data that have field to extract. fh: file descriptor to print to, e.g. stdout. Description: Print ASN.1 structure in human readable form, typically for debugging purposes.

5.17 Error Handling

Most functions in ‘Libshishi’ are returning an error if they fail. For this reason, the application should always catch the error condition and take appropriate measures, for example by releasing the resources and passing the error up to the caller, or by displaying a descriptive message to the user and cancelling the operation.

Some error values do not indicate a system error or an error in the operation, but the result of an operation that failed properly.

5.17.1 Error Values

Errors are returned as an int. Except for the SHISHI OK case, an application should always use the constants instead of their numeric value. Applications are encouraged to use the constants even for SHISHI OK as it improves readability. Possible values are:

SHISHI_OK

This value indicates success. The value of this error is guaranteed to always be 0 so you may use it in boolean constructs.

SHISHI_ASN1_ERROR

Error in ASN.1 function. (corrupt data?)

SHISHI_FOPEN_ERROR

Could not open file.

SHISHI_IO_ERROR

File input/output error.

SHISHI_MALLOC_ERROR

Memory allocation error in shishi library.

SHISHI_BASE64_ERROR

Base64 encoding or decoding failed. Data corrupt?

SHISHI_REALM_MISMATCH

Client realm value differ between request and reply.

SHISHI_CNAME_MISMATCH

Client name value differ between request and reply.

SHISHI_NONCE_MISMATCH

Replay protection value (nonce) differ between request and reply.

SHISHI_TGSREP_BAD_KEYTYPE

Incorrect key type used in TGS reply.

SHISHI_KDCREP_BAD_KEYTYPE

Incorrect key type used in reply from KDC.

SHISHI_APREP_BAD_KEYTYPE

Incorrect key type used in AP reply.

SHISHI_APREP_VERIFY_FAILED

Failed verification of AP reply.

SHISHI_APREQ_BAD_KEYTYPE

Incorrect key type used in AP request.

SHISHI_TOO_SMALL_BUFFER

Provided buffer was too small.

SHISHI_DERIVEDKEY_TOO_SMALL

Derived key material is too short to be applicable.

SHISHI_KEY_TOO_LARGE

The key is too large to be usable.

SHISHI_CRYPTO_ERROR

Low-level cryptographic primitive failed. This usually indicates bad password or data corruption.

SHISHI_CRYPTO_INTERNAL_ERROR

Internal error in low-level crypto routines.

SHISHI_SOCKET_ERROR

The system call socket() failed. This usually indicates that your system does not support the socket type.

SHISHI_BIND_ERROR

The system call bind() failed. This usually indicates insufficient permissions.

SHISHI_SENDTO_ERROR

The system call sendto() failed.

SHISHI_RECVFROM_ERROR

Error receiving data from server.

SHISHI_CLOSE_ERROR

The system call close() failed.

SHISHI_KDC_TIMEOUT

Timed out talking to KDC. This usually indicates a network or KDC address problem.

SHISHI_KDC_NOT_KNOWN_FOR_REALM

No KDC known for given realm.

SHISHI_TTY_ERROR

No TTY assigned to process.

SHISHI_GOT_KRBERROR

Server replied to the request with an error message.

SHISHI_HANDLE_ERROR

Failure to use handle. Missing handle, or misconfigured.

SHISHI_INVALID_TKTS

Ticket set not initialized. This usually indicates an internal application error.

SHISHI_TICKET_BAD_KEYTYPE

Key type used to encrypt ticket doesn’t match provided key. This usually indicates an internal application error.

SHISHI_INVALID_KEY

Reference to invalid encryption key.

SHISHI_APREQ_DECRYPT_FAILED

Could not decrypt AP-REQ using provided key. This usually indicates an internal application error.

SHISHI_TICKET_DECRYPT_FAILED

Could not decrypt Ticket using provided key. This usually indicates an internal application error.

SHISHI_INVALID_TICKET

Invalid ticked passed in call.

SHISHI_OUT_OF_RANGE

Argument lies outside of valid range.

SHISHI_ASN1_NO_ELEMENT

The ASN.1 structure does not contain the indicated element.

SHISHI_SAFE_BAD_KEYTYPE

Attempted access to non-existent key type.

SHISHI_SAFE_VERIFY_FAILED

Verification failed on either side.

SHISHI_PKCS5_INVALID_PRF

Invalid PKCS5 descriptor.

SHISHI_PKCS5_INVALID_ITERATION_COUNT

Invalid claim of iteration count in PKCS5 descriptor.

SHISHI_PKCS5_INVALID_DERIVED_KEY_LENGTH

Derived key length is incorrect for PKCS5 descriptor.

SHISHI_PKCS5_DERIVED_KEY_TOO_LONG

Derived key is too long for PKCS5 descriptor.

SHISHI_INVALID_PRINCIPAL_NAME

Principal name syntax error.

SHISHI_INVALID_ARGUMENT

Invalid argument passed in call. Wrong or unknown value.

SHISHI_ASN1_NO_VALUE

The indicated ASN.1 element does not carry a value.

SHISHI_CONNECT_ERROR

Connection attempt failed. Try again, or check availability.

SHISHI_VERIFY_FAILED

Verification failed on either side.

SHISHI_PRIV_BAD_KEYTYPE

The private key uses an incompatible encryption type.

SHISHI_FILE_ERROR

The desired file could not be accessed. Check permissions.

SHISHI_ENCAPREPPART_BAD_KEYTYPE

The present AP reply specifies an inpermissible key type.

SHISHI_GETTIMEOFDAY_ERROR

A request for present time of day has failed. This is usually internal, but a valid time is imperative for us.

SHISHI_KEYTAB_ERROR

Failed to parse keytab file.

SHISHI_CCACHE_ERROR

Failed to parse credential cache file.

5.17.2 Error Functions

shishi strerror

[Function] const char * shishi_strerror (int err)

err: shishi error code.

Description: Converts the return code in err to a human readable string.

Return value: Returns a pointer to a statically allocated string containing a description of the error with code err. This string can be used to output a diagnostic message to the user.

shishi error

[Function] const char * shishi_error (Shishi * handle)

handle: shishi handle as allocated by shishi_init().

Description: Extracts detailed information on the most recently occurred error condition. Note that memory is managed by the Shishi library, so the returned string must not be deallocated. Return value: Returns a pointer to a string describing an error. The string must not be deallocated by the caller.

shishi error clear

[Function] void shishi_error_clear (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Clears the internal error description. See shishi_error() on how to

access the error string, and shishi_error_set() as well as shishi_error_printf()

on how to set the error string. This function is mostly for Shishi’s internal use, but if you develop an extension of Shishi, it may be useful to support the same error handling infrastructure.

shishi error set

[Function] void shishi_error_set (Shishi * handle, const char * errstr)

handle: shishi handle as allocated by shishi_init(). errstr: A null-terminated character string holding a description, or NULL to clear the internal error string. Description: Sets the error description to the content of errstr. The string is copied into the Shishi internal structure, so you can deallocate any string passed to this function. This function is mostly for Shishi’s internal use, but if you develop an extension of Shishi, it may be useful to support the same error handling infrastructure.

shishi error printf

[Function] void shishi_error_printf (Shishi * handle, const char * format,
    ...)

handle: shishi handle as allocated by shishi_init(). format: printf style format string. ...: printf style arguments. Description: Sets the internal error description to a printf(3) formatted string. This function is mostly for Shishi’s internal use, but if you develop an extension of Shishi, it may be useful to support the same infrastructure for error handling.

shishi error outputtype

[Function] int shishi_error_outputtype (Shishi * handle)

handle: shishi handle as allocated by shishi_init(). Description: Reports the current output type used in message logging. Return value: Returns the output type. SHISHI_OUTPUTTYPE_NULL means no out-

put. SHISHI_OUTPUTTYPE_STDERR and SHISHI_OUTPUTTYPE_SYSLOG direct text to

the console, or to the syslog system.

shishi error set outputtype

[Function] void shishi_error_set_outputtype (Shishi * handle, int type)

handle: shishi handle as allocated by shishi_init(). type: output type, of enum type Shishi_outputtype.

Description: Sets the output type (NULL, stderr or syslog) used for information and warning messages. Intended values are SHISHI_OUTPUTTYPE_NULL, for no output at

all, SHISHI_OUTPUTTYPE_STDERR for output to the console, and SHISHI_OUTPUTTYPE_

SYSLOG for syslog messaging. The first value covers everything different from the latter two values.

shishi info

[Function] void shishi_info (Shishi * handle, const char * format, ...)

handle: shishi handle as allocated by shishi_init(). format: printf style format string. ...: printf style arguments.

Description: Prints an informational message, composed from the arguments, to the output stream set in handle.

shishi warn

[Function] void shishi_warn (Shishi * handle, const char * format, ...)

handle: shishi handle as allocated by shishi_init(). format: printf style format string. ...: printf style arguments.

Description: Prints a warning, composed from the arguments, to the output stream set in handle.

shishi verbose

[Function] void shishi_verbose (Shishi * handle, const char * format, ...)

handle: shishi handle as allocated by shishi_init(). format: printf style format string. ...: printf style arguments.

Description: Prints a diagnostic message, composed from the arguments, to the output stream set in handle. The current verbosity setting determines whether the message is actually printed, or is suppressed due to low significance.

5.18 Examples

This section will be extended to contain walk-throughs of example code that demonstrate how ‘Shishi’ is used to write your own applications that support Kerberos 5. The rest of the current section consists of some crude hints for the example client/server applications that is part of Shishi, taken from an email but saved here for lack of a better place to put it.

There are two programs: ’client’ and ’server’ in src/.

The client output an AP-REQ, waits for an AP-REP, and then simply reads data from stdin.

The server waits for an AP-REQ, parses it and prints an AP-REP, and then read data from stdin.

Both programs accept a Kerberos server name as the first command line argument. Your KDC must know this server, since the client tries to get a ticket for it (first it gets a ticket granting ticket for the default username), and you must write the key for the server into /usr/local/etc/shishi.keys on the Shishi format, e.g.:

     -----BEGIN SHISHI KEY-----
     Keytype: 16 (des3-cbc-sha1-kd)
     Principal: sample/latte.josefsson.org
     Realm: JOSEFSSON.ORG
     8W0VrQQBpxlACPQEqN91EHxbvFFo2ltt
     -----END SHISHI KEY-----

You must extract the proper encryption key from the KDC in some way. (This part will be easier when Shishi include a KDC, a basic one isn’t far away, give me a week or to.)

The intention is that the data read, after the authentication phase, should be protected using KRB SAFE (see RFC) but I haven’t added this yet.

5.19 Kerberos Database Functions

Shisa is a separate and standalone library from Shishi (see Section 3.1 [Introduction to Shisa], page 18). If you only wish to manipulate the information stored in the Kerberos user database used by Shishi, you do not need to link or use the Shishi library at all. However, you may find it useful to combine the two libraries.

For two real world examples on using the Shisa library, refer to src/shisa.c (Shisa command line tool) and src/kdc.c (part of Shishid server).

Shisa uses two ‘struct’s to carry information. The first, Shisa_principal, is used to hold information about principals. The struct does not contain pointers to strings etc, so the library assumes the caller is responsible for allocating and deallocating the struct itself. Each such struct is (uniquely) identified by the combination of principal name and realm name.

     struct Shisa_principal
     {
       int isdisabled;
       uint32_t kvno;
       time_t notusedbefore;
       time_t lastinitialtgt;      /* time of last initial request for a TGT */
       time_t lastinitialrequest;  /* time of last initial request */
       time_t lasttgt;             /* time of issue for the newest TGT used */
       time_t lastrenewal;         /* time of the last renewal */
       time_t passwordexpire;      /* time when the password will expire */
       time_t accountexpire;       /* time when the account will expire. */
     };
     typedef struct Shisa_principal Shisa_principal;

The second structure is called Shisa_key and hold information about cryptographic keys. Because the struct contain pointers, and the caller cannot know how many keys a principal have, the Shisa library manages memory for the struct. The library allocate the structs, and the pointers within them. The caller may deallocate them, but it is recommended to use shisa_key_free or shisa_keys_free instead. Note that each principal may have multiple keys.

struct Shisa_key
{
  uint32_t kvno;
  int32_t etype;
  int priority;
  char *key;
  size_t keylen;
  char *salt;
  size_t saltlen;
  char *str2keyparam;
  size_t str2keyparamlen;
  char *password;
};
typedef struct Shisa_key Shisa_key;

Shisa is typically initialized by calling shisa_init, and deinitialized (when the application no longer need to use Shisa, typically when it shuts down) by calling shisa_done, but here are the complete (de)initialization interface functions.

shisa

[Function] Shisa * shisa ()

Description: Initializes the Shisa library. If this function fails, it may print diagnostic errors to standard error. Return value: Returns a Shisa library handle, or NULL on error.

shisa done

[Function] void shisa_done (Shisa * dbh)

dbh: Shisa handle as allocated by shisa(). Description: Deallocates the shisa library handle. The handle must not be used in calls to any shisa function after the completion of this call.

shisa init

[Function] int shisa_init (Shisa ** dbh)

dbh: Returned pointer to a created Shisa library handle. Description: Creates a Shisa library handle, using shisa(), reading the system configuration file from its default location. The path to the default system configuration file is decided at compile time (sysconfdir/shisa.conf). The handle is allocated regardless of return value, the only exception being SHISA_ INIT_ERROR, which indicates a problem in allocating the handle. Other error conditions arise while reading a file. Return value: Returns SHISA_OK, or an error code. The value SHISA_INIT_ERROR indicates a failure to create the handle.

shisa init with paths

[Function] int shisa_init_with_paths (Shisa ** dbh, const char * file)

dbh: Returned pointer to a created Shisa library handle. file: Filename of system configuration, or NULL.

Description: Creates a Shisa library handle, using shisa(), but reading the system configuration file at the location file, or at the default location, should file be NULL. The path to the default system configuration file is decided at compile time (sysconfdir/shisa.conf).

The handle is allocated regardless of return value, the only exception being SHISA_ INIT_ERROR, which indicates a problem in allocating the handle. Other error conditions arise while reading a file.

Return value: Returns SHISA_OK, or an error code. The value SHISA_INIT_ERROR indicates a failure to create the handle.

The default configuration file is typically read automatically by calling shisa_init, but if you wish to manually access the Shisa configuration file functions, here is the complete interface.

shisa cfg db

[Function] int shisa_cfg_db (Shisa * dbh, const char * value)

dbh: Shisa library handle created by shisa(). value: String containing database definition.

Description: Sets up and opens a new database. The syntax of the parameter value is "TYPE[ LOCATION[ PARAMETER]]", where TYPE is one of the supported database types, typically "file".

The above substrings LOCATION and PARAMETER are optional strings passed on verbatim to the database during initialization. Neither TYPE nor LOCATION may contain embedded spaces, but PARAMETER may do so.

Return value: Returns SHISA_OK if a database was parsed and opened successfully.

shisa cfg

[Function] int shisa_cfg (Shisa * dbh, const char * option)

dbh: Shisa library handle created by shisa(). option: String with options to prime the Shisa library.

Description: Configures the Shisa library from the specification option. This call expects a string declaration of the form "db=VALUE", or "db VALUE". Here VALUE is the same declaration as used by shisa_cfg_db(), i.e., of the form "TYPE[ LO-CATION[ PARAMETER]]".

The prefix "db", mandatory in option, makes shisa_cfg() suited to a syntax with key-value pairs also in PARAMETER.

Return value: Returns SHISA_OK if option is valid.

shisa cfg from file

[Function] int shisa_cfg_from_file (Shisa * dbh, const char * cfg)

dbh: Shisa library handle created by shisa(). cfg: File name where to read configuration.

Description: Configures the Shisa library using a configuration file located at cfg.

Return value: Returns SHISA_OK if successful. Typically returns SHISA_CFG_NO_FILE in response to a misnamed file.

shisa cfg default systemfile

[Function] const char * shisa_cfg_default_systemfile (Shisa * dbh)

dbh: Shisa library handle created by shisa().

Description: Fetches information on the installed configuration.

Return value: Returns file name of the active system configuration.

The core part of the Shisa interface follows. The typical procedure is to use shisa_ principal_find to verify that a specific principal exists, and to extract some information about it, and then use shisa_keys_find to get the cryptographic keys for the principal, usually suppliying some hints as to which of all keys you are interested in (e.g., key version number and encryption algorithm number).

shisa enumerate realms

[Function] int shisa_enumerate_realms (Shisa * dbh, char *** realms,
    size t * nrealms)

dbh: Shisa library handle created by shisa(). realms: Returned pointer to a newly allocated array of also allocated and nullterminated UTF-8 strings with realm names. nrealms: Pointer to a number which is updated with the number of just allocated and returned realm strings.

Description: Extracts a list of all realm names in backend, as null-terminated UTF-8 strings. The caller is responsible for deallocating all strings as well as the array *realms.

Return value: Returns SHISA_OK on success, or an error code.

shisa enumerate principals

[Function] int shisa_enumerate_principals (Shisa * dbh,
   const char * realm, char *** principals, size t * nprincipals)

dbh: Shisa library handle created by shisa(). realm: Name of realm, as null-terminated UTF-8 string. principals: Returned pointer to newly allocated array of just allocated null-terminated UTF-8 strings with principal names. nprincipals: Pointer to an integer updated with the number of just allocated and returned principal names.

Description: Extracts a list of all principal names in backend belonging to the realm realm, as null-terminated UTF-8 strings. The caller is responsible for deallocating all strings and the array *principals.

Return value: Returns SHISA_OK on success, SHISA_NO_REALM if the specified realm does not exist, or an error code otherwise.

shisa principal find

[Function] int shisa_principal_find (Shisa * dbh, const char * realm,

const char * principal, Shisa principal * ph)

dbh: Shisa library handle created by shisa(). realm: Name of the realm the principal belongs to. principal: Name of principal to get information about. ph: Pointer to a previously allocated principal structure where information about the principal is to be stored.

Description: Extracts information about given the PRINCIPAL@REALM pair selected by principal and realm.

Return value: Returns SHISA_OK if successful, SHISA_NO_REALM if the indicated realm does not exist, SHISA_NO_PRINCIPAL if the indicated principal does not exist, or an error code otherwise.

shisa principal update

[Function] int shisa_principal_update (Shisa * dbh, const char * realm,

const char * principal, const Shisa principal * ph)

dbh: Shisa library handle created by shisa(). realm: Name of the realm the principal belongs to. principal: Name of principal to get information about. ph: Pointer to an existing principal structure containing information to store in the database.

Description: Modifies information stored about the given principal PRINCI-PAL@REALM. Note that it is usually a good idea to set in ph only the fields that are to be updated.

It is generally suggested to first call shisa_principal_find(), to get the current information, then to modify one field and call shisa_principal_update().

Modifying several values is not recommended in general, as this will 1) overwrite any modifications made to other fields between the two calls (by other processes) and 2) will cause all values to be written again, which may generate more overhead.

Return value: Returns SHISA_OK if successful, SHISA_NO_REALM if the indicated realm does not exist, SHISA_NO_PRINCIPAL if the indicated principal does not exist, or an error code otherwise.

shisa principal add

[Function] int shisa_principal_add (Shisa * dbh, const char * realm,

const char * principal, const Shisa principal * ph,

const Shisa key * key) dbh: Shisa library handle created by shisa(). realm: Name of the realm the principal belongs to. principal: Name of principal to add. When set to NULL, only the realm realm is created. ph: Pointer to a principal structure with information to store in the database. key: Pointer to a key structure with information to store in the database.

Description: Inserts the given information into the database for the principal PRIN-CIPAL@REALM. In case principal is NULL, the parameters ph and key are not used, so only the realm is added to the database.

Return value: Returns SHISA_OK if the information was successfully added, or an error code otherwise.

shisa principal remove

[Function] int shisa_principal_remove (Shisa * dbh, const char * realm,

const char * principal) dbh: Shisa library handle created by shisa(). realm: Name of the realm the principal belongs to. principal: Name of the principal to remove. Set to NULL, only the realm realm is removed.

Description: Removes all information stored in the database for the given principal PRINCIPAL@REALM. When principal is NULL, then the realm realm is itself removed, but this can only succeed if the realm is already empty of principals.

Return value: Returns SHISA_OK if successful, or an error code.

shisa keys find

[Function] int shisa_keys_find (Shisa * dbh, const char * realm, const char * principal, const Shisa key * hint, Shisa key *** keys,

size t * nkeys) dbh: Shisa library handle created by shisa(). realm: Name of the realm the principal belongs to. principal: Name of the principal whose keys are examined. hint: Pointer to a Shisa key structure with hints on matching criteria for relevant keys. NULL matches all keys. keys: Returned pointer to a newly allocated array of Shisa key structures. nkeys: Pointer to an integer updated with the number of allocated Shisa key structures in *keys.

Description: Iterates through the set of keys belonging to PRINCIPAL@REALM, as selected by principal and realm. Then extracts any keys that match the criteria in hint.

Not all elements of hint need to be filled in. Set only the fields you are interested in. For example, if you want to extract all keys of etype 3, i.e., DES-CBC-MD5, then set the field key->etype to 3, and all other fields to zero. Return value: Returns SHISA_OK if successful, or an error code.

shisa key add

[Function] int shisa_key_add (Shisa * dbh, const char * realm,

const char * principal, const Shisa key * key)

dbh: Shisa library handle created by shisa(). realm: Name of the realm the principal belongs to. principal: Name of the principal to add a new key for. key: Pointer to a Shisa key structure with the new key. Description: Adds a complete key key to the database entry belonging to the principal PRINCIPAL@REALM, as set by principal and realm. Return value: Returns SHISA_OK if successful, or an error code.

shisa key update

[Function] int shisa_key_update (Shisa * dbh, const char * realm, const char * principal, const Shisa key * oldkey,

const Shisa key * newkey) dbh: Shisa library handle created by shisa(). realm: Name of the realm the principal belongs to. principal: Name of the principal needing an updated key. oldkey: Pointer to a Shisa key structure giving matching criteria for locating the key to be updated. newkey: Pointer to a complete Shisa key structure, in which all fields are used for the new key. Note that oldkey normally has far fewer fields filled-in. Description: Modifies data about a key stored in the database, a key belonging to the principal selected by principal and realm. First oldkey is used to locate the key to update, as does shisa_keys_find(). Then the found key is modified to carry whatever information is stored in newkey. Not all elements of oldkey need to be filled out, only sufficiently many so as to uniquely identify the desired key. For example, if you want to modify the information stored about a unique key of etype 3, i.e., DES-CBC-MD5, then set the field key->etype to 3, leaving all other fields as zero. Return value: Returns SHISA_OK on success, SHISA_NO_KEY if no key could be located, SHISA_MULTIPLE_KEY_MATCH if more than a single key matched the given criteria, or an error code otherwise.

shisa key remove

[Function] int shisa_key_remove (Shisa * dbh, const char * realm,

const char * principal, const Shisa key * key)

dbh: Shisa library handle created by shisa(). realm: Name of the realm the principal belongs to. principal: Name of the principal whose key is to be removed. key: Pointer to a Shisa key structure with hints on matching criteria for the key to select. Description: Removes from the Shisa database a key, matching the hints in key, for the user PRINCIPAL@REALM. Not all elements of key need to be filled in, only those relevant to locate the key uniquely. For example, if you want to remove the only key of etype 3, i.e., DES-CBC-MD5, then set the field key->etype to 3, and all other fields to zero. Return value: Returns SHISA_OK on success, SHISA_NO_KEY if no key could be located, SHISA_MULTIPLE_KEY_MATCH if more than one key matched the given criteria, or an error code otherwise.

shisa key free

[Function] void shisa_key_free (Shisa * dbh, Shisa key * key)

dbh: Shisa library handle created by shisa(). key: Pointer to a Shisa key structure to deallocate. Description: Deallocates the fields of a Shisa key structure, as well as the structure itself.

shisa keys free

[Function] void shisa_keys_free (Shisa * dbh, Shisa key ** keys,

size t nkeys) dbh: Shisa library handle created by shisa(). keys: Pointer to an array of Shisa key structures. nkeys: Number of key elements in the array keys. Description: Deallocates each key element in the array keys of Shisa database keys, using repeated calls to shisa_key_free().

Error handling is similar to that for Shishi in general (see Section 5.17 [Error Handling], page 219), i.e., you invoke shisa_strerror on the integer return value received by some function, if the return value is non-zero. Below is the complete interface.

shisa strerror

[Function] const char * shisa_strerror (int err)

err: Shisa error code. Description: Explains verbally an error status err. The returned string can be used to compose a diagnostic message of benefit to a user. Return value: Returns a pointer to a statically allocated string, containing a description of the error given as input argument.

shisa info

[Function] void shisa_info (Shisa * dbh, const char * format, ...)

dbh: Shisa library handle created by shisa(). format: printf style format string. ...: printf style arguments.

Description: Prints an informational message to standard error. The text is composed from the arguments, like printf(3).

5.20 Generic Security Service

As an alternative to the native Shishi programming API, it is possible to program Shishi through the Generic Security Services (GSS) API. The advantage of using GSS-API in your security application, instead of the native Shishi API, is that it will be easier to port your application between different Kerberos 5 implementations, and even beyond Kerberos 5 to different security systems, that support GSS-API. In the free software world, however, almost the only widely used security system that supports GSS-API is Kerberos 5, so the last advantage is somewhat academic. But if you are porting applications using GSS-API for other Kerberos 5 implementations, or want a more mature and stable API than the native Shishi API, you may find using Shishi’s GSS-API interface compelling. Note that GSS-API only offer basic services, for more advanced uses you must use the native API.

Since the GSS-API is not specific to Shishi, it is distributed independently from Shishi. Further information on the GSS project can be found at http://www.gnu.org/software/

gss/.

6 Acknowledgements

Shishi uses Libtasn1 by Fabio Fiorina, Libgcrypt and Libgpg-error by Werner Koch, Libidn by Simon Josefsson, cvs2cl by Karl Fogel, and gdoc by Michael Zucchi.

Several GNU packages simplified development considerably, those packages include Autoconf, Automake, Libtool, Gnulib, Gettext, Indent, CVS, Texinfo, Help2man and Emacs.

Several people reported bugs, sent patches or suggested improvements, see the file THANKS.

Nicolas Pouvesle wrote the section about the Kerberos rsh/rlogin protocol.

This manual borrows text from the Kerberos 5 specification.

Appendix A Criticism of Kerberos

The intention with this section is to discuss various problems with Kerberos 5, so you can form a conscious decision how to deploy and use Shishi correctly in your organization. Currently the issues below are condensed, and mostly serve as a reminder for the author to elaborate on them.

No encryption scheme with security proof.

No standardized API, and GSS mechanism lack important functionality.

Lack of authorization system. (krb5 kuserok())

Host to realm mapping relies on insecure DNS or static configuration files.

Informational model and user database administration.

Non-formal specification. Unclear on the etype to use for session keys (etype in request or database?). Unclear on how to populate some “evident” fields (e.g., cname in tickets for AS-REQ, or crealm, cname, realm, sname, ctime and cusec in KRB-ERROR). Unclear error code semantics (e.g., logic for when to use S PRINCIPAL UNKNOWN absent). Some KRB-ERROR fields are required, but can’t be usefully populated in some situations, and no guidance is given on what they should contain.

RFC 1510/1510bis incompatibilities. NULL enctype removed without discussion, and it is still used by some 1964 GSSAPI implementations. KRB SAFE text (3.4.1) says the checksum is generated using the session or sub-session key, which contradicts itself (compare section 3.2.6) and also RFC 1510, which both allow the application to define the key. Verification of KRB SAFE now require the key to be compatible with the (sub-)session key, in 1510 the only requirement was that it was collision proof.

Problems with RFC 1510bis. Uses bignum INTEGER for TYPED-DATA and AD-AND-OR.

Problems with crypto specification. It uses the word “random” many times, but there is no discussion on the randomness requirements. Practical experience indicate it is impossible to use true randomness for all “random” fields, and no implementation does this. A post by Don Davis on the ietf-krb-wg list tried to provide insight, but the information was never added to the specification.

Appendix B Protocol Extensions

This appendix specifies the non-standard protocol elements implemented by Shishi. By nature of being non-standard, everything described here is experimental. Comments and feedback is appreciated.

B.1 STARTTLS protected KDC exchanges

Shishi is able to “upgrade” TCP communications with the KDC to use the Transport Layer Security (TLS) protocol. The TLS protocol offers integrity and privacy protected exchanges. TLS also offers authentication using username and passwords, X.509 certificates, or OpenPGP certificates. Kerberos 5 claims to offer some of these features, although it is not as rich as the TLS protocol. An inconclusive list of the motivation for using TLS is given below.

• Server authentication of the KDC to the client. In traditional Kerberos 5, KDC authentication is only proved as a side effect that the KDC knows your encryption key (i.e., your password).

• Client authentication against KDC. Kerberos 5 assume the user knows a key (usually in the form of a password). Sometimes external factors make this hard to fulfill. In some situations, users are equipped with smart cards with a RSA authentication key. In others, users have a OpenPGP client on their desktop, with a public OpenPGP key known to the server. In some situations, the policy may be that password authentication may only be done through SRP.

• Kerberos exchanges are privacy protected. Part of many Kerberos packets are transfered without privacy protection (i.e., encryption). That part contains information, such as the client principal name, the server principal name, the encryption types supported by the client, the lifetime of tickets, etc. Revealing such information is, in some threat models, considered a problem. Thus, this enables “anonymity”.

• Prevents downgrade attacks affecting encryption types. The encryption type of the ticket in KDC-REQ are sent in the clear in Kerberos 5. This allows an attacker to replace the encryption type with a compromised mechanisms, e.g. 56-bit DES. Since clients in general cannot know the encryption types other servers support, it is difficult for the client to detect if there was a man-in-the-middle or if the remote server simply did not support a stronger mechanism. Clients may chose to refuse 56-bit DES altogether, but in some environments this leads to operational difficulties.

• TLS is well-proved and the protocol is studied by many parties. This is an advantage in network design, where TLS is often already assumed as part of the solution since it is used to protect HTTP, IMAP, SMTP etc. In some threat models, the designer prefer to reduce the number of protocols that can hurt the overall system security if they are compromised.

Other reasons for using TLS exists.

B.1.1 TCP/IP transport with TLS upgrade (STARTTLS)

RFC 1510bis requires Kerberos servers (KDCs) to accept TCP requests. Each request and response is prefixed by a 4 octet integer in network byte order, indicating the length of the packet. The high bit of the length was reserved for future expansion, and servers that do not understand how to interpret a set high bit must return a KRB-ERROR with a

KRB_ERR_FIELD_TOOLONG and close the TCP stream.

The TCP/IP transport with TLS upgrade (STARTTLS) uses this reserved bit as follows. First we define a new extensible typed hole for Kerberos 5 messages, because we used the only reserved bit. It is thus prudent to offer future extensions on our proposal. Secondly we reserve two values in this new typed hole, and described how they are used to implement STARTTLS.

B.1.2 Extensible typed hole based on reserved high bit

When the high bit is set, the remaining 31 bits of the 4 octets are treated as an extensible typed hole, and thus form a 31 bit integer enumerating various extensions. Each of the values indicate a specific extended operation mode, two of which are used and defined here, and the rest are left for others to use. If the KDC do not understand a requested extension, it MUST return a KRB-ERROR with a KRB_ERR_FIELD_TOOLONG value (prefixed by the 4 octet length integer, with the high bit clear, as usual) and close the TCP stream.

Meaning of the 31 lower bits in the 4 octet field, when the high bit is set:

 0               RESERVED.
 1               STARTTLS requested by client.
 2               STARTTLS request accepted by server.
 3...2147483647  AVAILABLE for registration (via bug-shishi@josefsson.org).
 2147483648      RESERVED.

B.1.3 STARTTLS requested by client (extension mode 1)

When this is sent by the client, the client is requesting the server to start TLS negotiation on the TCP stream. The client MUST NOT start TLS negotiation immediately. Instead, the client wait for either a KRB-ERROR (sent normally, prefixed by a 4 octet length integer) indicating the server do not understand the set high bit, or 4 octet which is to interpreted as an integer in network byte order, where the high bit is set and the remaining 31 bit are interpreted as an integer specifying the “STARTTLS request accepted by server”. In the first case, the client infer that the server do not understand (or wish to support) STARTTLS, and can re-try using normal TCP, if unprotected Kerberos 5 exchanges are acceptable to the client policy. In the latter case, it should invoke TLS negotiation on the stream. If any other data is received, the client MUST close the TCP stream.

B.1.4 STARTTLS request accepted by server (extension mode 2)

This 4 octet message should be sent by the server when it has received the previous 4 octet message. The message is an acknowledgment of the client’s request to initiate STARTTLS on the channel. The server MUST then invoke a TLS negotiation.

B.1.5 Proceeding after successful TLS negotiation

If the TLS negotiation ended successfully, possibly also considering client or server policies, the exchange within the TLS protected stream is performed like normal UDP Kerberos 5 exchanges, i.e., there is no TCP 4 octet length field before each packet. Instead each Kerberos packet MUST be sent within one TLS record, so the application can use the TLS record length as the Kerberos 5 packet length.

B.1.6 Proceeding after failed TLS negotiation

If the TLS negotiation fails, possibly due to client or server policy (e.g., inadequate support of encryption types in TLS, or lack of client or server authentication) the entity that detect the failure MUST disconnected the connection. It is expected that any error messages that explain the error condition is transfered by TLS.

B.1.7 Interaction with KDC addresses in DNS

Administrators for a KDC may announce the KDC address by placing SRV records in

DNS for the realm, as described in draft-ietf-krb-wg-krb-dns-locate-03.txt. That

document mention TLS, but do not reference any work that describe how KDCs uses TLS. Until further clarified, consider the TLS field in that document to refer to implementation supporting this STARTTLS protocol.

B.1.8 Using TLS authentication logic in Kerberos

The server MAY consider the authentication performed by the TLS exchange as sufficient to issue Kerberos 5 tickets to the client, without requiring, e.g., pre-authentication. However, it is not an error to require or use pre-authentication as well.

The client may also indicate that it wishes to use TLS both for authentication and data protection by using the ‘NULL’ encryption type in its request. The server can decide from its local policy whether or not issuing tickets based solely on TLS authentication, and whether ‘NULL’ encryption within TLS, is acceptable or not. This mode is currently under investigation.

B.1.9 Security considerations

Because the initial token is not protected, it is possible for an active attacker to make it appear to the client that the server do not support this extension. It is up to client configuration to disallow non-TLS connections, if this vulnerability is deemed unacceptable. For interoperability, we suggest the default behaviour should be to allow automatic fallback to TCP or UDP.

The security considerations of both TLS and Kerberos 5 are inherited. Using TLS for authentication and/or data protection together with Kerberos alter the authentication logic fundamentally. Thus, it may be that even if the TLS and Kerberos 5 protocols and implementations were secure, the combination of TLS and Kerberos 5 described here could be insecure.

No channel bindings are provided in the Kerberos messages. It is an open question whether, and how, this should be fixed.

B.2 Telnet encryption with AES-CCM

This appendix describe how Shishi use the Advanced Encryption Standard (AES) encryption algorithm in Counter with CBC-MAC mode (RFC 3610) with the telnet encryption option (RFC 2946).

B.2.1 Command Names and Codes

Encryption Type
      AES_CCM             12
Suboption Commands
      AES_CCM_INFO         1
      AES_CCM_INFO_OK      2
      AES_CCM_INFO_BAD     3

B.2.2 Command Meanings

   IAC SB ENCRYPT IS AES_CCM AES_CCM_INFO <M> <L> <nonce> IAC SE

The sender of this command selects desired M and L parameters, and nonce, as described in RFC 3610, and sends it to the other side of the connection. The parameters and the nonce are sent in clear text. Only the side of the connection that is WILL ENCRYPT may send the AES CCM INFO command.

   IAC SB ENCRYPT REPLY AES_CCM AES_CCM_INFO_BAD IAC SE

The sender of this command rejects the parameters received in the AES CCM INFO command. Only the side of the connection that is DO ENCRYPT may send the AES CCM INFO BAD command. The command MUST be sent if the nonce field length does not match the selected value of L. The command MAY be sent if the receiver does not accept the parameters for a reason such as policy. No capability is provided for negotiating these parameters.

   IAC SB ENCRYPT REPLY AES_CCM AES_CCM_INFO_OK IAC SE

The sender of this command accepts the parameters received in the AES CCM INFO command. Only the side of the connection that is DO ENCRYPT may send the AES CCM INFO BAD command. The command MUST NOT be sent if the nonce field length does not match the selected value of L.

B.2.3 Implementation Rules

Once an AES CCM INFO OK command has been received, the WILL ENCRYPT side of the connection should do keyid negotiation using the ENC KEYID command. Once the keyid negotiation has successfully identified a common keyid, then START and END commands may be sent by the side of the connection that is WILL ENCRYPT. Data will be encrypted using the AES-CCM algorithm, with the negotiated nonce and parameters M and L. After each successful encryption and decryption, the nonce is treated as an integer in network byte order, and is incremented by one.

If encryption (decryption) is turned off and back on again, and the same keyid is used when re-starting the encryption (decryption), then the intervening clear text must not change the state of the encryption (decryption) machine. In particular, the AES-CCM nonce must not have been re-set.

If a START command is sent (received) with a different keyid, the encryption (decryption) machine must be re-initialized immediately following the end of the START command with the new key and the parameters sent (received) in the last AES CCM INFO command.

If a new AES CCM INFO command is sent (received), and encryption (decryption) is enabled, the encryption (decryption) machine must be re-initialized immediately following the end of the AES CCM INFO command with the new nonce and parameters, and the keyid sent (received) in the last START command.

If encryption (decryption) is not enabled when an AES CCM INFO command is sent (received), the encryption (decryption) machine must be re- initialized after the next START command, with the keyid sent (received) in that START command, and the nonce and parameters sent (received) in this AES CCM INFO command.

At all times each peer MUST make sure that an AES-CCM nonce is not used twice with the same encryption key. The rules above help accomplish this in an interoperable way.

B.2.4 Integration with the AUTHENTICATION telnet option

As noted in the telnet ENCRYPTION option specifications, a keyid value of zero indicates the default encryption key, as might be derived from the telnet AUTHENTICATION option. If the default encryption key negotiated as a result of the telnet AUTHENTICATION option contains less than 32 bytes (corresponding to two 128 bit keys), then the AES CCM option MUST NOT be offered or used as a valid telnet encryption option. Furthermore, depending on policy for key lengths, the AES CCM option MAY be disabled if the default encryption key contain less than 48 bytes (for two 192 bit keys), or less than 64 bytes (for two 256 bit keys), as well.

The available encrypt key data is divided on two halves, where the first half is used to encrypt data sent from the server (decrypt data received by the client), and the second half is used to encrypt data sent from the client (decrypt data received by the server).

Note that the above algorithm assumes that the AUTHENTICATION mechanism generate keying material suitable for AES-CCM as used in this specification. This is not necessarily true in general, but we specify this behaviour as the default since it is true for most authentication systems in popular use today. New telnet AUTHENTICATION mechanisms may specify alternative methods for determining the keys to be used for this cipher suite in their specification, if the session key negotiated by that authentication mechanism is not a DES key and where this algorithm may not be safely used.

Kerberos 5 authentication clarification: The key used to encrypt data from the client to the server is taken from the sub-session key in the AP-REQ. The key used to decrypt data from the server to the client is taken from the sub-session key in the AP-REP. If mutual authentication is not negotiated, the key used to encrypt data from the client to the server is taken from the session key in the ticket, and the key used to decrypt data from the server to the client is taken from the sub-session key in the AP-REQ. Leaving the AP-REQ sub-key field empty MUST disable the AES CCM option.

B.2.5 Security Considerations

The protocol must be properly and securely implemented. For example, an implementation should not be vulnerable to various implementation-specific attacks such as buffer overflows or side-channel analysis.

We wish to repeat the suggestion from RFC 2946, to investigate a STARTTLS approach for Telnet encryption (and also authentication), when the security level provided by this specification is not adequate.

B.2.5.1 Telnet Encryption Protocol Security Considerations

The security consideration of the Telnet encryption protocol are inherited.

It should be noted that it is up to the authentication protocol used, if any, to bind the authenticity of the peers to a specific session.

The Telnet encryption protocol does not, in general, protect against possibly malicious downgrading to any mutually acceptable, but not preferred, encryption type. This places a requirement on each peer to only accept encryption types it trust fully. In other words, the Telnet encryption protocol do not guarantee that the strongest mutually acceptable encryption type is always selected.

B.2.5.2 AES-CCM Security Considerations

The integrity and privacy claims are inherited from AES-CCM. In particular, the implementation must make sure a nonce is not used more than once together with a single key.

Furthermore, the encryption key is assumed to be random, i.e., it should not be possible to guess it with probability of success higher than guessing any uniformly selected random key. RFC 1750 gives an overview of issues and recommendations related to randomness.

B.2.6 Acknowledgments

This document is based on the various Telnet Encryption RFCs (RFC 2946, RFC 2947, RFC 2948, RFC 2952 and RFC 2953).

B.3 Kerberized rsh and rlogin

This appendix describes the KCMDV0.2 protocol used in versions of inetutils patched for shishi. The KCMD protocol was developed by the MIT Kerberos team for the kerberized rsh and rlogin programs. The differences between rlogin and rsh are explained below, as are differences between protocol versions v0.1 and v0.2. Both remain in use resons of compatibility.

It is possible that some parts of this document are not in conformity with the original KCMD protocol, because there is no official specification of it. However, it seems that shishi’s implementation is compatible with MIT’s protocol.

Warning: If you are seriously considering using Kerberized rsh or rlogin, instead of more robust remote access protocols, such as SSH, you may first want to explore http://

www.cs.berkeley.edu/~hildrum/kerberos/ or the full paper at http://www.cs.
berkeley.edu/~hildrum/043.pdf.

B.3.1 Establish connection

Initially the client establishs a TCP connection to the server. Default ports are 543 (‘klogin’), 544 (‘kshell’), and 2105 (‘eklogin’). Here ‘eklogin’ is the same as ‘klogin’, but with encryption. There is no longer a separate ‘ekshell’ port, because encrypted and normal connection now use the same port ‘kshell’.

Normally ‘kshell’ needs a second connection for stderr. The client should send a null terminated string containing an ascii encoding of the port number to be used for this second connection. Since ‘klogin’ and ‘eklogin’ do not use a second connection for stderr, the client just sends an additional null byte to the server, which can be thought of as an empty string. Contrary to classic rsh and rlogin, the server need not check if the client’s port lies in the range 0-1023.

B.3.2 Kerberos identification

When a connection is being established, the first thing to do is to indicate that Kerberos authentication is desired. The client will send a string to indicate it will use Kerberos 5. Let us say “length-string” of strl, and mean the couple

  (length of the string strl, null-terminated string strl).

when strl itself is given. The string length is encoded as an int32 (32-bit integer) in MSB order, i.e., network byte order. So the client first sends an authentication message, the length-string of

   KRB5_SENDAUTH_V1.0

After that, the client must tell which version of the protocol it uses by sending a version message, consisting of a second length-string. This time based on

   KCMDV0.2

or on “KCMDV0.1”, for the older version.

If the client’s indentification is acceptable, the server will respond with a null byte (0x00). Otherwise, if the authentication message was incorrect, then the server responds with the single byte 0x01, while if the protocol version message was unacceptable, then the response is a single 0x02.

B.3.3 Kerberos authentication

When the client is indentified, Kerberos authentication can begin. The client must send an AP-REQ to the server; an AP-REQ authenticator must contain a subkey (only for KCMDV0.2) and a checksum. The authenticator checksum is calculated from the following string.

     "serverport":"terminaltype""remoteusername"

Observe the mandatory colon, serving as a delimiter to the terminal type. Here remoteusername is whatever identity the client desires to use at the remote end. The simple example

     543:linux/38400user

demonstrates that the terminal type expects a terminfo name and a speed as decimal number.

Once the AP-REQ has been updated with the checksum, it is ready for transmission to the server. Its length (as int32) is first sent in network order (MSB), followed by the DER encoded AP-REQ itself.

If all is acceptable, the server reponds with an int32 of value null. (In MSB order, but as it is null, order is irrelevant!). The KCMD protocol uses mutual authentication, so the server must also send an AP-REP back:

     (int32-length in MSB of DER encoded AP-REP)
     (DER encoded AP-REP)

Now server and client are provisionally authenticated.

B.3.4 Extended authentication

The client next sends three different null terminated strings, without length : • remote user name (user identity on remote machine)

• terminal type for rlogin or command for rsh

• local user name (user identity on client machine)

An example for rsh :

     "rname\0"
     "cat /usr/local/etc/shishi.conf"
     "lname\0"

The server must next verify that the checksum delivered in the AP-REQ authenticator is correct, by computing a new hash like the client has previously done.

The server must also verify that the requesting principal (found in AP-REQ) has the right to log in to the remote user account. In the ‘basic’ authorization, this is done by checking whether the remote user name is identical to the principal’s name. The alternative is ‘k5login’ mode, which is discussed in [kerberos authorization], page 246.

If all is correct, the server sends a null byte, i.e., an empty string, or else an error message string (null terminated). The remote client reads the first byte. If it is equal to zero, authentication is successful and the user is logged in at the remote host. In any other case, the client has the error message available, as sent by the server, to help the user understand the failure.

B.3.5 Window size

In the rlogin protocol, when authentication completes, the server can optionally send a message asking for terminal window size suitable for the user. The client can respond, but it is not obliged to do so.

Under KCMDV0.1, the server sends a single byte, an urgent TCP message (MSG_OOB) :

     TIOCPKT_WINDOW = 0x80

With KCMDV0.2, the server does not send an urgent message at all, but writes five bytes to the socket :

     '\377', '\377', 'o', 'o', TIOCPKT_WINDOW

If encryption is enabled (‘eklogin’) the server must send those five bytes in encrypted form.

The client answers alike in both protocol versions :

     '\377', '\377', 's', 's', "struct winsize"

The winsize structure is filled in with the settings from the client’s terminal. With encryption enabled, this answer must be sent encrypted.

B.3.6 End of authentication

The legacy exchange supported by rsh and rlogin continues from this point onwards.

B.3.7 Encryption

Encryption mode is always used when connecting via the port ‘eklogin’. The port ‘kshell’ also supports encryption. Previously, there was a specific port ‘ekshell’ for that purpose, a use which is now extinct. Instead, whenever an encrypted exchange is desired via the port ‘kshell’, client must prefix the string “-x ” to the command string, when it is sent inbetween the remote user name and the local user name. In contrast, when the client computes the checksum for the AP-REQ authenticator, the string “-x ” must not be included.

Encryption porcedure under version KCMDV0.2 differs from that in the older protocol version. Under version KCMDV0.1, the ticket session key is put to use as encryption key, and only standard Kerberos encryption functions are usable. Thus it supports only ‘des-cbc-crc’, ‘des-cbc-md4’, ‘des-cbc-md5’, and does not allow use of initialisation vectors.

As an example of encryption/decryption calls, the following Kerberos function prototype should be used :

     kerberos_encrypt (key, keyusage, in, out)  (or kerberos_decrypt)

To contrast, KCMDV0.2 can be used with all Kerberos encryption modes, i.e., ‘des’, ‘3des’, ‘aes’, ‘arcfour’, and it uses an AP-REQ authenticator subkey. In opposition to KCMDV0.1, initialisation vectors are used. All encryption and descryption must be made using a cryptographic context. A typical coding example updates the context with an iv, then executes an encryption call :

     kerberos_init(ctx, iv, key, keyusage)
     kerberos_encrypt (ctx, in, out)

For both protocols, default keyusage identity allowing ‘des-cbc-md5’, ‘des-cbc-md4’, ‘des-cbc-crc’, and ‘des3-cbc-sha1’ (the latter applicable only to KCMDV0.2) is identical :

     keyusage = 1026

KCMDV0.2 encryption modes, other than the four named above, specify distinct values for keyusage, unique to each encryption/decryption mode.

In conclusion, ‘eklogin’ uses a single socket. It encrypts data “output 1” prior to sending, and it decrypts “input 1” received data.

‘kshell’ uses two sockets: one for transmitting data, and one for a side channel carrying stderr. Thus there are four streams available :

   transmit  : input  1
               output 1
   stderr    : input  2
               output 2

There is a key usage set for each mode. The values of each keyusage must be compatible between client and server side.

An example with ‘klogin’, shows the client’s “input 1” key usage to be identical to the server’s “output 1” usage.

I/O  Client  Server
input 1  1028  1030
output 1  1030  1028
input 2  1032  1034
output 2  1034  1032

The stated key usages are for AES and ARCFOUR modes.

KCMDV0.2 uses an IV (initialisation vector) with AES. Like for key usage, the client IV must correspond to the server IV. The IV size is equal to the blocksize of the chosen key type. All bytes of IV must be initialised too :

I/O  Client  Server
input 1  0  1
output 1  1  0
input 2  2  3
output 2  3  2

ARCFOUR mode does not use an IV. However, like mentioned before, a context must be used to keep the updating the sbox.

A normal message for ‘klogin’ or ‘kshell’ is set up like this :

     (int32-length of message in MSB order)
     (message)

In encrypted mode, the format is only slightly different :

     (int32-length of unencrypted message in MSB order)
     (encrypted message)

Under KCMDV0.2, the encrypted message is create like this :

     encrypt (

     (int32-length of message in MSB order)
     (message)
     )

A check on message size can be made in second version of the protocol.

B.3.8 KCMDV0.3

This part only gives possible ways to extend the KCMD protocol. They should not be understood as some kind of “must have” in a future KCMD implementation.

Extensions to KCMV0.2 can be imagined. For example, ‘kshell’ assumes there are no files of names like “-x *”. I think the same thing can be assumed of terminal names for ‘klogin’. So the client could add “-x ” to the terminal type it transmits to the server, in order to indicate the desire to use encryption.

Under this provision, there need only be one port shared by ‘klogin’ and ‘eklogin’, namely the IANA defined port number 543.

Before encryption begins, ‘kshell’ passes the intended commands in the clear through the network. This could be considered insecure, once the user has decided to use encryption. It is not really a problem with ‘klogin’, because it just tells which terminal type to target.

One could imagine, when the client intends to use encrypted mode, that the client side initially transmits a mundane “-x” and nothing else to either of ‘klogin’ and ‘kshell’, in place of either a command or a terminal type, respectively. Once encryption is in place, the client could send terminal type or command in a second, now encrypted exchange. The server could respond with a single null byte, saying that all is well, or respond with an error message, which already enjoys the added benefit of being encrypted.

B.3.9 MIT/Heimdal authorization

This short part describes how MIT/Heimdal versions of Kerberos 5 check authorization of any user seeking to log in to a remote machine.

The authorization stage begins by testing whether the file .k5login exists in the home directory of the remote user. If this file does not exist, then a valid authorization demands that the remote user’s name must be the same as the name of the principal contained in the request AP-REQ. (This is the legacy ‘basic’ authorization.) Else, if the file is present, the serve first verifies that the remote user, or root, is the owner of .k5login, then goes on to verify that the file is not readable by group, nor by world. If either fails, then the check fails entirely.

If all is good so far, then each line of that file is examined, and each name read in is compared to the principal. If the principal is found listed somewhere in .k5login, then authorization is successful. In the contrary case, the user is not admitted to the remote host as the requested remote user, a name that could have been derived to be the very same as in the principal itself.

So someone, “user1” say, can remotely log in to the account of “user2”, if the file .k5login is present in the home directory of “user2”, and it is owned by “user2” or by root, and at the same time the name “user1” is listed in this file.

B.4 Key as initialization vector

The des-cbc-crc algorithm (see Section 1.4 [Cryptographic Overview], page 5) uses the DES key as the initialization vector. This is problematic in general (see below^{1}), but may be mitigated in Kerberos by the CRC checksum that is also included.

From daw@espresso.CS.Berkeley.EDU Fri Mar  1 13:32:34 PST 1996
Article: 50440 of sci.crypt
Path: agate!daw
From: daw@espresso.CS.Berkeley.EDU (David A Wagner)
Newsgroups: sci.crypt
Subject: Re: DES-CBC and Initialization Vectors
Date: 29 Feb 1996 21:48:16 GMT
Organization: University of California, Berkeley
Lines: 31
Message-ID: <4h56v0$3no@agate.berkeley.edu>
References: <4h39li$33o@gaia.ns.utk.edu>
NNTP-Posting-Host: espresso.cs.berkeley.edu
In article <4h39li$33o@gaia.ns.utk.edu>,
Nair Venugopal <venu@mars.utcc.utk.edu> wrote:
> Is there anything wrong in using the key as the I.V. in DES-CBC mode?
Yes, you're open to a chosen-ciphertext attack which recovers the key.
Alice is sending stuff DES-CBC encrypted with key K to Bob.  Mary is an
active adversary in the middle.  Suppose Alice encrypts some plaintext
blocks P_1, P_2, P_3, ... in DES-CBC mode with K as the IV, and sends off
the resulting ciphertext
        A->B: C_1, C_2, C_3, ...
where each C_j is a 8-byte DES ciphertext block.  Mary wants to discover
the key K, but doesn't even know any of the P_j's.  She replaces the above
message by
        M->B: C_1, 0, C_1
where 0 is the 8-byte all-zeros block.  Bob will decrypt under DES-CBC,
recovering the blocks
        Q_1, Q_2, Q_3
where
        Q_1 = DES-decrypt(K, C_1) xor K = P_1
        Q_2 = DES-decrypt(K, C_2) xor C_1 = (some unimportant junk)
        Q_3 = DES-decrypt(K, C_1) xor 0 = P_1 xor K
Bob gets this garbage-looking message Q_1,Q_2,Q_3 which Mary recovers
(under the chosen-ciphertext assumption: this is like a known-plaintext
attack, which isn't too implausible).  Notice that Mary can recover K by
        K = Q_1 xor Q_3;
so after this one simple active attack, Mary gets the key back!

1 The post is copyrighted by David Wagner, included here with permission, the canonical location is

http://www.cs.berkeley.edu/~daw/my-posts/key-as-iv-broken

So, if you must use a fixed IV, don't use the key-- use 0 or something
like that.  Even better, don't use a fixed IV-- use the DES encryption
of a counter, or something like that.

B.5 The Keytab Binary File Format

The keytab file format is described in the file keytab.txt, here included verbatim.

The Kerberos Keytab Binary File Format
Copyright (C) 2006 Michael B Allen <mba2000 ioplex.com>
http://www.ioplex.com/utilities/keytab.txt
Last updated: Fri May  5 13:39:40 EDT 2006
The MIT keytab binary format is not a standard format, nor is it
documented anywhere in detail. The format has evolved and may continue
to. It is however understood by several Kerberos implementations including
Heimdal and of course MIT and keytab files are created by the ktpass.exe
utility from Windows. So it has established itself as the defacto format
for storing Kerberos keys.
The following C-like structure definitions illustrate the MIT keytab
file format. All values are in network byte order. All text is ASCII.
  keytab {
      uint16_t file_format_version;                    /* 0x502 */
      keytab_entry entries[*];
  };
  keytab_entry {
      int32_t size;
      uint16_t num_components;    /* sub 1 if version 0x501 */
      counted_octet_string realm;
      counted_octet_string components[num_components];
      uint32_t name_type;   /* not present if version 0x501 */
      uint32_t timestamp;
      uint8_t vno8;
      keyblock key;
      uint32_t vno; /* only present if >= 4 bytes left in entry */
  };
  counted_octet_string {
      uint16_t length;
      uint8_t data[length];
  };
  keyblock {
      uint16_t type;

      counted_octet_string;
  };
The keytab file format begins with the 16 bit file_format_version which
at the time this document was authored is 0x502. The format of older
keytabs is described at the end of this document.
The file_format_version is immediately followed by an array of
keytab_entry structures which are prefixed with a 32 bit size indicating
the number of bytes that follow in the entry. Note that the size should be
evaluated as signed. This is because a negative value indicates that the
entry is in fact empty (e.g. it has been deleted) and that the negative
value of that negative value (which is of course a positive value) is
the offset to the next keytab_entry. Based on these size values alone
the entire keytab file can be traversed.
The size is followed by a 16 bit num_components field indicating the
number of counted_octet_string components in the components array.
The num_components field is followed by a counted_octet_string
representing the realm of the principal.
A counted_octet_string is simply an array of bytes prefixed with a 16
bit length. For the realm and name components, the counted_octet_string
bytes are ASCII encoded text with no zero terminator.
Following the realm is the components array that represents the name of
the principal. The text of these components may be joined with slashs
to construct the typical SPN representation. For example, the service
principal HTTP/www.foo.net@FOO.NET would consist of name components
"HTTP" followed by "www.foo.net".
Following the components array is the 32 bit name_type (e.g. 1 is
KRB5_NT_PRINCIPAL, 2 is KRB5_NT_SRV_INST, 5 is KRB5_NT_UID, etc). In
practice the name_type is almost certainly 1 meaning KRB5_NT_PRINCIPAL.
The 32 bit timestamp indicates the time the key was established for that
principal. The value represents the number of seconds since Jan 1, 1970.
The 8 bit vno8 field is the version number of the key. This value is
overridden by the 32 bit vno field if it is present.
The keyblock structure consists of a 16 bit value indicating the keytype
(e.g. 3 is des-cbc-md5, 23 is arcfour-hmac-md5, 16 is des3-cbc-sha1,
etc). This is followed by a counted_octet_string containing the key.
The last field of the keytab_entry structure is optional. If the size of

the keytab_entry indicates that there are at least 4 bytes remaining,
a 32 bit value representing the key version number is present. This
value supersedes the 8 bit vno8 value preceeding the keyblock.
Older keytabs with a file_format_version of 0x501 are different in
three ways:
  1) All integers are in host byte order [1].
  2) The num_components field is 1 too large (i.e. after decoding,
     decrement by 1).
  3) The 32 bit name_type field is not present.
[1] The file_format_version field should really be treated as two
    separate 8 bit quantities representing the major and minor version
    number respectively.
Permission to copy, modify, and distribute this document, with or
without modification, for any purpose and without fee or royalty is
hereby granted, provided that you include this copyright notice in ALL
copies of the document or portions thereof, including modifications.

B.6 The Credential Cache Binary File Format

The credential cache file format is described in the file keytab.txt, included verbatim.

The Kerberos Credential Cache Binary File Format
Copyright (C) 2006-2022 Simon Josefsson <simon josefsson.org>
http://josefsson.org/shishi/ccache.txt
Last updated: Sat Sep 23 12:04:11 CEST 2006
Like the MIT keytab binary format (see Michael B Allen's reverse
engineered description in keytab.txt), the credential cache format is
not standard nor documented anywhere.
In C style notation, the MIT credential cache file format is as
follows.  All values are in network byte order.  All text is ASCII.
ccache {
          uint16_t file_format_version; /* 0x0504 */
          uint16_t headerlen;           /* only if version is 0x0504 */
          header headers[];             /* only if version is 0x0504 */
          principal primary_principal;
          credential credentials[*];
};
header {
       uint16_t tag;                    /* 1 = DeltaTime */
       uint16_t taglen;

       uint8_t tagdata[taglen]
};
The ccache.taglen and ccache.tags fields are only present in 0x0504
versions, not in earlier.  Both MIT and Heimdal appear to correctly
ignore unknown tags, so it appears safe to add them (although there is
no central place to "register" tags).
Currently only one tag is widely implemented, DeltaTime (0x0001).  Its
taglen is always 8, and tagdata will contain:
DeltaTime {
       uint32_t time_offset;
       uint32_t usec_offset;
};
After reading the file_format_version, header tags, and default
principal, a list of credentials follow.  You deduce from the file
length when there are no more credentials.
credential {
           principal client;
           principal server;
           keyblock key;
           times    time;
           uint8_t  is_skey;            /* 1 if skey, 0 otherwise */
           uint32_t tktflags;           /* stored in reversed byte order */
           uint32_t num_address;
           address  addrs[num_address];
           uint32_t num_authdata;
           authdata authdata[num_authdata];
           countet_octet_string ticket;
           countet_octet_string second_ticket;
};
keyblock {
         uint16_t keytype;
         uint16_t etype;                /* only present if version 0x0503 */
         uint16_t keylen;
         uint8_t keyvalue[keylen];
};
times {
      uint32_t  authtime;
      uint32_t  starttime;
      uint32_t  endtime;
      uint32_t  renew_till;

};
address {
        uint16_t addrtype;
        counted_octet_string addrdata;
};
authdata {
         uint16_t authtype;
         counted_octet_string authdata;
};
principal {
          uint32_t name_type;           /* not present if version 0x0501 */
          uint32_t num_components;      /* sub 1 if version 0x501 */
          counted_octet_string realm;
          counted_octet_string components[num_components];
};
counted_octet_string {
    uint32_t length;
    uint8_t data[length];
};
Permission to copy, modify, and distribute this document, with or
without modification, for any purpose and without fee or royalty is
hereby granted, provided that you include this copyright notice in ALL
copies of the document or portions thereof, including modifications.

Appendix C Copying Information

C.1 GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 https://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text. The “publisher” means any person or entity that distributes copies of the Document to the public. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING

You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3.

You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY

If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects.

If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages.

If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public.

It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS

You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version:

A. Use in the Title Page (and on the covers, if any) a title distinct from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See https://www.gnu.org/licenses/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Function and Data Index

A

ago in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 am in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

D

day in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

F

first in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 fortnight in date strings . . . . . . . . . . . . . . . . . . . . . . 53

H

hour in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

L

last day. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 last in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

M

midnight in date strings . . . . . . . . . . . . . . . . . . . . . . . 52 minute in date strings. . . . . . . . . . . . . . . . . . . . . . . . . . 53 month in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

N

next day. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 next in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 noon in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 now in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

P

parse_datetime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 pm in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

S

shisa. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 shisa_cfg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 shisa_cfg_db . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227

shisa_cfg_default_systemfile . . . . . . . . . . . . . . 228

shisa_cfg_from_file . . . . . . . . . . . . . . . . . . . . . . . . 228 shisa_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226

shisa_enumerate_principals . . . . . . . . . . . . . . . . 228

shisa_enumerate_realms . . . . . . . . . . . . . . . . . . . . . 228 shisa_info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 shisa_init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 shisa_init_with_paths . . . . . . . . . . . . . . . . . . . . . . 227 shisa_key_add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 shisa_key_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 shisa_key_remove . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 shisa_key_update . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 shisa_keys_find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 shisa_keys_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 shisa_principal_add . . . . . . . . . . . . . . . . . . . . . . . . 230 shisa_principal_find . . . . . . . . . . . . . . . . . . . . . . . 229 shisa_principal_remove . . . . . . . . . . . . . . . . . . . . . 230 shisa_principal_update . . . . . . . . . . . . . . . . . . . . . 229 shisa_strerror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 shishi. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 shishi_3des. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 shishi_aes_cts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 shishi_ap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

shishi_ap_authenticator . . . . . . . . . . . . . . . . . . . . . 77
shishi_ap_authenticator_cksumdata . . . . . . . . . 75
shishi_ap_authenticator_cksumdata_set . . . . . 76
shishi_ap_authenticator_cksumraw_set . . . . . . 76
shishi_ap_authenticator_cksumtype . . . . . . . . . 76
shishi_ap_authenticator_cksumtype_set . . . . . 77
shishi_ap_authenticator_set . . . . . . . . . . . . . . . . 77

shishi_ap_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 shishi_ap_encapreppart . . . . . . . . . . . . . . . . . . . . . . 81

shishi_ap_encapreppart_set. . . . . . . . . . . . . . . . . . 81

shishi_ap_etype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

shishi_ap_etype_tktoptionsdata . . . . . . . . . . . . . 74

shishi_ap_key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 shishi_ap_nosubkey . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

shishi_ap_option2string . . . . . . . . . . . . . . . . . . . . . 81

shishi_ap_rep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 shishi_ap_rep_asn1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 shishi_ap_rep_build. . . . . . . . . . . . . . . . . . . . . . . . . . 80 shishi_ap_rep_der . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 shishi_ap_rep_der_set . . . . . . . . . . . . . . . . . . . . . . . 80 shishi_ap_rep_set . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 shishi_ap_rep_verify . . . . . . . . . . . . . . . . . . . . . . . . 80

shishi_ap_rep_verify_asn1. . . . . . . . . . . . . . . . . . . 81
shishi_ap_rep_verify_der . . . . . . . . . . . . . . . . . . . . 80

Function and Data Index 262

shishi_ap_req . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 shishi_ap_req_asn1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 shishi_ap_req_build. . . . . . . . . . . . . . . . . . . . . . . . . . 78 shishi_ap_req_decode . . . . . . . . . . . . . . . . . . . . . . . . 78 shishi_ap_req_der . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 shishi_ap_req_der_set . . . . . . . . . . . . . . . . . . . . . . . 78 shishi_ap_req_process . . . . . . . . . . . . . . . . . . . . . . . 78

shishi_ap_req_process_keyusage . . . . . . . . . . . . . 78

shishi_ap_req_set . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

shishi_ap_set_tktoptions . . . . . . . . . . . . . . . . . . . . 72
shishi_ap_set_tktoptionsasn1usage . . . . . . . . . 73
shishi_ap_set_tktoptionsdata . . . . . . . . . . . . . . . 72
shishi_ap_set_tktoptionsraw . . . . . . . . . . . . . . . . 73
shishi_ap_string2option . . . . . . . . . . . . . . . . . . . . . 81

shishi_ap_tkt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 shishi_ap_tkt_set . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 shishi_ap_tktoptions . . . . . . . . . . . . . . . . . . . . . . . . 73

shishi_ap_tktoptionsasn1usage . . . . . . . . . . . . . . 75
shishi_ap_tktoptionsdata . . . . . . . . . . . . . . . . . . . . 74
shishi_ap_tktoptionsraw . . . . . . . . . . . . . . . . . . . . . 74

shishi_aprep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 shishi_aprep_from_file . . . . . . . . . . . . . . . . . . . . . . 87

shishi_aprep_get_enc_part_etype . . . . . . . . . . . . 87

shishi_aprep_parse . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 shishi_aprep_print . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 shishi_aprep_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 shishi_aprep_save . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 shishi_aprep_to_file . . . . . . . . . . . . . . . . . . . . . . . . 86 shishi_apreq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

shishi_apreq_add_authenticator . . . . . . . . . . . . . 84

shishi_apreq_from_file . . . . . . . . . . . . . . . . . . . . . . 83

shishi_apreq_get_authenticator_etype . . . . . . 85
shishi_apreq_get_ticket . . . . . . . . . . . . . . . . . . . . . 86
shishi_apreq_mutual_required_p . . . . . . . . . . . . . 85

shishi_apreq_options . . . . . . . . . . . . . . . . . . . . . . . . 84

shishi_apreq_options_add . . . . . . . . . . . . . . . . . . . . 85
shishi_apreq_options_remove . . . . . . . . . . . . . . . . 85
shishi_apreq_options_set . . . . . . . . . . . . . . . . . . . . 85

shishi_apreq_parse . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 shishi_apreq_print . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 shishi_apreq_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 shishi_apreq_save . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

shishi_apreq_set_authenticator . . . . . . . . . . . . . 83
shishi_apreq_set_ticket . . . . . . . . . . . . . . . . . . . . . 84

shishi_apreq_to_file . . . . . . . . . . . . . . . . . . . . . . . . 82

shishi_apreq_use_session_key_p . . . . . . . . . . . . . 84

shishi_arcfour . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 shishi_as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 shishi_as_check_cname . . . . . . . . . . . . . . . . . . . . . . 133 shishi_as_check_crealm . . . . . . . . . . . . . . . . . . . . . 132 shishi_as_derive_salt . . . . . . . . . . . . . . . . . . . . . . 131 shishi_as_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 shishi_as_krberror. . . . . . . . . . . . . . . . . . . . . . . . . . 116 shishi_as_krberror_der . . . . . . . . . . . . . . . . . . . . . 116 shishi_as_krberror_set . . . . . . . . . . . . . . . . . . . . . 117 shishi_as_process . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 shishi_as_rep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 shishi_as_rep_build . . . . . . . . . . . . . . . . . . . . . . . . 115 shishi_as_rep_der . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 shishi_as_rep_der_set . . . . . . . . . . . . . . . . . . . . . . 116 shishi_as_rep_process . . . . . . . . . . . . . . . . . . . . . . 115 shishi_as_rep_set . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 shishi_as_req . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 shishi_as_req_build . . . . . . . . . . . . . . . . . . . . . . . . 114 shishi_as_req_der . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 shishi_as_req_der_set . . . . . . . . . . . . . . . . . . . . . . 115 shishi_as_req_set . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 shishi_as_sendrecv. . . . . . . . . . . . . . . . . . . . . . . . . . 117

                                                shishi_as_sendrecv_hint . . . . . . . . . . . . . . . . . . . . 117

shishi_as_tkt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 shishi_as_tkt_set . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 shishi_asn1_aprep . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 shishi_asn1_apreq . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 shishi_asn1_asrep . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 shishi_asn1_asreq . . . . . . . . . . . . . . . . . . . . . . . . . . . 210

                                                shishi_asn1_authenticator. . . . . . . . . . . . . . . . . . 211

shishi_asn1_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209

                                                shishi_asn1_encapreppart. . . . . . . . . . . . . . . . . . . 211
                                                shishi_asn1_encasreppart. . . . . . . . . . . . . . . . . . . 212
                                                shishi_asn1_enckdcreppart. . . . . . . . . . . . . . . . . . 211
                                                shishi_asn1_encprivpart . . . . . . . . . . . . . . . . . . . . 212
                                                shishi_asn1_encrypteddata. . . . . . . . . . . . . . . . . . 209
                                                shishi_asn1_encticketpart. . . . . . . . . . . . . . . . . . 211

shishi_asn1_etype_info . . . . . . . . . . . . . . . . . . . . . 210

                                                shishi_asn1_etype_info2 . . . . . . . . . . . . . . . . . . . . 210

shishi_asn1_krberror . . . . . . . . . . . . . . . . . . . . . . . 212 shishi_asn1_krbsafe . . . . . . . . . . . . . . . . . . . . . . . . 212 shishi_asn1_methoddata . . . . . . . . . . . . . . . . . . . . . 210 shishi_asn1_msgtype . . . . . . . . . . . . . . . . . . . . . . . . 213

                                                shishi_asn1_pa_enc_ts_enc. . . . . . . . . . . . . . . . . . 209

shishi_asn1_padata. . . . . . . . . . . . . . . . . . . . . . . . . . 209 shishi_asn1_print . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 shishi_asn1_priv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 shishi_asn1_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208

                                                shishi_asn1_read_inline . . . . . . . . . . . . . . . . . . . . 208
                                                shishi_asn1_read_optional. . . . . . . . . . . . . . . . . . 209

shishi_asn1_tgsrep. . . . . . . . . . . . . . . . . . . . . . . . . . 210 shishi_asn1_tgsreq. . . . . . . . . . . . . . . . . . . . . . . . . . 210 shishi_asn1_ticket. . . . . . . . . . . . . . . . . . . . . . . . . . 211 shishi_asn1_to_der. . . . . . . . . . . . . . . . . . . . . . . . . . 213

                                                shishi_asn1_to_der_field. . . . . . . . . . . . . . . . . . . 212

shishi_asrep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 shishi_asreq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

                                                shishi_asreq_clientrealm. . . . . . . . . . . . . . . . . . . 137

shishi_authenticator . . . . . . . . . . . . . . . . . . . . . . . 152

                                                shishi_authenticator_add_

authorizationdata . . . . . . . . . . . . . . . . . . . . . . . . 159

                                                shishi_authenticator_add_cksum . . . . . . . . . . . . 158
                                                shishi_authenticator_add_cksum_type . . . . . . 158
                                                shishi_authenticator_add_random_subkey. . . 160
                                                shishi_authenticator_add_

random_subkey_etype . . . . . . . . . . . . . . . . . . . . . . 160

                                                shishi_authenticator_add_subkey. . . . . . . . . . . 161
                                                shishi_authenticator_authorizationdata. . . 159
                                                shishi_authenticator_cksum . . . . . . . . . . . . . . . . 157
                                                shishi_authenticator_clear_

authorizationdata . . . . . . . . . . . . . . . . . . . . . . . . 159

                                                shishi_authenticator_client . . . . . . . . . . . . . . . 157
                                                shishi_authenticator_client_set. . . . . . . . . . . 155

shishi_authenticator_clientrealm . . . . . . . . . 157
shishi_authenticator_ctime . . . . . . . . . . . . . . . . 155
shishi_authenticator_ctime_set . . . . . . . . . . . . 155
shishi_authenticator_cusec_get . . . . . . . . . . . . 156
shishi_authenticator_cusec_set . . . . . . . . . . . . 156
shishi_authenticator_from_file . . . . . . . . . . . . 154
shishi_authenticator_get_subkey. . . . . . . . . . . 160
shishi_authenticator_parse . . . . . . . . . . . . . . . . 154
shishi_authenticator_print . . . . . . . . . . . . . . . . 153
shishi_authenticator_read. . . . . . . . . . . . . . . . . . 154
shishi_authenticator_remove_subkey . . . . . . . 159
shishi_authenticator_save. . . . . . . . . . . . . . . . . . 153
shishi_authenticator_seqnumber_get . . . . . . . 156
shishi_authenticator_seqnumber_remove . . . . 156
shishi_authenticator_seqnumber_set . . . . . . . 156
shishi_authenticator_set_cksum . . . . . . . . . . . . 158
shishi_authenticator_set_cname . . . . . . . . . . . . 155
shishi_authenticator_set_crealm. . . . . . . . . . . 154
shishi_authenticator_set_subkey. . . . . . . . . . . 160
shishi_authenticator_subkey . . . . . . . . . . . . . . . 153
shishi_authenticator_to_file . . . . . . . . . . . . . . 153
shishi_authorization_parse . . . . . . . . . . . . . . . . 205
shishi_authorize_k5login. . . . . . . . . . . . . . . . . . . 205
shishi_authorize_strcmp . . . . . . . . . . . . . . . . . . . . 204

shishi_authorized_p . . . . . . . . . . . . . . . . . . . . . . . . 205 shishi_cfg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

shishi_cfg_authorizationtype_set. . . . . . . . . . . 64
shishi_cfg_clientkdcetype. . . . . . . . . . . . . . . . . . . 63
shishi_cfg_clientkdcetype_fast . . . . . . . . . . . . . 63
shishi_cfg_clientkdcetype_set . . . . . . . . . . . . . . 64
shishi_cfg_default_systemfile . . . . . . . . . . . . . . 62
shishi_cfg_default_userdirectory. . . . . . . . . . . 62
shishi_cfg_default_userfile . . . . . . . . . . . . . . . . 63

shishi_cfg_from_file . . . . . . . . . . . . . . . . . . . . . . . . 62 shishi_cfg_print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

shishi_cfg_userdirectory_file . . . . . . . . . . . . . . 63

shishi_check_version . . . . . . . . . . . . . . . . . . . . . . . . 57 shishi_checksum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184

shishi_checksum_cksumlen. . . . . . . . . . . . . . . . . . . 183

shishi_checksum_name . . . . . . . . . . . . . . . . . . . . . . . 182 shishi_checksum_parse . . . . . . . . . . . . . . . . . . . . . . 183

shishi_checksum_supported_p . . . . . . . . . . . . . . . 182
shishi_cipher_blocksize . . . . . . . . . . . . . . . . . . . . 181
shishi_cipher_confoundersize . . . . . . . . . . . . . . 181
shishi_cipher_defaultcksumtype . . . . . . . . . . . . 182

shishi_cipher_keylen . . . . . . . . . . . . . . . . . . . . . . . 182 shishi_cipher_name. . . . . . . . . . . . . . . . . . . . . . . . . . 181 shishi_cipher_parse . . . . . . . . . . . . . . . . . . . . . . . . 182

shishi_cipher_randomlen . . . . . . . . . . . . . . . . . . . . 182
shishi_cipher_supported_p. . . . . . . . . . . . . . . . . . 181

shishi_crc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 shishi_crypto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 shishi_crypto_close . . . . . . . . . . . . . . . . . . . . . . . . 193 shishi_crypto_decrypt . . . . . . . . . . . . . . . . . . . . . . 193 shishi_crypto_encrypt . . . . . . . . . . . . . . . . . . . . . . 193 shishi_ctime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 shishi_decrypt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 shishi_decrypt_etype . . . . . . . . . . . . . . . . . . . . . . . 189 shishi_decrypt_iv . . . . . . . . . . . . . . . . . . . . . . . . . . . 190

                                                shishi_decrypt_iv_etype . . . . . . . . . . . . . . . . . . . . 188
                                                shishi_decrypt_ivupdate . . . . . . . . . . . . . . . . . . . . 190
                                                shishi_decrypt_ivupdate_etype . . . . . . . . . . . . . 188

shishi_der_msgtype. . . . . . . . . . . . . . . . . . . . . . . . . . 213 shishi_der2asn1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 shishi_der2asn1_aprep . . . . . . . . . . . . . . . . . . . . . . 218 shishi_der2asn1_apreq . . . . . . . . . . . . . . . . . . . . . . 218 shishi_der2asn1_asrep . . . . . . . . . . . . . . . . . . . . . . 215 shishi_der2asn1_asreq . . . . . . . . . . . . . . . . . . . . . . 215

                                                shishi_der2asn1_authenticator . . . . . . . . . . . . . 217
                                                shishi_der2asn1_encapreppart . . . . . . . . . . . . . . 218
                                                shishi_der2asn1_encasreppart . . . . . . . . . . . . . . 216
                                                shishi_der2asn1_enckdcreppart . . . . . . . . . . . . . 216
                                                shishi_der2asn1_encprivpart . . . . . . . . . . . . . . . 218
                                                shishi_der2asn1_enctgsreppart . . . . . . . . . . . . . 216
                                                shishi_der2asn1_encticketpart . . . . . . . . . . . . . 215
                                                shishi_der2asn1_etype_info . . . . . . . . . . . . . . . . 214
                                                shishi_der2asn1_etype_info2 . . . . . . . . . . . . . . . 214

shishi_der2asn1_kdcrep . . . . . . . . . . . . . . . . . . . . . 216 shishi_der2asn1_kdcreq . . . . . . . . . . . . . . . . . . . . . 218

                                                shishi_der2asn1_krberror. . . . . . . . . . . . . . . . . . . 217
                                                shishi_der2asn1_krbsafe . . . . . . . . . . . . . . . . . . . . 217
                                                shishi_der2asn1_methoddata . . . . . . . . . . . . . . . . 214

shishi_der2asn1_padata . . . . . . . . . . . . . . . . . . . . . 214 shishi_der2asn1_priv . . . . . . . . . . . . . . . . . . . . . . . 217 shishi_der2asn1_tgsrep . . . . . . . . . . . . . . . . . . . . . 216 shishi_der2asn1_tgsreq . . . . . . . . . . . . . . . . . . . . . 215 shishi_der2asn1_ticket . . . . . . . . . . . . . . . . . . . . . 214

                                                shishi_derive_default_salt . . . . . . . . . . . . . . . . 204

shishi_des . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 shishi_des_cbc_mac. . . . . . . . . . . . . . . . . . . . . . . . . . 195 shishi_dk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 shishi_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 shishi_dr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 shishi_encapreppart. . . . . . . . . . . . . . . . . . . . . . . . . . 88

                                                shishi_encapreppart_ctime. . . . . . . . . . . . . . . . . . . 89
                                                shishi_encapreppart_ctime_set . . . . . . . . . . . . . . 90
                                                shishi_encapreppart_cusec_get . . . . . . . . . . . . . . 90
                                                shishi_encapreppart_cusec_set . . . . . . . . . . . . . . 90
                                                shishi_encapreppart_from_file . . . . . . . . . . . . . . 89
                                                shishi_encapreppart_get_key . . . . . . . . . . . . . . . . 89
                                                shishi_encapreppart_parse. . . . . . . . . . . . . . . . . . . 88
                                                shishi_encapreppart_print. . . . . . . . . . . . . . . . . . . 88
                                                shishi_encapreppart_read . . . . . . . . . . . . . . . . . . . . 89
                                                shishi_encapreppart_save . . . . . . . . . . . . . . . . . . . . 88
                                                shishi_encapreppart_seqnumber_get . . . . . . . . . 90
                                                shishi_encapreppart_seqnumber_remove . . . . . . 90
                                                shishi_encapreppart_seqnumber_set . . . . . . . . . 91
                                                shishi_encapreppart_time_copy . . . . . . . . . . . . . . 91
                                                shishi_encapreppart_to_file . . . . . . . . . . . . . . . . 88
                                                shishi_enckdcreppart_authtime_set . . . . . . . . 150
                                                shishi_enckdcreppart_endtime_set . . . . . . . . . 151
                                                shishi_enckdcreppart_flags_set . . . . . . . . . . . . 150
                                                shishi_enckdcreppart_get_key . . . . . . . . . . . . . . 150
                                                shishi_enckdcreppart_key_set . . . . . . . . . . . . . . 150
                                                shishi_enckdcreppart_nonce_set . . . . . . . . . . . . 150

shishi_enckdcreppart_populate_

encticketpart . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152

shishi_enckdcreppart_renew_till_set . . . . . . 151
shishi_enckdcreppart_sname_set . . . . . . . . . . . . 152
shishi_enckdcreppart_srealm_set. . . . . . . . . . . 151
shishi_enckdcreppart_starttime_set . . . . . . . 151
shishi_encprivpart_set_user_data . . . . . . . . . 101
shishi_encprivpart_user_data . . . . . . . . . . . . . . 101

shishi_encrypt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 shishi_encrypt_etype . . . . . . . . . . . . . . . . . . . . . . . 186 shishi_encrypt_iv . . . . . . . . . . . . . . . . . . . . . . . . . . . 187

shishi_encrypt_iv_etype . . . . . . . . . . . . . . . . . . . . 185
shishi_encrypt_ivupdate . . . . . . . . . . . . . . . . . . . . 186
shishi_encrypt_ivupdate_etype . . . . . . . . . . . . . 184
shishi_encticketpart_authtime_set . . . . . . . . 128
shishi_encticketpart_client . . . . . . . . . . . . . . . 128
shishi_encticketpart_clientrealm . . . . . . . . . 129
shishi_encticketpart_cname_set . . . . . . . . . . . . 127
shishi_encticketpart_crealm_set. . . . . . . . . . . 127
shishi_encticketpart_endtime_set . . . . . . . . . 128
shishi_encticketpart_flags_set . . . . . . . . . . . . 127
shishi_encticketpart_get_key . . . . . . . . . . . . . . 126
shishi_encticketpart_key_set . . . . . . . . . . . . . . 127
shishi_encticketpart_transited_set . . . . . . . 128

shishi_error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 shishi_error_clear. . . . . . . . . . . . . . . . . . . . . . . . . . 223

shishi_error_outputtype . . . . . . . . . . . . . . . . . . . . 223

shishi_error_printf . . . . . . . . . . . . . . . . . . . . . . . . 223 shishi_error_set . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

shishi_error_set_outputtype . . . . . . . . . . . . . . . 223
shishi_generalize_ctime . . . . . . . . . . . . . . . . . . . . 206

shishi_generalize_now . . . . . . . . . . . . . . . . . . . . . . 206 shishi_generalize_time . . . . . . . . . . . . . . . . . . . . . 205 shishi_hmac_md5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 shishi_hmac_sha1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195

shishi_hostkeys_default_file . . . . . . . . . . . . . . 179
shishi_hostkeys_default_file_set . . . . . . . . . 180
shishi_hostkeys_for_localservice . . . . . . . . . 181
shishi_hostkeys_for_localservicerealm . . . . 180
shishi_hostkeys_for_server . . . . . . . . . . . . . . . . 180
shishi_hostkeys_for_serverrealm. . . . . . . . . . . 180

shishi_info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 shishi_init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 shishi_init_server . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

shishi_init_server_with_paths . . . . . . . . . . . . . . 61

shishi_init_with_paths . . . . . . . . . . . . . . . . . . . . . . 61 shishi_kdc_check_nonce . . . . . . . . . . . . . . . . . . . . . 133 shishi_kdc_copy_cname . . . . . . . . . . . . . . . . . . . . . . 132 shishi_kdc_copy_crealm . . . . . . . . . . . . . . . . . . . . . 132 shishi_kdc_copy_nonce . . . . . . . . . . . . . . . . . . . . . . 133 shishi_kdc_process. . . . . . . . . . . . . . . . . . . . . . . . . . 134

shishi_kdcrep_add_enc_part . . . . . . . . . . . . . . . . 149
shishi_kdcrep_clear_padata . . . . . . . . . . . . . . . . 149
shishi_kdcrep_client_set. . . . . . . . . . . . . . . . . . . 148
shishi_kdcrep_cname_set . . . . . . . . . . . . . . . . . . . . 147
shishi_kdcrep_crealm_set. . . . . . . . . . . . . . . . . . . 147
shishi_kdcrep_from_file . . . . . . . . . . . . . . . . . . . . 147
shishi_kdcrep_get_enc_part_etype . . . . . . . . . 148

                                                shishi_kdcrep_get_ticket. . . . . . . . . . . . . . . . . . . 148

shishi_kdcrep_parse . . . . . . . . . . . . . . . . . . . . . . . . 147 shishi_kdcrep_print . . . . . . . . . . . . . . . . . . . . . . . . 146 shishi_kdcrep_read. . . . . . . . . . . . . . . . . . . . . . . . . . 147 shishi_kdcrep_save. . . . . . . . . . . . . . . . . . . . . . . . . . 146

                                                shishi_kdcrep_set_enc_part . . . . . . . . . . . . . . . . 149
                                                shishi_kdcrep_set_ticket. . . . . . . . . . . . . . . . . . . 148

shishi_kdcrep_to_file . . . . . . . . . . . . . . . . . . . . . . 146

                                                shishi_kdcreq_add_padata. . . . . . . . . . . . . . . . . . . 145
                                                shishi_kdcreq_add_padata_preauth . . . . . . . . . 145
                                                shishi_kdcreq_add_padata_tgs . . . . . . . . . . . . . . 145
                                                shishi_kdcreq_allow_postdate_p . . . . . . . . . . . . 141
                                                shishi_kdcreq_clear_padata . . . . . . . . . . . . . . . . 144

shishi_kdcreq_client . . . . . . . . . . . . . . . . . . . . . . . 137

                                                shishi_kdcreq_disable_

transited_check_p . . . . . . . . . . . . . . . . . . . . . . . . 142

                                                shishi_kdcreq_enc_tkt_in_skey_p. . . . . . . . . . . 143

shishi_kdcreq_etype . . . . . . . . . . . . . . . . . . . . . . . . 139

                                                shishi_kdcreq_forwardable_p . . . . . . . . . . . . . . . 140
                                                shishi_kdcreq_forwarded_p. . . . . . . . . . . . . . . . . . 140
                                                shishi_kdcreq_from_file . . . . . . . . . . . . . . . . . . . . 136
                                                shishi_kdcreq_get_padata. . . . . . . . . . . . . . . . . . . 144
                                                shishi_kdcreq_get_padata_tgs . . . . . . . . . . . . . . 144
                                                shishi_kdcreq_nonce_set . . . . . . . . . . . . . . . . . . . . 136

shishi_kdcreq_options . . . . . . . . . . . . . . . . . . . . . . 140

                                                shishi_kdcreq_options_add. . . . . . . . . . . . . . . . . . 144
                                                shishi_kdcreq_options_set. . . . . . . . . . . . . . . . . . 143

shishi_kdcreq_parse . . . . . . . . . . . . . . . . . . . . . . . . 136

                                                shishi_kdcreq_postdated_p. . . . . . . . . . . . . . . . . . 141

shishi_kdcreq_print . . . . . . . . . . . . . . . . . . . . . . . . 135

                                                shishi_kdcreq_proxiable_p. . . . . . . . . . . . . . . . . . 140

shishi_kdcreq_proxy_p . . . . . . . . . . . . . . . . . . . . . . 141 shishi_kdcreq_read. . . . . . . . . . . . . . . . . . . . . . . . . . 136 shishi_kdcreq_realm . . . . . . . . . . . . . . . . . . . . . . . . 138 shishi_kdcreq_renew_p . . . . . . . . . . . . . . . . . . . . . . 143

                                                shishi_kdcreq_renewable_ok_p . . . . . . . . . . . . . . 142
                                                shishi_kdcreq_renewable_p. . . . . . . . . . . . . . . . . . 142

shishi_kdcreq_save. . . . . . . . . . . . . . . . . . . . . . . . . . 135 shishi_kdcreq_sendrecv . . . . . . . . . . . . . . . . . . . . . 131

                                                shishi_kdcreq_sendrecv_hint . . . . . . . . . . . . . . . 131

shishi_kdcreq_server . . . . . . . . . . . . . . . . . . . . . . . 138

                                                shishi_kdcreq_set_cname . . . . . . . . . . . . . . . . . . . . 137
                                                shishi_kdcreq_set_etype . . . . . . . . . . . . . . . . . . . . 139
                                                shishi_kdcreq_set_realm . . . . . . . . . . . . . . . . . . . . 138
                                                shishi_kdcreq_set_sname . . . . . . . . . . . . . . . . . . . . 138

shishi_kdcreq_till. . . . . . . . . . . . . . . . . . . . . . . . . . 139 shishi_kdcreq_tillc . . . . . . . . . . . . . . . . . . . . . . . . 139 shishi_kdcreq_to_file . . . . . . . . . . . . . . . . . . . . . . 135

                                                shishi_kdcreq_validate_p. . . . . . . . . . . . . . . . . . . 143

shishi_key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 shishi_key_copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 shishi_key_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 shishi_key_from_base64 . . . . . . . . . . . . . . . . . . . . . 175 shishi_key_from_name . . . . . . . . . . . . . . . . . . . . . . . 176 shishi_key_from_random . . . . . . . . . . . . . . . . . . . . . 176 shishi_key_from_string . . . . . . . . . . . . . . . . . . . . . 176 shishi_key_from_value . . . . . . . . . . . . . . . . . . . . . . 175 shishi_key_length . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 shishi_key_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 shishi_key_principal . . . . . . . . . . . . . . . . . . . . . . . 172

shishi_key_principal_set. . . . . . . . . . . . . . . . . . . 172

shishi_key_random . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 shishi_key_realm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 shishi_key_realm_set . . . . . . . . . . . . . . . . . . . . . . . 172 shishi_key_timestamp . . . . . . . . . . . . . . . . . . . . . . . 174

shishi_key_timestamp_set. . . . . . . . . . . . . . . . . . . 174

shishi_key_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 shishi_key_type_set . . . . . . . . . . . . . . . . . . . . . . . . 173 shishi_key_value . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 shishi_key_value_set . . . . . . . . . . . . . . . . . . . . . . . 173 shishi_key_version. . . . . . . . . . . . . . . . . . . . . . . . . . 173 shishi_key_version_set . . . . . . . . . . . . . . . . . . . . . 173 shishi_keys. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 shishi_keys_add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 shishi_keys_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177

shishi_keys_for_
localservicerealm_in_file . . . . . . . . . . . . . . . 179
shishi_keys_for_server_in_file . . . . . . . . . . . . 179
shishi_keys_for_serverrealm_in_file . . . . . . 178

shishi_keys_from_file . . . . . . . . . . . . . . . . . . . . . . 178 shishi_keys_nth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 shishi_keys_print . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 shishi_keys_remove. . . . . . . . . . . . . . . . . . . . . . . . . . 177 shishi_keys_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 shishi_keys_to_file . . . . . . . . . . . . . . . . . . . . . . . . 178 shishi_krberror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 shishi_krberror_build . . . . . . . . . . . . . . . . . . . . . . 163 shishi_krberror_client . . . . . . . . . . . . . . . . . . . . . 164

shishi_krberror_client_set . . . . . . . . . . . . . . . . 165

shishi_krberror_crealm . . . . . . . . . . . . . . . . . . . . . 163 shishi_krberror_ctime . . . . . . . . . . . . . . . . . . . . . . 166

shishi_krberror_ctime_set. . . . . . . . . . . . . . . . . . 167

shishi_krberror_cusec . . . . . . . . . . . . . . . . . . . . . . 167

shishi_krberror_cusec_set. . . . . . . . . . . . . . . . . . 167

shishi_krberror_der . . . . . . . . . . . . . . . . . . . . . . . . 163 shishi_krberror_edata . . . . . . . . . . . . . . . . . . . . . . 170

shishi_krberror_errorcode. . . . . . . . . . . . . . . . . . 169
shishi_krberror_errorcode_fast . . . . . . . . . . . . 169
shishi_krberror_errorcode_message . . . . . . . . 171
shishi_krberror_errorcode_set . . . . . . . . . . . . . 169

shishi_krberror_etext . . . . . . . . . . . . . . . . . . . . . . 169

shishi_krberror_from_file. . . . . . . . . . . . . . . . . . 163
shishi_krberror_message . . . . . . . . . . . . . . . . . . . . 171
shishi_krberror_methoddata . . . . . . . . . . . . . . . . 170

shishi_krberror_parse . . . . . . . . . . . . . . . . . . . . . . 162

shishi_krberror_pretty_print . . . . . . . . . . . . . . 171

shishi_krberror_print . . . . . . . . . . . . . . . . . . . . . . 161 shishi_krberror_read . . . . . . . . . . . . . . . . . . . . . . . 162 shishi_krberror_realm . . . . . . . . . . . . . . . . . . . . . . 165

shishi_krberror_remove_cname . . . . . . . . . . . . . . 165
shishi_krberror_remove_crealm . . . . . . . . . . . . . 164
shishi_krberror_remove_ctime . . . . . . . . . . . . . . 167
shishi_krberror_remove_cusec . . . . . . . . . . . . . . 167
shishi_krberror_remove_edata . . . . . . . . . . . . . . 171
shishi_krberror_remove_etext . . . . . . . . . . . . . . 170
shishi_krberror_remove_sname . . . . . . . . . . . . . . 166

shishi_krberror_save . . . . . . . . . . . . . . . . . . . . . . . 162 shishi_krberror_server . . . . . . . . . . . . . . . . . . . . . 165

shishi_krberror_server_set . . . . . . . . . . . . . . . . 166
shishi_krberror_set_cname. . . . . . . . . . . . . . . . . . 164
shishi_krberror_set_crealm . . . . . . . . . . . . . . . . 164
shishi_krberror_set_edata. . . . . . . . . . . . . . . . . . 170
shishi_krberror_set_etext. . . . . . . . . . . . . . . . . . 169
shishi_krberror_set_realm. . . . . . . . . . . . . . . . . . 165
shishi_krberror_set_sname. . . . . . . . . . . . . . . . . . 166

shishi_krberror_stime . . . . . . . . . . . . . . . . . . . . . . 168

shishi_krberror_stime_set. . . . . . . . . . . . . . . . . . 168

shishi_krberror_susec . . . . . . . . . . . . . . . . . . . . . . 168

shishi_krberror_susec_set. . . . . . . . . . . . . . . . . . 168
shishi_krberror_to_file . . . . . . . . . . . . . . . . . . . . 162

shishi_md4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 shishi_md5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 shishi_n_fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 shishi_parse_name . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 shishi_pbkdf2_sha1. . . . . . . . . . . . . . . . . . . . . . . . . . 197

shishi_principal_default. . . . . . . . . . . . . . . . . . . 202
shishi_principal_default_guess . . . . . . . . . . . . 202
shishi_principal_default_set . . . . . . . . . . . . . . 202

shishi_principal_name . . . . . . . . . . . . . . . . . . . . . . 203

shishi_principal_name_realm . . . . . . . . . . . . . . . 203
shishi_principal_name_set. . . . . . . . . . . . . . . . . . 203

shishi_principal_set . . . . . . . . . . . . . . . . . . . . . . . 204 shishi_priv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 shishi_priv_build . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 shishi_priv_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

shishi_priv_enc_part_etype . . . . . . . . . . . . . . . . 100
shishi_priv_encprivpart . . . . . . . . . . . . . . . . . . . . . 98
shishi_priv_encprivpart_der . . . . . . . . . . . . . . . . 98
shishi_priv_encprivpart_der_set . . . . . . . . . . . . 98
shishi_priv_encprivpart_set . . . . . . . . . . . . . . . . 98

shishi_priv_from_file . . . . . . . . . . . . . . . . . . . . . . 100 shishi_priv_key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 shishi_priv_key_set. . . . . . . . . . . . . . . . . . . . . . . . . . 97 shishi_priv_parse . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 shishi_priv_print . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 shishi_priv_priv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 shishi_priv_priv_der . . . . . . . . . . . . . . . . . . . . . . . . 97

shishi_priv_priv_der_set . . . . . . . . . . . . . . . . . . . . 98

shishi_priv_priv_set . . . . . . . . . . . . . . . . . . . . . . . . 97 shishi_priv_process . . . . . . . . . . . . . . . . . . . . . . . . 101 shishi_priv_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 shishi_priv_save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

shishi_priv_set_enc_part. . . . . . . . . . . . . . . . . . . 100

shishi_priv_to_file. . . . . . . . . . . . . . . . . . . . . . . . . . 99 shishi_prompt_password . . . . . . . . . . . . . . . . . . . . . 207

shishi_prompt_password_callback_get . . . . . . 207
shishi_prompt_password_callback_set . . . . . . 207

shishi_random_to_key . . . . . . . . . . . . . . . . . . . . . . . 183 shishi_randomize . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 shishi_realm_default . . . . . . . . . . . . . . . . . . . . . . . 200

shishi_realm_default_guess . . . . . . . . . . . . . . . . 200
shishi_realm_default_set. . . . . . . . . . . . . . . . . . . 200
shishi_realm_for_server . . . . . . . . . . . . . . . . . . . . 201
shishi_realm_for_server_dns . . . . . . . . . . . . . . . 201

shishi_realm_for_server_file . . . . . . . . . . . . . . 201

shishi_resolv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 shishi_resolv_free. . . . . . . . . . . . . . . . . . . . . . . . . . 208 shishi_safe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 shishi_safe_build . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 shishi_safe_cksum . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 shishi_safe_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 shishi_safe_from_file . . . . . . . . . . . . . . . . . . . . . . . 94 shishi_safe_key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 shishi_safe_key_set. . . . . . . . . . . . . . . . . . . . . . . . . . 92 shishi_safe_parse . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 shishi_safe_print . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 shishi_safe_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 shishi_safe_safe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 shishi_safe_safe_der . . . . . . . . . . . . . . . . . . . . . . . . 92

shishi_safe_safe_der_set . . . . . . . . . . . . . . . . . . . . 93

shishi_safe_safe_set . . . . . . . . . . . . . . . . . . . . . . . . 92 shishi_safe_save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 shishi_safe_set_cksum . . . . . . . . . . . . . . . . . . . . . . . 95

shishi_safe_set_user_data. . . . . . . . . . . . . . . . . . . 95

shishi_safe_to_file. . . . . . . . . . . . . . . . . . . . . . . . . . 93 shishi_safe_user_data . . . . . . . . . . . . . . . . . . . . . . . 95 shishi_safe_verify . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 shishi_server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

shishi_server_for_local_service. . . . . . . . . . . 204

shishi_strerror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 shishi_string_to_key . . . . . . . . . . . . . . . . . . . . . . . 183 shishi_tgs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 shishi_tgs_ap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 shishi_tgs_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 shishi_tgs_krberror . . . . . . . . . . . . . . . . . . . . . . . . 122

shishi_tgs_krberror_der . . . . . . . . . . . . . . . . . . . . 122
shishi_tgs_krberror_set . . . . . . . . . . . . . . . . . . . . 122

shishi_tgs_process. . . . . . . . . . . . . . . . . . . . . . . . . . 133 shishi_tgs_rep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 shishi_tgs_rep_build . . . . . . . . . . . . . . . . . . . . . . . 122 shishi_tgs_rep_der. . . . . . . . . . . . . . . . . . . . . . . . . . 121 shishi_tgs_rep_process . . . . . . . . . . . . . . . . . . . . . 121 shishi_tgs_req . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 shishi_tgs_req_build . . . . . . . . . . . . . . . . . . . . . . . 121 shishi_tgs_req_der. . . . . . . . . . . . . . . . . . . . . . . . . . 120 shishi_tgs_req_der_set . . . . . . . . . . . . . . . . . . . . . 120 shishi_tgs_req_process . . . . . . . . . . . . . . . . . . . . . 121 shishi_tgs_req_set. . . . . . . . . . . . . . . . . . . . . . . . . . 120 shishi_tgs_sendrecv . . . . . . . . . . . . . . . . . . . . . . . . 123

shishi_tgs_sendrecv_hint. . . . . . . . . . . . . . . . . . . 123

shishi_tgs_set_realm . . . . . . . . . . . . . . . . . . . . . . . 123

shishi_tgs_set_realmserver . . . . . . . . . . . . . . . . 124

shishi_tgs_set_server . . . . . . . . . . . . . . . . . . . . . . 123 shishi_tgs_tgtkt . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 shishi_tgs_tgtkt_set . . . . . . . . . . . . . . . . . . . . . . . 119 shishi_tgs_tkt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 shishi_tgs_tkt_set. . . . . . . . . . . . . . . . . . . . . . . . . . 123 shishi_tgsrep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 shishi_tgsreq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 shishi_ticket . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

shishi_ticket_add_enc_part . . . . . . . . . . . . . . . . 126
shishi_ticket_get_enc_part_etype . . . . . . . . . 126

                                                shishi_ticket_realm_get . . . . . . . . . . . . . . . . . . . . 125
                                                shishi_ticket_realm_set . . . . . . . . . . . . . . . . . . . . 125

shishi_ticket_server . . . . . . . . . . . . . . . . . . . . . . . 125

                                                shishi_ticket_set_enc_part . . . . . . . . . . . . . . . . 126
                                                shishi_ticket_sname_set . . . . . . . . . . . . . . . . . . . . 125

shishi_time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 shishi_tkt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 shishi_tkt_authctime . . . . . . . . . . . . . . . . . . . . . . . 111 shishi_tkt_client . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 shishi_tkt_client_p . . . . . . . . . . . . . . . . . . . . . . . . 104 shishi_tkt_clientrealm . . . . . . . . . . . . . . . . . . . . . 104

                                                shishi_tkt_clientrealm_p. . . . . . . . . . . . . . . . . . . 105

shishi_tkt_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

                                                shishi_tkt_enckdcreppart. . . . . . . . . . . . . . . . . . . 103
                                                shishi_tkt_enckdcreppart_set . . . . . . . . . . . . . . 103
                                                shishi_tkt_encticketpart. . . . . . . . . . . . . . . . . . . 103
                                                shishi_tkt_encticketpart_set . . . . . . . . . . . . . . 103

shishi_tkt_endctime . . . . . . . . . . . . . . . . . . . . . . . . 111 shishi_tkt_expired_p . . . . . . . . . . . . . . . . . . . . . . . 112 shishi_tkt_flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 shishi_tkt_flags_add . . . . . . . . . . . . . . . . . . . . . . . 106 shishi_tkt_flags_set . . . . . . . . . . . . . . . . . . . . . . . 106

                                                shishi_tkt_forwardable_p. . . . . . . . . . . . . . . . . . . 106

shishi_tkt_forwarded_p . . . . . . . . . . . . . . . . . . . . . 106

                                                shishi_tkt_hw_authent_p . . . . . . . . . . . . . . . . . . . . 109

shishi_tkt_initial_p . . . . . . . . . . . . . . . . . . . . . . . 108 shishi_tkt_invalid_p . . . . . . . . . . . . . . . . . . . . . . . 108 shishi_tkt_kdcrep . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 shishi_tkt_key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 shishi_tkt_key_set. . . . . . . . . . . . . . . . . . . . . . . . . . 104 shishi_tkt_keytype. . . . . . . . . . . . . . . . . . . . . . . . . . 110

                                                shishi_tkt_keytype_fast . . . . . . . . . . . . . . . . . . . . 110

shishi_tkt_keytype_p . . . . . . . . . . . . . . . . . . . . . . . 110

                                                shishi_tkt_lastreq_pretty_print. . . . . . . . . . . 112

shishi_tkt_lastreqc . . . . . . . . . . . . . . . . . . . . . . . . 111 shishi_tkt_match_p . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

                                                shishi_tkt_may_postdate_p. . . . . . . . . . . . . . . . . . 107
                                                shishi_tkt_ok_as_delegate_p . . . . . . . . . . . . . . . 110

shishi_tkt_postdated_p . . . . . . . . . . . . . . . . . . . . . 107

                                                shishi_tkt_pre_authent_p. . . . . . . . . . . . . . . . . . . 109
                                                shishi_tkt_pretty_print . . . . . . . . . . . . . . . . . . . . 112

shishi_tkt_proxiable_p . . . . . . . . . . . . . . . . . . . . . 107 shishi_tkt_proxy_p. . . . . . . . . . . . . . . . . . . . . . . . . . 107 shishi_tkt_realm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 shishi_tkt_renew_tillc . . . . . . . . . . . . . . . . . . . . . 111 shishi_tkt_renewable_p . . . . . . . . . . . . . . . . . . . . . 108 shishi_tkt_server . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 shishi_tkt_server_p . . . . . . . . . . . . . . . . . . . . . . . . 105 shishi_tkt_startctime . . . . . . . . . . . . . . . . . . . . . . 111 shishi_tkt_ticket . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 shishi_tkt_ticket_set . . . . . . . . . . . . . . . . . . . . . . 102

                                                shishi_tkt_transited_policy_checked_p . . . . 109
                                                shishi_tkt_valid_at_time_p . . . . . . . . . . . . . . . . 111

shishi_tkt_valid_now_p . . . . . . . . . . . . . . . . . . . . . 112 shishi_tkt2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 shishi_tkts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 shishi_tkts_add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 shishi_tkts_default. . . . . . . . . . . . . . . . . . . . . . . . . . 65

shishi_tkts_default_file . . . . . . . . . . . . . . . . . . . . 65
shishi_tkts_default_file_guess . . . . . . . . . . . . . 64
shishi_tkts_default_file_set . . . . . . . . . . . . . . . 65

shishi_tkts_done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 shishi_tkts_expire . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 shishi_tkts_find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

shishi_tkts_find_for_clientserver . . . . . . . . . 69
shishi_tkts_find_for_server . . . . . . . . . . . . . . . . 69

shishi_tkts_from_file . . . . . . . . . . . . . . . . . . . . . . . 67 shishi_tkts_get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

shishi_tkts_get_for_clientserver. . . . . . . . . . . 70
shishi_tkts_get_for_localservicepasswd . . . . 70
shishi_tkts_get_for_server. . . . . . . . . . . . . . . . . . 70

shishi_tkts_get_tgs. . . . . . . . . . . . . . . . . . . . . . . . . . 69 shishi_tkts_get_tgt. . . . . . . . . . . . . . . . . . . . . . . . . . 69 shishi_tkts_new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 shishi_tkts_nth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 shishi_tkts_print . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

shishi_tkts_print_for_service . . . . . . . . . . . . . . 67

shishi_tkts_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 shishi_tkts_remove . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 shishi_tkts_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 shishi_tkts_to_file. . . . . . . . . . . . . . . . . . . . . . . . . . 67 shishi_tkts_write . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 shishi_verbose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 shishi_verify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 shishi_warn. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

                                                shishi_x509ca_default_file . . . . . . . . . . . . . . . . 198
                                                shishi_x509ca_default_file_guess . . . . . . . . . 198
                                                shishi_x509ca_default_file_set . . . . . . . . . . . . 198
                                                shishi_x509cert_default_file . . . . . . . . . . . . . . 199
                                                shishi_x509cert_default_file_guess . . . . . . . 198
                                                shishi_x509cert_default_file_set . . . . . . . . . 199
                                                shishi_x509key_default_file . . . . . . . . . . . . . . . 199
                                                shishi_x509key_default_file_guess . . . . . . . . 199
                                                shishi_x509key_default_file_set. . . . . . . . . . . 199

T

this in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 today in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 tomorrow in date strings . . . . . . . . . . . . . . . . . . . . . . . 53

W

week in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

Y

year in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 yesterday in date strings . . . . . . . . . . . . . . . . . . . . . . 53

Concept Index

3

3DES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

A

abbreviations for months . . . . . . . . . . . . . . . . . . . . . . . 51 AES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 AES-CCM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 AIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 anonymous tls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 Application Programming Interface (API) . . . . . . 57 ARCFOUR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 authenticated tls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Authentication. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Authentication header . . . . . . . . . . . . . . . . . . . . . . . . . 35 Authentication path . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Authenticator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Authorization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 authors of parse_datetime. . . . . . . . . . . . . . . . . . . . . 56 Autoconf tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

B

basic authorization. . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 beginning of time, for POSIX . . . . . . . . . . . . . . . . . . 54 Bellovin, Steven M. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Berets, Jim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Berry, K. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

C

calendar date item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Capability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 case, ignored in dates . . . . . . . . . . . . . . . . . . . . . . . . . . 50 certificate authority (CA) . . . . . . . . . . . . . . . . . . . . . . 28 Ciphertext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Client . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 client authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 combined date and time of day item. . . . . . . . . . . . 52 comments, in dates . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Compiling your application. . . . . . . . . . . . . . . . . . . . . 58 concurrent writers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 configuration file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Configure tests. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Contributing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 copying information. . . . . . . . . . . . . . . . . . . . . . . . . . . 253 credential cache, file format of. . . . . . . . . . . . . . . . . 250 Credentials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

D

database definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Database interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 date and time of day format, ISO 8601 . . . . . . . . . 52 date format, ISO 8601 . . . . . . . . . . . . . . . . . . . . . . . . . 51 date input formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 day of week item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Debian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 9 DES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Diffie Hellman key exchange. . . . . . . . . . . . . . . . . . . . 27 displacement of dates . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Download. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

E

Eggert, Paul . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 eklogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 Encryption Type (etype) . . . . . . . . . . . . . . . . . . . . . . . 36 End-user Shishi usage . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Epoch, for POSIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Error Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

F

fail over. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 FDL, GNU Free Documentation License . . . . . . 253 FreeBSD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

G

general date syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Generic Security Service . . . . . . . . . . . . . . . . . . . . . . 233 GNUTLS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 GSS-API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 GSSLib . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233

H

Hacking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 High Availability. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 HP-UX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

I

Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 IPSEC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 IRIX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ISO 8601 date and time of day format . . . . . . . . . . 52 ISO 8601 date format . . . . . . . . . . . . . . . . . . . . . . . . . . 51 items in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

Concept Index 269

K

k5login authorization . . . . . . . . . . . . . . . . . . . . . . . . . 246 KCMDV0.2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 KCMDV0.3. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 KDC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Kerberos. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Kerberos Ticket. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Key Version Number (kvno) . . . . . . . . . . . . . . . . . . . 36 keytab, file format of. . . . . . . . . . . . . . . . . . . . . . . . . . 248

L

language, in dates. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 LDAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 leap seconds . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50, 51, 55

M

MacKenzie, David . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 MacOS X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Mandrake . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 master server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Meyering, Jim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 minutes, time zone correction by . . . . . . . . . . . . . . . 52 month names in date strings . . . . . . . . . . . . . . . . . . . 51 months, written-out. . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Motorola Coldfire . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

N

NetBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 NFS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 numbers, written-out. . . . . . . . . . . . . . . . . . . . . . . . . . . 50

O

OpenBSD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ordinal numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

P

Pinard, F. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Plaintext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Principal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Principal identifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Protocol Extensions. . . . . . . . . . . . . . . . . . . . . . . . . . . 236 pure numbers in date strings . . . . . . . . . . . . . . . . . . . 54

R

RedHat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 RedHat Advanced Server. . . . . . . . . . . . . . . . . . . . . . . . 9 relative items in date strings . . . . . . . . . . . . . . . . . . . 53 remote databases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Reporting Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 rsh and rlogin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 rsync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

S

Salz, Rich. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Seal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 secondary server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Secret key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 server authentication. . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Session key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Shisa API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 Solaris . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 specifying user database. . . . . . . . . . . . . . . . . . . . . . . . 44 SQL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 STARTTLS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Sub-session key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 SuSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 SuSE Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

T

Telnet encryption . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 Ticket . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 time of day item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 time zone correction . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 time zone item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50, 52 tls resume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 TLS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Tru64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

U

uClibc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 uClinux. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 user database definition . . . . . . . . . . . . . . . . . . . . . . . . 44

X

X.509 authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

Short Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 User Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3 Administration Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4 Reference Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 5 Programming Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 6 Acknowledgements. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 A Criticism of Kerberos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 B Protocol Extensions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 C Copying Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 Function and Data Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Getting Started. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Features and Status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.4 Cryptographic Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.5 Supported Platforms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.6 Getting help. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.7 Downloading and Installing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.8 Bug Reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 1.9 Contributing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

2 User Manual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

2.1 Proxiable and Proxy Tickets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.2 Forwardable and Forwarded Tickets. . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

3 Administration Manual . . . . . . . . . . . . . . . . . . . . . . . . 18

3.1 Introduction to Shisa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.2 Configuring Shisa. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.3 Using Shisa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.4 Starting Shishid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.5 Configuring DNS for KDC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

3.5.1 DNS vs. Kerberos - Case Sensitivity of Realm Names . . . . . 25 3.5.2 Overview - KDC location information . . . . . . . . . . . . . . . . . . . . . 25 3.5.3 Example - KDC location information . . . . . . . . . . . . . . . . . . . . . 26 3.5.4 Security considerations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

3.6 Kerberos via TLS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6.1 Setting up TLS resume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6.2 Setting up Anonymous TLS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.6.3 Setting up X.509 authenticated TLS . . . . . . . . . . . . . . . . . . . . . . 28

3.6.3.1 Create a Kerberos Certificate Authority . . . . . . . . . . . . . . 28 3.6.3.2 Create a Kerberos KDC Certificate . . . . . . . . . . . . . . . . . . 29 3.6.3.3 Create a Kerberos Client Certificate. . . . . . . . . . . . . . . . . . 30 3.6.3.4 Starting KDC with X.509 authentication support. . . . . 31

3.7 Multiple servers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 3.8 Developer information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

4 Reference Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

4.1 Environmental Assumptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.2 Glossary of terms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.3 Realm and Principal Naming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

4.3.1 Realm Names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.3.2 Principal Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.3.2.1 Name of server principals . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.3.2.2 Name of the TGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

4.3.3 Choosing a principal with which to communicate . . . . . . . . . . 40 4.3.4 Principal Name Form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

4.4 Shishi Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.4.1 ‘default-realm’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.4.2 ‘default-principal’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.4.3 ‘client-kdc-etypes’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

4.4.4 ‘verbose’, ‘verbose-asn1’, ‘verbose-noise’,
  ‘verbose-crypto’, ‘verbose-crypto-noise’ . . . . . . . . . . . . . . . . . . . 42

4.4.5 ‘realm-kdc’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.4.6 ‘server-realm’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.4.7 ‘kdc-timeout’, ‘kdc-retries’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.4.8 ‘stringprocess’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.4.9 ‘ticket-life’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.4.10 ‘renew-life’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

4.5 Shisa Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

4.5.1 ‘db’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 4.6 Parameters for shishi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 4.7 Parameters for shishid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 4.8 Parameters for shisa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 4.9 Environment variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.10 Date input formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

4.10.1 General date syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.10.2 Calendar date items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 4.10.3 Time of day items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 4.10.4 Time zone items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 4.10.5 Combined date and time of day items . . . . . . . . . . . . . . . . . . . . 52 4.10.6 Day of week items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 4.10.7 Relative items in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 4.10.8 Pure numbers in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.10.9 Seconds since the Epoch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.10.10 Specifying time zone rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 4.10.11 Authors of parse_datetime . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

5 Programming Manual . . . . . . . . . . . . . . . . . . . . . . . . . 57

5.1 Preparation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.1.1 Header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.1.2 Initialization. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.1.3 Version Check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.1.4 Building the source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.1.5 Autoconf tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

5.1.5.1 Autoconf test via ‘pkg-config’ . . . . . . . . . . . . . . . . . . . . . . 58 5.1.5.2 Standalone Autoconf test using Libtool . . . . . . . . . . . . . . 59 5.1.5.3 Standalone Autoconf test . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

5.2 Initialization Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 5.3 Ticket Set Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 5.4 AP-REQ and AP-REP Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 5.5 SAFE and PRIV Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 5.6 Ticket Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 5.7 AS Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 5.8 TGS Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 5.9 Ticket (ASN.1) Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 5.10 AS/TGS Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 5.11 Authenticator Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 5.12 KRB-ERROR Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 5.13 Cryptographic Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 5.14 X.509 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 5.15 Utility Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 5.16 ASN.1 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 5.17 Error Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

5.17.1 Error Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 5.17.2 Error Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

5.18 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 5.19 Kerberos Database Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 5.20 Generic Security Service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233

6 Acknowledgements. . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

Appendix A Criticism of Kerberos . . . . . . . . . . . . 235

Appendix B Protocol Extensions . . . . . . . . . . . . . . 236

B.1 STARTTLS protected KDC exchanges . . . . . . . . . . . . . . . . . . . . . . . 236 B.1.1 TCP/IP transport with TLS upgrade (STARTTLS) . . . . . 236 B.1.2 Extensible typed hole based on reserved high bit. . . . . . . . . 237 B.1.3 STARTTLS requested by client (extension mode 1) . . . . . . 237 B.1.4 STARTTLS request accepted by server (extension mode 2) . . 237 B.1.5 Proceeding after successful TLS negotiation . . . . . . . . . . . . . 237 B.1.6 Proceeding after failed TLS negotiation. . . . . . . . . . . . . . . . . . 238 B.1.7 Interaction with KDC addresses in DNS . . . . . . . . . . . . . . . . . 238 B.1.8 Using TLS authentication logic in Kerberos. . . . . . . . . . . . . . 238 B.1.9 Security considerations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238

B.2 Telnet encryption with AES-CCM. . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 B.2.1 Command Names and Codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 B.2.2 Command Meanings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 B.2.3 Implementation Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 B.2.4 Integration with the AUTHENTICATION telnet option . . 240 B.2.5 Security Considerations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241

B.2.5.1 Telnet Encryption Protocol Security Considerations . . 241 B.2.5.2 AES-CCM Security Considerations . . . . . . . . . . . . . . . . . 241

B.2.6 Acknowledgments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 B.3 Kerberized rsh and rlogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242

B.3.1 Establish connection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 B.3.2 Kerberos identification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 B.3.3 Kerberos authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 B.3.4 Extended authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 B.3.5 Window size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 B.3.6 End of authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 B.3.7 Encryption. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 B.3.8 KCMDV0.3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 B.3.9 MIT/Heimdal authorization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246

B.4 Key as initialization vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 B.5 The Keytab Binary File Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 B.6 The Credential Cache Binary File Format . . . . . . . . . . . . . . . . . . . . 250

Appendix C Copying Information . . . . . . . . . . . . . 253

C.1 GNU Free Documentation License. . . . . . . . . . . . . . . . . . . . . . . . . . . . 253

Function and Data Index . . . . . . . . . . . . . . . . . . . . . . . . 261

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268